Effective AWK Programming

A User’s Guide for GNU Awk Edition 1.0.3 February 1997

Arnold D. Robbins

“To boldly go where no man has gone before” is a Registered Trademark of Paramount Pictures Corporation.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991, 92, 93, 96, 97 Free Software Foundation, Inc.

This is Edition 1.0.3 of Effective AWK Programming, for the 3.0.3 (or later) version of the GNU implementation of AWK.

Published jointly by:

Specialized Systems Consultants, Inc. (SSC) Free Software Foundation
PO Box 55549  59 Temple Place — Suite 330
Seattle, WA 98155 USA  Boston, MA 02111-1307 USA
Phone: +1-206-782-7733  Phone: +1-617-542-5942
Fax: +1-206-782-7191  Fax: +1-617-542-2652

E-mail: sales@ssc.com  E-mail: gnu@prep.ai.mit.edu
URL: http://www.ssc.com/  URL: http://www.fsf.org/

ISBN 1-57831-000-8

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

Cover art by Amy Wells Wood.

To Miriam, for making me complete. To Chana, for the joy you bring us. To Rivka, for the exponential increase. To Nachum, for the added dimension.

Preface 1

Preface

This book teaches you about the awk language and how you can use it effectively. You should already be familiar with basic system commands, such as cat and ls,^{1} and basic shell facilities, such as Input/Output (I/O) redirection and pipes.

Implementations of the awk language are available for many different computing environments. This book, while describing the awk language in general, also describes a particular implementation of awk called gawk (which stands for “GNU Awk”). gawk runs on a broad range of Unix systems, ranging from 80386 PC-based computers, up through large scale systems, such as Crays. gawk has also been ported to MS-DOS and OS/2 PC’s, Atari and Amiga micro-computers, and VMS.

History of awk and gawk

The name awk comes from the initials of its designers: Alfred V. Aho, Peter J. Weinberger, and Brian W. Kernighan. The original version of awk was written in 1977 at AT&T Bell Laboratories. In 1985 a new version made the programming language more powerful, introducing user-defined functions, multiple input streams, and computed regular expressions. This new version became generally available with Unix System V Release 3.1. The version in System V Release 4 added some new features and also cleaned up the behavior in some of the “dark corners” of the language. The specification for awk in the POSIX Command Language and Utilities standard further clarified the language based on feedback from both the gawk designers, and the original Bell Labs awk designers.

The GNU implementation, gawk, was written in 1986 by Paul Rubin and Jay Fenlason, with advice from Richard Stallman. John Woods contributed parts of the code as well. In 1988 and 1989, David Trueman, with help from Arnold Robbins, thoroughly reworked gawk for compatibility with the newer awk. Current development focuses on bug fixes, performance improvements, standards compliance, and occasionally, new features.

The GNU Project and This Book

The Free Software Foundation (FSF) is a non-profit organization dedicated to the production and distribution of freely distributable software.

1 These commands are available on POSIX compliant systems, as well as on traditional Unix based systems. If you are using some other operating system, you still need to be familiar with the ideas of I/O redirection and pipes.

It was founded by Richard M. Stallman, the author of the original Emacs editor. GNU Emacs is the most widely used version of Emacs today.

The GNU project is an on-going effort on the part of the Free Software Foundation to create a complete, freely distributable, POSIX compliant computing environment. (GNU stands for “GNU’s not Unix”.) The FSF uses the “GNU General Public License” (or GPL) to ensure that source code for their software is always available to the end user. A copy of the GPL is included for your reference (see [GNU GENERAL PUBLIC LICENSE], page 307). The GPL applies to the C language source code for gawk.

A shell, an editor (Emacs), highly portable optimizing C, C++, and Objective-C compilers, a symbolic debugger, and dozens of large and small utilities (such as gawk), have all been completed and are freely available. As of this writing (early 1997), the GNU operating system kernel (the HURD), has been released, but is still in an early stage of development.

Until the GNU operating system is more fully developed, you should consider using Linux, a freely distributable, Unix-like operating system for 80386, DEC Alpha, Sun SPARC and other systems. There are many books on Linux. One freely available one is Linux Installation and Getting Started, by Matt Welsh. Many Linux distributions are available, often in computer stores or bundled on CD-ROM with books about Linux. (There are three other freely available, Unix-like operating systems for 80386 and other systems, NetBSD, FreeBSD,and OpenBSD. All are based on the 4.4-Lite Berkeley Software Distribution, and they use recent versions of gawk for their versions of awk.)

This book you are reading now is actually free. The information in it is freely available to anyone, the machine readable source code for the book comes with gawk, and anyone may take this book to a copying machine and make as many copies of it as they like. (Take a moment to check the copying permissions on the Copyright page.)

If you paid money for this book, what you actually paid for was the book’s nice printing and binding, and the publisher’s associated costs to produce it. We have made an effort to keep these costs reasonable; most people would prefer a bound book to over 330 pages of photo-copied text that would then have to be held in a loose-leaf binder (not to mention the time and labor involved in doing the copying). The same is true of producing this book from the machine readable source; the retail price is only slightly more than the cost per page of printing it on a laser printer.

This book itself has gone through several previous, preliminary editions. I started working on a preliminary draft of The GAWK Manual, by Diane Close, Paul Rubin, and Richard Stallman in the fall of 1988. It was around 90 pages long, and barely described the original, “old” version of awk. After substantial revision, the first version of the The GAWK Manual to be released was Edition 0.11 Beta in October of 1989. The manual then underwent more substantial revision for Edition 0.13 of December 1991. David Trueman, Pat Rankin, and Michal Jaegermann contributed sections of the manual for Edition 0.13. That edition was published by the FSF as a bound book early in 1992. Since then there have been several minor revisions, notably Edition 0.14 of November 1992 that was published by the FSF in January of 1993, and Edition 0.16 of August 1993.

Edition 1.0 of Effective AWK Programming represents a significant reworking of The GAWK Manual, with much additional material. The FSF and I agree that I am now the primary author. I also felt that it needed a more descriptive title.

Effective AWK Programming will undoubtedly continue to evolve. An electronic version comes with the gawk distribution from the FSF. If you find an error in this book, please report it! See Section B.7 [Reporting Problems and Bugs], page 290, for information on submitting problem reports electronically, or write to me in care of the FSF.

Acknowledgements

I would like to acknowledge Richard M. Stallman, for his vision of a better world, and for his courage in founding the FSF and starting the GNU project.

The initial draft of The GAWK Manual had the following acknowledgements:

Many people need to be thanked for their assistance in producing this manual. Jay Fenlason contributed many ideas and sample programs. Richard Mlynarik and Robert Chassell gave helpful comments on drafts of this manual. The paper A Supplemental Document for awk by John W. Pierce of the Chemistry Department at UC San Diego, pinpointed several issues relevant both to awk implementation and to this manual, that would otherwise have escaped us. The following people provided many helpful comments on Edition 0.13 of

The GAWK Manual: Rick Adams, Michael Brennan, Rich Burridge, Diane Close, Christopher (“Topher”) Eliot, Michael Lijewski, Pat Rankin, Miriam Robbins, and Michal Jaegermann.

The following people provided many helpful comments for Edition 1.0 of Effective AWK Programming: Karl Berry, Michael Brennan, Darrel Hankerson, Michal Jaegermann, Michael Lijewski, and Miriam Robbins. Pat Rankin, Michal Jaegermann, Darrel Hankerson and Scott Deifik updated their respective sections for Edition 1.0.

Robert J. Chassell provided much valuable advice on the use of Texinfo. He also deserves special thanks for convincing me not to title this book How To Gawk Politely. Karl Berry helped significantly with the TEX part of Texinfo.

David Trueman deserves special credit; he has done a yeoman job of evolving gawk so that it performs well, and without bugs. Although he is no longer involved with gawk, working with him on this project was a significant pleasure.

Scott Deifik, Darrel Hankerson, Kai Uwe Rommel, Pat Rankin, and Michal Jaegermann (in no particular order) are long time members of the gawk “crack portability team.” Without their hard work and help, gawk would not be nearly the fine program it is today. It has been and continues to be a pleasure working with this team of fine people.

Jeffrey Friedl provided invaluable help in tracking down a number of last minute problems with regular expressions in gawk 3.0.

David and I would like to thank Brian Kernighan of Bell Labs for invaluable assistance during the testing and debugging of gawk, and for help in clarifying numerous points about the language. We could not have done nearly as good a job on either gawk or its documentation without his help.

I would like to thank Marshall and Elaine Hartholz of Seattle, and Dr. Bert and Rita Schreiber of Detroit for large amounts of quiet vacation time in their homes, which allowed me to make significant progress on this book and on gawk itself. Phil Hughes of SSC contributed in a very important way by loaning me his laptop Linux system, not once, but twice, allowing me to do a lot of work while away from home.

Finally, I must thank my wonderful wife, Miriam, for her patience through the many versions of this project, for her proof-reading, and for sharing me with the computer. I would like to thank my parents for their love, and for the grace with which they raised and educated me. I also must acknowledge my gratitude to G-d, for the many opportunities He has sent my way, as well as for the gifts He has given me with which to take advantage of those opportunities.

Arnold Robbins Atlanta, Georgia February, 1997

Chapter 1: Introduction 5

1 Introduction

If you are like many computer users, you would frequently like to make changes in various text files wherever certain patterns appear, or extract data from parts of certain lines while discarding the rest. To write a program to do this in a language such as C or Pascal is a time-consuming inconvenience that may take many lines of code. The job may be easier with awk.

The awk utility interprets a special-purpose programming language that makes it possible to handle simple data-reformatting jobs with just a few lines of code.

The GNU implementation of awk is called gawk; it is fully upward compatible with the System V Release 4 version of awk. gawk is also upward compatible with the POSIX specification of the awk language. This means that all properly written awk programs should work with gawk. Thus, we usually don’t distinguish between gawk and other awk implementations.

Using awk you can:

• manage small, personal databases

• generate reports

• validate data

• produce indexes, and perform other document preparation tasks

• even experiment with algorithms that can be adapted later to other computer languages

1.1 Using This Book

The term awk refers to a particular program, and to the language you use to tell this program what to do. When we need to be careful, we call the program “the awk utility” and the language “the awk language.” The term gawk refers to a version of awk developed as part the GNU project. The purpose of this book is to explain both the awk language and how to run the awk utility.

The main purpose of the book is to explain the features of awk, as defined in the POSIX standard. It does so in the context of one particular implementation, gawk. While doing so, it will also attempt to describe important differences between gawk and other awk implementations. Finally, any gawk features that are not in the POSIX standard for awk will be noted.

This book has the difficult task of being both tutorial and reference. If you are a novice, feel free to skip over details that seem too complex. You should also ignore the many cross references; they are for the expert user, and for the on-line Info version of the document.

The term awk program refers to a program written by you in the awk programming language.

See Chapter 2 [Getting Started with awk], page 9, for the bare essentials you need to know to start using awk.

Some useful “one-liners” are included to give you a feel for the awk language (see Chapter 3 [Useful One Line Programs], page 21).

Many sample awk programs have been provided for you (see Chapter 15 [A Library of awk Functions], page 167; also see Chapter 16 [Practical awk Programs], page 203).

The entire awk language is summarized for quick reference in Appendix A [gawk Summary], page 257. Look there if you just need to refresh your memory about a particular feature.

If you find terms that you aren’t familiar with, try looking them up in the glossary (see Appendix D [Glossary], page 299).

Most of the time complete awk programs are used as examples, but in some of the more advanced sections, only the part of the awk program that illustrates the concept being described is shown.

While this book is aimed principally at people who have not been exposed to awk, there is a lot of information here that even the awk expert should find useful. In particular, the description of POSIX awk, and the example programs in Chapter 15 [A Library of awk Functions], page 167, and Chapter 16 [Practical awk Programs], page 203, should be of interest.

Dark Corners

Who opened that window shade?!?

Count Dracula

Until the POSIX standard (and The Gawk Manual), many features of awk were either poorly documented, or not documented at all. Descriptions of such features (often called “dark corners”) are noted in this book with “(d.c.)”. They also appear in the index under the heading “dark corner.”

1.2 Typographical Conventions

This book is written using Texinfo, the GNU documentation formatting language. A single Texinfo source file is used to produce both the printed and on-line versions of the documentation. Because of this, the typographical conventions are slightly different than in other books you may have read.

Examples you would type at the command line are preceded by the common shell primary and secondary prompts, ‘$’ and ‘>’. Output from the command is preceded by the glyph “ ⊣ ”. This typically represents the command’s standard output. Error messages, and other output on the command’s standard error, are preceded by the glyph “ error ”. For example:

    $ echo hi on stdout
    ⊣ hi on stdout
    $ echo hello on stderr 1>&2
     error hello on stderr

In the text, command names appear in this font, while code segments appear in the same font and quoted, ‘like this’. Some things will be emphasized like this, and if a point needs to be made strongly, it will be done like this. The first occurrence of a new term is usually its definition, and appears in the same font as the previous occurrence of “definition” in this sentence. File names are indicated like this: ‘/path/to/ourfile’.

Characters that you type at the keyboard look like this. In particular, there are special characters called “control characters.” These are characters that you type by holding down both the CONTROL key and another key, at the same time. For example, a Control-d is typed by first pressing and holding the CONTROL key, next pressing the d key, and finally releasing both keys.

1.3 Data Files for the Examples

Many of the examples in this book take their input from two sample data files. The first, called ‘BBS-list’, represents a list of computer bulletin board systems together with information about those systems. The second data file, called ‘inventory-shipped’, contains information about shipments on a monthly basis. In both files, each line is considered to be one record.

In the file ‘BBS-list’, each record contains the name of a computer bulletin board, its phone number, the board’s baud rate(s), and a code for the number of hours it is operational. An ‘A’ in the last column means the board operates 24 hours a day. A ‘B’ in the last column means the board operates evening and weekend hours, only. A ‘C’ means the board operates only on weekends.

    aardvark     555-5553     1200/300          B
    alpo-net     555-3412     2400/1200/300     A
    barfly       555-7685     1200/300          A
    bites        555-1675     2400/1200/300     A
    camelot      555-0542     300               C
    core         555-2912     1200/300          C
    fooey        555-1234     2400/1200/300     B
    foot         555-6699     1200/300          B
    macfoo       555-6480     1200/300          A
    sdace        555-3430     2400/1200/300     A
    sabafoo      555-2127     1200/300          C

The second data file, called ‘inventory-shipped’, represents information about shipments during the year. Each record contains the month of the year, the number of green crates shipped, the number of red boxes shipped, the number of orange bags shipped, and the number of blue packages shipped, respectively. There are 16 entries, covering the 12 months of one year and four months of the next year.

Jan  13  25  15 115
Feb  15  32  24 226
Mar  15  24  34 228
Apr  31  52  63 420
May  16  34  29 208
Jun  31  42  75 492
Jul  24  34  67 436
Aug  15  34  47 316
Sep  13  55  37 277
Oct  29  54  68 525
Nov  20  87  82 577
Dec  17  35  61 401
Jan  21  36  64 620
Feb  26  58  80 652
Mar  24  75  70 495
Apr  21  70  74 514

Chapter 2: Getting Started with awk 9

2 Getting Started with awk

The basic function of awk is to search files for lines (or other units of text) that contain certain patterns. When a line matches one of the patterns, awk performs specified actions on that line. awk keeps processing input lines in this way until the end of the input files are reached.

Programs in awk are different from programs in most other languages, because awk programs are data-driven; that is, you describe the data you wish to work with, and then what to do when you find it. Most other languages are procedural; you have to describe, in great detail, every step the program is to take. When working with procedural languages, it is usually much harder to clearly describe the data your program will process. For this reason, awk programs are often refreshingly easy to both write and read.

When you run awk, you specify an awk program that tells awk what to do. The program consists of a series of rules. (It may also contain function definitions, an advanced feature which we will ignore for now. See Chapter 13 [User-defined Functions], page 151.) Each rule specifies one pattern to search for, and one action to perform when that pattern is found.

Syntactically, a rule consists of a pattern followed by an action. The action is enclosed in curly braces to separate it from the pattern. Rules are usually separated by newlines. Therefore, an awk program looks like this:

pattern { action } pattern { action } …

2.1 A Rose By Any Other Name

The awk language has evolved over the years. Full details are provided in Chapter 17 [The Evolution of the awk Language], page 251. The language described in this book is often referred to as “new awk.”

Because of this, many systems have multiple versions of awk. Some systems have an awk utility that implements the original version of the awk language, and a nawk utility for the new version. Others have an oawk for the “old awk” language, and plain awk for the new one. Still others only have one version, usually the new one.^{1}

All in all, this makes it difficult for you to know which version of awk you should run when writing your programs. The best advice we can give here is to check your local documentation. Look for awk, oawk, and nawk, as well as for gawk. Chances are, you will have some version of new awk on your

1 Often, these systems use gawk for their awk implementation! system, and that is what you should use when running your programs. (Of course, if you’re reading this book, chances are good that you have gawk!)

Throughout this book, whenever we refer to a language feature that should be available in any complete implementation of POSIX awk, we simply use the term awk. When referring to a feature that is specific to the GNU implementation, we use the term gawk.

2.2 How to Run awk Programs

There are several ways to run an awk program. If the program is short, it is easiest to include it in the command that runs awk, like this:

awk 'program' input-file1 input-file2 … where program consists of a series of patterns and actions, as described earlier. (The reason for the single quotes is described below, in Section 2.2.1 [One-shot Throw-away awk Programs], page 10.)

When the program is long, it is usually more convenient to put it in a file and run it with a command like this:

awk -f program-file input-file1 input-file2 …

2.2.1 One-shot Throw-away awk Programs

Once you are familiar with awk, you will often type in simple programs the moment you want to use them. Then you can write the program as the first argument of the awk command, like this:

awk 'program' input-file1 input-file2 … where program consists of a series of patterns and actions, as described earlier.

This command format instructs the shell, or command interpreter, to start awk and use the program to process records in the input file(s). There are single quotes around program so that the shell doesn’t interpret any awk characters as special shell characters. They also cause the shell to treat all of program as a single argument for awk and allow program to be more than one line long.

This format is also useful for running short or medium-sized awk programs from shell scripts, because it avoids the need for a separate file for the awk program. A self-contained shell script is more reliable since there are no other files to misplace.

Chapter 3 [Useful One Line Programs], page 21, presents several short, self-contained programs.

As an interesting side point, the command

    awk '/foo/' files . . .

is essentially the same as

    egrep foo files . . .

2.2.2 Running awk without Input Files

You can also run awk without any input files. If you type the command line:

awk 'program' then awk applies the program to the standard input, which usually means whatever you type on the terminal. This continues until you indicate end-of-file by typing Control-d. (On other operating systems, the end-of-file character may be different. For example, on OS/2 and MS-DOS, it is Controlz.)

For example, the following program prints a friendly piece of advice (from Douglas Adams’ The Hitchhiker’s Guide to the Galaxy), to keep you from worrying about the complexities of computer programming (‘BEGIN’ is a feature we haven’t discussed yet).

    $ awk "BEGIN { print \"Don't Panic!\" }"
    ⊣ Don't Panic!

This program does not read any input. The ‘\’ before each of the inner double quotes is necessary because of the shell’s quoting rules, in particular because it mixes both single quotes and double quotes.

This next simple awk program emulates the cat utility; it copies whatever you type at the keyboard to its standard output. (Why this works is explained shortly.)

    $ awk '{ print }'
    Now is the time for all good men
    ⊣ Now is the time for all good men
    to come to the aid of their country.
    ⊣ to come to the aid of their country.
    Four score and seven years ago, ...
    ⊣ Four score and seven years ago, ...
    What, me worry?
    ⊣ What, me worry?
    Control-d

2.2.3 Running Long Programs

Sometimes your awk programs can be very long. In this case it is more convenient to put the program into a separate file. To tell awk to use that file for its program, you type:

awk -f source-file input-file1 input-file2 …

The ‘-f’ instructs the awk utility to get the awk program from the file source-file. Any file name can be used for source-file. For example, you could put the program:

    BEGIN { print "Don't Panic!" }

into the file ‘advice’. Then this command:

    awk -f advice

does the same thing as this one:

    awk "BEGIN { print \"Don't Panic!\" }"

which was explained earlier (see Section 2.2.2 [Running awk without Input Files], page 11). Note that you don’t usually need single quotes around the file name that you specify with ‘-f’, because most file names don’t contain any of the shell’s special characters. Notice that in ‘advice’, the awk program did not have single quotes around it. The quotes are only needed for programs that are provided on the awk command line.

If you want to identify your awk program files clearly as such, you can add the extension ‘.awk’ to the file name. This doesn’t affect the execution of the awk program, but it does make “housekeeping” easier.

2.2.4 Executable awk Programs

Once you have learned awk, you may want to write self-contained awk scripts, using the ‘#!’ script mechanism. You can do this on many Unix systems^{2} (and someday on the GNU system).

For example, you could update the file ‘advice’ to look like this:

    #! /bin/awk -f
    BEGIN    { print "Don't Panic!" }

After making this file executable (with the chmod utility), you can simply type ‘advice’ at the shell, and the system will arrange to run awk^{3} as if you

had typed ‘awk -f advice’.
    $ advice
    ⊣ Don't Panic!

2 The ‘#!’ mechanism works on Linux systems, Unix systems derived from Berkeley Unix, System V Release 4, and some System V Release 3 systems.

3 The line beginning with ‘#!’ lists the full file name of an interpreter to be run, and an optional initial command line argument to pass to that interpreter. The operating system then runs the interpreter with the given argument and the full argument list of the executed program. The first argument in the list is the full file name of the awk program. The rest of the argument list will either be options to awk, or data files, or both.

Self-contained awk scripts are useful when you want to write a program which users can invoke without their having to know that the program is written

in awk.

Some older systems do not support the ‘#!’ mechanism. You can get a similar effect using a regular shell script. It would look something like this:

    : The colon ensures execution by the standard shell.
    awk 'program' "$@"

Using this technique, it is vital to enclose the program in single quotes to protect it from interpretation by the shell. If you omit the quotes, only a shell wizard can predict the results.

The "$@" causes the shell to forward all the command line arguments to the awk program, without interpretation. The first line, which starts with a colon, is used so that this shell script will work even if invoked by a user who uses the C shell. (Not all older systems obey this convention, but many do.)

2.2.5 Comments in awk Programs

A comment is some text that is included in a program for the sake of human readers; it is not really part of the program. Comments can explain what the program does, and how it works. Nearly all programming languages have provisions for comments, because programs are typically hard to understand without their extra help.

In the awk language, a comment starts with the sharp sign character, ‘#’, and continues to the end of the line. The ‘#’ does not have to be the first character on the line. The awk language ignores the rest of a line following a sharp sign. For example, we could have put the following into ‘advice’:

    # This program prints a nice friendly message.  It helps
    # keep novice users from being afraid of the computer.
    BEGIN    { print "Don't Panic!" }

You can put comment lines into keyboard-composed throw-away awk programs also, but this usually isn’t very useful; the purpose of a comment is to help you or another person understand the program at a later time.

2.3 A Very Simple Example

The following command runs a simple awk program that searches the input file ‘BBS-list’ for the string of characters: ‘foo’. (A string of characters is usually called a string. The term string is perhaps based on similar usage in English, such as “a string of pearls,” or, “a string of cars in a train.”)

    awk '/foo/ { print $0 }' BBS-list

When lines containing ‘foo’ are found, they are printed, because ‘print $0’ means print the current line. (Just ‘print’ by itself means the same thing, so we could have written that instead.)

You will notice that slashes, ‘/’, surround the string ‘foo’ in the awk program. The slashes indicate that ‘foo’ is a pattern to search for. This type of pattern is called a regular expression, and is covered in more detail later (see Chapter 4 [Regular Expressions], page 23). The pattern is allowed to match parts of words. There are single-quotes around the awk program so that the shell won’t interpret any of it as special shell characters.

Here is what this program prints:

    $ awk '/foo/ { print $0 }' BBS-list
    ⊣ fooey        555-1234     2400/1200/300     B
    ⊣ foot         555-6699     1200/300          B
    ⊣ macfoo       555-6480     1200/300          A
    ⊣ sabafoo      555-2127     1200/300          C

In an awk rule, either the pattern or the action can be omitted, but not both. If the pattern is omitted, then the action is performed for every input line. If the action is omitted, the default action is to print all lines that match the pattern.

Thus, we could leave out the action (the print statement and the curly braces) in the above example, and the result would be the same: all lines matching the pattern ‘foo’ would be printed. By comparison, omitting the print statement but retaining the curly braces makes an empty action that does nothing; then no lines would be printed.

2.4 An Example with Two Rules

The awk utility reads the input files one line at a time. For each line, awk tries the patterns of each of the rules. If several patterns match then several actions are run, in the order in which they appear in the awk program. If no patterns match, then no actions are run.

After processing all the rules (perhaps none) that match the line, awk reads the next line (however, see Section 9.7 [The next Statement], page 111, and also see Section 9.8 [The nextfile Statement], page 112). This continues until the end of the file is reached.

For example, the awk program:

    /12/  { print $0 }
    /21/  { print $0 }

contains two rules. The first rule has the string ‘12’ as the pattern and ‘print $0’ as the action. The second rule has the string ‘21’ as the pattern and also has ‘print $0’ as the action. Each rule’s action is enclosed in its own pair of braces.

This awk program prints every line that contains the string ‘12’ or the string ‘21’. If a line contains both strings, it is printed twice, once by each rule.

This is what happens if we run this program on our two sample data files,

‘BBS-list’ and ‘inventory-shipped’, as shown here:
    $ awk '/12/ { print $0 }
    >      /21/ { print $0 }' BBS-list inventory-shipped
    ⊣ aardvark     555-5553     1200/300          B
    ⊣ alpo-net     555-3412     2400/1200/300     A
    ⊣ barfly       555-7685     1200/300          A
    ⊣ bites        555-1675     2400/1200/300     A
    ⊣ core         555-2912     1200/300          C
    ⊣ fooey        555-1234     2400/1200/300     B
    ⊣ foot         555-6699     1200/300          B
    ⊣ macfoo       555-6480     1200/300          A
    ⊣ sdace        555-3430     2400/1200/300     A
    ⊣ sabafoo      555-2127     1200/300          C
    ⊣ sabafoo      555-2127     1200/300          C
    ⊣ Jan  21  36  64 620
    ⊣ Apr  21  70  74 514

Note how the line in ‘BBS-list’ beginning with ‘sabafoo’ was printed twice, once for each rule.

2.5 A More Complex Example

Here is an example to give you an idea of what typical awk programs do. This example shows how awk can be used to summarize, select, and rearrange the output of another utility. It uses features that haven’t been covered yet, so don’t worry if you don’t understand all the details.

    ls -lg | awk '$6 == "Nov" { sum += $5 }
                 END { print sum }'

This command prints the total number of bytes in all the files in the current directory that were last modified in November (of any year). (In the C shell you would need to type a semicolon and then a backslash at the end of the first line; in a POSIX-compliant shell, such as the Bourne shell or Bash, the GNU Bourne-Again shell, you can type the example as shown.)

The ‘ls -lg’ part of this example is a system command that gives you a listing of the files in a directory, including file size and the date the file was last modified. Its output looks like this:

    -rw-r--r--  1 arnold   user   1933 Nov  7 13:05 Makefile
    -rw-r--r--  1 arnold   user  10809 Nov  7 13:03 gawk.h
    -rw-r--r--  1 arnold   user    983 Apr 13 12:14 gawk.tab.h

    -rw-r--r--  1 arnold   user  31869 Jun 15 12:20 gawk.y
    -rw-r--r--  1 arnold   user  22414 Nov  7 13:03 gawk1.c
    -rw-r--r--  1 arnold   user  37455 Nov  7 13:03 gawk2.c
    -rw-r--r--  1 arnold   user  27511 Dec  9 13:07 gawk3.c
    -rw-r--r--  1 arnold   user   7989 Nov  7 13:03 gawk4.c

The first field contains read-write permissions, the second field contains the number of links to the file, and the third field identifies the owner of the file. The fourth field identifies the group of the file. The fifth field contains the size of the file in bytes. The sixth, seventh and eighth fields contain the month, day, and time, respectively, that the file was last modified. Finally, the ninth field contains the name of the file.

The ‘$6 == "Nov"’ in our awk program is an expression that tests whether the sixth field of the output from ‘ls -lg’ matches the string ‘Nov’. Each time a line has the string ‘Nov’ for its sixth field, the action ‘sum += $5’ is performed. This adds the fifth field (the file size) to the variable sum. As a result, when awk has finished reading all the input lines, sum is the sum of the sizes of files whose lines matched the pattern. (This works because awk variables are automatically initialized to zero.)

After the last line of output from ls has been processed, the END rule is executed, and the value of sum is printed. In this example, the value of sum would be 80600.

These more advanced awk techniques are covered in later sections (see Section 8.2 [Overview of Actions], page 102). Before you can move on to more advanced awk programming, you have to know how awk interprets your input and displays your output. By manipulating fields and using print statements, you can produce some very useful and impressive looking reports.

2.6 awk Statements Versus Lines

Most often, each line in an awk program is a separate statement or separate rule, like this:

    awk '/12/  { print $0 }
         /21/  { print $0 }' BBS-list inventory-shipped

However, gawk will ignore newlines after any of the following:

    ,    {    ?    :    ||    &&    do    else

A newline at any other point is considered the end of the statement. (Splitting lines after ‘?’ and ‘:’ is a minor gawk extension. The ‘?’ and ‘:’ referred to here is the three operand conditional expression described in Section 7.12 [Conditional Expressions], page 92.)

If you would like to split a single statement into two lines at a point where a newline would terminate it, you can continue it by ending the first line with a backslash character, ‘\’. The backslash must be the final character on the line to be recognized as a continuation character. This is allowed absolutely anywhere in the statement, even in the middle of a string or regular expression. For example:

    awk '/This regular expression is too long, so continue it\
     on the next line/ { print $1 }'

We have generally not used backslash continuation in the sample programs in this book. Since in gawk there is no limit on the length of a line, it is never strictly necessary; it just makes programs more readable. For this same reason, as well as for clarity, we have kept most statements short in the sample programs presented throughout the book. Backslash continuation is most useful when your awk program is in a separate source file, instead of typed in on the command line. You should also note that many awk implementations are more particular about where you may use backslash continuation. For example, they may not allow you to split a string constant using backslash continuation. Thus, for maximal portability of your awk programs, it is best not to split your lines in the middle of a regular expression or a string.

Caution: backslash continuation does not work as described above with the C shell. Continuation with backslash works for awk programs in files, and also for one-shot programs provided you are using a POSIX-compliant shell, such as the Bourne shell or Bash, the GNU Bourne-Again shell. But the C shell (csh) behaves differently! There, you must use two backslashes in a row, followed by a newline. Note also that when using the C shell, every newline in your awk program must be escaped with a backslash. To illustrate:

    % awk 'BEGIN { \
    ?   print \\
    ?       "hello, world" \
    ? }'
    ⊣ hello, world

Here, the ‘%’ and ‘?’ are the C shell’s primary and secondary prompts, analogous to the standard shell’s ‘$’ and ‘>’.

awk is a line-oriented language. Each rule’s action has to begin on the same line as the pattern. To have the pattern and action on separate lines, you must use backslash continuation—there is no other way.

Note that backslash continuation and comments do not mix. As soon as awk sees the ‘#’ that starts a comment, it ignores everything on the rest of the line. For example:

    $ gawk 'BEGIN { print "dont panic" # a friendly \
    >                                    BEGIN rule
    > }'
     error gawk: cmd. line:2:                BEGIN rule
     error gawk: cmd. line:2:                ^parse error

Here, it looks like the backslash would continue the comment onto the next line. However, the backslash-newline combination is never even noticed, since it is “hidden” inside the comment. Thus, the ‘BEGIN’ is noted as a syntax error.

When awk statements within one rule are short, you might want to put more than one of them on a line. You do this by separating the statements with a semicolon, ‘;’.

This also applies to the rules themselves. Thus, the previous program could have been written:

    /12/ { print $0 } ; /21/ { print $0 }

Note: the requirement that rules on the same line must be separated with a semicolon was not in the original awk language; it was added for consistency with the treatment of statements within an action.

2.7 Other Features of awk

The awk language provides a number of predefined, or built-in variables, which your programs can use to get information from awk. There are other variables your program can set to control how awk processes your data.

In addition, awk provides a number of built-in functions for doing common computational and string related operations.

As we develop our presentation of the awk language, we introduce most of the variables and many of the functions. They are defined systematically in Chapter 10 [Built-in Variables], page 115, and Chapter 12 [Built-in Functions], page 133.

2.8 When to Use awk

You might wonder how awk might be useful for you. Using utility programs, advanced patterns, field separators, arithmetic statements, and other selection criteria, you can produce much more complex output. The awk language is very useful for producing reports from large amounts of raw data, such as summarizing information from the output of other utility programs like ls. (See Section 2.5 [A More Complex Example], page 15.)

Programs written with awk are usually much smaller than they would be in other languages. This makes awk programs easy to compose and use. Often, awk programs can be quickly composed at your terminal, used once, and thrown away. Since awk programs are interpreted, you can avoid the (usually lengthy) compilation part of the typical edit-compile-test-debug cycle of software development.

Complex programs have been written in awk, including a complete retargetable assembler for eight-bit microprocessors (see Appendix D [Glossary], page 299, for more information) and a microcode assembler for a special purpose Prolog computer. However, awk’s capabilities are strained by tasks of such complexity.

If you find yourself writing awk scripts of more than, say, a few hundred lines, you might consider using a different programming language. Emacs Lisp is a good choice if you need sophisticated string or pattern matching capabilities. The shell is also good at string and pattern matching; in addition, it allows powerful use of the system utilities. More conventional languages, such as C, C++, and Lisp, offer better facilities for system programming and for managing the complexity of large programs. Programs in these languages may require more lines of source code than the equivalent awk programs, but they are easier to maintain and usually run more efficiently.

20 Effective AWK Programming

Chapter 3: Useful One Line Programs 21

3 Useful One Line Programs

Many useful awk programs are short, just a line or two. Here is a collection of useful, short programs to get you started. Some of these programs contain constructs that haven’t been covered yet. The description of the program will give you a good idea of what is going on, but please read the rest of the book to become an awk expert!

Most of the examples use a data file named ‘data’. This is just a placeholder; if you were to use these programs yourself, you would substitute your own file names for ‘data’.

awk '{ if (length($0) > max) max = length($0) }
END { print max }' data

This program prints the length of the longest input line.

awk 'length($0) > 80' data

This program prints every line that is longer than 80 characters. The sole rule has a relational expression as its pattern, and has no action (so the default action, printing the record, is used).

expand data | awk '{ if (x < length()) x = length() }
         END { print "maximum line length is " x }'

This program prints the length of the longest line in ‘data’. The input is processed by the expand program to change tabs into spaces, so the widths compared are actually the right-margin columns.

awk 'NF > 0' data

This program prints every line that has at least one field. This is an easy way to delete blank lines from a file (or rather, to create a new file similar to the old file but from which the blank lines have been deleted).

awk 'BEGIN { for (i = 1; i <= 7; i++)
     print int(101 * rand()) }'

This program prints seven random numbers from zero to 100, inclusive.

ls -lg files | awk '{ x += $5 } ; END { print "total bytes: " x }'

This program prints the total number of bytes used by files.

ls -lg files | awk '{ x += $5 }
       END { print "total K-bytes: " (x + 1023)/1024 }'

This program prints the total number of kilobytes used by files.

awk -F: '{ print $1 }' /etc/passwd | sort

This program prints a sorted list of the login names of all users.

awk 'END { print NR }' data

This program counts lines in a file.

awk 'NR % 2 == 0' data

This program prints the even numbered lines in the data file. If you were to use the expression ‘NR % 2 == 1’ instead, it would print the odd numbered lines.

Chapter 4: Regular Expressions 23

4 Regular Expressions

A regular expression, or regexp, is a way of describing a set of strings. Because regular expressions are such a fundamental part of awk programming, their format and use deserve a separate chapter.

A regular expression enclosed in slashes (‘/’) is an awk pattern that matches every input record whose text belongs to that set.

The simplest regular expression is a sequence of letters, numbers, or both. Such a regexp matches any string that contains that sequence. Thus, the regexp ‘foo’ matches any string containing ‘foo’. Therefore, the pattern /foo/ matches any input record containing the three characters ‘foo’, anywhere in the record. Other kinds of regexps let you specify more complicated classes of strings.

Initially, the examples will be simple. As we explain more about how regular expressions work, we will present more complicated examples.

4.1 How to Use Regular Expressions

A regular expression can be used as a pattern by enclosing it in slashes. Then the regular expression is tested against the entire text of each record. (Normally, it only needs to match some part of the text in order to succeed.) For example, this prints the second field of each record that contains the three characters ‘foo’ anywhere in it:

    $ awk '/foo/ { print $2 }' BBS-list
    ⊣ 555-1234
    ⊣ 555-6699
    ⊣ 555-6480
    ⊣ 555-2127

Regular expressions can also be used in matching expressions. These expressions allow you to specify the string to match against; it need not be the entire current input record. The two operators, ‘~’ and ‘!~’, perform regular expression comparisons. Expressions using these operators can be used as patterns or in if, while, for, and do statements.

exp ~ /regexp/ This is true if the expression exp (taken as a string) is matched by regexp. The following example matches, or selects, all input records with the upper-case letter ‘J’ somewhere in the first field:

              $ awk '$1 ~ /J/' inventory-shipped
              ⊣ Jan  13  25  15 115
              ⊣ Jun  31  42  75 492
              ⊣ Jul  24  34  67 436
              ⊣ Jan  21  36  64 620

So does this:

              awk '{ if ($1 ~ /J/) print }' inventory-shipped

exp !~ /regexp/ This is true if the expression exp (taken as a character string) is not matched by regexp. The following example matches, or selects, all input records whose first field does not contain the upper-case letter ‘J’:

              $ awk '$1 !~ /J/' inventory-shipped
              ⊣ Feb  15  32  24 226
              ⊣ Mar  15  24  34 228
              ⊣ Apr  31  52  63 420
              ⊣ May  16  34  29 208

…

When a regexp is written enclosed in slashes, like /foo/, we call it a regexp constant, much like 5.27 is a numeric constant, and "foo" is a string constant.

4.2 Escape Sequences

Some characters cannot be included literally in string constants ("foo") or regexp constants (/foo/). You represent them instead with escape sequences, which are character sequences beginning with a backslash (‘\’).

One use of an escape sequence is to include a double-quote character in a string constant. Since a plain double-quote would end the string, you must use ‘\"’ to represent an actual double-quote character as a part of the string. For example:

    $ awk 'BEGIN { print "He said \"hi!\" to her." }'
    ⊣ He said "hi!" to her.

The backslash character itself is another character that cannot be included normally; you write ‘\\’ to put one backslash in the string or regexp. Thus, the string whose contents are the two characters ‘"’ and ‘\’ must be written "\"\\".

Another use of backslash is to represent unprintable characters such as tab or newline. While there is nothing to stop you from entering most unprintable characters directly in a string constant or regexp constant, they may look ugly.

Here is a table of all the escape sequences used in awk, and what they represent. Unless noted otherwise, all of these escape sequences apply to both string constants and regexp constants.

\\ A literal backslash, ‘\’.

\a The “alert” character, Control-g, ASCII code 7 (BEL).

\b Backspace, Control-h, ASCII code 8 (BS).

\f Formfeed, Control-l, ASCII code 12 (FF).

\n Newline, Control-j, ASCII code 10 (LF).

\r Carriage return, Control-m, ASCII code 13 (CR).

\t Horizontal tab, Control-i, ASCII code 9 (HT).

\v Vertical tab, Control-k, ASCII code 11 (VT).

\nnn The octal value nnn, where nnn are one to three digits between ‘0’ and ‘7’. For example, the code for the ASCII ESC (escape) character is ‘\033’.

\xhh… The hexadecimal value hh, where hh are hexadecimal digits (‘0’ through ‘9’ and either ‘A’ through ‘F’ or ‘a’ through ‘f’). Like the same construct in ANSI C, the escape sequence continues until the first non-hexadecimal digit is seen. However, using more than two hexadecimal digits produces undefined results. (The ‘\x’ escape sequence is not allowed in POSIX awk.)

\/ A literal slash (necessary for regexp constants only). You use this when you wish to write a regexp constant that contains a slash. Since the regexp is delimited by slashes, you need to escape the slash that is part of the pattern, in order to tell awk to keep processing the rest of the regexp.

\" A literal double-quote (necessary for string constants only). You use this when you wish to write a string constant that contains a double-quote. Since the string is delimited by double-quotes, you need to escape the quote that is part of the string, in order to tell awk to keep processing the rest of the string.

In gawk, there are additional two character sequences that begin with backslash that have special meaning in regexps. See Section 4.4 [Additional Regexp Operators Only in gawk], page 31.

In a string constant, what happens if you place a backslash before something that is not one of the characters listed above? POSIX awk purposely leaves this case undefined. There are two choices.

• Strip the backslash out. This is what Unix awk and gawk both do. For example, "a\qc" is the same as "aqc".

• Leave the backslash alone. Some other awk implementations do this. In such implementations, "a\qc" is the same as if you had typed "a\\qc".

In a regexp, a backslash before any character that is not in the above table, and not listed in Section 4.4 [Additional Regexp Operators Only in gawk], page 31, means that the next character should be taken literally, even if it would normally be a regexp operator. E.g., /a\+b/ matches the three characters ‘a+b’.

For complete portability, do not use a backslash before any character not listed in the table above.

Another interesting question arises. Suppose you use an octal or hexadecimal escape to represent a regexp metacharacter (see Section 4.3 [Regular Expression Operators], page 26). Does awk treat the character as literal character, or as a regexp operator?

It turns out that historically, such characters were taken literally (d.c.). However, the POSIX standard indicates that they should be treated as real metacharacters, and this is what gawk does. However, in compatibility mode (see Section 14.1 [Command Line Options], page 159), gawk treats the characters represented by octal and hexadecimal escape sequences literally when used in regexp constants. Thus, /a\52b/ is equivalent to /a\*b/.

To summarize:

1. The escape sequences in the table above are always processed first, for both string constants and regexp constants. This happens very early, as soon as awk reads your program.

2. gawk processes both regexp constants and dynamic regexps (see Section 4.7 [Using Dynamic Regexps], page 35), for the special operators listed in Section 4.4 [Additional Regexp Operators Only in gawk], page 31.

3. A backslash before any other character means to treat that character literally.

4.3 Regular Expression Operators

You can combine regular expressions with the following characters, called regular expression operators, or metacharacters, to increase the power and versatility of regular expressions.

The escape sequences described above in Section 4.2 [Escape Sequences], page 24, are valid inside a regexp. They are introduced by a ‘\’. They are recognized and converted into the corresponding real characters as the very first step in processing regexps.

Here is a table of metacharacters. All characters that are not escape sequences and that are not listed in the table stand for themselves.

\ This is used to suppress the special meaning of a character when

matching. For example:

              \$

matches the character ‘$’.

^This matches the beginning of a string. For example:

    ^@chapter

matches the ‘@chapter’ at the beginning of a string, and can be used to identify chapter beginnings in Texinfo source files. The ‘^’ is known as an anchor, since it anchors the pattern to matching only at the beginning of the string. It is important to realize that ‘^’ does not match the beginning of a line embedded in a string. In this example the condition is not true:

    if ("line1\nLINE 2" ~ /^L/) . . .

$ This is similar to ‘^’, but it matches only at the end of a string.

For example:

    p$

matches a record that ends with a ‘p’. The ‘$’ is also an anchor, and also does not match the end of a line embedded in a string. In this example the condition is not true:

    if ("line1\nLINE 2" ~ /1$/) . . .

. The period, or dot, matches any single character, including the

newline character. For example:

    .P

matches any single character followed by a ‘P’ in a string. Using concatenation we can make a regular expression like ‘U.A’, which matches any three-character sequence that begins with ‘U’ and ends with ‘A’. In strict POSIX mode (see Section 14.1 [Command Line Options], page 159), ‘.’ does not match the nul character, which is a character with all bits equal to zero. Otherwise, nul is just another character. Other versions of awk may not be able to match the nul character.

[…] This is called a character list. It matches any one of the char-

acters that are enclosed in the square brackets. For example:

    [MVX]

matches any one of the characters ‘M’, ‘V’, or ‘X’ in a string. Ranges of characters are indicated by using a hyphen between the beginning and ending characters, and enclosing the whole thing in brackets. For example:

    [0-9]

matches any digit. Multiple ranges are allowed. E.g., the list [A-Za-z0-9] is a common way to express the idea of “all alphanumeric characters.”

To include one of the characters ‘\’, ‘]’, ‘-’ or ‘^’ in a character list, put a ‘\’ in front of it. For example:

    [d\]]

matches either ‘d’, or ‘]’. This treatment of ‘\’ in character lists is compatible with other awk implementations, and is also mandated by POSIX. The regular expressions in awk are a superset of the POSIX specification for Extended Regular Expressions (EREs). POSIX EREs are based on the regular expressions accepted by the traditional egrep utility. Character classes are a new feature introduced in the POSIX standard. A character class is a special notation for describing lists of characters that have a specific attribute, but where the actual characters themselves can vary from country to country and/or from character set to character set. For example, the notion of what is an alphabetic character differs in the USA and in France. A character class is only valid in a regexp inside the brackets of a character list. Character classes consist of ‘[:’, a keyword denoting the class, and ‘:]’. Here are the character classes defined by the POSIX standard.

[:alnum:]

Alphanumeric characters.

[:alpha:]

Alphabetic characters.

[:blank:]

Space and tab characters.

[:cntrl:]

Control characters.

[:digit:]

Numeric characters.

[:graph:]

Characters that are printable and are also visible. (A space is printable, but not visible, while an ‘a’ is both.)

[:lower:]

Lower-case alphabetic characters.

[:print:]

Printable characters (characters that are not control characters.)

[:punct:]

Punctuation characters (characters that are not letter, digits, control characters, or space characters).

[:space:]

Space characters (such as space, tab, and formfeed, to name a few).

[:upper:]

Upper-case alphabetic characters.

[:xdigit:]

Characters that are hexadecimal digits.

For example, before the POSIX standard, to match alphanumeric characters, you had to write /[A-Za-z0-9]/. If your character set had other alphabetic characters in it, this would not match them. With the POSIX character classes, you can write /[[:alnum:]]/, and this will match all the alphabetic and numeric characters in your character set. Two additional special sequences can appear in character lists. These apply to non-ASCII character sets, which can have single symbols (called collating elements) that are represented with more than one character, as well as several characters that are equivalent for collating, or sorting, purposes. (E.g., in French, a plain “e” and a grave-accented “è” are equivalent.)

Collating Symbols A collating symbol is a multi-character collating element enclosed in ‘[.’ and ‘.]’. For example, if ‘ch’ is a collating element, then [[.ch.]] is a regexp that matches this collating element, while [ch] is a regexp that matches either ‘c’ or ‘h’.

Equivalence Classes An equivalence class is a locale-specific name for a list of characters that are equivalent. The name is enclosed in ‘[=’ and ‘=]’. For example, the name ‘e’ might be used to represent all of “e,” “è,” and “é.” In this case, [[=e]] is a regexp that matches any of ‘e’, ‘é’, or ‘è’.

These features are very valuable in non-English speaking locales. Caution: The library functions that gawk uses for regular expression matching currently only recognize POSIX character classes; they do not recognize collating symbols or equivalence classes.

[^…] This is a complemented character list. The first character after the ‘[’ must be a ‘^’. It matches any characters except those in the square brackets. For example:

    [^0-9]

matches any character that is not a digit.

| This is the alternation operator, and it is used to specify alter-

natives. For example:

    ^P|[0-9]

matches any string that matches either ‘^P’ or ‘[0-9]’. This means it matches any string that starts with ‘P’ or contains a digit. The alternation applies to the largest possible regexps on either side. In other words, ‘|’ has the lowest precedence of all the regular expression operators.

(…) Parentheses are used for grouping in regular expressions as in arithmetic. They can be used to concatenate regular expressions containing the alternation operator, ‘|’. For exam-

ple, ‘@(samp|code)\{[^}]+\}’ matches both ‘@code{foo}’ and

‘@samp{bar}’. (These are Texinfo formatting control sequences.)

* This symbol means that the preceding regular expression is to be repeated as many times as necessary to find a match. For example:

    ph*

applies the ‘*’ symbol to the preceding ‘h’ and looks for matches of one ‘p’ followed by any number of ‘h’s. This will also match just ‘p’ if no ‘h’s are present. The ‘*’ repeats the smallest possible preceding expression. (Use parentheses if you wish to repeat a larger expression.) It finds as many repetitions as possible. For example:

    awk '/\(c[ad][ad]*r x\)/ { print }' sample

prints every record in ‘sample’ containing a string of the form ‘(car x)’, ‘(cdr x)’, ‘(cadr x)’, and so on. Notice the escaping of the parentheses by preceding them with backslashes.

+ This symbol is similar to ‘*’, but the preceding expression must

be matched at least once. This means that:

    wh+y

would match ‘why’ and ‘whhy’ but not ‘wy’, whereas ‘wh*y’ would match all three of these strings. This is a simpler way of writing the last ‘*’ example:

    awk '/\(c[ad]+r x\)/ { print }' sample

? This symbol is similar to ‘*’, but the preceding expression can be matched either once or not at all. For example:

    fe?d

will match ‘fed’ and ‘fd’, but nothing else.

{n}
{n,}

{n,m} One or two numbers inside braces denote an interval expression. If there is one number in the braces, the preceding regexp is repeated n times. If there are two numbers separated by a comma, the preceding regexp is repeated n to m times. If there is one number followed by a comma, then the preceding regexp is repeated at least n times.

wh{3}y matches ‘whhhy’ but not ‘why’ or ‘whhhhy’.

wh{3,5}y matches ‘whhhy’ or ‘whhhhy’ or ‘whhhhhy’, only.

wh{2,}y matches ‘whhy’ or ‘whhhy’, and so on.

Interval expressions were not traditionally available in awk. As part of the POSIX standard they were added, to make awk and egrep consistent with each other. However, since old programs may use ‘{’ and ‘}’ in regexp constants, by default gawk does not match interval expressions in regexps. If either ‘--posix’ or ‘--re-interval’ are specified (see Section 14.1 [Command Line Options], page 159), then interval expressions are allowed in regexps.

In regular expressions, the ‘*’, ‘+’, and ‘?’ operators, as well as the braces ‘{’ and ‘}’, have the highest precedence, followed by concatenation, and finally by ‘|’. As in arithmetic, parentheses can change how operators are grouped.

If gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), character classes and interval expressions are not available in regular expressions.

The next section discusses the GNU-specific regexp operators, and provides more detail concerning how command line options affect the way gawk interprets the characters in regular expressions.

4.4 Additional Regexp Operators Only in gawk

GNU software that deals with regular expressions provides a number of additional regexp operators. These operators are described in this section, and are specific to gawk; they are not available in other awk implementations.

Most of the additional operators are for dealing with word matching. For our purposes, a word is a sequence of one or more letters, digits, or underscores (‘_’).

\w This operator matches any word-constituent character, i.e. any letter, digit, or underscore. Think of it as a short-hand for

          [[:alnum:]_].

\W This operator matches any character that is not wordconstituent. Think of it as a short-hand for [^[:alnum:]_].

\< This operator matches the empty string at the beginning of a word. For example, /\<away/ matches ‘away’, but not

          ‘stowaway’.

\> This operator matches the empty string at the end of a word. For example, /stow\>/ matches ‘stow’, but not ‘stowaway’.

\y This operator matches the empty string at either the beginning or the end of a word (the word boundary). For example, ‘\yballs?\y’ matches either ‘ball’ or ‘balls’ as a separate word.

\B This operator matches the empty string within a word. In other words, ‘\B’ matches the empty string that occurs between two word-constituent characters. For example, /\Brat\B/ matches ‘crate’, but it does not match ‘dirty rat’. ‘\B’ is essentially the opposite of ‘\y’.

There are two other operators that work on buffers. In Emacs, a buffer is, naturally, an Emacs buffer. For other programs, the regexp library routines that gawk uses consider the entire string to be matched as the buffer.

For awk, since ‘^’ and ‘$’ always work in terms of the beginning and end of strings, these operators don’t add any new capabilities. They are provided for compatibility with other GNU software.

\` This operator matches the empty string at the beginning of the buffer.

\' This operator matches the empty string at the end of the buffer.

In other GNU software, the word boundary operator is ‘\b’. However, that conflicts with the awk language’s definition of ‘\b’ as backspace, so gawk uses a different letter.

An alternative method would have been to require two backslashes in the GNU operators, but this was deemed to be too confusing, and the current method of using ‘\y’ for the GNU ‘\b’ appears to be the lesser of two evils.

The various command line options (see Section 14.1 [Command Line Options], page 159) control how gawk interprets characters in regexps.

No options In the default case, gawk provide all the facilities of POSIX regexps and the GNU regexp operators described above. However, interval expressions are not supported.

--posix Only POSIX regexps are supported, the GNU operators are not special (e.g., ‘\w’ matches a literal ‘w’). Interval expressions are allowed.

--traditional

Traditional Unix awk regexps are matched. The GNU operators are not special, interval expressions are not available, and neither are the POSIX character classes ([[:alnum:]] and so on). Characters described by octal and hexadecimal escape sequences are treated literally, even if they represent regexp metacharacters.

--re-interval

Allow interval expressions in regexps, even if ‘--traditional’ has been provided.

4.5 Case-sensitivity in Matching

Case is normally significant in regular expressions, both when matching ordinary characters (i.e. not metacharacters), and inside character sets. Thus a ‘w’ in a regular expression matches only a lower-case ‘w’ and not an uppercase ‘W’.

The simplest way to do a case-independent match is to use a character list: ‘[Ww]’. However, this can be cumbersome if you need to use it often; and it can make the regular expressions harder to read. There are two alternatives that you might prefer.

One way to do a case-insensitive match at a particular point in the program is to convert the data to a single case, using the tolower or toupper built-in string functions (which we haven’t discussed yet; see Section 12.3 [Built-in Functions for String Manipulation], page 135). For example:

    tolower($1) ~ /foo/  { . . . }

converts the first field to lower-case before matching against it. This will work in any POSIX-compliant implementation of awk.

Another method, specific to gawk, is to set the variable IGNORECASE to a non-zero value (see Chapter 10 [Built-in Variables], page 115). When IGNORECASE is not zero, all regexp and string operations ignore case. Changing the value of IGNORECASE dynamically controls the case sensitivity of your program as it runs. Case is significant by default because IGNORECASE (like most variables) is initialized to zero.

    x = "aB"
    if (x ~ /ab/) . . . # this test will fail
    IGNORECASE = 1
    if (x ~ /ab/) . . . # now it will succeed

In general, you cannot use IGNORECASE to make certain rules caseinsensitive and other rules case-sensitive, because there is no way to set IGNORECASE just for the pattern of a particular rule. To do this, you must use character lists or tolower. However, one thing you can do only with IGNORECASE is turn case-sensitivity on or off dynamically for all the rules at once.

IGNORECASE can be set on the command line, or in a BEGIN rule (see Section 14.2 [Other Command Line Arguments], page 163; also see Section 8.1.5.1 [Startup and Cleanup Actions], page 100). Setting IGNORECASE from the command line is a way to make a program case-insensitive without having to edit it.

Prior to version 3.0 of gawk, the value of IGNORECASE only affected regexp operations. It did not affect string comparison with ‘==’, ‘!=’, and so on. Beginning with version 3.0, both regexp and string comparison operations are affected by IGNORECASE.

Beginning with version 3.0 of gawk, the equivalences between upper-case and lower-case characters are based on the ISO-8859-1 (ISO Latin-1) character set. This character set is a superset of the traditional 128 ASCII characters, that also provides a number of characters suitable for use with European languages.

The value of IGNORECASE has no effect if gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159). Case is always significant in compatibility mode.

4.6 How Much Text Matches?

Consider the following example:

    echo aaaabcd | awk '{ sub(/a+/, "<A>"); print }'

This example uses the sub function (which we haven’t discussed yet, see Section 12.3 [Built-in Functions for String Manipulation], page 135) to make a change to the input record. Here, the regexp /a+/ indicates “one or more ‘a’ characters,” and the replacement text is ‘<A>’.

The input contains four ‘a’ characters. What will the output be? In other words, how many is “one or more”—will awk match two, three, or all four ‘a’ characters?

The answer is, awk (and POSIX) regular expressions always match the leftmost, longest sequence of input characters that can match. Thus, in this example, all four ‘a’ characters are replaced with ‘<A>’.

    $ echo aaaabcd | awk '{ sub(/a+/, "<A>"); print }'
    ⊣ <A>bcd

For simple match/no-match tests, this is not so important. But when doing regexp-based field and record splitting, and text matching and substitutions with the match, sub, gsub, and gensub functions, it is very important. Understanding this principle is also important for regexp-based record and field splitting (see Section 5.1 [How Input is Split into Records], page 37, and also see Section 5.5 [Specifying How Fields are Separated], page 44).

4.7 Using Dynamic Regexps

The right hand side of a ‘~’ or ‘!~’ operator need not be a regexp constant (i.e. a string of characters between slashes). It may be any expression. The expression is evaluated, and converted if necessary to a string; the contents of the string are used as the regexp. A regexp that is computed in this way is called a dynamic regexp. For example:

    BEGIN { identifier_regexp = "[A-Za-z_][A-Za-z_0-9]+" }
    $0 ~ identifier_regexp    { print }

sets identifier_regexp to a regexp that describes awk variable names, and tests if the input record matches this regexp.

Caution: When using the ‘~’ and ‘!~’ operators, there is a difference between a regexp constant enclosed in slashes, and a string constant enclosed in double quotes. If you are going to use a string constant, you have to understand that the string is in essence scanned twice; the first time when awk reads your program, and the second time when it goes to match the string on the left-hand side of the operator with the pattern on the right. This is true of any string valued expression (such as identifier_regexp above), not just string constants.

What difference does it make if the string is scanned twice? The answer has to do with escape sequences, and particularly with backslashes. To get a backslash into a regular expression inside a string, you have to type two backslashes.

For example, /\*/ is a regexp constant for a literal ‘*’. Only one backslash is needed. To do the same thing with a string, you would have to type "\\*". The first backslash escapes the second one, so that the string actually contains the two characters ‘\’ and ‘*’.

Given that you can use both regexp and string constants to describe regular expressions, which should you use? The answer is “regexp constants,” for several reasons.

1. String constants are more complicated to write, and more difficult to read. Using regexp constants makes your programs less error-prone. Not understanding the difference between the two kinds of constants is a common source of errors.

2. It is also more efficient to use regexp constants: awk can note that you have supplied a regexp and store it internally in a form that makes pattern matching more efficient. When using a string constant, awk must first convert the string into this internal form, and then perform the pattern matching.

3. Using regexp constants is better style; it shows clearly that you intend a regexp match.

Chapter 5: Reading Input Files 37

5 Reading Input Files

In the typical awk program, all input is read either from the standard input (by default the keyboard, but often a pipe from another command) or from files whose names you specify on the awk command line. If you specify input files, awk reads them in order, reading all the data from one before going on to the next. The name of the current input file can be found in the built-in variable FILENAME (see Chapter 10 [Built-in Variables], page 115).

The input is read in units called records, and processed by the rules of your program one record at a time. By default, each record is one line. Each record is automatically split into chunks called fields. This makes it more convenient for programs to work on the parts of a record.

On rare occasions you will need to use the getline command. The getline command is valuable, both because it can do explicit input from any number of files, and because the files used with it do not have to be named on the awk command line (see Section 5.8 [Explicit Input with getline], page 53).

5.1 How Input is Split into Records

The awk utility divides the input for your awk program into records and fields. Records are separated by a character called the record separator. By default, the record separator is the newline character. This is why records are, by default, single lines. You can use a different character for the record separator by assigning the character to the built-in variable RS.

You can change the value of RS in the awk program, like any other variable, with the assignment operator, ‘=’ (see Section 7.7 [Assignment Expressions], page 84). The new record-separator character should be enclosed in quotation marks, which indicate a string constant. Often the right time to do this is at the beginning of execution, before any input has been processed, so that the very first record will be read with the proper separator. To do this, use the special BEGIN pattern (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100). For example:

    awk 'BEGIN { RS = "/" } ; { print $0 }' BBS-list

changes the value of RS to "/", before reading any input. This is a string whose first character is a slash; as a result, records are separated by slashes. Then the input file is read, and the second rule in the awk program (the action with no pattern) prints each record. Since each print statement adds a newline at the end of its output, the effect of this awk program is to copy the input with each slash changed to a newline. Here are the results of running the program on ‘BBS-list’:

$ awk 'BEGIN { RS = "/" } ; { print $0 }' BBS-list
⊣ aardvark     555-5553     1200
⊣ 300          B
⊣ alpo-net     555-3412     2400
⊣ 1200
⊣ 300     A
⊣ barfly       555-7685     1200
⊣ 300          A
⊣ bites        555-1675     2400
⊣ 1200
⊣ 300     A
⊣ camelot      555-0542     300               C
⊣ core         555-2912     1200
⊣ 300          C
⊣ fooey        555-1234     2400
⊣ 1200
⊣ 300     B
⊣ foot         555-6699     1200
⊣ 300          B
⊣ macfoo       555-6480     1200
⊣ 300          A
⊣ sdace        555-3430     2400
⊣ 1200
⊣ 300     A
⊣ sabafoo      555-2127     1200
⊣ 300          C

⊣ Note that the entry for the ‘camelot’ BBS is not split. In the original data file (see Section 1.3 [Data Files for the Examples], page 7), the line looks like this:

camelot      555-0542     300               C

It only has one baud rate; there are no slashes in the record.

Another way to change the record separator is on the command line, using the variable-assignment feature (see Section 14.2 [Other Command Line Arguments], page 163).

awk '{ print $0 }' RS="/" BBS-list

This sets RS to ‘/’ before processing ‘BBS-list’.

Using an unusual character such as ‘/’ for the record separator produces correct behavior in the vast majority of cases. However, the following (extreme) pipeline prints a surprising ‘1’. There is one field, consisting of a newline. The value of the built-in variable NF is the number of fields in the current record.

    $ echo | awk 'BEGIN { RS = "a" } ; { print NF }'
    ⊣ 1

Reaching the end of an input file terminates the current input record, even if the last character in the file is not the character in RS (d.c.).

The empty string, "" (a string of no characters), has a special meaning as the value of RS: it means that records are separated by one or more blank lines, and nothing else. See Section 5.7 [Multiple-Line Records], page 51, for more details.

If you change the value of RS in the middle of an awk run, the new value is used to delimit subsequent records, but the record currently being processed (and records already processed) are not affected.

After the end of the record has been determined, gawk sets the variable RT to the text in the input that matched RS.

The value of RS is in fact not limited to a one-character string. It can be any regular expression (see Chapter 4 [Regular Expressions], page 23). In general, each record ends at the next string that matches the regular expression; the next record starts at the end of the matching string. This general rule is actually at work in the usual case, where RS contains just a newline: a record ends at the beginning of the next matching string (the next newline in the input) and the following record starts just after the end of this string (at the first character of the following line). The newline, since it matches RS, is not part of either record.

When RS is a single character, RT will contain the same single character. However, when RS is a regular expression, then RT becomes more useful; it contains the actual input text that matched the regular expression.

The following example illustrates both of these features. It sets RS equal to a regular expression that matches either a newline, or a series of one or more upper-case letters with optional leading and/or trailing white space (see Chapter 4 [Regular Expressions], page 23).

    $ echo record 1 AAAA record 2 BBBB record 3 |
    > gawk 'BEGIN { RS = "\n|( *[[:upper:]]+ *)" }
    >             { print "Record =", $0, "and RT =", RT }'
    ⊣ Record = record 1 and RT =  AAAA
    ⊣ Record = record 2 and RT =  BBBB
    ⊣ Record = record 3 and RT =

⊣ The final line of output has an extra blank line. This is because the value of RT is a newline, and then the print statement supplies its own terminating newline.

See Section 16.2.8 [A Simple Stream Editor], page 240, for a more useful example of RS as a regexp and RT.

The use of RS as a regular expression and the RT variable are gawk extensions; they are not available in compatibility mode (see Section 14.1 [Command Line Options], page 159). In compatibility mode, only the first character of the value of RS is used to determine the end of the record.

The awk utility keeps track of the number of records that have been read so far from the current input file. This value is stored in a built-in variable called FNR. It is reset to zero when a new file is started. Another built-in variable, NR, is the total number of input records read so far from all data files. It starts at zero but is never automatically reset to zero.

5.2 Examining Fields

When awk reads an input record, the record is automatically separated or parsed by the interpreter into chunks called fields. By default, fields are separated by whitespace, like words in a line. Whitespace in awk means any string of one or more spaces, tabs or newlines;^{1} other characters such as formfeed, and so on, that are considered whitespace by other languages are not considered whitespace by awk.

The purpose of fields is to make it more convenient for you to refer to these pieces of the record. You don’t have to use them—you can operate on the whole record if you wish—but fields are what make simple awk programs so powerful.

To refer to a field in an awk program, you use a dollar-sign, ‘$’, followed by the number of the field you want. Thus, $1 refers to the first field, $2 to the second, and so on. For example, suppose the following is a line of input:

    This seems like a pretty nice example.

Here the first field, or $1, is ‘This’; the second field, or $2, is ‘seems’; and so on. Note that the last field, $7, is ‘example.’. Because there is no space between the ‘e’ and the ‘.’, the period is considered part of the seventh field.

NF is a built-in variable whose value is the number of fields in the current record. awk updates the value of NF automatically, each time a record is read.

No matter how many fields there are, the last field in a record can be represented by $NF. So, in the example above, $NF would be the same as $7, which is ‘example.’. Why this works is explained below (see Section 5.3 [Non-constant Field Numbers], page 41). If you try to reference a field beyond the last one, such as $8 when the record has only seven fields, you get the empty string.

1 In POSIX awk, newlines are not considered whitespace for separating fields.

$0, which looks like a reference to the “zeroth” field, is a special case: it represents the whole input record. $0 is used when you are not interested in fields.

Here are some more examples:

    $ awk '$1 ~ /foo/ { print $0 }' BBS-list
    ⊣ fooey        555-1234     2400/1200/300     B
    ⊣ foot         555-6699     1200/300          B
    ⊣ macfoo       555-6480     1200/300          A
    ⊣ sabafoo      555-2127     1200/300          C

This example prints each record in the file ‘BBS-list’ whose first field contains the string ‘foo’. The operator ‘~’ is called a matching operator (see Section 4.1 [How to Use Regular Expressions], page 23); it tests whether a string (here, the field $1) matches a given regular expression.

By contrast, the following example looks for ‘foo’ in the entire record and prints the first field and the last field for each input record containing a match.

    $ awk '/foo/ { print $1, $NF }' BBS-list
    ⊣ fooey B
    ⊣ foot B
    ⊣ macfoo A
    ⊣ sabafoo C

5.3 Non-constant Field Numbers

The number of a field does not need to be a constant. Any expression in the awk language can be used after a ‘$’ to refer to a field. The value of the expression specifies the field number. If the value is a string, rather than a number, it is converted to a number. Consider this example:

    awk '{ print $NR }'

Recall that NR is the number of records read so far: one in the first record, two in the second, etc. So this example prints the first field of the first record, the second field of the second record, and so on. For the twentieth record, field number 20 is printed; most likely, the record has fewer than 20 fields, so this prints a blank line.

Here is another example of using expressions as field numbers:

    awk '{ print $(2*2) }' BBS-list

awk must evaluate the expression ‘(2*2)’ and use its value as the number of the field to print. The ‘*’ sign represents multiplication, so the expression ‘2*2’ evaluates to four. The parentheses are used so that the multiplication is done before the ‘$’ operation; they are necessary whenever there is a binary operator in the field-number expression. This example, then, prints the hours of operation (the fourth field) for every line of the file ‘BBS-list’.

(All of the awk operators are listed, in order of decreasing precedence, in Section 7.14 [Operator Precedence (How Operators Nest)], page 94.)

If the field number you compute is zero, you get the entire record. Thus, $(2-2) has the same value as $0. Negative field numbers are not allowed; trying to reference one will usually terminate your running awk program. (The POSIX standard does not define what happens when you reference a negative field number. gawk will notice this and terminate your program. Other awk implementations may behave differently.)

As mentioned in Section 5.2 [Examining Fields], page 40, the number of fields in the current record is stored in the built-in variable NF (also see Chapter 10 [Built-in Variables], page 115). The expression $NF is not a special feature: it is the direct consequence of evaluating NF and using its value as a field number.

5.4 Changing the Contents of a Field

You can change the contents of a field as seen by awk within an awk program; this changes what awk perceives as the current input record. (The actual input is untouched; awk never modifies the input file.)

Consider this example and its output:

    $ awk '{ $3 = $2 - 10; print $2, $3 }' inventory-shipped
    ⊣ 13 3
    ⊣ 15 5
    ⊣ 15 5

…

The ‘-’ sign represents subtraction, so this program reassigns field three, $3, to be the value of field two minus ten, ‘$2 - 10’. (See Section 7.5 [Arithmetic Operators], page 82.) Then field two, and the new value for field three, are printed.

In order for this to work, the text in field $2 must make sense as a number; the string of characters must be converted to a number in order for the computer to do arithmetic on it. The number resulting from the subtraction is converted back to a string of characters which then becomes field three. See Section 7.4 [Conversion of Strings and Numbers], page 81.

When you change the value of a field (as perceived by awk), the text of the input record is recalculated to contain the new field where the old one was. Therefore, $0 changes to reflect the altered field. Thus, this program prints a copy of the input file, with 10 subtracted from the second field of each line.

    $ awk '{ $2 = $2 - 10; print $0 }' inventory-shipped
    ⊣ Jan 3 25 15 115
    ⊣ Feb 5 32 24 226
    ⊣ Mar 5 24 34 228

…

You can also assign contents to fields that are out of range. For example:

    $ awk '{ $6 = ($5 + $4 + $3 + $2)
    >        print $6 }' inventory-shipped
    ⊣ 168
    ⊣ 297
    ⊣ 301

…

We’ve just created $6, whose value is the sum of fields $2, $3, $4, and $5. The ‘+’ sign represents addition. For the file ‘inventory-shipped’, $6 represents the total number of parcels shipped for a particular month.

Creating a new field changes awk’s internal copy of the current input record—the value of $0. Thus, if you do ‘print $0’ after adding a field, the record printed includes the new field, with the appropriate number of field separators between it and the previously existing fields.

This recomputation affects and is affected by NF (the number of fields; see Section 5.2 [Examining Fields], page 40), and by a feature that has not been discussed yet, the output field separator, OFS, which is used to separate the fields (see Section 6.3 [Output Separators], page 63). For example, the value of NF is set to the number of the highest field you create.

Note, however, that merely referencing an out-of-range field does not change the value of either $0 or NF. Referencing an out-of-range field only produces an empty string. For example:

    if ($(NF+1) != "")
        print "can't happen"
    else
        print "everything is normal"

should print ‘everything is normal’, because NF+1 is certain to be out of range. (See Section 9.1 [The if-else Statement], page 105, for more information about awk’s if-else statements. See Section 7.10 [Variable Typing and Comparison Expressions], page 88, for more information about the ‘!=’ operator.)

It is important to note that making an assignment to an existing field will change the value of $0, but will not change the value of NF, even when you assign the empty string to a field. For example:

    $ echo a b c d | awk '{ OFS = ":"; $2 = ""
    >                       print $0; print NF }'
    ⊣ a::c:d
    ⊣ 4

The field is still there; it just has an empty value. You can tell because there are two colons in a row.

This example shows what happens if you create a new field.

    $ echo a b c d | awk '{ OFS = ":"; $2 = ""; $6 = "new"
    >                       print $0; print NF }'
    ⊣ a::c:d::new
    ⊣ 6

The intervening field, $5 is created with an empty value (indicated by the second pair of adjacent colons), and NF is updated with the value six.

Finally, decrementing NF will lose the values of the fields after the new value of NF, and $0 will be recomputed. Here is an example:

    $ echo a b c d e f | ../gawk '{ print "NF =", NF;
    >                               NF = 3; print $0 }'
    ⊣ NF = 6
    ⊣ a b c

5.5 Specifying How Fields are Separated

This section is rather long; it describes one of the most fundamental operations in awk.

5.5.1 The Basics of Field Separating

The field separator, which is either a single character or a regular expression, controls the way awk splits an input record into fields. awk scans the input record for character sequences that match the separator; the fields themselves are the text between the matches.

In the examples below, we use the bullet symbol “•” to represent spaces in the output.

If the field separator is ‘oo’, then the following line:

    moo goo gai pan

would be split into three fields: ‘m’, ‘•g’ and ‘•gai•pan’. Note the leading spaces in the values of the second and third fields.

The field separator is represented by the built-in variable FS. Shell programmers take note! awk does not use the name IFS which is used by the POSIX compatible shells (such as the Bourne shell, sh, or the GNU Bourne-Again Shell, Bash).

You can change the value of FS in the awk program with the assignment operator, ‘=’ (see Section 7.7 [Assignment Expressions], page 84). Often the right time to do this is at the beginning of execution, before any input has been processed, so that the very first record will be read with the proper separator. To do this, use the special BEGIN pattern (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100). For example, here we set the value of FS to the string ",":

    awk 'BEGIN { FS = "," } ; { print $2 }'

Given the input line,

    John Q. Smith, 29 Oak St., Walamazoo, MI 42139

this awk program extracts and prints the string ‘•29•Oak•St.’.

Sometimes your input data will contain separator characters that don’t separate fields the way you thought they would. For instance, the person’s name in the example we just used might have a title or suffix attached, such as ‘John Q. Smith, LXIX’. From input containing such a name:

    John Q. Smith, LXIX, 29 Oak St., Walamazoo, MI 42139

the above program would extract ‘•LXIX’, instead of ‘•29•Oak•St.’. If you were expecting the program to print the address, you would be surprised. The moral is: choose your data layout and separator characters carefully to prevent such problems.

As you know, normally, fields are separated by whitespace sequences (spaces, tabs and newlines), not by single spaces: two spaces in a row do not delimit an empty field. The default value of the field separator FS is a string containing a single space, " ". If this value were interpreted in the usual way, each space character would separate fields, so two spaces in a row would make an empty field between them. The reason this does not happen is that a single space as the value of FS is a special case: it is taken to specify the default manner of delimiting fields.

If FS is any other single character, such as ",", then each occurrence of that character separates two fields. Two consecutive occurrences delimit an empty field. If the character occurs at the beginning or the end of the line, that too delimits an empty field. The space character is the only single character which does not follow these rules.

5.5.2 Using Regular Expressions to Separate Fields

The previous subsection discussed the use of single characters or simple strings as the value of FS. More generally, the value of FS may be a string containing any regular expression. In this case, each match in the record for the regular expression separates fields. For example, the assignment:

    FS = ", \t"

makes every area of an input line that consists of a comma followed by a space and a tab, into a field separator. (‘\t’ is an escape sequence that stands for a tab; see Section 4.2 [Escape Sequences], page 24, for the complete list of similar escape sequences.)

For a less trivial example of a regular expression, suppose you want single spaces to separate fields the way single commas were used above. You can set FS to "[ ]" (left bracket, space, right bracket). This regular expression matches a single space and nothing else (see Chapter 4 [Regular Expressions], page 23).

There is an important difference between the two cases of ‘FS = " "’ (a single space) and ‘FS = "[ \t\n]+"’ (left bracket, space, backslash, “t”, backslash, “n”, right bracket, which is a regular expression matching one or more spaces, tabs, or newlines). For both values of FS, fields are separated by runs of spaces, tabs and/or newlines. However, when the value of FS is " ", awk will first strip leading and trailing whitespace from the record, and then decide where the fields are.

For example, the following pipeline prints ‘b’:

    $ echo ' a b c d ' | awk '{ print $2 }'
    ⊣ b

However, this pipeline prints ‘a’ (note the extra spaces around each letter):

    $ echo ' a  b  c  d ' | awk 'BEGIN { FS = "[ \t]+" }
    >                                  { print $2 }'
    ⊣ a

In this case, the first field is null, or empty.

The stripping of leading and trailing whitespace also comes into play whenever $0 is recomputed. For instance, study this pipeline:

    $ echo '   a b c d' | awk '{ print; $2 = $2; print }'
    ⊣  a b c d
    ⊣ a b c d

The first print statement prints the record as it was read, with leading whitespace intact. The assignment to $2 rebuilds $0 by concatenating $1 through $NF together, separated by the value of OFS. Since the leading whitespace was ignored when finding $1, it is not part of the new $0. Finally, the last print statement prints the new $0.

5.5.3 Making Each Character a Separate Field

There are times when you may want to examine each character of a record separately. In gawk, this is easy to do, you simply assign the null string ("") to FS. In this case, each individual character in the record will become a separate field. Here is an example:

    $ echo a b | gawk 'BEGIN { FS = "" }
    >                  {
    >                      for (i = 1; i <= NF; i = i + 1)
    >                          print "Field", i, "is", $i
    >                  }'
    ⊣ Field 1 is a
    ⊣ Field 2 is
    ⊣ Field 3 is b

Traditionally, the behavior for FS equal to "" was not defined. In this case, Unix awk would simply treat the entire record as only having one field (d.c.). In compatibility mode (see Section 14.1 [Command Line Options], page 159), if FS is the null string, then gawk will also behave this way.

5.5.4 Setting FS from the Command Line

FS can be set on the command line. You use the ‘-F’ option to do so. For example:

awk -F, 'program' input-files sets FS to be the ‘,’ character. Notice that the option uses a capital ‘F’. Contrast this with ‘-f’, which specifies a file containing an awk program. Case is significant in command line options: the ‘-F’ and ‘-f’ options have nothing to do with each other. You can use both options at the same time to set the FS variable and get an awk program from a file.

The value used for the argument to ‘-F’ is processed in exactly the same way as assignments to the built-in variable FS. This means that if the field separator contains special characters, they must be escaped appropriately. For example, to use a ‘\’ as the field separator, you would have to type:

    # same as FS = "\\"
    awk -F\\\\ '. . .' files . . .

Since ‘\’ is used for quoting in the shell, awk will see ‘-F\\’. Then awk processes the ‘\\’ for escape characters (see Section 4.2 [Escape Sequences], page 24), finally yielding a single ‘\’ to be used for the field separator.

As a special case, in compatibility mode (see Section 14.1 [Command Line Options], page 159), if the argument to ‘-F’ is ‘t’, then FS is set to the tab character. This is because if you type ‘-F\t’ at the shell, without any quotes, the ‘\’ gets deleted, so awk figures that you really want your fields to be separated with tabs, and not ‘t’s. Use ‘-v FS="t"’ on the command line if you really do want to separate your fields with ‘t’s (see Section 14.1 [Command Line Options], page 159).

For example, let’s use an awk program file called ‘baud.awk’ that contains the pattern /300/, and the action ‘print $1’. Here is the program:

    /300/   { print $1 }

Let’s also set FS to be the ‘-’ character, and run the program on the file ‘BBS-list’. The following command prints a list of the names of the bulletin boards that operate at 300 baud and the first three digits of their phone numbers:

    $ awk -F- -f baud.awk BBS-list
    ⊣ aardvark     555
    ⊣ alpo
    ⊣ barfly       555

…

Note the second line of output. In the original file (see Section 1.3 [Data Files for the Examples], page 7), the second line looked like this:

    alpo-net     555-3412     2400/1200/300     A

The ‘-’ as part of the system’s name was used as the field separator, instead of the ‘-’ in the phone number that was originally intended. This demonstrates why you have to be careful in choosing your field and record separators.

On many Unix systems, each user has a separate entry in the system password file, one line per user. The information in these lines is separated by colons. The first field is the user’s logon name, and the second is the user’s encrypted password. A password file entry might look like this:

    arnold:xyzzy:2076:10:Arnold Robbins:/home/arnold:/bin/sh

The following program searches the system password file, and prints the entries for users who have no password:

    awk -F: '$2 == ""' /etc/passwd

5.5.5 Field Splitting Summary

According to the POSIX standard, awk is supposed to behave as if each record is split into fields at the time that it is read. In particular, this means that you can change the value of FS after a record is read, and the value of the fields (i.e. how they were split) should reflect the old value of FS, not the new one.

However, many implementations of awk do not work this way. Instead, they defer splitting the fields until a field is actually referenced. The fields will be split using the current value of FS! (d.c.) This behavior can be difficult to diagnose. The following example illustrates the difference between the two methods. (The sed^{2} command prints just the first line of

‘/etc/passwd’.)
    sed 1q /etc/passwd | awk '{ FS = ":" ; print $1 }'

2 The sed utility is a “stream editor.” Its behavior is also defined by the POSIX standard. will usually print

    root

on an incorrect implementation of awk, while gawk will print something like

    root:nSijPlPhZZwgE:0:0:Root:/:

The following table summarizes how fields are split, based on the value of FS. (‘==’ means “is equal to.”)

FS == " " Fields are separated by runs of whitespace. Leading and trailing whitespace are ignored. This is the default.

FS == any other single character Fields are separated by each occurrence of the character. Multiple successive occurrences delimit empty fields, as do leading and trailing occurrences. The character can even be a regexp metacharacter; it does not need to be escaped.

FS == regexp Fields are separated by occurrences of characters that match regexp. Leading and trailing matches of regexp delimit empty fields.

FS == "" Each individual character in the record becomes a separate field.

5.6 Reading Fixed-width Data

(This section discusses an advanced, experimental feature. If you are a novice awk user, you may wish to skip it on the first reading.)

gawk version 2.13 introduced a new facility for dealing with fixed-width fields with no distinctive field separator. Data of this nature arises, for example, in the input for old FORTRAN programs where numbers are run together; or in the output of programs that did not anticipate the use of their output as input for other programs.

An example of the latter is a table where all the columns are lined up by the use of a variable number of spaces and empty fields are just spaces. Clearly, awk’s normal field splitting based on FS will not work well in this case. Although a portable awk program can use a series of substr calls on $0 (see Section 12.3 [Built-in Functions for String Manipulation], page 135), this is awkward and inefficient for a large number of fields.

The splitting of an input record into fixed-width fields is specified by assigning a string containing space-separated numbers to the built-in variable FIELDWIDTHS. Each number specifies the width of the field including columns between fields. If you want to ignore the columns between fields, you can specify the width as a separate field that is subsequently ignored.

The following data is the output of the Unix w utility. It is useful to illustrate the use of FIELDWIDTHS.

 10:06pm  up 21 days, 14:04,  23 users
User     tty       login  idle   JCPU   PCPU  what
hzuo     ttyV0     8:58pm            9      5  vi p24.tex
hzang    ttyV3     6:37pm    50                -csh
eklye    ttyV5     9:53pm            7      1  em thes.tex
dportein ttyV6     8:17pm  1:47                -csh
gierd    ttyD3    10:00pm     1                elm
dave     ttyD4     9:47pm            4      4  w
brent    ttyp0    26Jun91  4:46  26:46   4:41  bash
dave     ttyq4    26Jun9115days     46     46  wnewmail

The following program takes the above input, converts the idle time to number of seconds and prints out the first two fields and the calculated idle time. (This program uses a number of awk features that haven’t been introduced yet.)

BEGIN  { FIELDWIDTHS = "9 6 10 6 7 7 35" }
NR > 2 {
    idle = $4
    sub(/^*/, "", idle)   # strip leading spaces
    if (idle == "")
        idle = 0
    if (idle ~ /:/) {
        split(idle, t, ":")
        idle = t[1] * 60 + t[2]
    }
    if (idle ~ /days/)
        idle *= 24 * 60 * 60
    print $1, $2, idle
}

Here is the result of running the program on the data:

hzuo      ttyV0  0
hzang     ttyV3  50
eklye     ttyV5  0
dportein  ttyV6  107
gierd     ttyD3  1
dave      ttyD4  0
brent     ttyp0  286
dave      ttyq4  1296000

Another (possibly more practical) example of fixed-width input data would be the input from a deck of balloting cards. In some parts of the United States, voters mark their choices by punching holes in computer cards. These cards are then processed to count the votes for any particular candidate or on any particular issue. Since a voter may choose not to vote on some issue, any column on the card may be empty. An awk program for processing such data could use the FIELDWIDTHS feature to simplify reading the data. (Of course, getting gawk to run on a system with card readers is another story!)

Assigning a value to FS causes gawk to return to using FS for field splitting. Use ‘FS = FS’ to make this happen, without having to know the current value of FS.

This feature is still experimental, and may evolve over time. Note that in particular, gawk does not attempt to verify the sanity of the values used in the value of FIELDWIDTHS.

5.7 Multiple-Line Records

In some data bases, a single line cannot conveniently hold all the information in one entry. In such cases, you can use multi-line records.

The first step in doing this is to choose your data format: when records are not defined as single lines, how do you want to define them? What should separate records?

One technique is to use an unusual character or string to separate records. For example, you could use the formfeed character (written ‘\f’ in awk, as in C) to separate them, making each record a page of the file. To do this, just set the variable RS to "\f" (a string containing the formfeed character). Any other character could equally well be used, as long as it won’t be part of the data in a record.

Another technique is to have blank lines separate records. By a special dispensation, an empty string as the value of RS indicates that records are separated by one or more blank lines. If you set RS to the empty string, a record always ends at the first blank line encountered. And the next record doesn’t start until the first non-blank line that follows—no matter how many blank lines appear in a row, they are considered one record-separator.

You can achieve the same effect as ‘RS = ""’ by assigning the string "\n\n+" to RS. This regexp matches the newline at the end of the record, and one or more blank lines after the record. In addition, a regular expression always matches the longest possible sequence when there is a choice (see Section 4.6 [How Much Text Matches?], page 34). So the next record doesn’t start until the first non-blank line that follows—no matter how many blank lines appear in a row, they are considered one record-separator.

There is an important difference between ‘RS = ""’ and ‘RS = "\n\n+"’. In the first case, leading newlines in the input data file are ignored, and if a file ends without extra blank lines after the last record, the final newline is removed from the record. In the second case, this special processing is not done (d.c.).

Now that the input is separated into records, the second step is to separate the fields in the record. One way to do this is to divide each of the lines into fields in the normal manner. This happens by default as the result of a special feature: when RS is set to the empty string, the newline character always acts as a field separator. This is in addition to whatever field separations result from FS.

The original motivation for this special exception was probably to provide useful behavior in the default case (i.e. FS is equal to " "). This feature can be a problem if you really don’t want the newline character to separate fields, since there is no way to prevent it. However, you can work around this by using the split function to break up the record manually (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

Another way to separate fields is to put each field on a separate line: to do this, just set the variable FS to the string "\n". (This simple regular expression matches a single newline.)

A practical example of a data file organized this way might be a mailing list, where each entry is separated by blank lines. If we have a mailing list in a file named ‘addresses’, that looks like this:

    Jane Doe
    123 Main Street
    Anywhere, SE 12345-6789
    John Smith
    456 Tree-lined Avenue
    Smallville, MW 98765-4321

…

A simple program to process this file would look like this:

    # addrs.awk --- simple mailing list program
    # Records are separated by blank lines.
    # Each line is one field.
    BEGIN { RS = "" ; FS = "\n" }
    {
          print "Name is:", $1
          print "Address is:", $2
          print "City and State are:", $3
          print ""
    }

Running the program produces the following output:

    $ awk -f addrs.awk addresses
    ⊣ Name is: Jane Doe
    ⊣ Address is: 123 Main Street
    ⊣ City and State are: Anywhere, SE 12345-6789

⊣

    ⊣ Name is: John Smith
    ⊣ Address is: 456 Tree-lined Avenue
    ⊣ City and State are: Smallville, MW 98765-4321

⊣ … See Section 16.2.4 [Printing Mailing Labels], page 231, for a more realistic

program that deals with address lists.

The following table summarizes how records are split, based on the value of RS. (‘==’ means “is equal to.”)

RS == "\n"

Records are separated by the newline character (‘\n’). In effect, every line in the data file is a separate record, including blank lines. This is the default.

RS == any single character Records are separated by each occurrence of the character. Multiple successive occurrences delimit empty records.

RS == "" Records are separated by runs of blank lines. The newline character always serves as a field separator, in addition to whatever value FS may have. Leading and trailing newlines in a file are ignored.

RS == regexp Records are separated by occurrences of characters that match regexp. Leading and trailing matches of regexp delimit empty records.

In all cases, gawk sets RT to the input text that matched the value specified by RS.

5.8 Explicit Input with getline

So far we have been getting our input data from awk’s main input stream—either the standard input (usually your terminal, sometimes the output from another program) or from the files specified on the command line. The awk language has a special built-in command called getline that can be used to read input under your explicit control.

5.8.1 Introduction to getline

This command is used in several different ways, and should not be used by beginners. It is covered here because this is the chapter on input. The examples that follow the explanation of the getline command include material that has not been covered yet. Therefore, come back and study the getline command after you have reviewed the rest of this book and have a good knowledge of how awk works.

getline returns one if it finds a record, and zero if the end of the file is encountered. If there is some error in getting a record, such as a file that cannot be opened, then getline returns −1. In this case, gawk sets the variable ERRNO to a string describing the error that occurred.

In the following examples, command stands for a string value that represents a shell command.

5.8.2 Using getline with No Arguments

The getline command can be used without arguments to read input from the current input file. All it does in this case is read the next input record and split it up into fields. This is useful if you’ve finished processing the current record, but you want to do some special processing right now on the next record. Here’s an example:

    awk '{
         if ((t = index($0, "/*")) != 0) {
              # value will be "" if t is 1
              tmp = substr($0, 1, t - 1)
              u = index(substr($0, t + 2), "*/")
              while (u == 0) {
                   if (getline <= 0) {
                        m = "unexpected EOF or error"
                        m = (m ": " ERRNO)
                        print m > "/dev/stderr"
                        exit
                   }
                   t = -1
                   u = index($0, "*/")
              }
              # substr expression will be "" if */
              # occurred at end of line
              $0 = tmp substr($0, t + u + 3)
         }
         print $0
    }'

This awk program deletes all C-style comments, ‘/* … */’, from the input. By replacing the ‘print $0’ with other statements, you could perform more complicated processing on the decommented input, like searching for matches of a regular expression. This program has a subtle problem—it does not work if one comment ends and another begins on the same line.

This form of the getline command sets NF (the number of fields; see Section 5.2 [Examining Fields], page 40), NR (the number of records read so far; see Section 5.1 [How Input is Split into Records], page 37), FNR (the number of records read from this input file), and the value of $0.

Note: the new value of $0 is used in testing the patterns of any subsequent rules. The original value of $0 that triggered the rule which executed getline is lost (d.c.). By contrast, the next statement reads a new record but immediately begins processing it normally, starting with the first rule in the program. See Section 9.7 [The next Statement], page 111.

5.8.3 Using getline Into a Variable

You can use ‘getline var’ to read the next record from awk’s input into the variable var. No other processing is done.

For example, suppose the next line is a comment, or a special string, and you want to read it, without triggering any rules. This form of getline allows you to read that line and store it in a variable so that the main read-a-line-and-check-each-rule loop of awk never sees it.

The following example swaps every two lines of input. For example, given:

    wan
    tew
    free
    phore

it outputs:

    tew
    wan
    phore
    free

Here’s the program:

    awk '{
         if ((getline tmp) > 0) {
              print tmp
              print $0
         } else
              print $0
    }'

The getline command used in this way sets only the variables NR and FNR (and of course, var). The record is not split into fields, so the values of the fields (including $0) and the value of NF do not change.

5.8.4 Using getline from a File

Use ‘getline < file’ to read the next record from the file file. Here file is a string-valued expression that specifies the file name. ‘< file’ is called a redirection since it directs input to come from a different place.

For example, the following program reads its input record from the file ‘secondary.input’ when it encounters a first field with a value equal to 10 in the current input file.

    awk '{
        if ($1 == 10) {
             getline < "secondary.input"
             print
        } else
             print
    }'

Since the main input stream is not used, the values of NR and FNR are not changed. But the record read is split into fields in the normal manner, so the values of $0 and other fields are changed. So is the value of NF.

According to POSIX, ‘getline < expression’ is ambiguous if expression contains unparenthesized operators other than ‘$’; for example, ‘getline < dir "/" file’ is ambiguous because the concatenation operator is not parenthesized, and you should write it as ‘getline < (dir "/" file)’ if you want your program to be portable to other awk implementations.

5.8.5 Using getline Into a Variable from a File

Use ‘getline var < file’ to read input the file file and put it in the variable var. As above, file is a string-valued expression that specifies the file from which to read.

In this version of getline, none of the built-in variables are changed, and the record is not split into fields. The only variable changed is var.

For example, the following program copies all the input files to the output, except for records that say ‘@include filename’. Such a record is replaced by the contents of the file filename.

    awk '{
         if (NF == 2 && $1 == "@include") {
              while ((getline line < $2) > 0)
                   print line
              close($2)
         } else
              print
    }'

Note here how the name of the extra input file is not built into the program; it is taken directly from the data, from the second field on the

‘@include’ line.

The close function is called to ensure that if two identical ‘@include’ lines appear in the input, the entire specified file is included twice. See Section 6.8 [Closing Input and Output Files and Pipes], page 74.

One deficiency of this program is that it does not process nested ‘@include’ statements (‘@include’ statements in included files) the way a true macro preprocessor would. See Section 16.2.9 [An Easy Way to Use Library Functions], page 242, for a program that does handle nested ‘@include’ statements.

5.8.6 Using getline from a Pipe

You can pipe the output of a command into getline, using ‘command | getline’. In this case, the string command is run as a shell command and its output is piped into awk to be used as input. This form of getline reads one record at a time from the pipe.

For example, the following program copies its input to its output, except for lines that begin with ‘@execute’, which are replaced by the output produced by running the rest of the line as a shell command:

    awk '{
         if ($1 == "@execute") {
              tmp = substr($0, 10)
              while ((tmp | getline) > 0)
                   print
              close(tmp)
         } else
              print
    }'

The close function is called to ensure that if two identical ‘@execute’ lines appear in the input, the command is run for each one. See Section 6.8 [Closing Input and Output Files and Pipes], page 74.

Given the input:

    foo
    bar
    baz
    @execute who
    bletch

the program might produce:

    foo
    bar
    baz
    arnold     ttyv0   Jul 13 14:22
    miriam     ttyp0   Jul 13 14:23     (murphy:0)
    bill       ttyp1   Jul 13 14:23     (murphy:0)
    bletch

Notice that this program ran the command who and printed the result. (If you try this program yourself, you will of course get different results, showing you who is logged in on your system.)

This variation of getline splits the record into fields, sets the value of NF and recomputes the value of $0. The values of NR and FNR are not changed.

According to POSIX, ‘expression | getline’ is ambiguous if expression contains unparenthesized operators other than ‘$’; for example, ‘"echo " "date" | getline’ is ambiguous because the concatenation operator is not parenthesized, and you should write it as ‘("echo " "date") | getline’ if you want your program to be portable to other awk implementations.

5.8.7 Using getline Into a Variable from a Pipe

When you use ‘command | getline var’, the output of the command command is sent through a pipe to getline and into the variable var. For example, the following program reads the current date and time into the variable current_time, using the date utility, and then prints it.

    awk 'BEGIN {
         "date" | getline current_time
         close("date")
         print "Report printed on " current_time
    }'

In this version of getline, none of the built-in variables are changed, and the record is not split into fields.

5.8.8 Summary of getline Variants

With all the forms of getline, even though $0 and NF, may be updated, the record will not be tested against all the patterns in the awk program, in the way that would happen if the record were read normally by the main processing loop of awk. However the new record is tested against any subsequent rules.

Many awk implementations limit the number of pipelines an awk program may have open to just one! In gawk, there is no such limit. You can open as many pipelines as the underlying operating system will permit.

An interesting side-effect occurs if you use getline (without a redirection) inside a BEGIN rule. Since an unredirected getline reads from the command line data files, the first getline command causes awk to set the value of FILENAME. Normally, FILENAME does not have a value inside BEGIN rules, since you have not yet started to process the command line data files (d.c.). (See Section 8.1.5 [The BEGIN and END Special Patterns], page 100, also see Section 10.2 [Built-in Variables that Convey Information], page 117.)

The following table summarizes the six variants of getline, listing which built-in variables are set by each one.

getline sets $0, NF, FNR, and NR.
getline var

sets var, FNR, and NR.

getline < file

sets $0, and NF.

getline var < file

sets var.

command | getline

sets $0, and NF.

command | getline var sets var.

60 Effective AWK Programming

Chapter 6: Printing Output 61

6 Printing Output

One of the most common actions is to print, or output, some or all of the input. You use the print statement for simple output. You use the printf statement for fancier formatting. Both are described in this chapter.

6.1 The print Statement

The print statement does output with simple, standardized formatting. You specify only the strings or numbers to be printed, in a list separated by commas. They are output, separated by single spaces, followed by a newline. The statement looks like this:

print item1, item2, …

The entire list of items may optionally be enclosed in parentheses. The parentheses are necessary if any of the item expressions uses the ‘>’ relational operator; otherwise it could be confused with a redirection (see Section 6.6 [Redirecting Output of print and printf], page 70).

The items to be printed can be constant strings or numbers, fields of the current record (such as $1), variables, or any awk expressions. Numeric values are converted to strings, and then printed.

The print statement is completely general for computing what values to print. However, with two exceptions, you cannot specify how to print them—how many columns, whether to use exponential notation or not, and so on. (For the exceptions, see Section 6.3 [Output Separators], page 63, and Section 6.4 [Controlling Numeric Output with print], page 64.) For that, you need the printf statement (see Section 6.5 [Using printf Statements for Fancier Printing], page 64).

The simple statement ‘print’ with no items is equivalent to ‘print $0’: it prints the entire current record. To print a blank line, use ‘print ""’, where "" is the empty string.

To print a fixed piece of text, use a string constant such as "Don't Panic" as one item. If you forget to use the double-quote characters, your text will be taken as an awk expression, and you will probably get an error. Keep in mind that a space is printed between any two items.

Each print statement makes at least one line of output. But it isn’t limited to one line. If an item value is a string that contains a newline, the newline is output along with the rest of the string. A single print can make any number of lines this way.

6.2 Examples of print Statements

Here is an example of printing a string that contains embedded newlines (the ‘\n’ is an escape sequence, used to represent the newline character; see Section 4.2 [Escape Sequences], page 24):

    $ awk 'BEGIN { print "line one\nline two\nline three" }'
    ⊣ line one
    ⊣ line two
    ⊣ line three

Here is an example that prints the first two fields of each input record, with a space between them:

    $ awk '{ print $1, $2 }' inventory-shipped
    ⊣ Jan 13
    ⊣ Feb 15
    ⊣ Mar 15

…

A common mistake in using the print statement is to omit the comma between two items. This often has the effect of making the items run together in the output, with no space. The reason for this is that juxtaposing two string expressions in awk means to concatenate them. Here is the same program, without the comma:

    $ awk '{ print $1 $2 }' inventory-shipped
    ⊣ Jan13
    ⊣ Feb15
    ⊣ Mar15

…

To someone unfamiliar with the file ‘inventory-shipped’, neither example’s output makes much sense. A heading line at the beginning would make it clearer. Let’s add some headings to our table of months ($1) and green crates shipped ($2). We do this using the BEGIN pattern (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100) to force the headings to be printed only once:

    awk 'BEGIN {  print "Month Crates"
                  print "----- ------" }
               {  print $1, $2 }' inventory-shipped

Did you already guess what happens? When run, the program prints the following:

    Month Crates
    ----- ------
    Jan 13
    Feb 15
    Mar 15

… The headings and the table data don’t line up! We can fix this by printing some spaces between the two fields:

    awk 'BEGIN { print "Month Crates"
                 print "----- ------" }
               { print $1, "     ", $2 }' inventory-shipped

You can imagine that this way of lining up columns can get pretty complicated when you have many columns to fix. Counting spaces for two or three columns can be simple, but more than this and you can get lost quite easily. This is why the printf statement was created (see Section 6.5 [Using printf Statements for Fancier Printing], page 64); one of its specialties is lining up columns of data.

As a side point, you can continue either a print or printf statement simply by putting a newline after any comma (see Section 2.6 [awk Statements Versus Lines], page 16).

6.3 Output Separators

As mentioned previously, a print statement contains a list of items, separated by commas. In the output, the items are normally separated by single spaces. This need not be the case; a single space is only the default. You can specify any string of characters to use as the output field separator by setting the built-in variable OFS. The initial value of this variable is the string " ", that is, a single space.

The output from an entire print statement is called an output record. Each print statement outputs one output record and then outputs a string called the output record separator. The built-in variable ORS specifies this string. The initial value of ORS is the string "\n", i.e. a newline character; thus, normally each print statement makes a separate line.

You can change how output fields and records are separated by assigning new values to the variables OFS and/or ORS. The usual place to do this is in the BEGIN rule (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100), so that it happens before any input is processed. You may also do this with assignments on the command line, before the names of your input files, or using the ‘-v’ command line option (see Section 14.1 [Command Line Options], page 159).

The following example prints the first and second fields of each input record separated by a semicolon, with a blank line added after each line:

    $ awk 'BEGIN { OFS = ";"; ORS = "\n\n" }
    >            { print $1, $2 }' BBS-list
    ⊣ aardvark;555-5553

⊣

    ⊣ alpo-net;555-3412

⊣

    ⊣ barfly;555-7685

…

If the value of ORS does not contain a newline, all your output will be run together on a single line, unless you output newlines some other way.

6.4 Controlling Numeric Output with print

When you use the print statement to print numeric values, awk internally converts the number to a string of characters, and prints that string. awk uses the sprintf function to do this conversion (see Section 12.3 [Built-in Functions for String Manipulation], page 135). For now, it suffices to say that the sprintf function accepts a format specification that tells it how to format numbers (or strings), and that there are a number of different ways in which numbers can be formatted. The different format specifications are discussed more fully in Section 6.5.2 [Format-Control Letters], page 65.

The built-in variable OFMT contains the default format specification that print uses with sprintf when it wants to convert a number to a string for printing. The default value of OFMT is "%.6g". By supplying different format specifications as the value of OFMT, you can change how print will print your numbers. As a brief example:

    $ awk 'BEGIN {
    >   OFMT = "%.0f"  # print numbers as integers (rounds)
    >   print 17.23 }'
    ⊣ 17

According to the POSIX standard, awk’s behavior will be undefined if OFMT contains anything but a floating point conversion specification (d.c.).

6.5 Using printf Statements for Fancier Printing

If you want more precise control over the output format than print gives you, use printf. With printf you can specify the width to use for each item, and you can specify various formatting choices for numbers (such as what radix to use, whether to print an exponent, whether to print a sign, and how many digits to print after the decimal point). You do this by supplying a string, called the format string, which controls how and where to print the other arguments.

6.5.1 Introduction to the printf Statement

The printf statement looks like this:

printf format, item1, item2, … The entire list of arguments may optionally be enclosed in parentheses. The parentheses are necessary if any of the item expressions use the ‘>’ relational operator; otherwise it could be confused with a redirection (see Section 6.6 [Redirecting Output of print and printf], page 70).

The difference between printf and print is the format argument. This is an expression whose value is taken as a string; it specifies how to output each of the other arguments. It is called the format string.

The format string is very similar to that in the ANSI C library function printf. Most of format is text to be output verbatim. Scattered among this text are format specifiers, one per item. Each format specifier says to output the next item in the argument list at that place in the format.

The printf statement does not automatically append a newline to its output. It outputs only what the format string specifies. So if you want a newline, you must include one in the format string. The output separator variables OFS and ORS have no effect on printf statements. For example:

    BEGIN {
       ORS = "\nOUCH!\n"; OFS = "!"
       msg = "Don't Panic!"; printf "%s\n", msg
    }

This program still prints the familiar ‘Don't Panic!’ message.

6.5.2 Format-Control Letters

A format specifier starts with the character ‘%’ and ends with a formatcontrol letter; it tells the printf statement how to output one item. (If you actually want to output a ‘%’, write ‘%%’.) The format-control letter specifies what kind of value to print. The rest of the format specifier is made up of optional modifiers which are parameters to use, such as the field width.

Here is a list of the format-control letters:

c This prints a number as an ASCII character. Thus, ‘printf "%c", 65’ outputs the letter ‘A’. The output for a string value is the first character of the string.

d

i These are equivalent. They both print a decimal integer. The ‘%i’ specification is for compatibility with ANSI C.

e

E This prints a number in scientific (exponential) notation. For example,

              printf "%4.3e\n", 1950

prints ‘1.950e+03’, with a total of four significant figures of which three follow the decimal point. The ‘4.3’ are modifiers, discussed below. ‘%E’ uses ‘E’ instead of ‘e’ in the output.

f This prints a number in floating point notation. For example,

              printf "%4.3f", 1950

prints ‘1950.000’, with a total of four significant figures of which three follow the decimal point. The ‘4.3’ are modifiers, discussed below.

g

G This prints a number in either scientific notation or floating point notation, whichever uses fewer characters. If the result is printed in scientific notation, ‘%G’ uses ‘E’ instead of ‘e’.

o This prints an unsigned octal integer. (In octal, or base-eight notation, the digits run from ‘0’ to ‘7’; the decimal number eight is represented as ‘10’ in octal.)

s This prints a string.

x

X This prints an unsigned hexadecimal integer. (In hexadecimal, or base-16 notation, the digits are ‘0’ through ‘9’ and ‘a’ through ‘f’. The hexadecimal digit ‘f’ represents the decimal number 15.) ‘%X’ uses the letters ‘A’ through ‘F’ instead of ‘a’ through ‘f’.

% This isn’t really a format-control letter, but it does have a meaning when used after a ‘%’: the sequence ‘%%’ outputs one ‘%’. It does not consume an argument, and it ignores any modifiers.

When using the integer format-control letters for values that are outside the range of a C long integer, gawk will switch to the ‘%g’ format specifier. Other versions of awk may print invalid values, or do something else entirely (d.c.).

6.5.3 Modifiers for printf Formats

A format specification can also include modifiers that can control how much of the item’s value is printed and how much space it gets. The modifiers come between the ‘%’ and the format-control letter. In the examples below, we use the bullet symbol “•” to represent spaces in the output. Here are the possible modifiers, in the order in which they may appear:

- The minus sign, used before the width modifier (see below), says to left-justify the argument within its specified width. Normally the argument is printed right-justified in the specified width. Thus,

    printf "%-4s", "foo"

prints ‘foo•’.

space For numeric conversions, prefix positive values with a space, and negative values with a minus sign.

+ The plus sign, used before the width modifier (see below), says to always supply a sign for numeric conversions, even if the data to be formatted is positive. The ‘+’ overrides the space modifier.

# Use an “alternate form” for certain control letters. For ‘%o’, supply a leading zero. For ‘%x’, and ‘%X’, supply a leading ‘0x’ or ‘0X’ for a non-zero result. For ‘%e’, ‘%E’, and ‘%f’, the result will always contain a decimal point. For ‘%g’, and ‘%G’, trailing zeros are not removed from the result.

0 A leading ‘0’ (zero) acts as a flag, that indicates output should be padded with zeros instead of spaces. This applies even to non-numeric output formats (d.c.). This flag only has an effect when the field width is wider than the value to be printed.

width This is a number specifying the desired minimum width of a field. Inserting any number between the ‘%’ sign and the format control character forces the field to be expanded to this width. The default way to do this is to pad with spaces on the left. For example,

    printf "%4s", "foo"

prints ‘•foo’. The value of width is a minimum width, not a maximum. If the item value requires more than width characters, it can be as wide as necessary. Thus,

    printf "%4s", "foobar"

prints ‘foobar’. Preceding the width with a minus sign causes the output to be padded with spaces on the right, instead of on the left.

.prec This is a number that specifies the precision to use when printing. For the ‘e’, ‘E’, and ‘f’ formats, this specifies the number of digits you want printed to the right of the decimal point. For the ‘g’, and ‘G’ formats, it specifies the maximum number of significant digits. For the ‘d’, ‘o’, ‘i’, ‘u’, ‘x’, and ‘X’ formats, it specifies the minimum number of digits to print. For a string, it specifies the maximum number of characters from the string that should be printed. Thus,

              printf "%.4s", "foobar"

prints ‘foob’.

The C library printf’s dynamic width and prec capability (for example, "%*.*s") is supported. Instead of supplying explicit width and/or prec values in the format string, you pass them in the argument list. For example:

    w = 5
    p = 3
    s = "abcdefg"
    printf "%*.*s\n", w, p, s

is exactly equivalent to

    s = "abcdefg"
    printf "%5.3s\n", s

Both programs output ‘••abc’.

Earlier versions of awk did not support this capability. If you must use such a version, you may simulate this feature by using concatenation to build up the format string, like so:

    w = 5
    p = 3
    s = "abcdefg"
    printf "%" w "." p "s\n", s

This is not particularly easy to read, but it does work.

C programmers may be used to supplying additional ‘l’ and ‘h’ flags in printf format strings. These are not valid in awk. Most awk implementations silently ignore these flags. If ‘--lint’ is provided on the command line (see Section 14.1 [Command Line Options], page 159), gawk will warn about their use. If ‘--posix’ is supplied, their use is a fatal error.

6.5.4 Examples Using printf

Here is how to use printf to make an aligned table:

    awk '{ printf "%-10s %s\n", $1, $2 }' BBS-list

prints the names of bulletin boards ($1) of the file ‘BBS-list’ as a string of 10 characters, left justified. It also prints the phone numbers ($2) afterward on the line. This produces an aligned two-column table of names and phone numbers:

    $ awk '{ printf "%-10s %s\n", $1, $2 }' BBS-list
    ⊣ aardvark   555-5553
    ⊣ alpo-net   555-3412
    ⊣ barfly     555-7685
    ⊣ bites      555-1675
    ⊣ camelot    555-0542
    ⊣ core       555-2912
    ⊣ fooey      555-1234
    ⊣ foot       555-6699
    ⊣ macfoo     555-6480
    ⊣ sdace      555-3430
    ⊣ sabafoo    555-2127

Did you notice that we did not specify that the phone numbers be printed as numbers? They had to be printed as strings because the numbers are separated by a dash. If we had tried to print the phone numbers as numbers, all we would have gotten would have been the first three digits, ‘555’. This would have been pretty confusing.

We did not specify a width for the phone numbers because they are the last things on their lines. We don’t need to put spaces after them.

We could make our table look even nicer by adding headings to the tops of the columns. To do this, we use the BEGIN pattern (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100) to force the header to be printed only once, at the beginning of the awk program:

    awk 'BEGIN { print "Name      Number"
                 print "----      ------" }
         { printf "%-10s %s\n", $1, $2 }' BBS-list

Did you notice that we mixed print and printf statements in the above example? We could have used just printf statements to get the same results:

    awk 'BEGIN { printf "%-10s %s\n", "Name", "Number"
                 printf "%-10s %s\n", "----", "------" }
         { printf "%-10s %s\n", $1, $2 }' BBS-list

By printing each column heading with the same format specification used for the elements of the column, we have made sure that the headings are aligned just like the columns.

The fact that the same format specification is used three times can be emphasized by storing it in a variable, like this:

    awk 'BEGIN { format = "%-10s %s\n"
                 printf format, "Name", "Number"
                 printf format, "----", "------" }
         { printf format, $1, $2 }' BBS-list

See if you can use the printf statement to line up the headings and table data for our ‘inventory-shipped’ example covered earlier in the section on the print statement (see Section 6.1 [The print Statement], page 61).

6.6 Redirecting Output of print and printf

So far we have been dealing only with output that prints to the standard output, usually your terminal. Both print and printf can also send their output to other places. This is called redirection.

A redirection appears after the print or printf statement. Redirections in awk are written just like redirections in shell commands, except that they are written inside the awk program.

There are three forms of output redirection: output to a file, output appended to a file, and output through a pipe to another command. They are all shown for the print statement, but they work identically for printf also.

print items > output-file This type of redirection prints the items into the output file output-file. The file name output-file can be any expression. Its value is changed to a string and then used as a file name (see Chapter 7 [Expressions], page 77). When this type of redirection is used, the output-file is erased before the first output is written to it. Subsequent writes to the same output-file do not erase output-file, but append to it. If output-file does not exist, then it is created. For example, here is how an awk program can write a list of BBS names to a file ‘name-list’ and a list of phone numbers to a file ‘phone-list’. Each output file contains one name or number per line.

    $ awk '{ print $2 > "phone-list"
    >        print $1 > "name-list" }' BBS-list
    $ cat phone-list
    ⊣ 555-5553
    ⊣ 555-3412

…

    $ cat name-list
    ⊣ aardvark
    ⊣ alpo-net

…

print items >> output-file This type of redirection prints the items into the pre-existing output file output-file. The difference between this and the single-‘>’ redirection is that the old contents (if any) of outputfile are not erased. Instead, the awk output is appended to the file. If output-file does not exist, then it is created.

print items | command It is also possible to send output to another program through a pipe instead of into a file. This type of redirection opens a pipe to command and writes the values of items through this pipe, to another process created to execute command. The redirection argument command is actually an awk expression. Its value is converted to a string, whose contents give the shell command to be run. For example, this produces two files, one unsorted list of BBS names and one list sorted in reverse alphabetical order:

    awk '{ print $1 > "names.unsorted"
           command = "sort -r > names.sorted"
           print $1 | command }' BBS-list

Here the unsorted list is written with an ordinary redirection while the sorted list is written by piping through the sort utility. This example uses redirection to mail a message to a mailing list ‘bug-system’. This might be useful when trouble is encountered in an awk script run periodically for system maintenance.

    report = "mail bug-system"
    print "Awk script failed:", $0 | report
    m = ("at record number " FNR " of " FILENAME)
    print m | report
    close(report)

The message is built using string concatenation and saved in the variable m. It is then sent down the pipeline to the mail program. We call the close function here because it’s a good idea to close the pipe as soon as all the intended output has been sent to it. See Section 6.8 [Closing Input and Output Files and Pipes], page 74, for more information on this. This example also illustrates the use of a variable to represent a file or command: it is not necessary to always use a string constant. Using a variable is generally a good idea, since awk requires you to spell the string value identically every time.

Redirecting output using ‘>’, ‘>>’, or ‘|’ asks the system to open a file or pipe only if the particular file or command you’ve specified has not already been written to by your program, or if it has been closed since it was last written to.

As mentioned earlier (see Section 5.8.8 [Summary of getline Variants], page 58), many awk implementations limit the number of pipelines an awk program may have open to just one! In gawk, there is no such limit. You can open as many pipelines as the underlying operating system will permit.

6.7 Special File Names in gawk

Running programs conventionally have three input and output streams already available to them for reading and writing. These are known as the standard input, standard output, and standard error output. These streams are, by default, connected to your terminal, but they are often redirected with the shell, via the ‘<’, ‘<<’, ‘>’, ‘>>’, ‘>&’ and ‘|’ operators. Standard error is typically used for writing error messages; the reason we have two separate streams, standard output and standard error, is so that they can be redirected separately.

In other implementations of awk, the only way to write an error message to standard error in an awk program is as follows:

    print "Serious error detected!" | "cat 1>&2"

This works by opening a pipeline to a shell command which can access the standard error stream which it inherits from the awk process. This is far from elegant, and is also inefficient, since it requires a separate process. So people writing awk programs often neglect to do this. Instead, they send the error messages to the terminal, like this:

    print "Serious error detected!" > "/dev/tty"

This usually has the same effect, but not always: although the standard error stream is usually the terminal, it can be redirected, and when that happens, writing to the terminal is not correct. In fact, if awk is run from a background job, it may not have a terminal at all. Then opening ‘/dev/tty’ will fail.

gawk provides special file names for accessing the three standard streams. When you redirect input or output in gawk, if the file name matches one of these special names, then gawk directly uses the stream it stands for.

‘/dev/stdin’

The standard input (file descriptor 0).

‘/dev/stdout’

The standard output (file descriptor 1).

‘/dev/stderr’

The standard error output (file descriptor 2).

‘/dev/fd/N’

The file associated with file descriptor N. Such a file must have been opened by the program initiating the awk execution (typically the shell). Unless you take special pains in the shell from which you invoke gawk, only descriptors 0, 1 and 2 are available.

The file names ‘/dev/stdin’, ‘/dev/stdout’, and ‘/dev/stderr’ are

aliases for ‘/dev/fd/0’, ‘/dev/fd/1’, and ‘/dev/fd/2’, respectively, but they are more self-explanatory.

The proper way to write an error message in a gawk program is to use ‘/dev/stderr’, like this:

print "Serious error detected!" > "/dev/stderr"

gawk also provides special file names that give access to information about the running gawk process. Each of these “files” provides a single record of information. To read them more than once, you must first close them with the close function (see Section 6.8 [Closing Input and Output Files and Pipes], page 74). The filenames are:

‘/dev/pid’

Reading this file returns the process ID of the current process, in decimal, terminated with a newline.

‘/dev/ppid’

Reading this file returns the parent process ID of the current process, in decimal, terminated with a newline.

‘/dev/pgrpid’

Reading this file returns the process group ID of the current process, in decimal, terminated with a newline.

‘/dev/user’

Reading this file returns a single record terminated with a newline. The fields are separated with spaces. The fields represent the following information: $1 The return value of the getuid system call (the real

user ID number).

$2 The return value of the geteuid system call (the effective user ID number).

$3 The return value of the getgid system call (the real group ID number).

$4 The return value of the getegid system call (the effective group ID number).

If there are any additional fields, they are the group IDs returned by getgroups system call. (Multiple groups may not be supported on all systems.)

These special file names may be used on the command line as data files, as well as for I/O redirections within an awk program. They may not be used as source files with the ‘-f’ option.

Recognition of these special file names is disabled if gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159).

Caution: Unless your system actually has a ‘/dev/fd’ directory (or any of the other above listed special files), the interpretation of these file names is done by gawk itself. For example, using ‘/dev/fd/4’ for output will actually write on file descriptor 4, and not on a new file descriptor that was dup’ed from file descriptor 4. Most of the time this does not matter; however, it is important to not close any of the files related to file descriptors 0, 1, and 2. If you do close one of these files, unpredictable behavior will result.

The special files that provide process-related information may disappear in a future version of gawk. See Section C.3 [Probable Future Extensions], page 297.

6.8 Closing Input and Output Files and Pipes

If the same file name or the same shell command is used with getline (see Section 5.8 [Explicit Input with getline], page 53) more than once during the execution of an awk program, the file is opened (or the command is executed) only the first time. At that time, the first record of input is read from that file or command. The next time the same file or command is used in getline, another record is read from it, and so on.

Similarly, when a file or pipe is opened for output, the file name or command associated with it is remembered by awk and subsequent writes to the same file or command are appended to the previous writes. The file or pipe stays open until awk exits.

This implies that if you want to start reading the same file again from the beginning, or if you want to rerun a shell command (rather than reading more output from the command), you must take special steps. What you must do is use the close function, as follows:

    close(filename)

or

    close(command)

The argument filename or command can be any expression. Its value must exactly match the string that was used to open the file or start the command (spaces and other “irrelevant” characters included). For example, if you open a pipe with this:

    "sort -r names" | getline foo

then you must close it with this:

    close("sort -r names")

Once this function call is executed, the next getline from that file or command, or the next print or printf to that file or command, will reopen the file or rerun the command.

Because the expression that you use to close a file or pipeline must exactly match the expression used to open the file or run the command, it is good practice to use a variable to store the file name or command. The previous example would become

sortcom = "sort -r names"
sortcom | getline foo

…

close(sortcom)

This helps avoid hard-to-find typographical errors in your awk programs. Here are some reasons why you might need to close an output file:

• To write a file and read it back later on in the same awk program. Close the file when you are finished writing it; then you can start reading it

with getline.

• To write numerous files, successively, in the same awk program. If you don’t close the files, eventually you may exceed a system limit on the number of open files in one process. So close each one when you are finished writing it.

• To make a command finish. When you redirect output through a pipe, the command reading the pipe normally continues to try to read input as long as the pipe is open. Often this means the command cannot really do its work until the pipe is closed. For example, if you redirect output to the mail program, the message is not actually sent until the pipe is closed.

• To run the same program a second time, with the same arguments. This is not the same thing as giving more input to the first run! For example, suppose you pipe output to the mail program. If you output several lines redirected to this pipe without closing it, they make a single message of several lines. By contrast, if you close the pipe after each line of output, then each line makes a separate message. close returns a value of zero if the close succeeded. Otherwise, the value

will be non-zero. In this case, gawk sets the variable ERRNO to a string describing the error that occurred.

If you use more files than the system allows you to have open, gawk will attempt to multiplex the available open files among your data files. gawk’s ability to do this depends upon the facilities of your operating system: it may not always work. It is therefore both good practice and good portability advice to always use close on your files when you are done with them.

76 Effective AWK Programming

Chapter 7: Expressions 77

7 Expressions

Expressions are the basic building blocks of awk patterns and actions. An expression evaluates to a value, which you can print, test, store in a variable or pass to a function. Additionally, an expression can assign a new value to a variable or a field, with an assignment operator.

An expression can serve as a pattern or action statement on its own. Most other kinds of statements contain one or more expressions which specify data on which to operate. As in other languages, expressions in awk include variables, array references, constants, and function calls, as well as combinations of these with various operators.

7.1 Constant Expressions

The simplest type of expression is the constant, which always has the same value. There are three types of constants: numeric constants, string constants, and regular expression constants.

7.1.1 Numeric and String Constants

A numeric constant stands for a number. This number can be an integer, a decimal fraction, or a number in scientific (exponential) notation.^{1} Here are some examples of numeric constants, which all have the same value:

    105
    1.05e+2
    1050e-1

A string constant consists of a sequence of characters enclosed in doublequote marks. For example:

    "parrot"

represents the string whose contents are ‘parrot’. Strings in gawk can be of any length and they can contain any of the possible eight-bit ASCII characters including ASCII NUL (character code zero). Other awk implementations may have difficulty with some character codes.

7.1.2 Regular Expression Constants

A regexp constant is a regular expression description enclosed in slashes, such as /^beginning and end$/. Most regexps used in awk programs are constant, but the ‘~’ and ‘!~’ matching operators can also match computed or “dynamic” regexps (which are just ordinary strings or variables that contain a regexp).

1 The internal representation uses double-precision floating point numbers. If you don’t know what that means, then don’t worry about it.

7.2 Using Regular Expression Constants

When used on the right hand side of the ‘~’ or ‘!~’ operators, a regexp constant merely stands for the regexp that is to be matched.

Regexp constants (such as /foo/) may be used like simple expressions. When a regexp constant appears by itself, it has the same meaning as if it appeared in a pattern, i.e. ‘($0 ~ /foo/)’ (d.c.) (see Section 8.1.3 [Expressions as Patterns], page 98). This means that the two code segments,

    if ($0 ~ /barfly/ || $0 ~ /camelot/)
        print "found"

and

    if (/barfly/ || /camelot/)
        print "found"

are exactly equivalent.

One rather bizarre consequence of this rule is that the following boolean expression is valid, but does not do what the user probably intended:

    # note that /foo/ is on the left of the ~
    if (/foo/ ~ $1) print "found foo"

This code is “obviously” testing $1 for a match against the regexp /foo/. But in fact, the expression ‘/foo/ ~ $1’ actually means ‘($0 ~ /foo/) ~ $1’. In other words, first match the input record against the regexp /foo/. The result will be either zero or one, depending upon the success or failure of the match. Then match that result against the first field in the record.

Since it is unlikely that you would ever really wish to make this kind of test, gawk will issue a warning when it sees this construct in a program.

Another consequence of this rule is that the assignment statement

    matches = /foo/

will assign either zero or one to the variable matches, depending upon the contents of the current input record.

This feature of the language was never well documented until the POSIX specification.

Constant regular expressions are also used as the first argument for the gensub, sub and gsub functions, and as the second argument of the match function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). Modern implementations of awk, including gawk, allow the third argument of split to be a regexp constant, while some older implementations do not (d.c.).

This can lead to confusion when attempting to use regexp constants as arguments to user defined functions (see Chapter 13 [User-defined Functions], page 151). For example:

    function mysub(pat, repl, str, global)
    {
        if (global)
            gsub(pat, repl, str)
        else
            sub(pat, repl, str)
        return str
    }
    {

…

        text = "hi! hi yourself!"
        mysub(/hi/, "howdy", text, 1)

…

    }

In this example, the programmer wishes to pass a regexp constant to the user-defined function mysub, which will in turn pass it on to either sub or gsub. However, what really happens is that the pat parameter will be either one or zero, depending upon whether or not $0 matches /hi/.

As it is unlikely that you would ever really wish to pass a truth value in this way, gawk will issue a warning when it sees a regexp constant used as a parameter to a user-defined function.

7.3 Variables

Variables are ways of storing values at one point in your program for use later in another part of your program. You can manipulate them entirely within your program text, and you can also assign values to them on the awk command line.

7.3.1 Using Variables in a Program

Variables let you give names to values and refer to them later. You have already seen variables in many of the examples. The name of a variable must be a sequence of letters, digits and underscores, but it may not begin with a digit. Case is significant in variable names; a and A are distinct variables.

A variable name is a valid expression by itself; it represents the variable’s current value. Variables are given new values with assignment operators, increment operators and decrement operators. See Section 7.7 [Assignment Expressions], page 84.

A few variables have special built-in meanings, such as FS, the field separator, and NF, the number of fields in the current input record. See Chapter 10 [Built-in Variables], page 115, for a list of them. These built-in variables can be used and assigned just like all other variables, but their values are also used or changed automatically by awk. All built-in variables names are entirely upper-case.

Variables in awk can be assigned either numeric or string values. By default, variables are initialized to the empty string, which is zero if converted to a number. There is no need to “initialize” each variable explicitly in awk, the way you would in C and in most other traditional languages.

7.3.2 Assigning Variables on the Command Line

You can set any awk variable by including a variable assignment among the arguments on the command line when you invoke awk (see Section 14.2 [Other Command Line Arguments], page 163). Such an assignment has this form:

variable=text With it, you can set a variable either at the beginning of the awk run or in between input files.

If you precede the assignment with the ‘-v’ option, like this:

-v variable=text then the variable is set at the very beginning, before even the BEGIN rules are run. The ‘-v’ option and its assignment must precede all the file name arguments, as well as the program text. (See Section 14.1 [Command Line Options], page 159, for more information about the ‘-v’ option.)

Otherwise, the variable assignment is performed at a time determined by its position among the input file arguments: after the processing of the preceding input file argument. For example:

    awk '{ print $n }' n=4 inventory-shipped n=2 BBS-list

prints the value of field number n for all input records. Before the first file is read, the command line sets the variable n equal to four. This causes the fourth field to be printed in lines from the file ‘inventory-shipped’. After the first file has finished, but before the second file is started, n is set to two, so that the second field is printed in lines from ‘BBS-list’.

    $ awk '{ print $n }' n=4 inventory-shipped n=2 BBS-list
    ⊣ 15
    ⊣ 24

…

    ⊣ 555-5553
    ⊣ 555-3412

…

Command line arguments are made available for explicit examination by the awk program in an array named ARGV (see Section 10.3 [Using ARGC and ARGV], page 120).

awk processes the values of command line assignments for escape sequences (d.c.) (see Section 4.2 [Escape Sequences], page 24).

7.4 Conversion of Strings and Numbers

Strings are converted to numbers, and numbers to strings, if the context of the awk program demands it. For example, if the value of either foo or bar in the expression ‘foo + bar’ happens to be a string, it is converted to a number before the addition is performed. If numeric values appear in string concatenation, they are converted to strings. Consider this:

    two = 2; three = 3
    print (two three) + 4

This prints the (numeric) value 27. The numeric values of the variables two and three are converted to strings and concatenated together, and the resulting string is converted back to the number 23, to which four is then added.

If, for some reason, you need to force a number to be converted to a string, concatenate the empty string, "", with that number. To force a string to be converted to a number, add zero to that string.

A string is converted to a number by interpreting any numeric prefix of the string as numerals: "2.5" converts to 2.5, "1e3" converts to 1000, and "25fix" has a numeric value of 25. Strings that can’t be interpreted as valid numbers are converted to zero.

The exact manner in which numbers are converted into strings is controlled by the awk built-in variable CONVFMT (see Chapter 10 [Built-in Variables], page 115). Numbers are converted using the sprintf function (see Section 12.3 [Built-in Functions for String Manipulation], page 135) with CONVFMT as the format specifier.

CONVFMT’s default value is "%.6g", which prints a value with at least six significant digits. For some applications you will want to change it to specify more precision. Double precision on most modern machines gives you 16 or 17 decimal digits of precision.

Strange results can happen if you set CONVFMT to a string that doesn’t tell sprintf how to format floating point numbers in a useful way. For example, if you forget the ‘%’ in the format, all numbers will be converted to the same constant string.

As a special case, if a number is an integer, then the result of converting it to a string is always an integer, no matter what the value of CONVFMT may be. Given the following code fragment:

    CONVFMT = "%2.2f"
    a = 12
    b = a ""

b has the value "12", not "12.00" (d.c.).

Prior to the POSIX standard, awk specified that the value of OFMT was used for converting numbers to strings. OFMT specifies the output format to use when printing numbers with print. CONVFMT was introduced in order to separate the semantics of conversion from the semantics of printing. Both CONVFMT and OFMT have the same default value: "%.6g". In the vast majority of cases, old awk programs will not change their behavior. However, this use of OFMT is something to keep in mind if you must port your program to other implementations of awk; we recommend that instead of changing your programs, you just port gawk itself! See Section 6.1 [The print Statement], page 61, for more information on the print statement.

7.5 Arithmetic Operators

The awk language uses the common arithmetic operators when evaluating expressions. All of these arithmetic operators follow normal precedence rules, and work as you would expect them to.

Here is a file ‘grades’ containing a list of student names and three test scores per student (it’s a small class):

    Pat   100 97 58
    Sandy  84 72 93
    Chris  72 92 89

This programs takes the file ‘grades’, and prints the average of the scores.

    $ awk '{ sum = $2 + $3 + $4 ; avg = sum / 3
    >        print $1, avg }' grades
    ⊣ Pat 85
    ⊣ Sandy 83
    ⊣ Chris 84.3333

This table lists the arithmetic operators in awk, in order from highest precedence to lowest:

- x Negation.

+ x Unary plus. The expression is converted to a number.

x ^y x ** y Exponentiation: x raised to the y power. ‘2 ^3’ has the value

eight. The character sequence ‘**’ is equivalent to ‘^’. (The POSIX standard only specifies the use of ‘^’ for exponentiation.)

x * y Multiplication.

x / y Division. Since all numbers in awk are real numbers, the result is not rounded to an integer: ‘3 / 4’ has the value 0.75. x % y Remainder. The quotient is rounded toward zero to an integer,

multiplied by y and this result is subtracted from x. This operation is sometimes known as “trunc-mod.” The following relation always holds:

              b * int(a / b) + (a % b) == a

One possibly undesirable effect of this definition of remainder is that x % y is negative if x is negative. Thus,

              -17 % 8 = -1

In other awk implementations, the signedness of the remainder may be machine dependent.

x + y Addition.

x - y Subtraction.

For maximum portability, do not use the ‘**’ operator.

Unary plus and minus have the same precedence, the multiplication operators all have the same precedence, and addition and subtraction have the same precedence.

7.6 String Concatenation

It seemed like a good idea at the time.

Brian Kernighan

There is only one string operation: concatenation. It does not have a specific operator to represent it. Instead, concatenation is performed by writing expressions next to one another, with no operator. For example:

    $ awk '{ print "Field number one: " $1 }' BBS-list
    ⊣ Field number one: aardvark
    ⊣ Field number one: alpo-net

…

Without the space in the string constant after the ‘:’, the line would run together. For example:

    $ awk '{ print "Field number one:" $1 }' BBS-list
    ⊣ Field number one:aardvark
    ⊣ Field number one:alpo-net

…

Since string concatenation does not have an explicit operator, it is often necessary to insure that it happens where you want it to by using parentheses to enclose the items to be concatenated. For example, the following code fragment does not concatenate file and name as you might expect:

    file = "file"
    name = "name"
    print "something meaningful" > file name

It is necessary to use the following:

    print "something meaningful" > (file name)

We recommend that you use parentheses around concatenation in all but the most common contexts (such as on the right-hand side of ‘=’).

7.7 Assignment Expressions

An assignment is an expression that stores a new value into a variable. For example, let’s assign the value one to the variable z:

    z = 1

After this expression is executed, the variable z has the value one. Whatever old value z had before the assignment is forgotten.

Assignments can store string values also. For example, this would store

the value "this food is good" in the variable message:
    thing = "food"
    predicate = "good"
    message = "this " thing " is " predicate

(This also illustrates string concatenation.)

The ‘=’ sign is called an assignment operator. It is the simplest assignment operator because the value of the right-hand operand is stored unchanged.

Most operators (addition, concatenation, and so on) have no effect except to compute a value. If you ignore the value, you might as well not use the operator. An assignment operator is different; it does produce a value, but even if you ignore the value, the assignment still makes itself felt through the alteration of the variable. We call this a side effect.

The left-hand operand of an assignment need not be a variable (see Section 7.3 [Variables], page 79); it can also be a field (see Section 5.4 [Changing the Contents of a Field], page 42) or an array element (see Chapter 11 [Arrays in awk], page 123). These are all called lvalues, which means they can appear on the left-hand side of an assignment operator. The right-hand operand may be any expression; it produces the new value which the assignment stores in the specified variable, field or array element. (Such values are called rvalues).

It is important to note that variables do not have permanent types. The type of a variable is simply the type of whatever value it happens to hold at the moment. In the following program fragment, the variable foo has a numeric value at first, and a string value later on:

    foo = 1 print foo foo = "bar" print foo

When the second assignment gives foo a string value, the fact that it previously had a numeric value is forgotten.

String values that do not begin with a digit have a numeric value of zero. After executing this code, the value of foo is five:

    foo = "a string"
    foo = foo + 5

(Note that using a variable as a number and then later as a string can be confusing and is poor programming style. The above examples illustrate how awk works, not how you should write your own programs!)

An assignment is an expression, so it has a value: the same value that is assigned. Thus, ‘z = 1’ as an expression has the value one. One consequence of this is that you can write multiple assignments together:

    x = y = z = 0

stores the value zero in all three variables. It does this because the value of ‘z = 0’, which is zero, is stored into y, and then the value of ‘y = z = 0’, which is zero, is stored into x.

You can use an assignment anywhere an expression is called for. For example, it is valid to write ‘x != (y = 1)’ to set y to one and then test whether x equals one. But this style tends to make programs hard to read; except in a one-shot program, you should not use such nesting of assignments.

Aside from ‘=’, there are several other assignment operators that do arithmetic with the old value of the variable. For example, the operator ‘+=’ computes a new value by adding the right-hand value to the old value of the variable. Thus, the following assignment adds five to the value of foo:

    foo += 5

This is equivalent to the following:

    foo = foo + 5

Use whichever one makes the meaning of your program clearer.

There are situations where using ‘+=’ (or any assignment operator) is not the same as simply repeating the left-hand operand in the right-hand expression. For example:

    # Thanks to Pat Rankin for this example
    BEGIN  {
        foo[rand()] += 5
        for (x in foo)
           print x, foo[x]
        bar[rand()] = bar[rand()] + 5
        for (x in bar)
           print x, bar[x]
    }

The indices of bar are guaranteed to be different, because rand will return different values each time it is called. (Arrays and the rand function haven’t been covered yet. See Chapter 11 [Arrays in awk], page 123, and see Section 12.2 [Numeric Built-in Functions], page 134, for more information). This example illustrates an important fact about the assignment operators: the left-hand expression is only evaluated once.

It is also up to the implementation as to which expression is evaluated first, the left-hand one or the right-hand one. Consider this example:

    i = 1
    a[i += 2] = i + 1

The value of a[3] could be either two or four.

Here is a table of the arithmetic assignment operators. In each case, the right-hand operand is an expression whose value is converted to a number.

lvalue += increment Adds increment to the value of lvalue to make the new value of lvalue.

lvalue -= decrement Subtracts decrement from the value of lvalue.

lvalue *= coefficient Multiplies the value of lvalue by coefficient.

lvalue /= divisor Divides the value of lvalue by divisor.

lvalue %= modulus Sets lvalue to its remainder by modulus.

lvalue ^= power lvalue **= power

Raises lvalue to the power power. (Only the ‘^=’ operator is specified by POSIX.)

For maximum portability, do not use the ‘**=’ operator.

7.8 Increment and Decrement Operators

Increment and decrement operators increase or decrease the value of a variable by one. You could do the same thing with an assignment operator, so the increment operators add no power to the awk language; but they are convenient abbreviations for very common operations.

The operator to add one is written ‘++’. It can be used to increment a variable either before or after taking its value.

To pre-increment a variable v, write ‘++v’. This adds one to the value of v and that new value is also the value of this expression. The assignment expression ‘v += 1’ is completely equivalent.

Writing the ‘++’ after the variable specifies post-increment. This increments the variable value just the same; the difference is that the value of the increment expression itself is the variable’s old value. Thus, if foo has the value four, then the expression ‘foo++’ has the value four, but it changes the value of foo to five.

The post-increment ‘foo++’ is nearly equivalent to writing ‘(foo += 1)

- 1’. It is not perfectly equivalent because all numbers in awk are floating point: in floating point, ‘foo + 1 - 1’ does not necessarily equal foo. But the difference is minute as long as you stick to numbers that are fairly small (less than 10e12).

Any lvalue can be incremented. Fields and array elements are incremented just like variables. (Use ‘$(i++)’ when you wish to do a field reference and a variable increment at the same time. The parentheses are necessary because of the precedence of the field reference operator, ‘$’.)

The decrement operator ‘--’ works just like ‘++’ except that it subtracts one instead of adding. Like ‘++’, it can be used before the lvalue to predecrement or after it to post-decrement.

Here is a summary of increment and decrement expressions.

++lvalue This expression increments lvalue and the new value becomes the value of the expression.

lvalue++ This expression increments lvalue, but the value of the expression is the old value of lvalue.

--lvalue Like ‘++lvalue’, but instead of adding, it subtracts. It decrements lvalue and delivers the value that results.

lvalue-- Like ‘lvalue++’, but instead of adding, it subtracts. It decrements lvalue. The value of the expression is the old value of lvalue.

7.9 True and False in awk

Many programming languages have a special representation for the concepts of “true” and “false.” Such languages usually use the special constants true and false, or perhaps their upper-case equivalents.

awk is different. It borrows a very simple concept of true and false from C. In awk, any non-zero numeric value, or any non-empty string value is true. Any other value (zero or the null string, "") is false. The following program will print ‘A strange truth value’ three times:

    BEGIN {
       if (3.1415927)
           print "A strange truth value"
       if ("Four Score And Seven Years Ago")
           print "A strange truth value"
       if (j = 57)
           print "A strange truth value"
    }

There is a surprising consequence of the “non-zero or non-null” rule: The string constant "0" is actually true, since it is non-null (d.c.).

7.10 Variable Typing and Comparison Expressions

The Guide is definitive. Reality is frequently inaccurate.

The Hitchhiker’s Guide to the Galaxy

Unlike other programming languages, awk variables do not have a fixed type. Instead, they can be either a number or a string, depending upon the value that is assigned to them.

The 1992 POSIX standard introduced the concept of a numeric string, which is simply a string that looks like a number, for example, " +2". This concept is used for determining the type of a variable.

The type of the variable is important, since the types of two variables determine how they are compared.

In gawk, variable typing follows these rules.

1. A numeric literal or the result of a numeric operation has the numeric attribute.

2. A string literal or the result of a string operation has the string attribute.

3. Fields, getline input, FILENAME, ARGV elements, ENVIRON elements and the elements of an array created by split that are numeric strings have the strnum attribute. Otherwise, they have the string attribute. Uninitialized variables also have the strnum attribute.

4. Attributes propagate across assignments, but are not changed by any use. The last rule is particularly important. In the following program, a has

numeric type, even though it is later used in a string operation.

    BEGIN {
             a = 12.345
             b = a " is a cute number"
             print b
    }

When two operands are compared, either string comparison or numeric comparison may be used, depending on the attributes of the operands, according to the following, symmetric, matrix:

STRING NUMERIC STRNUM
STRING  string  string  string
NUMERIC  string  numeric  numeric
STRNUM  string  numeric  numeric

The basic idea is that user input that looks numeric, and only user input, should be treated as numeric, even though it is actually made of characters, and is therefore also a string.

Comparison expressions compare strings or numbers for relationships such as equality. They are written using relational operators, which are a superset of those in C. Here is a table of them:

x < y True if x is less than y.

x <= y True if x is less than or equal to y.

x > y True if x is greater than y.

x >= y True if x is greater than or equal to y.

x == y True if x is equal to y.

x != y True if x is not equal to y.

x ~ y True if the string x matches the regexp denoted by y.

x !~ y True if the string x does not match the regexp denoted by y.

subscript in array True if the array array has an element with the subscript subscript.

Comparison expressions have the value one if true and zero if false.

When comparing operands of mixed types, numeric operands are converted to strings using the value of CONVFMT (see Section 7.4 [Conversion of Strings and Numbers], page 81).

Strings are compared by comparing the first character of each, then the second character of each, and so on. Thus "10" is less than "9". If there are two strings where one is a prefix of the other, the shorter string is less than the longer one. Thus "abc" is less than "abcd".

It is very easy to accidentally mistype the ‘==’ operator, and leave off one of the ‘=’s. The result is still valid awk code, but the program will not do what you mean:

    if (a = b)   # oops! should be a == b

…

    else

… Unless b happens to be zero or the null string, the if part of the test will always succeed. Because the operators are so similar, this kind of error is very difficult to spot when scanning the source code.

Here are some sample expressions, how gawk compares them, and what the result of the comparison is.

1.5 <= 2.0

numeric comparison (true)

"abc" >= "xyz"

string comparison (false)

1.5 != " +2"

string comparison (true)

"1e2" < "3"

string comparison (true)

a = 2; b = "2"

a == b string comparison (true)

a = 2; b = " +2"

a == b string comparison (false)

In this example,

    $ echo 1e2 3 | awk '{ print ($1 < $2) ? "true" : "false" }'
    ⊣ false

the result is ‘false’ since both $1 and $2 are numeric strings and thus both have the strnum attribute, dictating a numeric comparison.

The purpose of the comparison rules and the use of numeric strings is to attempt to produce the behavior that is “least surprising,” while still “doing the right thing.”

String comparisons and regular expression comparisons are very different. For example,

    x == "foo"

has the value of one, or is true, if the variable x is precisely ‘foo’. By contrast,

    x ~ /foo/

has the value one if x contains ‘foo’, such as "Oh, what a fool am I!".

The right hand operand of the ‘~’ and ‘!~’ operators may be either a regexp constant (/…/), or an ordinary expression, in which case the value of the expression as a string is used as a dynamic regexp (see Section 4.1 [How to Use Regular Expressions], page 23; also see Section 4.7 [Using Dynamic Regexps], page 35).

In recent implementations of awk, a constant regular expression in slashes by itself is also an expression. The regexp /regexp/ is an abbreviation for this comparison expression:

$0 ~ /regexp/ One special place where /foo/ is not an abbreviation for ‘$0 ~ /foo/’

is when it is the right-hand operand of ‘~’ or ‘!~’! See Section 7.2 [Using Regular Expression Constants], page 78, where this is discussed in more detail.

7.11 Boolean Expressions

A boolean expression is a combination of comparison expressions or matching expressions, using the boolean operators “or” (‘||’), “and” (‘&&’), and “not” (‘!’), along with parentheses to control nesting. The truth value of the boolean expression is computed by combining the truth values of the component expressions. Boolean expressions are also referred to as logical expressions. The terms are equivalent.

Boolean expressions can be used wherever comparison and matching expressions can be used. They can be used in if, while, do and for statements (see Chapter 9 [Control Statements in Actions], page 105). They have numeric values (one if true, zero if false), which come into play if the result of the boolean expression is stored in a variable, or used in arithmetic.

In addition, every boolean expression is also a valid pattern, so you can use one as a pattern to control the execution of rules.

Here are descriptions of the three boolean operators, with examples.

boolean1 && boolean2 True if both boolean1 and boolean2 are true. For example, the following statement prints the current input record if it contains both ‘2400’ and ‘foo’.

              if ($0 ~ /2400/ && $0 ~ /foo/) print

The subexpression boolean2 is evaluated only if boolean1 is true. This can make a difference when boolean2 contains expressions that have side effects: in the case of ‘$0 ~ /foo/ && ($2 == bar++)’, the variable bar is not incremented if there is no ‘foo’ in the record.

boolean1 || boolean2 True if at least one of boolean1 or boolean2 is true. For example, the following statement prints all records in the input that contain either ‘2400’ or ‘foo’, or both.

              if ($0 ~ /2400/ || $0 ~ /foo/) print

The subexpression boolean2 is evaluated only if boolean1 is false. This can make a difference when boolean2 contains expressions that have side effects.

! boolean True if boolean is false. For example, the following program prints all records in the input file ‘BBS-list’ that do not contain the string ‘foo’.

              awk '{ if (! ($0 ~ /foo/)) print }' BBS-list

The ‘&&’ and ‘||’ operators are called short-circuit operators because of the way they work. Evaluation of the full expression is “short-circuited” if the result can be determined part way through its evaluation.

You can continue a statement that uses ‘&&’ or ‘||’ simply by putting a newline after them. But you cannot put a newline in front of either of these operators without using backslash continuation (see Section 2.6 [awk Statements Versus Lines], page 16).

The actual value of an expression using the ‘!’ operator will be either one or zero, depending upon the truth value of the expression it is applied to.

The ‘!’ operator is often useful for changing the sense of a flag variable from false to true and back again. For example, the following program is one way to print lines in between special bracketing lines:

    $1 == "START"   { interested = ! interested }
    interested == 1 { print }
    $1 == "END"     { interested = ! interested }

The variable interested, like all awk variables, starts out initialized to zero, which is also false. When a line is seen whose first field is ‘START’, the value of interested is toggled to true, using ‘!’. The next rule prints lines as long as interested is true. When a line is seen whose first field is ‘END’, interested is toggled back to false.

7.12 Conditional Expressions

A conditional expression is a special kind of expression with three operands. It allows you to use one expression’s value to select one of two other expressions.

The conditional expression is the same as in the C language:

selector ? if-true-exp : if-false-exp There are three subexpressions. The first, selector, is always computed first. If it is “true” (not zero and not null) then if-true-exp is computed next and its value becomes the value of the whole expression. Otherwise, if-false-exp is computed next and its value becomes the value of the whole expression.

For example, this expression produces the absolute value of x:

    x > 0 ? x : -x

Each time the conditional expression is computed, exactly one of if-true-exp and if-false-exp is computed; the other is ignored. This is important when the expressions contain side effects. For example, this conditional expression examines element i of either array a or array b, and increments

i.
    x == y ? a[i++] : b[i++]

This is guaranteed to increment i exactly once, because each time only one of the two increment expressions is executed, and the other is not. See Chapter 11 [Arrays in awk], page 123, for more information about arrays.

As a minor gawk extension, you can continue a statement that uses ‘?:’ simply by putting a newline after either character. However, you cannot put a newline in front of either character without using backslash continuation (see Section 2.6 [awk Statements Versus Lines], page 16).

7.13 Function Calls

A function is a name for a particular calculation. Because it has a name, you can ask for it by name at any point in the program. For example, the function sqrt computes the square root of a number.

A fixed set of functions are built-in, which means they are available in every awk program. The sqrt function is one of these. See Chapter 12 [Builtin Functions], page 133, for a list of built-in functions and their descriptions. In addition, you can define your own functions for use in your program. See Chapter 13 [User-defined Functions], page 151, for how to do this.

The way to use a function is with a function call expression, which consists of the function name followed immediately by a list of arguments in parentheses. The arguments are expressions which provide the raw materials for the function’s calculations. When there is more than one argument, they are separated by commas. If there are no arguments, write just ‘()’ after the function name. Here are some examples:

    sqrt(x^2 + y^2)  one argument

atan2(y, x) two arguments rand() no arguments Do not put any space between the function name and the open-

parenthesis! A user-defined function name looks just like the name of a variable, and space would make the expression look like concatenation of a variable with an expression inside parentheses. Space before the parenthesis is harmless with built-in functions, but it is best not to get into the habit of using space to avoid mistakes with user-defined functions.

Each function expects a particular number of arguments. For example, the sqrt function must be called with a single argument, the number to take the square root of:

sqrt(argument)

Some of the built-in functions allow you to omit the final argument. If you do so, they use a reasonable default. See Chapter 12 [Built-in Functions], page 133, for full details. If arguments are omitted in calls to user-defined functions, then those arguments are treated as local variables, initialized to the empty string (see Chapter 13 [User-defined Functions], page 151).

Like every other expression, the function call has a value, which is computed by the function based on the arguments you give it. In this example, the value of ‘sqrt(argument)’ is the square root of argument. A function can also have side effects, such as assigning values to certain variables or doing I/O.

Here is a command to read numbers, one number per line, and print the square root of each one:

    $ awk '{ print "The square root of", $1, "is", sqrt($1) }'
    1
    ⊣ The square root of 1 is 1
    3
    ⊣ The square root of 3 is 1.73205
    5
    ⊣ The square root of 5 is 2.23607
    Control-d

7.14 Operator Precedence (How Operators Nest)

Operator precedence determines how operators are grouped, when different operators appear close by in one expression. For example, ‘*’ has higher precedence than ‘+’; thus, ‘a + b * c’ means to multiply b and c, and then add a to the product (i.e. ‘a + (b * c)’).

You can overrule the precedence of the operators by using parentheses. You can think of the precedence rules as saying where the parentheses are assumed to be if you do not write parentheses yourself. In fact, it is wise to always use parentheses whenever you have an unusual combination of operators, because other people who read the program may not remember what the precedence is in this case. You might forget, too; then you could make a mistake. Explicit parentheses will help prevent any such mistake.

When operators of equal precedence are used together, the leftmost operator groups first, except for the assignment, conditional and exponentiation operators, which group in the opposite order. Thus, ‘a - b + c’ groups as

‘(a - b) + c’, and ‘a = b = c’ groups as ‘a = (b = c)’.

The precedence of prefix unary operators does not matter as long as only unary operators are involved, because there is only one way to interpret them—innermost first. Thus, ‘$++i’ means ‘$(++i)’ and ‘++$x’ means ‘++($x)’. However, when another operator follows the operand, then the precedence of the unary operators can matter. Thus, ‘$x^2’ means ‘($x)^2’, but ‘-x^2’ means ‘-(x^2)’, because ‘-’ has lower precedence than ‘^’ while ‘$’ has higher precedence.

Here is a table of awk’s operators, in order from highest precedence to lowest:

(…) Grouping.

$ Field.

++ -- Increment, decrement.

^** Exponentiation. These operators group right-to-left. (The ‘**’ operator is not specified by POSIX.)

+ - ! Unary plus, minus, logical “not”.

* / % Multiplication, division, modulus.

+ - Addition, subtraction.

Concatenation No special token is used to indicate concatenation. The operands are simply written side by side.

< <= == !=

> >= >> | Relational, and redirection. The relational operators and the redirections have the same precedence level. Characters such as ‘>’ serve both as relationals and as redirections; the context distinguishes between the two meanings. Note that the I/O redirection operators in print and printf statements belong to the statement level, not to expressions. The redirection does not produce an expression which could be the operand of another operator. As a result, it does not make sense to use a redirection operator near another operator of lower precedence, without parentheses. Such combinations, for example ‘print foo > a ? b : c’, result in syntax errors. The correct way to write this statement is ‘print foo > (a ? b : c)’.

~ !~ Matching, non-matching.

in Array membership.

&& Logical “and”. || Logical “or”. ?: Conditional. This operator groups right-to-left.

= += -= *=
/= %= ^= **=

Assignment. These operators group right-to-left. (The ‘**=’ operator is not specified by POSIX.)

Chapter 8: Patterns and Actions 97

8 Patterns and Actions

As you have already seen, each awk statement consists of a pattern with an associated action. This chapter describes how you build patterns and actions.

8.1 Pattern Elements

Patterns in awk control the execution of rules: a rule is executed when its pattern matches the current input record. This section explains all about how to write patterns.

8.1.1 Kinds of Patterns

Here is a summary of the types of patterns supported in awk.

/regular expression/ A regular expression as a pattern. It matches when the text of the input record fits the regular expression. (See Chapter 4 [Regular Expressions], page 23.)

expression A single expression. It matches when its value is non-zero (if a number) or non-null (if a string). (See Section 8.1.3 [Expressions as Patterns], page 98.)

pat1, pat2 A pair of patterns separated by a comma, specifying a range of records. The range includes both the initial record that matches pat1, and the final record that matches pat2. (See Section 8.1.4 [Specifying Record Ranges with Patterns], page 99.)

BEGIN

END Special patterns for you to supply start-up or clean-up actions for your awk program. (See Section 8.1.5 [The BEGIN and END Special Patterns], page 100.)

empty The empty pattern matches every input record. (See Section 8.1.6 [The Empty Pattern], page 102.)

8.1.2 Regular Expressions as Patterns

We have been using regular expressions as patterns since our early examples. This kind of pattern is simply a regexp constant in the pattern part of a rule. Its meaning is ‘$0 ~ /pattern/’. The pattern matches when the input record matches the regexp. For example:

    /foo|bar|baz/  { buzzwords++ }
    END            { print buzzwords, "buzzwords seen" }

8.1.3 Expressions as Patterns

Any awk expression is valid as an awk pattern. Then the pattern matches if the expression’s value is non-zero (if a number) or non-null (if a string).

The expression is reevaluated each time the rule is tested against a new input record. If the expression uses fields such as $1, the value depends directly on the new input record’s text; otherwise, it depends only on what has happened so far in the execution of the awk program, but that may still be useful.

A very common kind of expression used as a pattern is the comparison expression, using the comparison operators described in Section 7.10 [Variable Typing and Comparison Expressions], page 88.

Regexp matching and non-matching are also very common expressions. The left operand of the ‘~’ and ‘!~’ operators is a string. The right operand is either a constant regular expression enclosed in slashes (/regexp/), or any expression, whose string value is used as a dynamic regular expression (see Section 4.7 [Using Dynamic Regexps], page 35).

The following example prints the second field of each input record whose first field is precisely ‘foo’.

    $ awk '$1 == "foo" { print $2 }' BBS-list

(There is no output, since there is no BBS site named “foo”.) Contrast this with the following regular expression match, which would accept any record with a first field that contains ‘foo’:

    $ awk '$1 ~ /foo/ { print $2 }' BBS-list
    ⊣ 555-1234
    ⊣ 555-6699
    ⊣ 555-6480
    ⊣ 555-2127

Boolean expressions are also commonly used as patterns. Whether the pattern matches an input record depends on whether its subexpressions match.

For example, the following command prints all records in ‘BBS-list’ that contain both ‘2400’ and ‘foo’.

    $ awk '/2400/ && /foo/' BBS-list
    ⊣ fooey        555-1234     2400/1200/300     B

The following command prints all records in ‘BBS-list’ that contain either ‘2400’ or ‘foo’, or both.

    $ awk '/2400/ || /foo/' BBS-list
    ⊣ alpo-net     555-3412     2400/1200/300     A
    ⊣ bites        555-1675     2400/1200/300     A
    ⊣ fooey        555-1234     2400/1200/300     B
    ⊣ foot         555-6699     1200/300          B
    ⊣ macfoo       555-6480     1200/300          A
    ⊣ sdace        555-3430     2400/1200/300     A
    ⊣ sabafoo      555-2127     1200/300          C

The following command prints all records in ‘BBS-list’ that do not contain the string ‘foo’.

    $ awk '! /foo/' BBS-list
    ⊣ aardvark     555-5553     1200/300          B
    ⊣ alpo-net     555-3412     2400/1200/300     A
    ⊣ barfly       555-7685     1200/300          A
    ⊣ bites        555-1675     2400/1200/300     A
    ⊣ camelot      555-0542     300               C
    ⊣ core         555-2912     1200/300          C
    ⊣ sdace        555-3430     2400/1200/300     A

The subexpressions of a boolean operator in a pattern can be constant regular expressions, comparisons, or any other awk expressions. Range patterns are not expressions, so they cannot appear inside boolean patterns. Likewise, the special patterns BEGIN and END, which never match any input record, are not expressions and cannot appear inside boolean patterns.

A regexp constant as a pattern is also a special case of an expression pattern. /foo/ as an expression has the value one if ‘foo’ appears in the current input record; thus, as a pattern, /foo/ matches any record containing ‘foo’.

8.1.4 Specifying Record Ranges with Patterns

A range pattern is made of two patterns separated by a comma, of the form ‘begpat, endpat’. It matches ranges of consecutive input records. The first pattern, begpat, controls where the range begins, and the second one, endpat, controls where it ends. For example,

    awk '$1 == "on", $1 == "off"'

prints every record between ‘on’/‘off’ pairs, inclusive.

A range pattern starts out by matching begpat against every input record; when a record matches begpat, the range pattern becomes turned on. The range pattern matches this record. As long as it stays turned on, it automatically matches every input record read. It also matches endpat against every input record; when that succeeds, the range pattern is turned off again for the following record. Then it goes back to checking begpat against each record.

The record that turns on the range pattern and the one that turns it off both match the range pattern. If you don’t want to operate on these records, you can write if statements in the rule’s action to distinguish them from the records you are interested in.

It is possible for a pattern to be turned both on and off by the same record, if the record satisfies both conditions. Then the action is executed for just that record.

For example, suppose you have text between two identical markers (say the ‘%’ symbol) that you wish to ignore. You might try to combine a range pattern that describes the delimited text with the next statement (not discussed yet, see Section 9.7 [The next Statement], page 111), which causes awk to skip any further processing of the current record and start over again with the next input record. Such a program would look like this:

    /^%$/,/^%$/    { next }
                   { print }

This program fails because the range pattern is both turned on and turned off by the first line with just a ‘%’ on it. To accomplish this task, you must write the program this way, using a flag:

    /^%$/     { skip = ! skip; next }
    skip == 1 { next } # skip lines with `skip' set

Note that in a range pattern, the ‘,’ has the lowest precedence (is evaluated last) of all the operators. Thus, for example, the following program attempts to combine a range pattern with another, simpler test.

    echo Yes | awk '/1/,/2/ || /Yes/'

The author of this program intended it to mean ‘(/1/,/2/) || /Yes/’. However, awk interprets this as ‘/1/, (/2/ || /Yes/)’. This cannot be changed or worked around; range patterns do not combine with other patterns.

8.1.5 The BEGIN and END Special Patterns

BEGIN and END are special patterns. They are not used to match input records. Rather, they supply start-up or clean-up actions for your awk script.

8.1.5.1 Startup and Cleanup Actions

A BEGIN rule is executed, once, before the first input record has been read. An END rule is executed, once, after all the input has been read. For example:

    $ awk '
    > BEGIN { print "Analysis of \"foo\"" }
    > /foo/ { ++n }
    > END   { print "\"foo\" appears " n " times." }' BBS-list
    ⊣ Analysis of "foo"
    ⊣ "foo" appears 4 times.

This program finds the number of records in the input file ‘BBS-list’ that contain the string ‘foo’. The BEGIN rule prints a title for the report. There is no need to use the BEGIN rule to initialize the counter n to zero, as awk does this automatically (see Section 7.3 [Variables], page 79).

The second rule increments the variable n every time a record containing the pattern ‘foo’ is read. The END rule prints the value of n at the end of the run.

The special patterns BEGIN and END cannot be used in ranges or with boolean operators (indeed, they cannot be used with any operators).

An awk program may have multiple BEGIN and/or END rules. They are executed in the order they appear, all the BEGIN rules at start-up and all the END rules at termination. BEGIN and END rules may be intermixed with other rules. This feature was added in the 1987 version of awk, and is included in the POSIX standard. The original (1978) version of awk required you to put the BEGIN rule at the beginning of the program, and the END rule at the end, and only allowed one of each. This is no longer required, but it is a good idea in terms of program organization and readability.

Multiple BEGIN and END rules are useful for writing library functions, since each library file can have its own BEGIN and/or END rule to do its own initialization and/or cleanup. Note that the order in which library functions are named on the command line controls the order in which their BEGIN and END rules are executed. Therefore you have to be careful to write such rules in library files so that the order in which they are executed doesn’t matter. See Section 14.1 [Command Line Options], page 159, for more information on using library functions. See Chapter 15 [A Library of awk Functions], page 167, for a number of useful library functions.

If an awk program only has a BEGIN rule, and no other rules, then the program exits after the BEGIN rule has been run. (The original version of awk used to keep reading and ignoring input until end of file was seen.) However, if an END rule exists, then the input will be read, even if there are no other rules in the program. This is necessary in case the END rule checks the FNR and NR variables (d.c.).

BEGIN and END rules must have actions; there is no default action for these rules since there is no current record when they run.

8.1.5.2 Input/Output from BEGIN and END Rules

There are several (sometimes subtle) issues involved when doing I/O from

a BEGIN or END rule.

The first has to do with the value of $0 in a BEGIN rule. Since BEGIN rules are executed before any input is read, there simply is no input record, and therefore no fields, when executing BEGIN rules. References to $0 and the fields yield a null string or zero, depending upon the context. One way to give $0 a real value is to execute a getline command without a variable (see Section 5.8 [Explicit Input with getline], page 53). Another way is to simply assign a value to it.

The second point is similar to the first, but from the other direction. Inside an END rule, what is the value of $0 and NF? Traditionally, due largely to implementation issues, $0 and NF were undefined inside an END rule. The POSIX standard specified that NF was available in an END rule, containing the number of fields from the last input record. Due most probably to an oversight, the standard does not say that $0 is also preserved, although logically one would think that it should be. In fact, gawk does preserve the value of $0 for use in END rules. Be aware, however, that Unix awk, and possibly other implementations, do not.

The third point follows from the first two. What is the meaning of ‘print’ inside a BEGIN or END rule? The meaning is the same as always, ‘print $0’. If $0 is the null string, then this prints an empty line. Many long time awk programmers use ‘print’ in BEGIN and END rules, to mean ‘print ""’, relying on $0 being null. While you might generally get away with this in BEGIN rules, in gawk at least, it is a very bad idea in END rules. It is also poor style, since if you want an empty line in the output, you should say so explicitly in your program.

8.1.6 The Empty Pattern

An empty (i.e. non-existent) pattern is considered to match every input record. For example, the program:

    awk '{ print $1 }' BBS-list

prints the first field of every record.

8.2 Overview of Actions

An awk program or script consists of a series of rules and function definitions, interspersed. (Functions are described later. See Chapter 13 [Userdefined Functions], page 151.)

A rule contains a pattern and an action, either of which (but not both) may be omitted. The purpose of the action is to tell awk what to do once a match for the pattern is found. Thus, in outline, an awk program generally looks like this:

[pattern] [{ action }]

[pattern] [{ action }] …

function name(args) { . . . }

… An action consists of one or more awk statements, enclosed in curly braces

(‘{’ and ‘}’). Each statement specifies one thing to be done. The statements are separated by newlines or semicolons.

The curly braces around an action must be used even if the action contains only one statement, or even if it contains no statements at all. However, if you omit the action entirely, omit the curly braces as well. An omitted action is equivalent to ‘{ print $0 }’.

/foo/  { }  # match foo, do nothing - empty action
/foo/       # match foo, print the record - omitted action

Here are the kinds of statements supported in awk:

• Expressions, which can call functions or assign values to variables (see Chapter 7 [Expressions], page 77). Executing this kind of statement simply computes the value of the expression. This is useful when the expression has side effects (see Section 7.7 [Assignment Expressions], page 84).

• Control statements, which specify the control flow of awk programs. The awk language gives you C-like constructs (if, for, while, and do) as well as a few special ones (see Chapter 9 [Control Statements in Actions], page 105).

• Compound statements, which consist of one or more statements enclosed in curly braces. A compound statement is used in order to put several statements together in the body of an if, while, do or for statement.

• Input statements, using the getline command (see Section 5.8 [Explicit Input with getline], page 53), the next statement (see Section 9.7 [The next Statement], page 111), and the nextfile statement (see Section 9.8 [The nextfile Statement], page 112).

• Output statements, print and printf. See Chapter 6 [Printing Output], page 61.

• Deletion statements, for deleting array elements. See Section 11.6 [The delete Statement], page 128. The next chapter covers control statements in detail.

104 Effective AWK Programming

Chapter 9: Control Statements in Actions 105

9 Control Statements in Actions

Control statements such as if, while, and so on control the flow of execution in awk programs. Most of the control statements in awk are patterned on similar statements in C.

All the control statements start with special keywords such as if and while, to distinguish them from simple expressions.

Many control statements contain other statements; for example, the if statement contains another statement which may or may not be executed. The contained statement is called the body. If you want to include more than one statement in the body, group them into a single compound statement with curly braces, separating them with newlines or semicolons.

9.1 The if-else Statement

The if-else statement is awk’s decision-making statement. It looks like this:

if (condition) then-body [else else-body] The condition is an expression that controls what the rest of the statement will do. If condition is true, then-body is executed; otherwise, else-body is executed. The else part of the statement is optional. The condition is considered false if its value is zero or the null string, and true otherwise.

Here is an example:

    if (x % 2 == 0)
        print "x is even"
    else
        print "x is odd"

In this example, if the expression ‘x % 2 == 0’ is true (that is, the value of x is evenly divisible by two), then the first print statement is executed, otherwise the second print statement is executed.

If the else appears on the same line as then-body, and then-body is not a compound statement (i.e. not surrounded by curly braces), then a semicolon must separate then-body from else. To illustrate this, let’s rewrite the previous example:

    if (x % 2 == 0) print "x is even"; else
            print "x is odd"

If you forget the ‘;’, awk won’t be able to interpret the statement, and you will get a syntax error.

We would not actually write this example this way, because a human reader might fail to see the else if it were not the first thing on its line.

9.2 The while Statement

In programming, a loop means a part of a program that can be executed two or more times in succession.

The while statement is the simplest looping statement in awk. It repeatedly executes a statement as long as a condition is true. It looks like this:

while (condition)

body Here body is a statement that we call the body of the loop, and condition is an expression that controls how long the loop keeps running.

The first thing the while statement does is test condition. If condition is true, it executes the statement body. After body has been executed, condition is tested again, and if it is still true, body is executed again. This process repeats until condition is no longer true. If condition is initially false, the body of the loop is never executed, and awk continues with the statement following the loop.

This example prints the first three fields of each record, one per line.

    awk '{ i = 1
           while (i <= 3) {
               print $i
               i++
           }
    }' inventory-shipped

Here the body of the loop is a compound statement enclosed in braces, containing two statements.

The loop works like this: first, the value of i is set to one. Then, the while tests whether i is less than or equal to three. This is true when i equals one, so the i-th field is printed. Then the ‘i++’ increments the value of i and the loop repeats. The loop terminates when i reaches four.

As you can see, a newline is not required between the condition and the body; but using one makes the program clearer unless the body is a compound statement or is very simple. The newline after the open-brace that begins the compound statement is not required either, but the program would be harder to read without it.

9.3 The do-while Statement

The do loop is a variation of the while looping statement. The do loop executes the body once, and then repeats body as long as condition is true. It looks like this:

    do

body while (condition) Even if condition is false at the start, body is executed at least once (and

only once, unless executing body makes condition true). Contrast this with the corresponding while statement:

while (condition)

body This statement does not execute body even once if condition is false to begin with.

Here is an example of a do statement:

    awk '{ i = 1
           do {
              print $0
              i++
           } while (i <= 10)
    }'

This program prints each input record ten times. It isn’t a very realistic example, since in this case an ordinary while would do just as well. But this reflects actual experience; there is only occasionally a real use for a do statement.

9.4 The for Statement

The for statement makes it more convenient to count iterations of a loop. The general form of the for statement looks like this:

for (initialization; condition; increment)

body The initialization, condition and increment parts are arbitrary awk expressions, and body stands for any awk statement.

The for statement starts by executing initialization. Then, as long as condition is true, it repeatedly executes body and then increment. Typically initialization sets a variable to either zero or one, increment adds one to it, and condition compares it against the desired number of iterations.

Here is an example of a for statement:

    awk '{ for (i = 1; i <= 3; i++)
              print $i
    }' inventory-shipped

This prints the first three fields of each input record, one field per line.

You cannot set more than one variable in the initialization part unless you use a multiple assignment statement such as ‘x = y = 0’, which is possible only if all the initial values are equal. (But you can initialize additional variables by writing their assignments as separate statements preceding the for loop.)

The same is true of the increment part; to increment additional variables, you must write separate statements at the end of the loop. The C compound expression, using C’s comma operator, would be useful in this context, but it is not supported in awk.

Most often, increment is an increment expression, as in the example above. But this is not required; it can be any expression whatever. For example, this statement prints all the powers of two between one and 100:

    for (i = 1; i <= 100; i *= 2)
      print i

Any of the three expressions in the parentheses following the for may be omitted if there is nothing to be done there. Thus, ‘for (; x > 0;)’ is equivalent to ‘while (x > 0)’. If the condition is omitted, it is treated as true, effectively yielding an infinite loop (i.e. a loop that will never terminate).

In most cases, a for loop is an abbreviation for a while loop, as shown here:

initialization

while (condition) { body increment

    }

The only exception is when the continue statement (see Section 9.6 [The continue Statement], page 110) is used inside the loop; changing a for statement to a while statement in this way can change the effect of the continue statement inside the loop.

There is an alternate version of the for loop, for iterating over all the indices of an array:

    for (i in array)

do something with array[i]

See Section 11.5 [Scanning All Elements of an Array], page 127, for more information on this version of the for loop.

The awk language has a for statement in addition to a while statement because often a for loop is both less work to type and more natural to think of. Counting the number of iterations is very common in loops. It can be easier to think of this counting as part of looping rather than as something to do inside the loop.

The next section has more complicated examples of for loops.

9.5 The break Statement

The break statement jumps out of the innermost for, while, or do loop that encloses it. The following example finds the smallest divisor of any integer, and also identifies prime numbers:

    awk '# find smallest divisor of num
         { num = $1
           for (div = 2; div*div <= num; div++)
             if (num % div == 0)
               break
           if (num % div == 0)
             printf "Smallest divisor of %d is %d\n", num, div
           else
             printf "%d is prime\n", num
         }'

When the remainder is zero in the first if statement, awk immediately breaks out of the containing for loop. This means that awk proceeds immediately to the statement following the loop and continues processing. (This is very different from the exit statement which stops the entire awk program. See Section 9.9 [The exit Statement], page 112.)

Here is another program equivalent to the previous one. It illustrates how the condition of a for or while could just as well be replaced with a break inside an if:

    awk '# find smallest divisor of num
         { num = $1
           for (div = 2; ; div++) {
             if (num % div == 0) {
               printf "Smallest divisor of %d is %d\n", num, div
               break
             }
             if (div*div > num) {
               printf "%d is prime\n", num
               break
             }
           }
    }'

As described above, the break statement has no meaning when used outside the body of a loop. However, although it was never documented, historical implementations of awk have treated the break statement outside of a loop as if it were a next statement (see Section 9.7 [The next Statement], page 111). Recent versions of Unix awk no longer allow this usage. gawk will support this use of break only if ‘--traditional’ has been specified on the command line (see Section 14.1 [Command Line Options], page 159).

Otherwise, it will be treated as an error, since the POSIX standard specifies that break should only be used inside the body of a loop (d.c.).

9.6 The continue Statement

The continue statement, like break, is used only inside for, while, and do loops. It skips over the rest of the loop body, causing the next cycle around the loop to begin immediately. Contrast this with break, which jumps out of the loop altogether.

The continue statement in a for loop directs awk to skip the rest of the body of the loop, and resume execution with the increment-expression of the for statement. The following program illustrates this fact:

    awk 'BEGIN {
         for (x = 0; x <= 20; x++) {
             if (x == 5)
                 continue
             printf "%d ", x
         }
         print ""
    }'

This program prints all the numbers from zero to 20, except for five, for which the printf is skipped. Since the increment ‘x++’ is not skipped, x does not remain stuck at five. Contrast the for loop above with this while loop:

    awk 'BEGIN {
         x = 0
         while (x <= 20) {
             if (x == 5)
                 continue
             printf "%d ", x
             x++
         }
         print ""
    }'

This program loops forever once x gets to five.

As described above, the continue statement has no meaning when used outside the body of a loop. However, although it was never documented, historical implementations of awk have treated the continue statement outside of a loop as if it were a next statement (see Section 9.7 [The next Statement], page 111). Recent versions of Unix awk no longer allow this usage. gawk will support this use of continue only if ‘--traditional’ has been specified on the command line (see Section 14.1 [Command Line Options], page 159). Otherwise, it will be treated as an error, since the POSIX standard specifies that continue should only be used inside the body of a loop (d.c.).

9.7 The next Statement

The next statement forces awk to immediately stop processing the current record and go on to the next record. This means that no further rules are executed for the current record. The rest of the current rule’s action is not executed either.

Contrast this with the effect of the getline function (see Section 5.8 [Explicit Input with getline], page 53). That too causes awk to read the next record immediately, but it does not alter the flow of control in any way. So the rest of the current action executes with a new input record.

At the highest level, awk program execution is a loop that reads an input record and then tests each rule’s pattern against it. If you think of this loop as a for statement whose body contains the rules, then the next statement is analogous to a continue statement: it skips to the end of the body of this implicit loop, and executes the increment (which reads another record).

For example, if your awk program works only on records with four fields, and you don’t want it to fail when given bad input, you might use this rule near the beginning of the program:

    NF != 4 {
      err = sprintf("%s:%d: skipped: NF != 4\n", FILENAME, FNR)
      print err > "/dev/stderr"
      next
    }

so that the following rules will not see the bad record. The error message is redirected to the standard error output stream, as error messages should be. See Section 6.7 [Special File Names in gawk], page 72.

According to the POSIX standard, the behavior is undefined if the next statement is used in a BEGIN or END rule. gawk will treat it as a syntax error. Although POSIX permits it, some other awk implementations don’t allow the next statement inside function bodies (see Chapter 13 [User-defined Functions], page 151). Just as any other next statement, a next inside a function body reads the next record and starts processing it with the first rule in the program.

If the next statement causes the end of the input to be reached, then the code in any END rules will be executed. See Section 8.1.5 [The BEGIN and END Special Patterns], page 100.

Caution: Some awk implementations generate a run-time error if you use the next statement inside a user-defined function (see Chapter 13 [Userdefined Functions], page 151). gawk does not have this problem.

9.8 The nextfile Statement

gawk provides the nextfile statement, which is similar to the next statement. However, instead of abandoning processing of the current record, the nextfile statement instructs gawk to stop processing the current data file.

Upon execution of the nextfile statement, FILENAME is updated to the name of the next data file listed on the command line, FNR is reset to one, ARGIND is incremented, and processing starts over with the first rule in the progam. See Chapter 10 [Built-in Variables], page 115.

If the nextfile statement causes the end of the input to be reached, then the code in any END rules will be executed. See Section 8.1.5 [The BEGIN and END Special Patterns], page 100.

The nextfile statement is a gawk extension; it is not (currently) available in any other awk implementation. See Section 15.2 [Implementing nextfile as a Function], page 168, for a user-defined function you can use to simulate the nextfile statement.

The nextfile statement would be useful if you have many data files to process, and you expect that you would not want to process every record in every file. Normally, in order to move on to the next data file, you would have to continue scanning the unwanted records. The nextfile statement accomplishes this much more efficiently.

Caution: Versions of gawk prior to 3.0 used two words (‘next file’) for the nextfile statement. This was changed in 3.0 to one word, since the treatment of ‘file’ was inconsistent. When it appeared after next, it was a keyword. Otherwise, it was a regular identifier. The old usage is still accepted. However, gawk will generate a warning message, and support for next file will eventually be discontinued in a future version of gawk.

9.9 The exit Statement

The exit statement causes awk to immediately stop executing the current rule and to stop processing input; any remaining input is ignored. It looks like this:

exit [return code] If an exit statement is executed from a BEGIN rule the program stops

processing everything immediately. No input records are read. However, if an END rule is present, it is executed (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100).

If exit is used as part of an END rule, it causes the program to stop immediately.

An exit statement that is not part of a BEGIN or END rule stops the execution of any further automatic rules for the current record, skips reading any remaining input records, and executes the END rule if there is one.

If you do not want the END rule to do its job in this case, you can set a variable to non-zero before the exit statement, and check that variable in the END rule. See Section 15.3 [Assertions], page 170, for an example that does this.

If an argument is supplied to exit, its value is used as the exit status code for the awk process. If no argument is supplied, exit returns status zero (success). In the case where an argument is supplied to a first exit statement, and then exit is called a second time with no argument, the previously supplied exit value is used (d.c.).

For example, let’s say you’ve discovered an error condition you really don’t know how to handle. Conventionally, programs report this by exiting with a non-zero status. Your awk program can do this using an exit statement with a non-zero argument. Here is an example:

    BEGIN {
           if (("date" | getline date_now) < 0) {
             print "Can't get system date" > "/dev/stderr"
             exit 1
           }
           print "current date is", date_now
           close("date")
    }

114 Effective AWK Programming

Chapter 10: Built-in Variables 115

10 Built-in Variables

Most awk variables are available for you to use for your own purposes; they never change except when your program assigns values to them, and never affect anything except when your program examines them. However, a few variables in awk have special built-in meanings. Some of them awk examines automatically, so that they enable you to tell awk how to do certain things. Others are set automatically by awk, so that they carry information from the internal workings of awk to your program.

This chapter documents all the built-in variables of gawk. Most of them are also documented in the chapters describing their areas of activity.

10.1 Built-in Variables that Control awk

This is an alphabetical list of the variables which you can change to control how awk does certain things. Those variables that are specific to gawk are marked with an asterisk, ‘*’.

CONVFMT This string controls conversion of numbers to strings (see Section 7.4 [Conversion of Strings and Numbers], page 81). It works by being passed, in effect, as the first argument to the sprintf function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). Its default value is "%.6g". CONVFMT was introduced by the POSIX standard.

FIELDWIDTHS *

This is a space separated list of columns that tells gawk how to split input with fixed, columnar boundaries. It is an experimental feature. Assigning to FIELDWIDTHS overrides the use of FS for field splitting. See Section 5.6 [Reading Fixed-width Data], page 49, for more information. If gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), then FIELDWIDTHS has no special meaning, and field splitting operations are done based exclusively on the value of FS.

FS FS is the input field separator (see Section 5.5 [Specifying How Fields are Separated], page 44). The value is a single-character string or a multi-character regular expression that matches the separations between fields in an input record. If the value is the null string (""), then each character in the record becomes a separate field. The default value is " ", a string consisting of a single space. As a special exception, this value means that any sequence of spaces, tabs, and/or newlines is a single separator.^{1} It also causes spaces, tabs, and newlines at the beginning and end of a record to be ignored. You can set the value of FS on the command line using the ‘-F’ option:

awk -F, 'program' input-files If gawk is using FIELDWIDTHS for field-splitting, assigning a value to FS will cause gawk to return to the normal, FS-based, field splitting. An easy way to do this is to simply say ‘FS = FS’, perhaps with an explanatory comment.

IGNORECASE *

If IGNORECASE is non-zero or non-null, then all string comparisons, and all regular expression matching are case-independent. Thus, regexp matching with ‘~’ and ‘!~’, and the gensub, gsub, index, match, split and sub functions, record termination with RS, and field splitting with FS all ignore case when doing their particular regexp operations. The value of IGNORECASE does not affect array subscripting. See Section 4.5 [Case-sensitivity in Matching], page 33. If gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), then IGNORECASE has no special meaning, and string and regexp operations are always case-sensitive.

OFMT This string controls conversion of numbers to strings (see Section 7.4 [Conversion of Strings and Numbers], page 81) for printing with the print statement. It works by being passed, in effect, as the first argument to the sprintf function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). Its default value is "%.6g". Earlier versions of awk also used OFMT to specify the format for converting numbers to strings in general expressions; this is now done by CONVFMT.

OFS This is the output field separator (see Section 6.3 [Output Separators], page 63). It is output between the fields output by a print statement. Its default value is " ", a string consisting of a single space.

ORS This is the output record separator. It is output at the end of every print statement. Its default value is "\n". (See Section 6.3 [Output Separators], page 63.)

RS This is awk’s input record separator. Its default value is a string containing a single newline character, which means that an input

1 In POSIX awk, newline does not count as whitespace. record consists of a single line of text. It can also be the null string, in which case records are separated by runs of blank lines, or a regexp, in which case records are separated by matches of the regexp in the input text. (See Section 5.1 [How Input is Split into Records], page 37.)

SUBSEP SUBSEP is the subscript separator. It has the default value of "\034", and is used to separate the parts of the indices of a multi-dimensional array. Thus, the expression foo["A", "B"] really accesses foo["A\034B"] (see Section 11.9 [Multi-dimensional Arrays], page 130).

10.2 Built-in Variables that Convey Information

This is an alphabetical list of the variables that are set automatically by awk on certain occasions in order to provide information to your program. Those variables that are specific to gawk are marked with an asterisk, ‘*’.

ARGC

ARGV The command-line arguments available to awk programs are stored in an array called ARGV. ARGC is the number of commandline arguments present. See Section 14.2 [Other Command Line Arguments], page 163. Unlike most awk arrays, ARGV is indexed from zero to ARGC − 1. For example:

    $ awk 'BEGIN {
    >        for (i = 0; i < ARGC; i++)
    >            print ARGV[i]
    >      }' inventory-shipped BBS-list
    ⊣ awk
    ⊣ inventory-shipped
    ⊣ BBS-list

In this example, ARGV[0] contains "awk", ARGV[1] contains

"inventory-shipped", and ARGV[2] contains "BBS-list". The

value of ARGC is three, one more than the index of the last element in ARGV, since the elements are numbered from zero. The names ARGC and ARGV, as well as the convention of indexing the array from zero to ARGC − 1, are derived from the C language’s method of accessing command line arguments. See Section 10.3 [Using ARGC and ARGV], page 120, for information about how awk uses these variables.

ARGIND * The index in ARGV of the current file being processed. Every time gawk opens a new data file for processing, it sets ARGIND to the index in ARGV of the file name. When gawk is processing the input files, it is always true that ‘FILENAME == ARGV[ARGIND]’.

This variable is useful in file processing; it allows you to tell how far along you are in the list of data files, and to distinguish between successive instances of the same filename on the command line. While you can change the value of ARGIND within your awk program, gawk will automatically set it to a new value when the next file is opened. This variable is a gawk extension. In other awk implementations, or if gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), it is not special.

ENVIRON An associative array that contains the values of the environment. The array indices are the environment variable names; the values are the values of the particular environment variables. For

example, ENVIRON["HOME"] might be ‘/home/arnold’. Chang-

ing this array does not affect the environment passed on to any programs that awk may spawn via redirection or the system function. (In a future version of gawk, it may do so.) Some operating systems may not have environment variables. On such systems, the ENVIRON array is empty (except for

ENVIRON["AWKPATH"]).

ERRNO * If a system error occurs either doing a redirection for getline, during a read for getline, or during a close operation, then ERRNO will contain a string describing the error. This variable is a gawk extension. In other awk implementations, or if gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), it is not special.

FILENAME This is the name of the file that awk is currently reading. When no data files are listed on the command line, awk reads from the standard input, and FILENAME is set to "-". FILENAME is changed each time a new file is read (see Chapter 5 [Reading Input Files], page 37). Inside a BEGIN rule, the value of FILENAME is "", since there are no input files being processed yet.^{2} (d.c.)

FNR FNR is the current record number in the current file. FNR is incremented each time a new record is read (see Section 5.8 [Explicit Input with getline], page 53). It is reinitialized to zero each time a new input file is started.

2 Some early implementations of Unix awk initialized FILENAME to "-", even if there were data files to be processed. This behavior was incorrect, and should not be relied upon in your programs.

NF NF is the number of fields in the current input record. NF is set each time a new record is read, when a new field is created, or when $0 changes (see Section 5.2 [Examining Fields], page 40).

NR This is the number of input records awk has processed since the beginning of the program’s execution (see Section 5.1 [How Input is Split into Records], page 37). NR is set each time a new record is read.

RLENGTH RLENGTH is the length of the substring matched by the match function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). RLENGTH is set by invoking the match function. Its value is the length of the matched string, or −1 if no match was found.

RSTART RSTART is the start-index in characters of the substring matched by the match function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). RSTART is set by invoking the match function. Its value is the position of the string where the matched substring starts, or zero if no match was found.

RT * RT is set each time a record is read. It contains the input text that matched the text denoted by RS, the record separator. This variable is a gawk extension. In other awk implementations, or if gawk is in compatibility mode (see Section 14.1 [Command Line Options], page 159), it is not special.

A side note about NR and FNR. awk simply increments both of these variables each time it reads a record, instead of setting them to the absolute value of the number of records read. This means that your program can change these variables, and their new values will be incremented for each record (d.c.). For example:

$ echo '1
> 2
> 3
> 4' | awk 'NR == 2 { NR = 17 }
> { print NR }'
⊣ 1
⊣ 17
⊣ 18
⊣ 19

Before FNR was added to the awk language (see Section 17.1 [Major Changes between V7 and SVR3.1], page 251), many awk programs used this feature to track the number of records in a file by resetting NR to zero when FILENAME changed.

10.3 Using ARGC and ARGV

In Section 10.2 [Built-in Variables that Convey Information], page 117, you saw this program describing the information contained in ARGC and ARGV:

    $ awk 'BEGIN {
    >        for (i = 0; i < ARGC; i++)
    >            print ARGV[i]
    >      }' inventory-shipped BBS-list
    ⊣ awk
    ⊣ inventory-shipped
    ⊣ BBS-list

In this example, ARGV[0] contains "awk", ARGV[1] contains "inventory-

shipped", and ARGV[2] contains "BBS-list".

Notice that the awk program is not entered in ARGV. The other special command line options, with their arguments, are also not entered. But variable assignments on the command line are treated as arguments, and do show up in the ARGV array.

Your program can alter ARGC and the elements of ARGV. Each time awk reaches the end of an input file, it uses the next element of ARGV as the name of the next input file. By storing a different string there, your program can change which files are read. You can use "-" to represent the standard input. By storing additional elements and incrementing ARGC you can cause additional files to be read.

If you decrease the value of ARGC, that eliminates input files from the end of the list. By recording the old value of ARGC elsewhere, your program can treat the eliminated arguments as something other than file names.

To eliminate a file from the middle of the list, store the null string ("") into ARGV in place of the file’s name. As a special feature, awk ignores file names that have been replaced with the null string. You may also use the delete statement to remove elements from ARGV (see Section 11.6 [The delete Statement], page 128).

All of these actions are typically done from the BEGIN rule, before actual processing of the input begins. See Section 16.1.4 [Splitting a Large File Into Pieces], page 215, and see Section 16.1.5 [Duplicating Output Into Multiple Files], page 217, for an example of each way of removing elements from ARGV.

The following fragment processes ARGV in order to examine, and then remove, command line options.

    BEGIN {
        for (i = 1; i < ARGC; i++) {
            if (ARGV[i] == "-v")
                verbose = 1
            else if (ARGV[i] == "-d")
                debug = 1
            else if (ARGV[i] ~ /^-?/) {
                e = sprintf("%s: unrecognized option -- %c",
                        ARGV[0], substr(ARGV[i], 1, ,1))
                print e > "/dev/stderr"
            } else
                break
            delete ARGV[i]
        }
    }

To actually get the options into the awk program, you have to end the awk options with ‘--’, and then supply your options, like so:

    awk -f myprog -- -v -d file1 file2 . . .

This is not necessary in gawk: Unless ‘--posix’ has been specified, gawk silently puts any unrecognized options into ARGV for the awk program to deal with.

As soon as it sees an unknown option, gawk stops looking for other options it might otherwise recognize. The above example with gawk would be:

    gawk -f myprog -d -v file1 file2 . . .

Since ‘-d’ is not a valid gawk option, the following ‘-v’ is passed on to the awk program.

122 Effective AWK Programming

Chapter 11: Arrays in awk 123

11 Arrays in awk

An array is a table of values, called elements. The elements of an array are distinguished by their indices. Indices may be either numbers or strings. awk maintains a single set of names that may be used for naming variables, arrays and functions (see Chapter 13 [User-defined Functions], page 151). Thus, you cannot have a variable and an array with the same name in the same awk program.

11.1 Introduction to Arrays

The awk language provides one-dimensional arrays for storing groups of related strings or numbers.

Every awk array must have a name. Array names have the same syntax as variable names; any valid variable name would also be a valid array name. But you cannot use one name in both ways (as an array and as a variable) in one awk program.

Arrays in awk superficially resemble arrays in other programming languages; but there are fundamental differences. In awk, you don’t need to specify the size of an array before you start to use it. Additionally, any number or string in awk may be used as an array index, not just consecutive integers.

In most other languages, you have to declare an array and specify how many elements or components it contains. In such languages, the declaration causes a contiguous block of memory to be allocated for that many elements. An index in the array usually must be a positive integer; for example, the index zero specifies the first element in the array, which is actually stored at the beginning of the block of memory. Index one specifies the second element, which is stored in memory right after the first element, and so on. It is impossible to add more elements to the array, because it has room for only as many elements as you declared. (Some languages allow arbitrary starting and ending indices, e.g., ‘15 .. 27’, but the size of the array is still fixed when the array is declared.)

A contiguous array of four elements might look like this, conceptually, if the element values are eight, "foo", "" and 30:

                8    "foo"    ""     30  value

0  1  2  3  index

Only the values are stored; the indices are implicit from the order of the values. Eight is the value at index zero, because eight appears in the position with zero elements before it.

Arrays in awk are different: they are associative. This means that each array is a collection of pairs: an index, and its corresponding array element value:

Element 4 Value 30 Element 2 Value "foo" Element 1 Value 8 Element 3 Value ""

We have shown the pairs in jumbled order because their order is irrelevant.

One advantage of associative arrays is that new pairs can be added at any time. For example, suppose we add to the above array a tenth element whose value is "number ten". The result is this:

    Element 10  Value "number ten"

Element 4 Value 30 Element 2 Value "foo" Element 1 Value 8 Element 3 Value ""

Now the array is sparse, which just means some indices are missing: it has elements 1–4 and 10, but doesn’t have elements 5, 6, 7, 8, or 9.

Another consequence of associative arrays is that the indices don’t have to be positive integers. Any number, or even a string, can be an index. For example, here is an array which translates words from English into French:

    Element "dog" Value "chien"

Element "cat" Value "chat" Element "one" Value "un" Element 1 Value "un"

Here we decided to translate the number one in both spelled-out and numeric form—thus illustrating that a single array can have both numbers and strings as indices. (In fact, array subscripts are always strings; this is discussed in more detail in Section 11.7 [Using Numbers to Subscript Arrays], page 129.)

The value of IGNORECASE has no effect upon array subscripting. You must use the exact same string value to retrieve an array element as you used to store it.

When awk creates an array for you, e.g., with the split built-in function, that array’s indices are consecutive integers starting at one. (See Section 12.3 [Built-in Functions for String Manipulation], page 135.)

11.2 Referring to an Array Element

The principal way of using an array is to refer to one of its elements. An array reference is an expression which looks like this:

array[index] Here, array is the name of an array. The expression index is the index of the element of the array that you want.

The value of the array reference is the current value of that array element. For example, foo[4.3] is an expression for the element of array foo at index

‘4.3’.

If you refer to an array element that has no recorded value, the value of the reference is "", the null string. This includes elements to which you have not assigned any value, and elements that have been deleted (see Section 11.6 [The delete Statement], page 128). Such a reference automatically creates that array element, with the null string as its value. (In some cases, this is unfortunate, because it might waste memory inside awk.)

You can find out if an element exists in an array at a certain index with the expression:

index in array This expression tests whether or not the particular index exists, without the side effect of creating that element if it is not present. The expression has the value one (true) if array[index] exists, and zero (false) if it does not exist.

For example, to test whether the array frequencies contains the index ‘2’, you could write this statement:

    if (2 in frequencies)
        print "Subscript 2 is present."

Note that this is not a test of whether or not the array frequencies contains an element whose value is two. (There is no way to do that except to scan all the elements.) Also, this does not create frequencies[2], while the following (incorrect) alternative would do so:

    if (frequencies[2] != "")
        print "Subscript 2 is present."

11.3 Assigning Array Elements

Array elements are lvalues: they can be assigned values just like awk variables:

array[subscript] = value Here array is the name of your array. The expression subscript is the index of the element of the array that you want to assign a value. The expression value is the value you are assigning to that element of the array.

11.4 Basic Array Example

The following program takes a list of lines, each beginning with a line number, and prints them out in order of line number. The line numbers are not in order, however, when they are first read: they are scrambled. This program sorts the lines by making an array using the line numbers as subscripts. It then prints out the lines in sorted order of their numbers. It is a very simple program, and gets confused if it encounters repeated numbers, gaps, or lines that don’t begin with a number.

{
  if ($1 > max)
    max = $1
  arr[$1] = $0
}
END {
  for (x = 1; x <= max; x++)
    print arr[x]
}

The first rule keeps track of the largest line number seen so far; it also stores each line into the array arr, at an index that is the line’s number.

The second rule runs after all the input has been read, to print out all the lines.

When this program is run with the following input:

5  I am the Five man
2  Who are you?  The new number two!
4  . . . And four on the floor
1  Who is number one?
3  I three you.

its output is this:

1  Who is number one?
2  Who are you?  The new number two!
3  I three you.
4  . . . And four on the floor
5  I am the Five man

If a line number is repeated, the last line with a given number overrides the others.

Gaps in the line numbers can be handled with an easy improvement to the program’s END rule:

END {
  for (x = 1; x <= max; x++)
    if (x in arr)
      print arr[x]
}

11.5 Scanning All Elements of an Array

In programs that use arrays, you often need a loop that executes once for each element of an array. In other languages, where arrays are contiguous and indices are limited to positive integers, this is easy: you can find all the valid indices by counting from the lowest index up to the highest. This technique won’t do the job in awk, since any number or string can be an array index. So awk has a special kind of for statement for scanning an array:

for (var in array)

body This loop executes body once for each index in array that your program has previously used, with the variable var set to that index.

Here is a program that uses this form of the for statement. The first rule scans the input records and notes which words appear (at least once) in the input, by storing a one into the array used with the word as index. The second rule scans the elements of used to find all the distinct words that appear in the input. It prints each word that is more than 10 characters long, and also prints the number of such words. See Section 12.3 [Built-in Functions for String Manipulation], page 135, for more information on the built-in function length.

    # Record a 1 for each word that is used at least once.
    {
        for (i = 1; i <= NF; i++)
            used[$i] = 1
    }
    # Find number of distinct words more than 10 characters long.
    END {
        for (x in used)
            if (length(x) > 10) {
                ++num_long_words
                print x
            }
        print num_long_words, "words longer than 10 characters"
    }

See Section 16.2.5 [Generating Word Usage Counts], page 233, for a more detailed example of this type.

The order in which elements of the array are accessed by this statement is determined by the internal arrangement of the array elements within awk and cannot be controlled or changed. This can lead to problems if new elements are added to array by statements in the loop body; you cannot predict whether or not the for loop will reach them. Similarly, changing var inside the loop may produce strange results. It is best to avoid such things.

11.6 The delete Statement

You can remove an individual element of an array using the delete statement:

delete array[index]

Once you have deleted an array element, you can no longer obtain any value the element once had. It is as if you had never referred to it and had never given it any value.

Here is an example of deleting elements in an array:

    for (i in frequencies)
      delete frequencies[i]

This example removes all the elements from the array frequencies.

If you delete an element, a subsequent for statement to scan the array will not report that element, and the in operator to check for the presence of that element will return zero (i.e. false):

    delete foo[4]
    if (4 in foo)
        print "This will never be printed"

It is important to note that deleting an element is not the same as assigning it a null value (the empty string, "").

    foo[4] = ""
    if (4 in foo)
      print "This is printed, even though foo[4] is empty"

It is not an error to delete an element that does not exist.

You can delete all the elements of an array with a single statement, by leaving off the subscript in the delete statement.

    delete array

This ability is a gawk extension; it is not available in compatibility mode (see Section 14.1 [Command Line Options], page 159).

Using this version of the delete statement is about three times more efficient than the equivalent loop that deletes each element one at a time.

The following statement provides a portable, but non-obvious way to clear out an array.

    # thanks to Michael Brennan for pointing this out
    split("", array)

The split function (see Section 12.3 [Built-in Functions for String Manipulation], page 135) clears out the target array first. This call asks it to split apart the null string. Since there is no data to split out, the function simply clears the array and then returns.

11.7 Using Numbers to Subscript Arrays

An important aspect of arrays to remember is that array subscripts are always strings. If you use a numeric value as a subscript, it will be converted to a string value before it is used for subscripting (see Section 7.4 [Conversion of Strings and Numbers], page 81).

This means that the value of the built-in variable CONVFMT can potentially affect how your program accesses elements of an array. For example:

    xyz = 12.153
    data[xyz] = 1
    CONVFMT = "%2.2f"
    if (xyz in data)
        printf "%s is in data\n", xyz
    else
        printf "%s is not in data\n", xyz

This prints ‘12.15 is not in data’. The first statement gives xyz a numeric value. Assigning to data[xyz] subscripts data with the string value "12.153" (using the default conversion value of CONVFMT, "%.6g"), and assigns one to data["12.153"]. The program then changes the value of CONVFMT. The test ‘(xyz in data)’ generates a new string value from xyz, this time "12.15", since the value of CONVFMT only allows two significant digits. This test fails, since "12.15" is a different string from "12.153".

According to the rules for conversions (see Section 7.4 [Conversion of Strings and Numbers], page 81), integer values are always converted to strings as integers, no matter what the value of CONVFMT may happen to be. So the usual case of:

    for (i = 1; i <= maxsub; i++)

do something with array[i]

will work, no matter what the value of CONVFMT.

Like many things in awk, the majority of the time things work as you would expect them to work. But it is useful to have a precise knowledge of the actual rules, since sometimes they can have a subtle effect on your programs.

11.8 Using Uninitialized Variables as Subscripts

Suppose you want to print your input data in reverse order. A reasonable attempt at a program to do so (with some test data) might look like this:

    $ echo 'line 1
    > line 2
    > line 3' | awk '{ l[lines] = $0; ++lines }
    > END {
    >     for (i = lines-1; i >= 0; --i)
    >        print l[i]
    > }'
    ⊣ line 3
    ⊣ line 2

Unfortunately, the very first line of input data did not come out in the output!

At first glance, this program should have worked. The variable lines is uninitialized, and uninitialized variables have the numeric value zero. So, the value of l[0] should have been printed.

The issue here is that subscripts for awk arrays are always strings. And uninitialized variables, when used as strings, have the value "", not zero. Thus, ‘line 1’ ended up stored in l[""].

The following version of the program works correctly:

    { l[lines++] = $0 }
    END {
        for (i = lines - 1; i >= 0; --i)
           print l[i]
    }

Here, the ‘++’ forces lines to be numeric, thus making the “old value” numeric zero, which is then converted to "0" as the array subscript.

As we have just seen, even though it is somewhat unusual, the null string ("") is a valid array subscript (d.c.). If ‘--lint’ is provided on the command line (see Section 14.1 [Command Line Options], page 159), gawk will warn about the use of the null string as a subscript.

11.9 Multi-dimensional Arrays

A multi-dimensional array is an array in which an element is identified by a sequence of indices, instead of a single index. For example, a twodimensional array requires two indices. The usual way (in most languages, including awk) to refer to an element of a two-dimensional array named grid is with grid[x,y].

Multi-dimensional arrays are supported in awk through concatenation of indices into one string. What happens is that awk converts the indices into strings (see Section 7.4 [Conversion of Strings and Numbers], page 81) and concatenates them together, with a separator between them. This creates a single string that describes the values of the separate indices. The combined string is used as a single index into an ordinary, one-dimensional array. The separator used is the value of the built-in variable SUBSEP.

For example, suppose we evaluate the expression ‘foo[5,12] = "value"’ when the value of SUBSEP is "@". The numbers five and 12 are converted to strings and concatenated with an ‘@’ between them, yielding "5@12"; thus, the array element foo["5@12"] is set to "value".

Once the element’s value is stored, awk has no record of whether it was stored with a single index or a sequence of indices. The two expressions ‘foo[5,12]’ and ‘foo[5 SUBSEP 12]’ are always equivalent.

The default value of SUBSEP is the string "\034", which contains a nonprinting character that is unlikely to appear in an awk program or in most input data.

The usefulness of choosing an unlikely character comes from the fact that index values that contain a string matching SUBSEP lead to combined strings that are ambiguous. Suppose that SUBSEP were "@"; then ‘foo["a@b", "c"]’ and ‘foo["a", "b@c"]’ would be indistinguishable because both would actually be stored as ‘foo["a@b@c"]’.

You can test whether a particular index-sequence exists in a “multidimensional” array with the same operator ‘in’ used for single dimensional arrays. Instead of a single index as the left-hand operand, write the whole sequence of indices, separated by commas, in parentheses:

(subscript1, subscript2, …) in array

The following example treats its input as a two-dimensional array of fields; it rotates this array 90 degrees clockwise and prints the result. It assumes that all lines have the same number of elements.

    awk '{
         if (max_nf < NF)
              max_nf = NF
         max_nr = NR
         for (x = 1; x <= NF; x++)
              vector[x, NR] = $x
    }
    END {
         for (x = 1; x <= max_nf; x++) {
              for (y = max_nr; y >= 1; --y)
                   printf("%s ", vector[x, y])
              printf("\n")
         }
    }'

When given the input:

    1 2 3 4 5 6 2 3 4 5 6 1 3 4 5 6 1 2 4 5 6 1 2 3

it produces:

    4 3 2 1 5 4 3 2 6 5 4 3 1 6 5 4 2 1 6 5 3 2 1 6

11.10 Scanning Multi-dimensional Arrays

There is no special for statement for scanning a “multi-dimensional” array; there cannot be one, because in truth there are no multi-dimensional arrays or elements; there is only a multi-dimensional way of accessing an array.

However, if your program has an array that is always accessed as multidimensional, you can get the effect of scanning it by combining the scanning for statement (see Section 11.5 [Scanning All Elements of an Array], page 127) with the split built-in function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). It works like this:

    for (combined in array) {
      split(combined, separate, SUBSEP)

…

    }

This sets combined to each concatenated, combined index in the array, and splits it into the individual indices by breaking it apart where the value of SUBSEP appears. The split-out indices become the elements of the array

separate.

Thus, suppose you have previously stored a value in array[1, "foo"]; then an element with index "1\034foo" exists in array. (Recall that the default value of SUBSEP is the character with code 034.) Sooner or later the for statement will find that index and do an iteration with combined set to "1\034foo". Then the split function is called as follows:

    split("1\034foo", separate, "\034")
The result of this is to set separate[1] to "1" and separate[2] to "foo".

Presto, the original sequence of separate indices has been recovered.

Chapter 12: Built-in Functions 133

12 Built-in Functions

Built-in functions are functions that are always available for your awk program to call. This chapter defines all the built-in functions in awk; some of them are mentioned in other sections, but they are summarized here for your convenience. (You can also define new functions yourself. See Chapter 13 [User-defined Functions], page 151.)

12.1 Calling Built-in Functions

To call a built-in function, write the name of the function followed by arguments in parentheses. For example, ‘atan2(y + z, 1)’ is a call to the function atan2, with two arguments.

Whitespace is ignored between the built-in function name and the openparenthesis, but we recommend that you avoid using whitespace there. Userdefined functions do not permit whitespace in this way, and you will find it easier to avoid mistakes by following a simple convention which always works: no whitespace after a function name.

Each built-in function accepts a certain number of arguments. In some cases, arguments can be omitted. The defaults for omitted arguments vary from function to function and are described under the individual functions. In some awk implementations, extra arguments given to built-in functions are ignored. However, in gawk, it is a fatal error to give extra arguments to a built-in function.

When a function is called, expressions that create the function’s actual parameters are evaluated completely before the function call is performed. For example, in the code fragment:

    i = 4
    j = sqrt(i++)

the variable i is set to five before sqrt is called with a value of four for its actual parameter.

The order of evaluation of the expressions used for the function’s parameters is undefined. Thus, you should not write programs that assume that parameters are evaluated from left to right or from right to left. For example,

    i = 5
    j = atan2(i++, i *= 2)

If the order of evaluation is left to right, then i first becomes six, and then 12, and atan2 is called with the two arguments six and 12. But if the order of evaluation is right to left, i first becomes 10, and then 11, and atan2 is called with the two arguments 11 and 10.

12.2 Numeric Built-in Functions

Here is a full list of built-in functions that work with numbers. Optional parameters are enclosed in square brackets (“[” and “]”).

int(x) This produces the nearest integer to x, located between x and zero, truncated toward zero. For example, int(3) is three, int(3.9) is three, int(-3.9) is −3, and int(-3) is −3 as well.

sqrt(x) This gives you the positive square root of x. It reports an error if x is negative. Thus, sqrt(4) is two.

exp(x) This gives you the exponential of x (e ^x), or reports an error if x is out of range. The range of values x can have depends on your machine’s floating point representation.

log(x) This gives you the natural logarithm of x, if x is positive; otherwise, it reports an error.

sin(x) This gives you the sine of x, with x in radians.

cos(x) This gives you the cosine of x, with x in radians.

atan2(y, x)

This gives you the arctangent of y / x in radians.

rand() This gives you a random number. The values of rand are uniformly-distributed between zero and one. The value is never zero and never one. Often you want random integers instead. Here is a user-defined function you can use to obtain a random non-negative integer less than n:

    function randint(n) {
         return int(n * rand())
    }

The multiplication produces a random real number greater than zero and less than n. We then make it an integer (using int) between zero and n − 1, inclusive. Here is an example where a similar function is used to produce random integers between one and n. This program prints a new random number for each input record.

    awk '
    # Function to roll a simulated die.
    function roll(n) { return 1 + int(rand() * n) }

    # Roll 3 six-sided dice and
    # print total number of points.
    {
          printf("%d points\n",
                 roll(6)+roll(6)+roll(6))
    }'

Caution: In most awk implementations, including gawk, rand starts generating numbers from the same starting number, or seed, each time you run awk. Thus, a program will generate the same results each time you run it. The numbers are random within one awk run, but predictable from run to run. This is convenient for debugging, but if you want a program to do different things each time it is used, you must change the seed to a value that will be different in each run. To do this, use srand.

srand([x])

The function srand sets the starting point, or seed, for generating random numbers to the value x. Each seed value leads to a particular sequence of random numbers.^{1} Thus, if you set the seed to the same value a second time, you will get the same sequence of random numbers again. If you omit the argument x, as in srand(), then the current date and time of day are used for a seed. This is the way to get random numbers that are truly unpredictable. The return value of srand is the previous seed. This makes it easy to keep track of the seeds for use in consistently reproducing sequences of random numbers.

12.3 Built-in Functions for String Manipulation

The functions in this section look at or change the text of one or more strings. Optional parameters are enclosed in square brackets (“[” and “]”).

index(in, find)

This searches the string in for the first occurrence of the string find, and returns the position in characters where that occurrence begins in the string in. For example:

    $ awk 'BEGIN { print index("peanut", "an") }'
    ⊣ 3

1 Computer generated random numbers really are not truly random. They are technically known as “pseudo-random.” This means that while the numbers in a sequence appear to be random, you can in fact generate the same sequence of random numbers over and over again.

If find is not found, index returns zero. (Remember that string indices in awk start at one.)

length([string])

This gives you the number of characters in string. If string is a number, the length of the digit string representing that number is returned. For example, length("abcde") is five. By contrast, length(15 * 35) works out to three. How? Well, 15 * 35 = 525, and 525 is then converted to the string "525", which has three characters. If no argument is supplied, length returns the length of $0. In older versions of awk, you could call the length function without any parentheses. Doing so is marked as “deprecated” in the POSIX standard. This means that while you can do this in your programs, it is a feature that can eventually be removed from a future version of the standard. Therefore, for maximal portability of your awk programs, you should always supply the parentheses.

match(string, regexp)

The match function searches the string, string, for the longest, leftmost substring matched by the regular expression, regexp. It returns the character position, or index, of where that substring begins (one, if it starts at the beginning of string). If no match is found, it returns zero. The match function sets the built-in variable RSTART to the index. It also sets the built-in variable RLENGTH to the length in characters of the matched substring. If no match is found, RSTART is set to zero, and RLENGTH to −1. For example:

    awk '{
           if ($1 == "FIND")
             regex = $2
           else {
             where = match($0, regex)
             if (where != 0)
               print "Match of", regex, "found at", \
                         where, "in", $0
           }
    }'

This program looks for lines that match the regular expression stored in the variable regex. This regular expression can be changed. If the first word on a line is ‘FIND’, regex is changed to be the second word on that line. Therefore, given:

    FIND ru+n
    My program runs
    but not very quickly
    FIND Melvin
    JF+KM
    This line is property of Reality Engineering Co.
    Melvin was here.

awk prints:

    Match of ru+n found at 12 in My program runs
    Match of Melvin found at 1 in Melvin was here.

split(string, array [, fieldsep]) This divides string into pieces separated by fieldsep, and stores the pieces in array. The first piece is stored in array[1], the second piece in array[2], and so forth. The string value of the third argument, fieldsep, is a regexp describing where to split string (much as FS can be a regexp describing where to split input records). If the fieldsep is omitted, the value of FS is used. split returns the number of elements created.

The split function splits strings into pieces in a manner similar to the way input lines are split into fields. For example:

    split("cul-de-sac", a, "-")

splits the string ‘cul-de-sac’ into three fields using ‘-’ as the separator. It sets the contents of the array a as follows:

    a[1] = "cul"
    a[2] = "de"
    a[3] = "sac"

The value returned by this call to split is three.

As with input field-splitting, when the value of fieldsep is " ", leading and trailing whitespace is ignored, and the elements are separated by runs of whitespace.

Also as with input field-splitting, if fieldsep is the null string, each individual character in the string is split into its own array element. (This is a gawk-specific extension.)

Recent implementations of awk, including gawk, allow the third argument to be a regexp constant (/abc/), as well as a string (d.c.). The POSIX standard allows this as well.

Before splitting the string, split deletes any previously existing elements in the array array (d.c.). sprintf(format, expression1,…)

This returns (without printing) the string that printf would have printed out with the same arguments (see Section 6.5 [Using printf Statements for Fancier Printing], page 64). For example:

    sprintf("pi = %.2f (approx.)", 22/7)
returns the string "pi = 3.14 (approx.)".

sub(regexp, replacement [, target]) The sub function alters the value of target. It searches this value, which is treated as a string, for the leftmost longest substring matched by the regular expression, regexp, extending this match as far as possible. Then the entire string is changed by replacing the matched text with replacement. The modified string becomes the new value of target. This function is peculiar because target is not simply used to compute a value, and not just any expression will do: it must be a variable, field or array element, so that sub can store a modified value there. If this argument is omitted, then the default is to use and alter $0. For example:

    str = "water, water, everywhere"
    sub(/at/, "ith", str)
sets str to "wither, water, everywhere", by replacing the

leftmost, longest occurrence of ‘at’ with ‘ith’. The sub function returns the number of substitutions made (either one or zero). If the special character ‘&’ appears in replacement, it stands for the precise substring that was matched by regexp. (If the regexp can match more than one string, then this precise substring may vary.) For example:

    awk '{ sub(/candidate/, "& and his wife"); print }'

changes the first occurrence of ‘candidate’ to ‘candidate and his wife’ on each input line. Here is another example:

    awk 'BEGIN {
            str = "daabaaa"
            sub(/a*/, "c&c", str)
            print str
    }'
    ⊣ dcaacbaaa

This shows how ‘&’ can represent a non-constant string, and also illustrates the “leftmost, longest” rule in regexp matching (see Section 4.6 [How Much Text Matches?], page 34). The effect of this special character (‘&’) can be turned off by putting a backslash before it in the string. As usual, to insert one backslash in the string, you must write two backslashes. Therefore, write ‘\\&’ in a string constant to include a literal ‘&’ in the replacement. For example, here is how to replace the first ‘|’ on each line with an ‘&’:

    awk '{ sub(/\|/, "\\&"); print }'

Note: As mentioned above, the third argument to sub must be a variable, field or array reference. Some versions of awk allow the third argument to be an expression which is not an lvalue. In such a case, sub would still search for the pattern and return zero or one, but the result of the substitution (if any) would be thrown away because there is no place to put it. Such versions of awk accept expressions like this:

    sub(/USA/, "United States", "the USA and Canada")

For historical compatibility, gawk will accept erroneous code, such as in the above example. However, using any other nonchangeable object as the third parameter will cause a fatal error, and your program will not run.

gsub(regexp, replacement [, target])

This is similar to the sub function, except gsub replaces all of the longest, leftmost, non-overlapping matching substrings it can find. The ‘g’ in gsub stands for “global,” which means replace everywhere. For example:

    awk '{ gsub(/Britain/, "United Kingdom"); print }'

replaces all occurrences of the string ‘Britain’ with ‘United Kingdom’ for all input records. The gsub function returns the number of substitutions made. If the variable to be searched and altered, target, is omitted, then the entire input record, $0, is used. As in sub, the characters ‘&’ and ‘\’ are special, and the third argument must be an lvalue.

gensub(regexp, replacement, how [, target]) gensub is a general substitution function. Like sub and gsub, it searches the target string target for matches of the regular expression regexp. Unlike sub and gsub, the modified string is returned as the result of the function, and the original target string is not changed. If how is a string beginning with ‘g’ or ‘G’, then it replaces all matches of regexp with replacement. Otherwise, how is a number indicating which match of regexp to replace. If no target is supplied, $0 is used instead. gensub provides an additional feature that is not available in sub or gsub: the ability to specify components of a regexp in the replacement text. This is done by using parentheses in the regexp to mark the components, and then specifying ‘\n’ in the replacement text, where n is a digit from one to nine. For example:

    $ gawk '
    > BEGIN {
    >      a = "abc def"
    >      b = gensub(/(.+) (.+)/, "\\2 \\1", "g", a)
    >      print b
    > }'
    ⊣ def abc

As described above for sub, you must type two backslashes in order to get one into the string. In the replacement text, the sequence ‘\0’ represents the entire matched text, as does the character ‘&’. This example shows how you can use the third argument to control which match of the regexp should be changed.

    $ echo a b c a b c |
    > gawk '{ print gensub(/a/, "AA", 2) }'
    ⊣ a b c AA b c

In this case, $0 is used as the default target string. gensub returns the new string as its result, which is passed directly to print for printing. If the how argument is a string that does not begin with ‘g’ or ‘G’, or if it is a number that is less than zero, only one substitution is performed. gensub is a gawk extension; it is not available in compatibility mode (see Section 14.1 [Command Line Options], page 159).

substr(string, start [, length]) This returns a length-character-long substring of string, starting at character number start. The first character of a string is character number one. For example, substr("washington",

5, 3) returns "ing".

If length is not present, this function returns the whole suffix of string that begins at character number start. For example,

substr("washington", 5) returns "ington". The whole suffix

is also returned if length is greater than the number of characters remaining in the string, counting from character number start. Note: The string returned by substr cannot be assigned to. Thus, it is a mistake to attempt to change a portion of a string, like this:

    string = "abcdef"
    # try to get "abCDEf", won't work
    substr(string, 3, 3) = "CDE"

or to use substr as the third agument of sub or gsub:

    gsub(/xyz/, "pdq", substr($0, 5, 20))  # WRONG
tolower(string)

This returns a copy of string, with each upper-case character in the string replaced with its corresponding lower-case character. Non-alphabetic characters are left unchanged. For example,

tolower("MiXeD cAsE 123") returns "mixed case 123".
toupper(string)

This returns a copy of string, with each lower-case character in the string replaced with its corresponding upper-case character. Non-alphabetic characters are left unchanged. For example,

toupper("MiXeD cAsE 123") returns "MIXED CASE 123".

More About ‘\’ and ‘&’ with sub, gsub and gensub

When using sub, gsub or gensub, and trying to get literal backslashes and ampersands into the replacement text, you need to remember that there are several levels of escape processing going on.

First, there is the lexical level, which is when awk reads your program, and builds an internal copy of your program that can be executed.

Then there is the run-time level, when awk actually scans the replacement string to determine what to generate.

At both levels, awk looks for a defined set of characters that can come after a backslash. At the lexical level, it looks for the escape sequences listed in Section 4.2 [Escape Sequences], page 24. Thus, for every ‘\’ that awk will process at the run-time level, you type two ‘\’s at the lexical level. When a character that is not valid for an escape sequence follows the ‘\’, Unix awk and gawk both simply remove the initial ‘\’, and put the following character into the string. Thus, for example, "a\qb" is treated as "aqb".

At the run-time level, the various functions handle sequences of ‘\’ and ‘&’ differently. The situation is (sadly) somewhat complex.

Historically, the sub and gsub functions treated the two character sequence ‘\&’ specially; this sequence was replaced in the generated text with a single ‘&’. Any other ‘\’ within the replacement string that did not precede an ‘&’ was passed through unchanged. To illustrate with a table:

You type  sub sees  sub generates
\&          &  the matched text
\\&         \&  a literal ‘&’
\\\&         \&  a literal ‘&’

   \\\\&        \\&  a literal ‘\&’
  \\\\\&        \\&  a literal ‘\&’
 \\\\\\&       \\\&  a literal ‘\\&’

\\q         \q  a literal ‘\q’

This table shows both the lexical level processing, where an odd number of backslashes becomes an even number at the run time level, and the run-time processing done by sub. (For the sake of simplicity, the rest of the tables below only show the case of even numbers of ‘\’s entered at the lexical level.)

The problem with the historical approach is that there is no way to get a literal ‘\’ followed by the matched text.

The 1992 POSIX standard attempted to fix this problem. The standard says that sub and gsub look for either a ‘\’ or an ‘&’ after the ‘\’. If either one follows a ‘\’, that character is output literally. The interpretation of ‘\’ and ‘&’ then becomes like this:

You type  sub sees  sub generates
&          &  the matched text
\\&         \&  a literal ‘&’
\\\\&        \\&  a literal ‘\’, then the matched text

 \\\\\\&       \\\&  a literal ‘\&’

This would appear to solve the problem. Unfortunately, the phrasing of the standard is unusual. It says, in effect, that ‘\’ turns off the special meaning of any following character, but that for anything other than ‘\’ and ‘&’, such special meaning is undefined. This wording leads to two problems.

1. Backslashes must now be doubled in the replacement string, breaking historical awk programs.

2. To make sure that an awk program is portable, every character in the replacement string must be preceded with a backslash.^{2}

2 This consequence was certainly unintended.

The POSIX standard is under revision.^{3} Because of the above problems, proposed text for the revised standard reverts to rules that correspond more closely to the original existing practice. The proposed rules have special cases that make it possible to produce a ‘\’ preceding the matched text.

You type  sub sees  sub generates

 \\\\\\&       \\\&  a literal ‘\&’

\\\\&        \\&  a literal ‘\’, followed by the matched text
\\&         \&  a literal ‘&’
\\q         \q  a literal ‘\q’

In a nutshell, at the run-time level, there are now three special sequences of characters, ‘\\\&’, ‘\\&’ and ‘\&’, whereas historically, there was only one. However, as in the historical case, any ‘\’ that is not part of one of these three sequences is not special, and appears in the output literally.

gawk 3.0 follows these proposed POSIX rules for sub and gsub. Whether these proposed rules will actually become codified into the standard is unknown at this point. Subsequent gawk releases will track the standard and implement whatever the final version specifies; this book will be updated as well.

The rules for gensub are considerably simpler. At the run-time level, whenever gawk sees a ‘\’, if the following character is a digit, then the text that matched the corresponding parenthesized subexpression is placed in the generated output. Otherwise, no matter what the character after the ‘\’ is, that character will appear in the generated text, and the ‘\’ will not.

You type  gensub sees  gensub generates
&             &  the matched text
\\&            \&  a literal ‘&’
\\\\            \\  a literal ‘\’
\\\\&           \\&  a literal ‘\’, then the matched text

 \\\\\\&          \\\&  a literal ‘\&’

\\q            \q  a literal ‘q’

Because of the complexity of the lexical and run-time level processing, and the special cases for sub and gsub, we recommend the use of gawk and gensub for when you have to do substitutions.

3 As of February 1997, with final approval and publication hopefully sometime in 1997.

12.4 Built-in Functions for Input/Output

The following functions are related to Input/Output (I/O). Optional parameters are enclosed in square brackets (“[” and “]”).

close(filename)

Close the file filename, for input or output. The argument may alternatively be a shell command that was used for redirecting to or from a pipe; then the pipe is closed. See Section 6.8 [Closing Input and Output Files and Pipes], page 74, for more information.

fflush([filename])

Flush any buffered output associated filename, which is either a file opened for writing, or a shell command for redirecting output to a pipe. Many utility programs will buffer their output; they save information to be written to a disk file or terminal in memory, until there is enough for it to be worthwhile to send the data to the ouput device. This is often more efficient than writing every little bit of information as soon as it is ready. However, sometimes it is necessary to force a program to flush its buffers; that is, write the information to its destination, even if a buffer is not full. This is the purpose of the fflush function; gawk too buffers its output, and the fflush function can be used to force gawk to flush its buffers. fflush is a recent (1994) addition to the Bell Labs research version of awk; it is not part of the POSIX standard, and will not be available if ‘--posix’ has been specified on the command line (see Section 14.1 [Command Line Options], page 159). gawk extends the fflush function in two ways. The first is to allow no argument at all. In this case, the buffer for the standard output is flushed. The second way is to allow the null string ("") as the argument. In this case, the buffers for all open output files and pipes are flushed. fflush returns zero if the buffer was successfully flushed, and nonzero otherwise.

system(command)

The system function allows the user to execute operating system commands and then return to the awk program. The system function executes the command given by the string command. It returns, as its value, the status returned by the command that was executed. For example, if the following fragment of code is put in your awk program:

    END {
         system("date | mail -s 'awk run done' root")
    }

the system administrator will be sent mail when the awk program finishes processing input and begins its end-of-input processing. Note that redirecting print or printf into a pipe is often enough to accomplish your task. However, if your awk program is interactive, system is useful for cranking up large self-contained programs, such as a shell or an editor. Some operating systems cannot implement the system function. system causes a fatal error if it is not supported.

Interactive vs. Non-Interactive Buffering

As a side point, buffering issues can be even more confusing depending upon whether or not your program is interactive, i.e., communicating with a user sitting at a keyboard.^{4}

Interactive programs generally line buffer their output; they write out every line. Non-interactive programs wait until they have a full buffer, which may be many lines of output.

Here is an example of the difference.

$ awk '{ print $1 + $2 }'
1 1
⊣ 2
2 3
⊣ 5
Control-d

Each line of output is printed immediately. Compare that behavior with this example.

$ awk '{ print $1 + $2 }' | cat
1 1
2 3
Control-d
⊣ 2
⊣ 5

Here, no output is printed until after the Control-d is typed, since it is all buffered, and sent down the pipe to cat in one shot.

4 A program is interactive if the standard output is connected to a terminal device.

Controlling Output Buffering with system

The fflush function provides explicit control over output buffering for individual files and pipes. However, its use is not portable to many other awk implementations. An alternative method to flush output buffers is by calling system with a null string as its argument:

    system("")   # flush output

gawk treats this use of the system function as a special case, and is smart enough not to run a shell (or other command interpreter) with the empty command. Therefore, with gawk, this idiom is not only useful, it is efficient. While this method should work with other awk implementations, it will not necessarily avoid starting an unnecessary shell. (Other implementations may only flush the buffer associated with the standard output, and not necessarily all buffered output.)

If you think about what a programmer expects, it makes sense that system should flush any pending output. The following program:

    BEGIN {
         print "first print"
         system("echo system echo")
         print "second print"
    }

must print

    first print
    system echo
    second print

and not

    system echo
    first print
    second print

If awk did not flush its buffers before calling system, the latter (undesirable) output is what you would see.

12.5 Functions for Dealing with Time Stamps

A common use for awk programs is the processing of log files containing time stamp information, indicating when a particular log record was written. Many programs log their time stamp in the form returned by the time system call, which is the number of seconds since a particular epoch. On POSIX systems, it is the number of seconds since Midnight, January 1, 1970, UTC.

In order to make it easier to process such log files, and to produce useful reports, gawk provides two functions for working with time stamps. Both of these are gawk extensions; they are not specified in the POSIX standard, nor are they in any other known version of awk.

Optional parameters are enclosed in square brackets (“[” and “]”).

systime()

This function returns the current time as the number of seconds since the system epoch. On POSIX systems, this is the number of seconds since Midnight, January 1, 1970, UTC. It may be a different number on other systems.

strftime([format [, timestamp]])

This function returns a string. It is similar to the function of the same name in ANSI C. The time specified by timestamp is used to produce a string, based on the contents of the format string. The timestamp is in the same format as the value returned by the systime function. If no timestamp argument is supplied, gawk will use the current time of day as the time stamp. If no format argument is supplied, strftime uses "%a %b %d %H:%M:%S %Z %Y". This format string produces output (almost) equivalent to that of the date utility. (Versions of gawk prior to 3.0 require the format argument.)

The systime function allows you to compare a time stamp from a log file with the current time of day. In particular, it is easy to determine how long ago a particular record was logged. It also allows you to produce log records using the “seconds since the epoch” format.

The strftime function allows you to easily turn a time stamp into human-readable information. It is similar in nature to the sprintf function (see Section 12.3 [Built-in Functions for String Manipulation], page 135), in that it copies non-format specification characters verbatim to the returned string, while substituting date and time values for format specifications in the format string.

strftime is guaranteed by the ANSI C standard to support the following date format specifications:

%a The locale’s abbreviated weekday name.

%A The locale’s full weekday name.

%b The locale’s abbreviated month name.

%B The locale’s full month name.

%c The locale’s “appropriate” date and time representation.

%d The day of the month as a decimal number (01–31).

%H The hour (24-hour clock) as a decimal number (00–23).

%I The hour (12-hour clock) as a decimal number (01–12).

%j The day of the year as a decimal number (001–366).

%m The month as a decimal number (01–12).

%M The minute as a decimal number (00–59).

%p The locale’s equivalent of the AM/PM designations associated with a 12-hour clock.

%S The second as a decimal number (00–60).^{5}

%U The week number of the year (the first Sunday as the first day of week one) as a decimal number (00–53).

%w The weekday as a decimal number (0–6). Sunday is day zero.

%W The week number of the year (the first Monday as the first day of week one) as a decimal number (00–53).

%x The locale’s “appropriate” date representation.

%X The locale’s “appropriate” time representation.

%y The year without century as a decimal number (00–99).

%Y The year with century as a decimal number (e.g., 1995).

%Z The time zone name or abbreviation, or no characters if no time zone is determinable.

%% A literal ‘%’.

If a conversion specifier is not one of the above, the behavior is undefined.^{6} Informally, a locale is the geographic place in which a program is meant

to run. For example, a common way to abbreviate the date September 4, 1991 in the United States would be “9/4/91”. In many countries in Europe, however, it would be abbreviated “4.9.91”. Thus, the ‘%x’ specification in a "US" locale might produce ‘9/4/91’, while in a "EUROPE" locale, it might produce ‘4.9.91’. The ANSI C standard defines a default "C" locale, which is an environment that is typical of what most C programmers are used to.

A public-domain C version of strftime is supplied with gawk for systems that are not yet fully ANSI-compliant. If that version is used to compile gawk (see Appendix B [Installing gawk], page 277), then the following additional format specifications are available:

5 Occasionally there are minutes in a year with a leap second, which is why

the seconds can go up to 60. 6 This is because ANSI C leaves the behavior of the C version of strftime

undefined, and gawk will use the system’s version of strftime if it’s there. Typically, the conversion specifier will either not appear in the returned string, or it will appear literally.

%D Equivalent to specifying ‘%m/%d/%y’.

%e The day of the month, padded with a space if it is only one digit.

%h Equivalent to ‘%b’, above.

%n A newline character (ASCII LF).

%r Equivalent to specifying ‘%I:%M:%S %p’.

%R Equivalent to specifying ‘%H:%M’.

%T Equivalent to specifying ‘%H:%M:%S’.

%t A tab character.

%k The hour (24-hour clock) as a decimal number (0-23). Single digit numbers are padded with a space.

%l The hour (12-hour clock) as a decimal number (1-12). Single digit numbers are padded with a space.

%C The century, as a number between 00 and 99.

%u The weekday as a decimal number [1 (Monday)–7].

%V The week number of the year (the first Monday as the first day of week one) as a decimal number (01–53). The method for determining the week number is as specified by ISO 8601 (to wit: if the week containing January 1 has four or more days in the new year, then it is week one, otherwise it is week 53 of the previous year and the next week is week one).

%G The year with century of the ISO week number, as a decimal number. For example, January 1, 1993, is in week 53 of 1992. Thus, the year of its ISO week number is 1992, even though its year is 1993. Similarly, December 31, 1973, is in week 1 of 1974. Thus, the year of its ISO week number is 1974, even though its year is 1973.

%g The year without century of the ISO week number, as a decimal number (00–99).

%Ec %EC %Ex %Ey %EY %Od %Oe %OH %OI
%Om %OM %OS %Ou %OU %OV %Ow %OW %Oy

These are “alternate representations” for the specifications that use only the second letter (‘%c’, ‘%C’, and so on). They are recognized, but their normal representations are used.^{7} (These facilitate compliance with the POSIX date utility.)

7 If you don’t understand any of this, don’t worry about it; these facilities are meant to make it easier to “internationalize” programs.

%v The date in VMS format (e.g., 20-JUN-1991).

%z The timezone offset in a +HHMM format (e.g., the format necessary to produce RFC-822/RFC-1036 date headers).

This example is an awk implementation of the POSIX date utility. Normally, the date utility prints the current date and time of day in a well known format. However, if you provide an argument to it that begins with a ‘+’, date will copy non-format specifier characters to the standard output, and will interpret the current time according to the format specifiers in the string. For example:

    $ date '+Today is %A, %B %d, %Y.'
    ⊣ Today is Thursday, July 11, 1991.

Here is the gawk version of the date utility. It has a shell “wrapper”, to handle the ‘-u’ option, which requires that date run as if the time zone was set to UTC.

    #! /bin/sh
    #
    # date --- approximate the P1003.2 'date' command
    case $1 in
    -u)  TZ=GMT0     # use UTC
         export TZ
         shift ;;
    esac
    gawk 'BEGIN  {
        format = "%a %b %d %H:%M:%S %Z %Y"
        exitval = 0
        if (ARGC > 2)
            exitval = 1
        else if (ARGC == 2) {
            format = ARGV[1]
            if (format ~ /^\+/)
                format = substr(format, 2)   # remove leading +
        }
        print strftime(format)
        exit exitval
    }' "$@"

Chapter 13: User-defined Functions 151

13 User-defined Functions

Complicated awk programs can often be simplified by defining your own functions. User-defined functions can be called just like built-in ones (see Section 7.13 [Function Calls], page 93), but it is up to you to define them—to tell awk what they should do.

13.1 Function Definition Syntax

Definitions of functions can appear anywhere between the rules of an awk program. Thus, the general form of an awk program is extended to include sequences of rules and user-defined function definitions. There is no need in awk to put the definition of a function before all uses of the function. This is because awk reads the entire program before starting to execute any of it.

The definition of a function named name looks like this:

function name(parameter-list)

    {

body-of-function

    }

name is the name of the function to be defined. A valid function name is like a valid variable name: a sequence of letters, digits and underscores, not starting with a digit. Within a single awk program, any particular name can only be used as a variable, array or function.

parameter-list is a list of the function’s arguments and local variable names, separated by commas. When the function is called, the argument names are used to hold the argument values given in the call. The local variables are initialized to the empty string. A function cannot have two parameters with the same name.

The body-of-function consists of awk statements. It is the most important part of the definition, because it says what the function should actually do. The argument names exist to give the body a way to talk about the arguments; local variables, to give the body places to keep temporary values.

Argument names are not distinguished syntactically from local variable names; instead, the number of arguments supplied when the function is called determines how many argument variables there are. Thus, if three argument values are given, the first three names in parameter-list are arguments, and the rest are local variables.

It follows that if the number of arguments is not the same in all calls to the function, some of the names in parameter-list may be arguments on some occasions and local variables on others. Another way to think of this is that omitted arguments default to the null string.

Usually when you write a function you know how many names you intend to use for arguments and how many you intend to use as local variables. It is conventional to place some extra space between the arguments and the local variables, to document how your function is supposed to be used.

During execution of the function body, the arguments and local variable values hide or shadow any variables of the same names used in the rest of the program. The shadowed variables are not accessible in the function definition, because there is no way to name them while their names have been taken away for the local variables. All other variables used in the awk program can be referenced or set normally in the function’s body.

The arguments and local variables last only as long as the function body is executing. Once the body finishes, you can once again access the variables that were shadowed while the function was running.

The function body can contain expressions which call functions. They can even call this function, either directly or by way of another function. When this happens, we say the function is recursive.

In many awk implementations, including gawk, the keyword function may be abbreviated func. However, POSIX only specifies the use of the keyword function. This actually has some practical implications. If gawk is in POSIX-compatibility mode (see Section 14.1 [Command Line Options], page 159), then the following statement will not define a function:

    func foo() { a = sqrt($1) ; print a }

Instead it defines a rule that, for each record, concatenates the value of the variable ‘func’ with the return value of the function ‘foo’. If the resulting string is non-null, the action is executed. This is probably not what was desired. (awk accepts this input as syntactically valid, since functions may be used before they are defined in awk programs.)

To ensure that your awk programs are portable, always use the keyword function when defining a function.

13.2 Function Definition Examples

Here is an example of a user-defined function, called myprint, that takes a number and prints it in a specific format.

    function myprint(num)
    {
         printf "%6.3g\n", num
    }

To illustrate, here is an awk rule which uses our myprint function:

    $3 > 0     { myprint($3) }

This program prints, in our special format, all the third fields that contain a positive number in our input. Therefore, when given:

     1.2   3.4    5.6   7.8
     9.10 11.12 -13.14 15.16
    17.18 19.20  21.22 23.24

this program, using our function to format the results, prints:

       5.6
      21.2

This function deletes all the elements in an array.

    function delarray(a,    i)
    {
        for (i in a)
           delete a[i]
    }

When working with arrays, it is often necessary to delete all the elements in an array and start over with a new list of elements (see Section 11.6 [The delete Statement], page 128). Instead of having to repeat this loop everywhere in your program that you need to clear out an array, your program can just call delarray.

Here is an example of a recursive function. It takes a string as an input parameter, and returns the string in backwards order.

    function rev(str, start)
    {
        if (start == 0)
            return ""
        return (substr(str, start, 1) rev(str, start - 1))
    }

If this function is in a file named ‘rev.awk’, we can test it this way:

    $ echo "Don't Panic!" |
    > gawk --source '{ print rev($0, length($0)) }' -f rev.awk
    ⊣ !cinaP t'noD

Here is an example that uses the built-in function strftime. (See Section 12.5 [Functions for Dealing with Time Stamps], page 146, for more information on strftime.) The C ctime function takes a timestamp and returns it in a string, formatted in a well known fashion. Here is an awk version:

    # ctime.awk
    #
    # awk version of C ctime(3) function
    function ctime(ts,    format)
    {
        format = "%a %b %d %H:%M:%S %Z %Y"
        if (ts == 0)
            ts = systime()       # use current time as default
        return strftime(format, ts)
    }

13.3 Calling User-defined Functions

Calling a function means causing the function to run and do its job. A function call is an expression, and its value is the value returned by the function.

A function call consists of the function name followed by the arguments in parentheses. What you write in the call for the arguments are awk expressions; each time the call is executed, these expressions are evaluated, and the values are the actual arguments. For example, here is a call to foo with three arguments (the first being a string concatenation):

    foo(x y, "lose", 4 * z)

Caution: whitespace characters (spaces and tabs) are not allowed between the function name and the open-parenthesis of the argument list. If you write whitespace by mistake, awk might think that you mean to concatenate a variable with an expression in parentheses. However, it notices that you used a function name and not a variable name, and reports an error.

When a function is called, it is given a copy of the values of its arguments. This is known as call by value. The caller may use a variable as the expression for the argument, but the called function does not know this: it only knows what value the argument had. For example, if you write this code:

    foo = "bar"
    z = myfunc(foo)

then you should not think of the argument to myfunc as being “the variable foo.” Instead, think of the argument as the string value, "bar".

If the function myfunc alters the values of its local variables, this has no effect on any other variables. Thus, if myfunc does this:

    function myfunc(str)
    {
      print str
      str = "zzz"
      print str
    }

to change its first argument variable str, this does not change the value of foo in the caller. The role of foo in calling myfunc ended when its value, "bar", was computed. If str also exists outside of myfunc, the function body cannot alter this outer value, because it is shadowed during the execution of myfunc and cannot be seen or changed from there.

However, when arrays are the parameters to functions, they are not copied. Instead, the array itself is made available for direct manipulation by the function. This is usually called call by reference. Changes made to an array parameter inside the body of a function are visible outside that function. This can be very dangerous if you do not watch what you are doing. For example:

    function changeit(array, ind, nvalue)
    {
         array[ind] = nvalue
    }
    BEGIN {
        a[1] = 1; a[2] = 2; a[3] = 3
        changeit(a, 2, "two")
        printf "a[1] = %s, a[2] = %s, a[3] = %s\n",
                a[1], a[2], a[3]
    }
This program prints ‘a[1] = 1, a[2] = two, a[3] = 3’, because changeit

stores "two" in the second element of a.

Some awk implementations allow you to call a function that has not been defined, and only report a problem at run-time when the program actually tries to call the function. For example:

    BEGIN {
        if (0)
            foo()
        else
            bar()
    }
    function bar() { . . . }
    # note that `foo' is not defined

Since the ‘if’ statement will never be true, it is not really a problem that foo has not been defined. Usually though, it is a problem if a program calls an undefined function.

If ‘--lint’ has been specified (see Section 14.1 [Command Line Options], page 159), gawk will report about calls to undefined functions.

Some awk implementations generate a run-time error if you use the next statement (see Section 9.7 [The next Statement], page 111) inside a userdefined function. gawk does not have this problem.

13.4 The return Statement

The body of a user-defined function can contain a return statement. This statement returns control to the rest of the awk program. It can also be used to return a value for use in the rest of the awk program. It looks like this:

return [expression]

The expression part is optional. If it is omitted, then the returned value is undefined and, therefore, unpredictable.

A return statement with no value expression is assumed at the end of every function definition. So if control reaches the end of the function body, then the function returns an unpredictable value. awk will not warn you if you use the return value of such a function.

Sometimes, you want to write a function for what it does, not for what it returns. Such a function corresponds to a void function in C or to a procedure in Pascal. Thus, it may be appropriate to not return any value; you should simply bear in mind that if you use the return value of such a function, you do so at your own risk.

Here is an example of a user-defined function that returns a value for the largest number among the elements of an array:

    function maxelt(vec,   i, ret)
    {
         for (i in vec) {
              if (ret == "" || vec[i] > ret)
                   ret = vec[i]
         }
         return ret
    }

You call maxelt with one argument, which is an array name. The local variables i and ret are not intended to be arguments; while there is nothing to stop you from passing two or three arguments to maxelt, the results would be strange. The extra space before i in the function parameter list indicates that i and ret are not supposed to be arguments. This is a convention that you should follow when you define functions.

Here is a program that uses our maxelt function. It loads an array, calls maxelt, and then reports the maximum number in that array:

awk '
function maxelt(vec,   i, ret)
{
     for (i in vec) {
          if (ret == "" || vec[i] > ret)
               ret = vec[i]
     }
     return ret
}
# Load all fields of each record into nums.
{
     for(i = 1; i <= NF; i++)
          nums[NR, i] = $i
}
END {
     print maxelt(nums)
}'

Given the following input:

 1 5 23 8 16
44 3 5 2 8 26
256 291 1396 2962 100
-6 467 998 1101
99385 11 0 225

our program tells us (predictably) that 99385 is the largest number in our array.

158 Effective AWK Programming

Chapter 14: Running awk 159

14 Running awk

There are two ways to run awk: with an explicit program, or with one or more program files. Here are templates for both of them; items enclosed in ‘[…]’ in these templates are optional.

Besides traditional one-letter POSIX-style options, gawk also supports GNU long options.

awk [options] -f progfile [--] file … awk [options] [--] 'program' file … It is possible to invoke awk with an empty program:

    $ awk '' datafile1 datafile2

Doing so makes little sense though; awk will simply exit silently when given an empty program (d.c.). If ‘--lint’ has been specified on the command line, gawk will issue a warning that the program is empty.

14.1 Command Line Options

Options begin with a dash, and consist of a single character. GNU style long options consist of two dashes and a keyword. The keyword can be abbreviated, as long the abbreviation allows the option to be uniquely identified. If the option takes an argument, then the keyword is either immediately followed by an equals sign (‘=’) and the argument’s value, or the keyword and the argument’s value are separated by whitespace. For brevity, the discussion below only refers to the traditional short options; however the long and short options are interchangeable in all contexts.

Each long option for gawk has a corresponding POSIX-style option. The options and their meanings are as follows:

-F fs
--field-separator fs

Sets the FS variable to fs (see Section 5.5 [Specifying How Fields are Separated], page 44).

-f source-file --file source-file

Indicates that the awk program is to be found in source-file instead of in the first non-option argument.

-v var=val

--assign var=val

Sets the variable var to the value val before execution of the program begins. Such variable values are available inside the BEGIN rule (see Section 14.2 [Other Command Line Arguments], page 163).

The ‘-v’ option can only set one variable, but you can use it more than once, setting another variable each time, like this:

‘awk -v foo=1 -v bar=2 . . .’.
-mf NNN

-mr NNN Set various memory limits to the value NNN. The ‘f’ flag sets the maximum number of fields, and the ‘r’ flag sets the maximum record size. These two flags and the ‘-m’ option are from the Bell Labs research version of Unix awk. They are provided for compatibility, but otherwise ignored by gawk, since gawk has no predefined limits.

-W gawk-opt Following the POSIX standard, options that are implementation specific are supplied as arguments to the ‘-W’ option. These options also have corresponding GNU style long options. See below.

-- Signals the end of the command line options. The following arguments are not treated as options even if they begin with ‘-’. This interpretation of ‘--’ follows the POSIX argument parsing conventions.

This is useful if you have file names that start with ‘-’, or in shell scripts, if you have file names that will be specified by the user which could start with ‘-’.

The following gawk-specific options are available:

-W traditional
-W compat
--traditional

--compat Specifies compatibility mode, in which the GNU extensions to the awk language are disabled, so that gawk behaves just like the Bell Labs research version of Unix awk. ‘--traditional’ is the preferred form of this option. See Section 17.5 [Extensions in gawk Not in POSIX awk], page 254, which summarizes the extensions. Also see Section C.1 [Downward Compatibility and Debugging], page 293.

-W copyleft
-W copyright
--copyleft
--copyright

Print the short version of the General Public License, and then exit. This option may disappear in a future version of gawk.

-W help
-W usage
--help

--usage Print a “usage” message summarizing the short and long style options that gawk accepts, and then exit.

-W lint

--lint Warn about constructs that are dubious or non-portable to other awk implementations. Some warnings are issued when gawk first reads your program. Others are issued at run-time, as your program executes.

-W lint-old
--lint-old

Warn about constructs that are not available in the original Version 7 Unix version of awk (see Section 17.1 [Major Changes between V7 and SVR3.1], page 251).

-W posix

--posix Operate in strict POSIX mode. This disables all gawk extensions (just like ‘--traditional’), and adds the following additional restrictions:

• \x escape sequences are not recognized (see Section 4.2 [Escape Sequences], page 24).

• Newlines do not act as whitespace to separate fields when FS is equal to a single space.

• The synonym func for the keyword function is not recognized (see Section 13.1 [Function Definition Syntax], page 151).

• The operators ‘**’ and ‘**=’ cannot be used in place of ‘^’ and ‘^=’ (see Section 7.5 [Arithmetic Operators], page 82, and also see Section 7.7 [Assignment Expressions], page 84).

• Specifying ‘-Ft’ on the command line does not set the value of FS to be a single tab character (see Section 5.5 [Specifying How Fields are Separated], page 44).

• The fflush built-in function is not supported (see Section 12.4 [Built-in Functions for Input/Output], page 144).

If you supply both ‘--traditional’ and ‘--posix’ on the command line, ‘--posix’ will take precedence. gawk will also issue a warning if both options are supplied.

-W re-interval
--re-interval

Allow interval expressions (see Section 4.3 [Regular Expression Operators], page 26), in regexps. Because interval expressions were traditionally not available in awk, gawk does not provide them by default. This prevents old awk programs from breaking.

-W source program-text --source program-text

Program source code is taken from the program-text. This option allows you to mix source code in files with source code that you enter on the command line. This is particularly useful when you have library functions that you wish to use from your command line programs (see Section 14.3 [The AWKPATH Environment Variable], page 164).

-W version
--version

Prints version information for this particular copy of gawk. This allows you to determine if your copy of gawk is up to date with respect to whatever the Free Software Foundation is currently distributing. It is also useful for bug reports (see Section B.7 [Reporting Problems and Bugs], page 290).

Any other options are flagged as invalid with a warning message, but are otherwise ignored.

In compatibility mode, as a special case, if the value of fs supplied to the ‘-F’ option is ‘t’, then FS is set to the tab character ("\t"). This is only true for ‘--traditional’, and not for ‘--posix’ (see Section 5.5 [Specifying How Fields are Separated], page 44).

The ‘-f’ option may be used more than once on the command line. If it is, awk reads its program source from all of the named files, as if they had been concatenated together into one big file. This is useful for creating libraries of awk functions. Useful functions can be written once, and then retrieved from a standard place, instead of having to be included into each individual program.

You can type in a program at the terminal and still use library functions, by specifying ‘-f /dev/tty’. awk will read a file from the terminal to use as part of the awk program. After typing your program, type Control-d (the end-of-file character) to terminate it. (You may also use ‘-f -’ to read program source from the standard input, but then you will not be able to also use the standard input as a source of data.)

Because it is clumsy using the standard awk mechanisms to mix source file and command line awk programs, gawk provides the ‘--source’ option. This does not require you to pre-empt the standard input for your source code, and allows you to easily mix command line and library source code (see Section 14.3 [The AWKPATH Environment Variable], page 164).

If no ‘-f’ or ‘--source’ option is specified, then gawk will use the first non-option command line argument as the text of the program source code.

If the environment variable POSIXLY_CORRECT exists, then gawk will behave in strict POSIX mode, exactly as if you had supplied the ‘--posix’ command line option. Many GNU programs look for this environment variable to turn on strict POSIX mode. If you supply ‘--lint’ on the command line, and gawk turns on POSIX mode because of POSIXLY_CORRECT, then it will print a warning message indicating that POSIX mode is in effect.

You would typically set this variable in your shell’s startup file. For a Bourne compatible shell (such as Bash), you would add these lines to the ‘.profile’ file in your home directory.

    POSIXLY_CORRECT=true
    export POSIXLY_CORRECT

For a csh compatible shell,^{1} you would add this line to the ‘.login’ file in your home directory.

    setenv POSIXLY_CORRECT true

14.2 Other Command Line Arguments

Any additional arguments on the command line are normally treated as input files to be processed in the order specified. However, an argument that has the form var=value, assigns the value value to the variable var—it does not specify a file at all.

All these arguments are made available to your awk program in the ARGV array (see Chapter 10 [Built-in Variables], page 115). Command line options and the program text (if present) are omitted from ARGV. All other arguments, including variable assignments, are included. As each element of ARGV is processed, gawk sets the variable ARGIND to the index in ARGV of the current element.

The distinction between file name arguments and variable-assignment arguments is made when awk is about to open the next input file. At that point in execution, it checks the “file name” to see whether it is really a variable assignment; if so, awk sets the variable instead of reading a file.

Therefore, the variables actually receive the given values after all previously specified files have been read. In particular, the values of variables assigned in this fashion are not available inside a BEGIN rule (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100), since such rules are run before awk begins scanning the argument list.

The variable values given on the command line are processed for escape sequences (d.c.) (see Section 4.2 [Escape Sequences], page 24).

In some earlier implementations of awk, when a variable assignment occurred before any file names, the assignment would happen before the BEGIN rule was executed. awk’s behavior was thus inconsistent; some command

1 Not recommended. line assignments were available inside the BEGIN rule, while others were not. However, some applications came to depend upon this “feature.” When awk was changed to be more consistent, the ‘-v’ option was added to accommodate applications that depended upon the old behavior.

The variable assignment feature is most useful for assigning to variables such as RS, OFS, and ORS, which control input and output formats, before scanning the data files. It is also useful for controlling state if multiple passes are needed over a data file. For example:

    awk 'pass == 1  { pass 1 stuff }
         pass == 2  { pass 2 stuff }' pass=1 mydata pass=2 mydata

Given the variable assignment feature, the ‘-F’ option for setting the value of FS is not strictly necessary. It remains for historical compatibility.

14.3 The AWKPATH Environment Variable

The previous section described how awk program files can be named on the command line with the ‘-f’ option. In most awk implementations, you must supply a precise path name for each program file, unless the file is in the current directory.

But in gawk, if the file name supplied to the ‘-f’ option does not contain a ‘/’, then gawk searches a list of directories (called the search path), one by one, looking for a file with the specified name.

The search path is a string consisting of directory names separated by colons. gawk gets its search path from the AWKPATH environment variable. If that variable does not exist, gawk uses a default path, which is ‘.:/usr/local/share/awk’.^{2} (Programs written for use by system administrators should use an AWKPATH variable that does not include the current directory, ‘.’.)

The search path feature is particularly useful for building up libraries of useful awk functions. The library files can be placed in a standard directory that is in the default path, and then specified on the command line with a short file name. Otherwise, the full file name would have to be typed for each file.

By using both the ‘--source’ and ‘-f’ options, your command line awk programs can use facilities in awk library files. See Chapter 15 [A Library of awk Functions], page 167.

2 Your version of gawk may use a directory that is different than ‘/usr/local/share/awk’; it will depend upon how gawk was built and installed. The actual directory will be the value of ‘$(datadir)’ generated when gawk was configured. You probably don’t need to worry about this though.

Path searching is not done if gawk is in compatibility mode. This is true for both ‘--traditional’ and ‘--posix’. See Section 14.1 [Command Line Options], page 159.

Note: if you want files in the current directory to be found, you must include the current directory in the path, either by including ‘.’ explicitly in the path, or by writing a null entry in the path. (A null entry is indicated by starting or ending the path with a colon, or by placing two colons next to each other (‘::’).) If the current directory is not included in the path, then files cannot be found in the current directory. This path search mechanism is identical to the shell’s.

Starting with version 3.0, if AWKPATH is not defined in the environment, gawk will place its default search path into ENVIRON["AWKPATH"]. This makes it easy to determine the actual search path gawk will use.

14.4 Obsolete Options and/or Features

This section describes features and/or command line options from previous releases of gawk that are either not available in the current version, or that are still supported but deprecated (meaning that they will not be in the next release).

For version 3.0.3 of gawk, there are no command line options or other deprecated features from the previous version of gawk. This section is thus essentially a place holder, in case some option becomes obsolete in a future version of gawk.

14.5 Undocumented Options and Features

Use the Source, Luke!

Obi-Wan

This section intentionally left blank.

14.6 Known Bugs in gawk

• The ‘-F’ option for changing the value of FS (see Section 14.1 [Command Line Options], page 159) is not necessary given the command line variable assignment feature; it remains only for backwards compatibility.

• If your system actually has support for ‘/dev/fd’ and the associated

    ‘/dev/stdin’, ‘/dev/stdout’, and ‘/dev/stderr’ files, you may get dif-

ferent output from gawk than you would get on a system without those files. When gawk interprets these files internally, it synchronizes output to the standard output with output to ‘/dev/stdout’, while on a system with those files, the output is actually to different open files (see Section 6.7 [Special File Names in gawk], page 72).

• Syntactically invalid single character programs tend to overflow the parse stack, generating a rather unhelpful message. Such programs are surprisingly difficult to diagnose in the completely general case, and the effort to do so really is not worth it.

Chapter 15: A Library of awk Functions 167

15 A Library of awk Functions

This chapter presents a library of useful awk functions. The sample programs presented later (see Chapter 16 [Practical awk Programs], page 203) use these functions. The functions are presented here in a progression from simple to complex.

Section 16.2.7 [Extracting Programs from Texinfo Source Files], page 236, presents a program that you can use to extract the source code for these example library functions and programs from the Texinfo source for this book. (This has already been done as part of the gawk distribution.)

If you have written one or more useful, general purpose awk functions, and would like to contribute them for a subsequent edition of this book, please contact the author. See Section B.7 [Reporting Problems and Bugs], page 290, for information on doing this. Don’t just send code, as you will be required to either place your code in the public domain, publish it under the GPL (see [GNU GENERAL PUBLIC LICENSE], page 307), or assign the copyright in it to the Free Software Foundation.

15.1 Simulating gawk-specific Features

The programs in this chapter and in Chapter 16 [Practical awk Programs], page 203, freely use features that are specific to gawk. This section briefly discusses how you can rewrite these programs for different implementations

of awk.

Diagnostic error messages are sent to ‘/dev/stderr’. Use ‘| "cat 1>&2"’ instead of ‘> "/dev/stderr"’, if your system does not have a ‘/dev/stderr’, or if you cannot use gawk.

A number of programs use nextfile (see Section 9.8 [The nextfile Statement], page 112), to skip any remaining input in the input file. Section 15.2 [Implementing nextfile as a Function], page 168, shows you how to write a function that will do the same thing.

Finally, some of the programs choose to ignore upper-case and lower-case distinctions in their input. They do this by assigning one to IGNORECASE. You can achieve the same effect by adding the following rule to the beginning of the program:

    # ignore case
    { $0 = tolower($0) }

Also, verify that all regexp and string constants used in comparisons only use lower-case letters.

15.2 Implementing nextfile as a Function

The nextfile statement presented in Section 9.8 [The nextfile Statement], page 112, is a gawk-specific extension. It is not available in other implementations of awk. This section shows two versions of a nextfile function that you can use to simulate gawk’s nextfile statement if you cannot use gawk.

Here is a first attempt at writing a nextfile function.

    # nextfile --- skip remaining records in current file
    # this should be read in before the "main" awk program
    function nextfile()    { _abandon_= FILENAME; next }
    _abandon_== FILENAME  { next }

This file should be included before the main program, because it supplies a rule that must be executed first. This rule compares the current data file’s name (which is always in the FILENAME variable) to a private variable named _abandon_. If the file name matches, then the action part of the rule executes a next statement, to go on to the next record. (The use of ‘_’ in the variable name is a convention. It is discussed more fully in Section 15.13 [Naming Library Function Global Variables], page 200.)

The use of the next statement effectively creates a loop that reads all the records from the current data file. Eventually, the end of the file is reached, and a new data file is opened, changing the value of FILENAME. Once this happens, the comparison of _abandon_to FILENAME fails, and execution continues with the first rule of the “real” program.

The nextfile function itself simply sets the value of _abandon_and then executes a next statement to start the loop going.^{1}

This initial version has a subtle problem. What happens if the same data file is listed twice on the command line, one right after the other, or even with just a variable assignment between the two occurrences of the file name?

In such a case, this code will skip right through the file, a second time, even though it should stop when it gets to the end of the first occurrence. Here is a second version of nextfile that remedies this problem.

1 Some implementations of awk do not allow you to execute next from within a function body. Some other work-around will be necessary if you use such a version.

    # nextfile --- skip remaining records in current file
    # correctly handle successive occurrences of the same file
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May, 1993
    # this should be read in before the "main" awk program
    function nextfile()   { _abandon_= FILENAME; next }
    _abandon_== FILENAME {
          if (FNR == 1)
              _abandon_= ""
          else
              next
    }

The nextfile function has not changed. It sets _abandon_equal to the current file name and then executes a next satement. The next statement reads the next record and increments FNR, so FNR is guaranteed to have a value of at least two. However, if nextfile is called for the last record in the file, then awk will close the current data file and move on to the next one. Upon doing so, FILENAME will be set to the name of the new file, and FNR will be reset to one. If this next file is the same as the previous one, _abandon_ will still be equal to FILENAME. However, FNR will be equal to one, telling us that this is a new occurrence of the file, and not the one we were reading when the nextfile function was executed. In that case, _abandon_is reset to the empty string, so that further executions of this rule will fail (until the next time that nextfile is called).

If FNR is not one, then we are still in the original data file, and the program executes a next statement to skip through it.

An important question to ask at this point is: “Given that the functionality of nextfile can be provided with a library file, why is it built into gawk?” This is an important question. Adding features for little reason leads to larger, slower programs that are harder to maintain.

The answer is that building nextfile into gawk provides significant gains in efficiency. If the nextfile function is executed at the beginning of a large data file, awk still has to scan the entire file, splitting it up into records, just to skip over it. The built-in nextfile can simply close the file immediately and proceed to the next one, saving a lot of time. This is particularly important in awk, since awk programs are generally I/O bound (i.e. they spend most of their time doing input and output, instead of performing computations).

15.3 Assertions

When writing large programs, it is often useful to be able to know that a condition or set of conditions is true. Before proceeding with a particular computation, you make a statement about what you believe to be the case. Such a statement is known as an “assertion.” The C language provides an <assert.h> header file and corresponding assert macro that the programmer can use to make assertions. If an assertion fails, the assert macro arranges to print a diagnostic message describing the condition that should have been true but was not, and then it kills the program. In C, using assert looks this:

#include <assert.h>
int myfunc(int a, double b)
{
     assert(a <= 5 && b >= 17);

…

}

If the assertion failed, the program would print a message similar to this:

prog.c:5: assertion failed: a <= 5 && b >= 17

The ANSI C language makes it possible to turn the condition into a string for use in printing the diagnostic message. This is not possible in awk, so this assert function also requires a string version of the condition that is being tested.

# assert --- assert that a condition is true. Otherwise exit.
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
# May, 1993
function assert(condition, string)
{
    if (! condition) {
        printf("%s:%d: assertion failed: %s\n",
            FILENAME, FNR, string) > "/dev/stderr"
        _assert_exit = 1
        exit 1
    }
}
END {
    if (_assert_exit)
        exit 1
}

The assert function tests the condition parameter. If it is false, it prints a message to standard error, using the string parameter to describe the failed condition. It then sets the variable _assert_exit to one, and executes the exit statement. The exit statement jumps to the END rule. If the END rules finds _assert_exit to be true, then it exits immediately.

The purpose of the END rule with its test is to keep any other END rules from running. When an assertion fails, the program should exit immediately. If no assertions fail, then _assert_exit will still be false when the END rule is run normally, and the rest of the program’s END rules will execute. For all of this to work correctly, ‘assert.awk’ must be the first source file read by

awk.

You would use this function in your programs this way:

    function myfunc(a, b)
    {
         assert(a <= 5 && b >= 17, "a <= 5 && b >= 17")

…

    }

If the assertion failed, you would see a message like this:

    mydata:1357: assertion failed: a <= 5 && b >= 17

There is a problem with this version of assert, that it may not be possible to work around. An END rule is automatically added to the program calling assert. Normally, if a program consists of just a BEGIN rule, the input files and/or standard input are not read. However, now that the program has an END rule, awk will attempt to read the input data files, or standard input (see Section 8.1.5.1 [Startup and Cleanup Actions], page 100), most likely causing the program to hang, waiting for input.

15.4 Rounding Numbers

The way printf and sprintf (see Section 6.5 [Using printf Statements for Fancier Printing], page 64) do rounding will often depend upon the system’s C sprintf subroutine. On many machines, sprintf rounding is “unbiased,” which means it doesn’t always round a trailing ‘.5’ up, contrary to naive expectations. In unbiased rounding, ‘.5’ rounds to even, rather than always up, so 1.5 rounds to 2 but 4.5 rounds to 4. The result is that if you are using a format that does rounding (e.g., "%.0f") you should check what your system does. The following function does traditional rounding; it might be useful if your awk’s printf does unbiased rounding.

    # round --- do normal rounding
    #
    # Arnold Robbins, arnold@gnu.ai.mit.edu, August, 1996
    # Public Domain

function round(x,   ival, aval, fraction)
{
ival = int(x)    # integer part, int() truncates
# see if fractional part
if (ival == x)   # no fraction
return x
if (x < 0) {
aval = -x     # absolute value
ival = int(aval)
fraction = aval - ival
if (fraction >= .5)
   return int(x) - 1   # -2.5 --> -3
else
   return int(x)       # -2.3 --> -2
} else {
fraction = x - ival
if (fraction >= .5)
   return ival + 1
else
   return ival
}
}
# test harness
{ print $0, round($0) }

15.5 Translating Between Characters and Numbers

One commercial implementation of awk supplies a built-in function, ord, which takes a character and returns the numeric value for that character in the machine’s character set. If the string passed to ord has more than one character, only the first one is used.

The inverse of this function is chr (from the function of the same name in Pascal), which takes a number and returns the corresponding character.

Both functions can be written very nicely in awk; there is no real reason to build them into the awk interpreter.

# ord.awk --- do ord and chr
#
# Global identifiers:
#    _ord_:        numerical values indexed by characters
#    _ord_init:    function to initialize _ord_
#
# Arnold Robbins
# arnold@gnu.ai.mit.edu
# Public Domain
# 16 January, 1992
# 20 July, 1992, revised
BEGIN    { _ord_init() }
function _ord_init(    low, high, i, t)
{
    low = sprintf("%c", 7) # BEL is ascii 7
    if (low == "\a") {    # regular ascii
        low = 0
        high = 127
    } else if (sprintf("%c", 128 + 7) == "\a") {
        # ascii, mark parity
        low = 128
        high = 255
    } else {        # ebcdic(!)
        low = 0
        high = 255
    }
    for (i = low; i <= high; i++) {
        t = sprintf("%c", i)
        _ord_[t] = i
    }
}

Some explanation of the numbers used by chr is worthwhile. The most prominent character set in use today is ASCII. Although an eight-bit byte can hold 256 distinct values (from zero to 255), ASCII only defines characters that use the values from zero to 127.^{2} At least one computer manufacturer that we know of uses ASCII, but with mark parity, meaning that the leftmost bit in the byte is always one. What this means is that on those systems,

2 ASCII has been extended in many countries to use the values from 128 to 255 for country-specific characters. If your system uses these extensions, you can simplify _ord_init to simply loop from zero to 255. characters have numeric values from 128 to 255. Finally, large mainframe systems use the EBCDIC character set, which uses all 256 values. While there are other character sets in use on some older systems, they are not really worth worrying about.

function ord(str,    c)
{
    # only first character is of interest
    c = substr(str, 1, 1)
    return _ord_[c]
}
function chr(c)
{
    # force c to be numeric by adding 0
    return sprintf("%c", c + 0)
}
#### test code ####
# BEGIN    \
# {
#    for (;;) {
#        printf("enter a character: ")
#        if (getline var <= 0)
#            break
#        printf("ord(%s) = %d\n", var, ord(var))
#    }
# }

An obvious improvement to these functions would be to move the code for the _ord_init function into the body of the BEGIN rule. It was written this way initially for ease of development.

There is a “test program” in a BEGIN rule, for testing the function. It is commented out for production use.

15.6 Merging an Array Into a String

When doing string processing, it is often useful to be able to join all the strings in an array into one long string. The following function, join, accomplishes this task. It is used later in several of the application programs (see Chapter 16 [Practical awk Programs], page 203).

Good function design is important; this function needs to be general, but it should also have a reasonable default behavior. It is called with an array and the beginning and ending indices of the elements in the array to be merged. This assumes that the array indices are numeric—a reasonable assumption since the array was likely created with split (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

    # join.awk --- join an array into a string
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    function join(array, start, end, sep,    result, i)
    {
        if (sep == "")
           sep = " "
        else if (sep == SUBSEP) # magic value
           sep = ""
        result = array[start]
        for (i = start + 1; i <= end; i++)
            result = result sep array[i]
        return result
    }

An optional additional argument is the separator to use when joining the strings back together. If the caller supplies a non-empty value, join uses it. If it is not supplied, it will have a null value. In this case, join uses a single blank as a default separator for the strings. If the value is equal to SUBSEP, then join joins the strings with no separator between them. SUBSEP serves as a “magic” value to indicate that there should be no separation between the component strings.

It would be nice if awk had an assignment operator for concatenation. The lack of an explicit operator for concatenation makes string operations more difficult than they really need to be.

15.7 Turning Dates Into Timestamps

The systime function built in to gawk returns the current time of day as a timestamp in “seconds since the Epoch.” This timestamp can be converted into a printable date of almost infinitely variable format using the built-in strftime function. (For more information on systime and strftime, see Section 12.5 [Functions for Dealing with Time Stamps], page 146.)

An interesting but difficult problem is to convert a readable representation of a date back into a timestamp. The ANSI C library provides a mktime function that does the basic job, converting a canonical representation of a date into a timestamp.

It would appear at first glance that gawk would have to supply a mktime built-in function that was simply a “hook” to the C language version. In fact though, mktime can be implemented entirely in awk.

Here is a version of mktime for awk. It takes a simple representation of the date and time, and converts it into a timestamp.

The code is presented here intermixed with explanatory prose. In Section 16.2.7 [Extracting Programs from Texinfo Source Files], page 236, you will see how the Texinfo source file for this book can be processed to extract the code into a single source file.

The program begins with a descriptive comment and a BEGIN rule that initializes a table _tm_months. This table is a two-dimensional array that has the lengths of the months. The first index is zero for regular years, and one for leap years. The values are the same for all the months in both kinds of years, except for February; thus the use of multiple assignment.

    # mktime.awk --- convert a canonical date representation
    #                into a timestamp
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    BEGIN    \
    {
        # Initialize table of month lengths
        _tm_months[0,1] = _tm_months[1,1] = 31
        _tm_months[0,2] = 28; _tm_months[1,2] = 29
        _tm_months[0,3] = _tm_months[1,3] = 31
        _tm_months[0,4] = _tm_months[1,4] = 30
        _tm_months[0,5] = _tm_months[1,5] = 31
        _tm_months[0,6] = _tm_months[1,6] = 30
        _tm_months[0,7] = _tm_months[1,7] = 31
        _tm_months[0,8] = _tm_months[1,8] = 31
        _tm_months[0,9] = _tm_months[1,9] = 30
        _tm_months[0,10] = _tm_months[1,10] = 31
        _tm_months[0,11] = _tm_months[1,11] = 30
        _tm_months[0,12] = _tm_months[1,12] = 31
    }

The benefit of merging multiple BEGIN rules (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100) is particularly clear when writing library files. Functions in library files can cleanly initialize their own private data and also provide clean-up actions in private END rules.

The next function is a simple one that computes whether a given year is or is not a leap year. If a year is evenly divisible by four, but not evenly divisible by 100, or if it is evenly divisible by 400, then it is a leap year. Thus, 1904 was a leap year, 1900 was not, but 2000 will be.

    # decide if a year is a leap year
    function _tm_isleap(year,    ret)
    {
        ret = (year % 4 == 0 && year % 100 != 0) ||
                (year % 400 == 0)
        return ret
    }

This function is only used a few times in this file, and its computation could have been written in-line (at the point where it’s used). Making it a separate function made the original development easier, and also avoids the possibility of typing errors when duplicating the code in multiple places.

The next function is more interesting. It does most of the work of generating a timestamp, which is converting a date and time into some number of seconds since the Epoch. The caller passes an array (rather imaginatively named a) containing six values: the year including century, the month as a number between one and 12, the day of the month, the hour as a number between zero and 23, the minute in the hour, and the seconds within the minute.

The function uses several local variables to precompute the number of seconds in an hour, seconds in a day, and seconds in a year. Often, similar C code simply writes out the expression in-line, expecting the compiler to do constant folding. E.g., most C compilers would turn ‘60 * 60’ into ‘3600’ at compile time, instead of recomputing it every time at run time. Precomputing these values makes the function more efficient.

    # convert a date into seconds
    function _tm_addup(a,    total, yearsecs, daysecs,
                             hoursecs, i, j)
    {
        hoursecs = 60 * 60
        daysecs = 24 * hoursecs
        yearsecs = 365 * daysecs
        total = (a[1] - 1970) * yearsecs
        # extra day for leap years
        for (i = 1970; i < a[1]; i++)
            if (_tm_isleap(i))
                total += daysecs
        j = _tm_isleap(a[1])
        for (i = 1; i < a[2]; i++)
            total += _tm_months[j, i] * daysecs

        total += (a[3] - 1) * daysecs
        total += a[4] * hoursecs
        total += a[5] * 60
        total += a[6]
        return total
    }

The function starts with a first approximation of all the seconds between Midnight, January 1, 1970,^{3} and the beginning of the current year. It then goes through all those years, and for every leap year, adds an additional day’s worth of seconds.

The variable j holds either one or zero, if the current year is or is not a leap year. For every month in the current year prior to the current month, it adds the number of seconds in the month, using the appropriate entry in

the _tm_months array.

Finally, it adds in the seconds for the number of days prior to the current day, and the number of hours, minutes, and seconds in the current day.

The result is a count of seconds since January 1, 1970. This value is not yet what is needed though. The reason why is described shortly.

The main mktime function takes a single character string argument. This string is a representation of a date and time in a “canonical” (fixed) form. This string should be "year month day hour minute second".

    # mktime --- convert a date into seconds,
    #            compensate for time zone
    function mktime(str,    res1, res2, a, b, i, j, t, diff)
    {
        i = split(str, a, " ")    # don't rely on FS
        if (i != 6)
            return -1
        # force numeric
        for (j in a)
            a[j] += 0

3 This is the Epoch on POSIX systems. It may be different on other systems.

        # validate
        if (a[1] < 1970 ||
            a[2] < 1 || a[2] > 12 ||
            a[3] < 1 || a[3] > 31 ||
            a[4] < 0 || a[4] > 23 ||
            a[5] < 0 || a[5] > 59 ||
            a[6] < 0 || a[6] > 60 )
                return -1
        res1 = _tm_addup(a)
        t = strftime("%Y %m %d %H %M %S", res1)
        if (_tm_debug)
            printf("(%s) -> (%s)\n", str, t) > "/dev/stderr"
        split(t, b, " ")
        res2 = _tm_addup(b)
        diff = res1 - res2
        if (_tm_debug)
            printf("diff = %d seconds\n", diff) > "/dev/stderr"
        res1 += diff
        return res1
    }

The function first splits the string into an array, using spaces and tabs as separators. If there are not six elements in the array, it returns an error, signaled as the value −1. Next, it forces each element of the array to be numeric, by adding zero to it. The following ‘if’ statement then makes sure that each element is within an allowable range. (This checking could be extended further, e.g., to make sure that the day of the month is within the correct range for the particular month supplied.) All of this is essentially preliminary set-up and error checking.

Recall that _tm_addup generated a value in seconds since Midnight, January 1, 1970. This value is not directly usable as the result we want, since the calculation does not account for the local timezone. In other words, the value represents the count in seconds since the Epoch, but only for UTC (Universal Coordinated Time). If the local timezone is east or west of UTC, then some number of hours should be either added to, or subtracted from the resulting timestamp.

For example, 6:23 p.m. in Atlanta, Georgia (USA), is normally five hours west of (behind) UTC. It is only four hours behind UTC if daylight savings time is in effect. If you are calling mktime in Atlanta, with the argument "1993 5 23 18 23 12", the result from _tm_addup will be for 6:23 p.m. UTC, which is only 2:23 p.m. in Atlanta. It is necessary to add another four hours worth of seconds to the result.

How can mktime determine how far away it is from UTC? This is surprisingly easy. The returned timestamp represents the time passed to mktime as UTC. This timestamp can be fed back to strftime, which will format it as a local time; i.e. as if it already had the UTC difference added in to it. This is done by giving "%Y %m %d %H %M %S" to strftime as the format argument. It returns the computed timestamp in the original string format. The result represents a time that accounts for the UTC difference. When the new time is converted back to a timestamp, the difference between the two timestamps is the difference (in seconds) between the local timezone and UTC. This difference is then added back to the original result. An example demonstrating this is presented below.

Finally, there is a “main” program for testing the function.

    BEGIN  {
        if (_tm_test) {
            printf "Enter date as yyyy mm dd hh mm ss: "
            getline _tm_test_date
            t = mktime(_tm_test_date)
            r = strftime("%Y %m %d %H %M %S", t)
            printf "Got back (%s)\n", r
        }
    }

The entire program uses two variables that can be set on the command line to control debugging output and to enable the test in the final BEGIN rule. Here is the result of a test run. (Note that debugging output is to standard error, and test output is to standard output.)

    $ gawk -f mktime.awk -v _tm_test=1 -v _tm_debug=1
    ⊣ Enter date as yyyy mm dd hh mm ss: 1993 5 23 15 35 10
     error (1993 5 23 15 35 10) -> (1993 05 23 11 35 10)
     error diff = 14400 seconds
    ⊣ Got back (1993 05 23 15 35 10)

The time entered was 3:35 p.m. (15:35 on a 24-hour clock), on May 23, 1993. The first line of debugging output shows the resulting time as UTC— four hours ahead of the local time zone. The second line shows that the difference is 14400 seconds, which is four hours. (The difference is only four hours, since daylight savings time is in effect during May.) The final line of test output shows that the timezone compensation algorithm works; the returned time is the same as the entered time.

This program does not solve the general problem of turning an arbitrary date representation into a timestamp. That problem is very involved. However, the mktime function provides a foundation upon which to build. Other software can convert month names into numeric months, and AM/PM times into 24-hour clocks, to generate the “canonical” format that mktime requires.

15.8 Managing the Time of Day

The systime and strftime functions described in Section 12.5 [Functions for Dealing with Time Stamps], page 146, provide the minimum functionality necessary for dealing with the time of day in human readable form. While strftime is extensive, the control formats are not necessarily easy to remember or intuitively obvious when reading a program.

The following function, gettimeofday, populates a user-supplied array with pre-formatted time information. It returns a string with the current time formatted in the same way as the date utility.

# gettimeofday --- get the time of day in a usable format
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain, May 1993
#
# Returns a string in the format of output of date(1)
# Populates the array argument time with individual values:
#    time["second"]       -- seconds (0 - 59)
#    time["minute"]       -- minutes (0 - 59)
#    time["hour"]         -- hours (0 - 23)
#    time["althour"]      -- hours (0 - 12)
#    time["monthday"]     -- day of month (1 - 31)
#    time["month"]        -- month of year (1 - 12)
#    time["monthname"]    -- name of the month
#    time["shortmonth"]   -- short name of the month
#    time["year"]         -- year within century (0 - 99)
#    time["fullyear"]     -- year with century (19xx or 20xx)
#    time["weekday"]      -- day of week (Sunday = 0)
#    time["altweekday"]   -- day of week (Monday = 0)
#    time["weeknum"]      -- week number, Sunday first day
#    time["altweeknum"]   -- week number, Monday first day
#    time["dayname"]      -- name of weekday
#    time["shortdayname"] -- short name of weekday
#    time["yearday"]      -- day of year (0 - 365)
#    time["timezone"]     -- abbreviation of timezone name
#    time["ampm"]         -- AM or PM designation

function gettimeofday(time,    ret, now, i)
{
# get time once, avoids unnecessary system calls
now = systime()
# return date(1)-style output
ret = strftime("%a %b %d %H:%M:%S %Z %Y", now)
# clear out target array
for (i in time)
    delete time[i]
# fill in values, force numeric values to be
# numeric by adding 0
time["second"]       = strftime("%S", now) + 0
time["minute"]       = strftime("%M", now) + 0
time["hour"]         = strftime("%H", now) + 0
time["althour"]      = strftime("%I", now) + 0
time["monthday"]     = strftime("%d", now) + 0
time["month"]        = strftime("%m", now) + 0
time["monthname"]    = strftime("%B", now)
time["shortmonth"]   = strftime("%b", now)
time["year"]         = strftime("%y", now) + 0
time["fullyear"]     = strftime("%Y", now) + 0
time["weekday"]      = strftime("%w", now) + 0
time["altweekday"]   = strftime("%u", now) + 0
time["dayname"]      = strftime("%A", now)
time["shortdayname"] = strftime("%a", now)
time["yearday"]      = strftime("%j", now) + 0
time["timezone"]     = strftime("%Z", now)
time["ampm"]         = strftime("%p", now)
time["weeknum"]      = strftime("%U", now) + 0
time["altweeknum"]   = strftime("%W", now) + 0
return ret
}

The string indices are easier to use and read than the various formats required by strftime. The alarm program presented in Section 16.2.2 [An Alarm Clock Program], page 226, uses this function.

The gettimeofday function is presented above as it was written. A more general design for this function would have allowed the user to supply an optional timestamp value that would have been used instead of the current time.

15.9 Noting Data File Boundaries

The BEGIN and END rules are each executed exactly once, at the beginning and end respectively of your awk program (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100). We (the gawk authors) once had a user who mistakenly thought that the BEGIN rule was executed at the beginning of each data file and the END rule was executed at the end of each data file. When informed that this was not the case, the user requested that we add new special patterns to gawk, named BEGIN_FILE and END_FILE, that would have the desired behavior. He even supplied us the code to do so.

However, after a little thought, I came up with the following library program. It arranges to call two user-supplied functions, beginfile and endfile, at the beginning and end of each data file. Besides solving the problem in only nine(!) lines of code, it does so portably; this will work with any implementation of awk.

    # transfile.awk
    #
    # Give the user a hook for filename transitions
    #
    # The user must supply functions beginfile() and endfile()
    # that each take the name of the file being started or
    # finished, respectively.
    #
    # Arnold Robbins, arnold@gnu.ai.mit.edu, January 1992
    # Public Domain
    FILENAME != _oldfilename \
    {
        if (_oldfilename != "")
            endfile(_oldfilename)
        _oldfilename = FILENAME
        beginfile(FILENAME)
    }
    END   { endfile(FILENAME) }

This file must be loaded before the user’s “main” program, so that the rule it supplies will be executed first.

This rule relies on awk’s FILENAME variable that automatically changes for each new data file. The current file name is saved in a private variable,

_oldfilename. If FILENAME does not equal _oldfilename, then a new data

file is being processed, and it is necessary to call endfile for the old file. Since endfile should only be called if a file has been processed, the program first checks to make sure that _oldfilename is not the null string.

The program then assigns the current file name to _oldfilename, and calls beginfile for the file. Since, like all awk variables, _oldfilename will be initialized to the null string, this rule executes correctly even for the first data file.

The program also supplies an END rule, to do the final processing for the last file. Since this END rule comes before any END rules supplied in the “main” program, endfile will be called first. Once again the value of multiple BEGIN and END rules should be clear.

This version has same problem as the first version of nextfile (see Section 15.2 [Implementing nextfile as a Function], page 168). If the same data file occurs twice in a row on command line, then endfile and beginfile will not be executed at the end of the first pass and at the beginning of the second pass. This version solves the problem.

    # ftrans.awk --- handle data file transitions
    #
    # user supplies beginfile() and endfile() functions
    #
    # Arnold Robbins, arnold@gnu.ai.mit.edu. November 1992
    # Public Domain
    FNR == 1 {
        if (_filename_!= "")
            endfile(_filename_)
        _filename_= FILENAME
        beginfile(FILENAME)
    }
    END  { endfile(_filename_) }

In Section 16.1.7 [Counting Things], page 223, you will see how this library function can be used, and how it simplifies writing the main program.

15.10 Processing Command Line Options

Most utilities on POSIX compatible systems take options or “switches” on the command line that can be used to change the way a program behaves. awk is an example of such a program (see Section 14.1 [Command Line Options], page 159). Often, options take arguments, data that the program needs to correctly obey the command line option. For example, awk’s ‘-F’ option requires a string to use as the field separator. The first occurrence on the command line of either ‘--’ or a string that does not begin with ‘-’ ends the options.

Most Unix systems provide a C function named getopt for processing command line arguments. The programmer provides a string describing the one letter options. If an option requires an argument, it is followed in the string with a colon. getopt is also passed the count and values of the command line arguments, and is called in a loop. getopt processes the command line arguments for option letters. Each time around the loop, it returns a single character representing the next option letter that it found, or ‘?’ if it found an invalid option. When it returns −1, there are no options left on the command line.

When using getopt, options that do not take arguments can be grouped together. Furthermore, options that take arguments require that the argument be present. The argument can immediately follow the option letter, or it can be a separate command line argument.

Given a hypothetical program that takes three command line options, ‘-a’, ‘-b’, and ‘-c’, and ‘-b’ requires an argument, all of the following are valid ways of invoking the program:

    prog -a -b foo -c data1 data2 data3
    prog -ac -bfoo -- data1 data2 data3
    prog -acbfoo data1 data2 data3

Notice that when the argument is grouped with its option, the rest of the command line argument is considered to be the option’s argument. In the above example, ‘-acbfoo’ indicates that all of the ‘-a’, ‘-b’, and ‘-c’ options were supplied, and that ‘foo’ is the argument to the ‘-b’ option.

getopt provides four external variables that the programmer can use.

optind The index in the argument value array (argv) where the first non-option command line argument can be found.

optarg The string value of the argument to an option.

opterr Usually getopt prints an error message when it finds an invalid option. Setting opterr to zero disables this feature. (An application might wish to print its own error message.)

optopt The letter representing the command line option. While not usually documented, most versions supply this variable.

The following C fragment shows how getopt might process command line arguments for awk.

    int
    main(int argc, char *argv[])
    {

…

        /* print our own message */
        opterr = 0;

while ((c = getopt(argc, argv, "v:f:F:W:")) != -1) {
switch (c) {
case 'f':    /* file */

…

break;
case 'F':    /* field separator */

…

break;
case 'v':    /* variable assignment */

…

break;
case 'W':    /* extension */

…

break;
case '?':
default:
usage();
break;
}
}

…

}

As a side point, gawk actually uses the GNU getopt_long function to process both normal and GNU-style long options (see Section 14.1 [Command Line Options], page 159).

The abstraction provided by getopt is very useful, and would be quite handy in awk programs as well. Here is an awk version of getopt. This function highlights one of the greatest weaknesses in awk, which is that it is very poor at manipulating single characters. Repeated calls to substr are necessary for accessing individual characters (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

The discussion walks through the code a bit at a time.

# getopt --- do C library getopt(3) function in awk
#
# arnold@gnu.ai.mit.edu
# Public domain
#
# Initial version: March, 1991
# Revised: May, 1993

# External variables:
#    Optind -- index of ARGV for first non-option argument
#    Optarg -- string value of argument to current option
#    Opterr -- if non-zero, print our own diagnostic
#    Optopt -- current option letter
# Returns
#    -1     at end of options
#    ?      for unrecognized option
#    <c>    a character representing the current option
# Private Data
#    _opti  index in multi-flag option, e.g., -abc

The function starts out with some documentation: who wrote the code, and when it was revised, followed by a list of the global variables it uses, what the return values are and what they mean, and any global variables that are “private” to this library function. Such documentation is essential for any program, and particularly for library functions.

function getopt(argc, argv, options,    optl, thisopt, i)
{
    optl = length(options)
    if (optl == 0)        # no options given
        return -1
    if (argv[Optind] == "--") {  # all done
        Optind++
        _opti = 0
        return -1
    } else if (argv[Optind] !~ /^-[^: \t\n\f\r\v\b]/) {
        _opti = 0
        return -1
    }

The function first checks that it was indeed called with a string of options (the options parameter). If options has a zero length, getopt immediately returns −1.

The next thing to check for is the end of the options. A ‘--’ ends the command line options, as does any command line argument that does not begin with a ‘-’. Optind is used to step through the array of command line arguments; it retains its value across calls to getopt, since it is a global variable.

The regexp used, /^-[^: \t\n\f\r\v\b]/, is perhaps a bit of overkill; it checks for a ‘-’ followed by anything that is not whitespace and not a colon.

If the current command line argument does not match this pattern, it is not an option, and it ends option processing.

        if (_opti == 0)
            _opti = 2
        thisopt = substr(argv[Optind], _opti, 1)
        Optopt = thisopt
        i = index(options, thisopt)
        if (i == 0) {
            if (Opterr)
                printf("%c -- invalid option\n",
                                      thisopt) > "/dev/stderr"
            if (_opti >= length(argv[Optind])) {
                Optind++
                _opti = 0
            } else
                _opti++
            return "?"
        }

The _opti variable tracks the position in the current command line argument (argv[Optind]). In the case that multiple options were grouped together with one ‘-’ (e.g., ‘-abx’), it is necessary to return them to the user one at a time.

If _opti is equal to zero, it is set to two, the index in the string of the next character to look at (we skip the ‘-’, which is at position one). The variable thisopt holds the character, obtained with substr. It is saved in Optopt for the main program to use.

If thisopt is not in the options string, then it is an invalid option. If Opterr is non-zero, getopt prints an error message on the standard error that is similar to the message from the C version of getopt.

Since the option is invalid, it is necessary to skip it and move on to the next option character. If _opti is greater than or equal to the length of the current command line argument, then it is necessary to move on to the next one, so Optind is incremented and _opti is reset to zero. Otherwise, Optind is left alone and _opti is merely incremented.

In any case, since the option was invalid, getopt returns ‘?’. The main program can examine Optopt if it needs to know what the invalid option letter actually was.

if (substr(options, i + 1, 1) == ":") {
# get option argument
if (length(substr(argv[Optind], _opti + 1)) > 0)
    Optarg = substr(argv[Optind], _opti + 1)
else
    Optarg = argv[++Optind]
_opti = 0
} else
Optarg = ""

If the option requires an argument, the option letter is followed by a colon in the options string. If there are remaining characters in the current command line argument (argv[Optind]), then the rest of that string is assigned to Optarg. Otherwise, the next command line argument is used (‘-xFOO’ vs. ‘-x FOO’). In either case, _opti is reset to zero, since there are no more characters left to examine in the current command line argument.

if (_opti == 0 || _opti >= length(argv[Optind])) {
Optind++
_opti = 0
} else
_opti++
return thisopt
}

Finally, if _opti is either zero or greater than the length of the current command line argument, it means this element in argv is through being processed, so Optind is incremented to point to the next element in argv. If neither condition is true, then only _opti is incremented, so that the next option letter can be processed on the next call to getopt.

BEGIN {
Opterr = 1    # default is to diagnose
Optind = 1    # skip ARGV[0]
# test program
if (_getopt_test) {
while ((_go_c = getopt(ARGC, ARGV, "ab:cd")) != -1)
    printf("c = <%c>, optarg = <%s>\n",
                               _go_c, Optarg)
printf("non-option arguments:\n")
for (; Optind < ARGC; Optind++)
    printf("\tARGV[%d] = <%s>\n",
                            Optind, ARGV[Optind])
}
}

The BEGIN rule initializes both Opterr and Optind to one. Opterr is set to one, since the default behavior is for getopt to print a diagnostic message upon seeing an invalid option. Optind is set to one, since there’s no reason to look at the program name, which is in ARGV[0].

The rest of the BEGIN rule is a simple test program. Here is the result of two sample runs of the test program.

    $ awk -f getopt.awk -v _getopt_test=1 -- -a -cbARG bax -x
    ⊣ c = <a>, optarg = <>
    ⊣ c = <c>, optarg = <>
    ⊣ c = <b>, optarg = <ARG>
    ⊣ non-option arguments:
    ⊣  ARGV[3] = <bax>
    ⊣  ARGV[4] = <-x>
    $ awk -f getopt.awk -v _getopt_test=1 -- -a -x -- xyz abc
    ⊣ c = <a>, optarg = <>
     error x -- invalid option
    ⊣ c = <?>, optarg = <>
    ⊣ non-option arguments:
    ⊣  ARGV[4] = <xyz>
    ⊣  ARGV[5] = <abc>

The first ‘--’ terminates the arguments to awk, so that it does not try to interpret the ‘-a’ etc. as its own options.

Several of the sample programs presented in Chapter 16 [Practical awk Programs], page 203, use getopt to process their arguments.

15.11 Reading the User Database

The ‘/dev/user’ special file (see Section 6.7 [Special File Names in gawk], page 72) provides access to the current user’s real and effective user and group id numbers, and if available, the user’s supplementary group set. However, since these are numbers, they do not provide very useful information to the average user. There needs to be some way to find the user information associated with the user and group numbers. This section presents a suite of functions for retrieving information from the user database. See Section 15.12 [Reading the Group Database], page 195, for a similar suite that retrieves information from the group database.

The POSIX standard does not define the file where user information is kept. Instead, it provides the <pwd.h> header file and several C language subroutines for obtaining user information. The primary function is getpwent, for “get password entry.” The “password” comes from the original user database file, ‘/etc/passwd’, which kept user information, along with the encrypted passwords (hence the name).

While an awk program could simply read ‘/etc/passwd’ directly (the format is well known), because of the way password files are handled on networked systems, this file may not contain complete information about the system’s set of users.

To be sure of being able to produce a readable, complete version of the user database, it is necessary to write a small C program that calls getpwent. getpwent is defined to return a pointer to a struct passwd. Each time it is called, it returns the next entry in the database. When there are no more entries, it returns NULL, the null pointer. When this happens, the C program should call endpwent to close the database. Here is pwcat, a C program that “cats” the password database.

    /*
     * pwcat.c
     *
     * Generate a printable version of the password database
     *
     * Arnold Robbins
     * arnold@gnu.ai.mit.edu
     * May 1993
     * Public Domain
     */
    #include <stdio.h>
    #include <pwd.h>
    int
    main(argc, argv)
    int argc;
    char **argv;
    {
        struct passwd *p;
        while ((p = getpwent()) != NULL)
            printf("%s:%s:%d:%d:%s:%s:%s\n",
                p->pw_name, p->pw_passwd, p->pw_uid,
                p->pw_gid, p->pw_gecos, p->pw_dir, p->pw_shell);
        endpwent();
        exit(0);
    }

If you don’t understand C, don’t worry about it. The output from pwcat is the user database, in the traditional ‘/etc/passwd’ format of colonseparated fields. The fields are:

Login name The user’s login name.

Encrypted password The user’s encrypted password. This may not be available on some systems.

User-ID The user’s numeric user-id number.

Group-ID The user’s numeric group-id number.

Full name The user’s full name, and perhaps other information associated with the user.

Home directory The user’s login, or “home” directory (familiar to shell programmers as $HOME).

Login shell The program that will be run when the user logs in. This is usually a shell, such as Bash (the Gnu Bourne-Again shell).

Here are a few lines representative of pwcat’s output.

    $ pwcat
    ⊣ root:3Ov02d5VaUPB6:0:1:Operator:/:/bin/sh
    ⊣ nobody:*:65534:65534::/:
    ⊣ daemon:*:1:1::/:
    ⊣ sys:*:2:2::/:/bin/csh
    ⊣ bin:*:3:3::/bin:
    ⊣ arnold:xyzzy:2076:10:Arnold Robbins:/home/arnold:/bin/sh
    ⊣ miriam:yxaay:112:10:Miriam Robbins:/home/miriam:/bin/sh
    ⊣ andy:abcca2:113:10:Andy Jacobs:/home/andy:/bin/sh

…

With that introduction, here is a group of functions for getting user information. There are several functions here, corresponding to the C functions of the same name.

    # passwd.awk --- access password file information
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    BEGIN {
        # tailor this to suit your system
        _pw_awklib = "/usr/local/libexec/awk/"
    }

function _pw_init(    oldfs, oldrs, olddol0, pwcat)
{
if (_pw_inited)
    return
oldfs = FS
oldrs = RS
olddol0 = $0
FS = ":"
RS = "\n"
pwcat = _pw_awklib "pwcat"
while ((pwcat | getline) > 0) {
    _pw_byname[$1] = $0
    _pw_byuid[$3] = $0
    _pw_bycount[++_pw_total] = $0
}
close(pwcat)
_pw_count = 0
_pw_inited = 1
FS = oldfs
RS = oldrs
$0 = olddol0
}

The BEGIN rule sets a private variable to the directory where pwcat is stored. Since it is used to help out an awk library routine, we have chosen to put it in ‘/usr/local/libexec/awk’. You might want it to be in a different directory on your system.

The function _pw_init keeps three copies of the user information in three associative arrays. The arrays are indexed by user name (_pw_byname), by user-id number (_pw_byuid), and by order of occurrence (_pw_bycount).

The variable _pw_inited is used for efficiency; _pw_init only needs to be called once.

Since this function uses getline to read information from pwcat, it first saves the values of FS, RS, and $0. Doing so is necessary, since these functions could be called from anywhere within a user’s program, and the user may have his or her own values for FS and RS.

The main part of the function uses a loop to read database lines, split the line into fields, and then store the line into each array as necessary. When the loop is done, _pw_init cleans up by closing the pipeline, setting _pw_inited to one, and restoring FS, RS, and $0. The use of _pw_count will be explained below.

    function getpwnam(name)
    {
        _pw_init()
        if (name in _pw_byname)
            return _pw_byname[name]
        return ""
    }

The getpwnam function takes a user name as a string argument. If that user is in the database, it returns the appropriate line. Otherwise it returns the null string.

    function getpwuid(uid)
    {
        _pw_init()
        if (uid in _pw_byuid)
            return _pw_byuid[uid]
        return ""
    }

Similarly, the getpwuid function takes a user-id number argument. If that user number is in the database, it returns the appropriate line. Otherwise it returns the null string.

    function getpwent()
    {
        _pw_init()
        if (_pw_count < _pw_total)
            return _pw_bycount[++_pw_count]
        return ""
    }

The getpwent function simply steps through the database, one entry at a time. It uses _pw_count to track its current position in the _pw_bycount array.

    function endpwent()
    {
        _pw_count = 0
    }

The endpwent function resets _pw_count to zero, so that subsequent calls to getpwent will start over again.

A conscious design decision in this suite is that each subroutine calls _pw_init to initialize the database arrays. The overhead of running a separate process to generate the user database, and the I/O to scan it, will only be incurred if the user’s main program actually calls one of these functions. If this library file is loaded along with a user’s program, but none of the routines are ever called, then there is no extra run-time overhead. (The alternative would be to move the body of _pw_init into a BEGIN rule, which would always run pwcat. This simplifies the code but runs an extra process that may never be needed.)

In turn, calling _pw_init is not too expensive, since the _pw_inited variable keeps the program from reading the data more than once. If you are worried about squeezing every last cycle out of your awk program, the check of _pw_inited could be moved out of _pw_init and duplicated in all the other functions. In practice, this is not necessary, since most awk programs are I/O bound, and it would clutter up the code.

The id program in Section 16.1.3 [Printing Out User Information], page 213, uses these functions.

15.12 Reading the Group Database

Much of the discussion presented in Section 15.11 [Reading the User Database], page 190, applies to the group database as well. Although there has traditionally been a well known file, ‘/etc/group’, in a well known format, the POSIX standard only provides a set of C library routines (<grp.h> and getgrent) for accessing the information. Even though this file may exist, it likely does not have complete information. Therefore, as with the user database, it is necessary to have a small C program that generates the group database as its output.

Here is grcat, a C program that “cats” the group database.

    /*
     * grcat.c
     *
     * Generate a printable version of the group database
     *
     * Arnold Robbins, arnold@gnu.ai.mit.edu
     * May 1993
     * Public Domain
     */
    #include <stdio.h>
    #include <grp.h>
    int
    main(argc, argv)
    int argc;
    char **argv;
    {
        struct group *g;
        int i;

while ((g = getgrent()) != NULL) {
  printf("%s:%s:%d:", g->gr_name, g->gr_passwd,
                                      g->gr_gid);
  for (i = 0; g->gr_mem[i] != NULL; i++) {
      printf("%s", g->gr_mem[i]);
      if (g->gr_mem[i+1] != NULL)
          putchar(',');
  }
  putchar('\n');
}
endgrent();
exit(0);
}

Each line in the group database represent one group. The fields are separated with colons, and represent the following information.

Group Name The name of the group.

Group Password The encrypted group password. In practice, this field is never used. It is usually empty, or set to ‘*’.

Group ID Number The numeric group-id number. This number should be unique within the file.

Group Member List A comma-separated list of user names. These users are members of the group. Most Unix systems allow users to be members of several groups simultaneously. If your system does, then reading ‘/dev/user’ will return those group-id numbers in $5 through $NF. (Note that ‘/dev/user’ is a gawk extension; see Section 6.7 [Special File Names in gawk], page 72.)

Here is what running grcat might produce:

$ grcat
⊣ wheel:*:0:arnold
⊣ nogroup:*:65534:
⊣ daemon:*:1:
⊣ kmem:*:2:
⊣ staff:*:10:arnold,miriam,andy
⊣ other:*:20:

…

Here are the functions for obtaining information from the group database. There are several, modeled after the C library functions of the same names.

# group.awk --- functions for dealing with the group file
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
# May 1993
BEGIN    \
{
# Change to suit your system
_gr_awklib = "/usr/local/libexec/awk/"
}
function _gr_init(    oldfs, oldrs, olddol0, grcat, n, a, i)
{
if (_gr_inited)
    return
oldfs = FS
oldrs = RS
olddol0 = $0
FS = ":"
RS = "\n"
grcat = _gr_awklib "grcat"
while ((grcat | getline) > 0) {
    if ($1 in _gr_byname)
        _gr_byname[$1] = _gr_byname[$1] "," $4
    else
        _gr_byname[$1] = $0
    if ($3 in _gr_bygid)
        _gr_bygid[$3] = _gr_bygid[$3] "," $4
    else
        _gr_bygid[$3] = $0
    n = split($4, a, "[ \t]*,[ \t]*")
    for (i = 1; i <= n; i++)
        if (a[i] in _gr_groupsbyuser)
            _gr_groupsbyuser[a[i]] = \
                _gr_groupsbyuser[a[i]] " " $1
        else
            _gr_groupsbyuser[a[i]] = $1
    _gr_bycount[++_gr_count] = $0
}

        close(grcat)
        _gr_count = 0
        _gr_inited++
        FS = oldfs
        RS = oldrs
        $0 = olddol0
    }

The BEGIN rule sets a private variable to the directory where grcat is stored. Since it is used to help out an awk library routine, we have chosen to put it in ‘/usr/local/libexec/awk’. You might want it to be in a different directory on your system.

These routines follow the same general outline as the user database routines (see Section 15.11 [Reading the User Database], page 190). The _gr_inited variable is used to ensure that the database is scanned no more than once. The _gr_init function first saves FS, RS, and $0, and then sets FS and RS to the correct values for scanning the group information.

The group information is stored is several associative arrays. The arrays are indexed by group name (_gr_byname), by group-id number (_gr_bygid), and by position in the database (_gr_bycount). There is an additional array indexed by user name (_gr_groupsbyuser), that is a space separated list of groups that each user belongs to.

Unlike the user database, it is possible to have multiple records in the database for the same group. This is common when a group has a large number of members. Such a pair of entries might look like:

    tvpeople:*:101:johny,jay,arsenio
    tvpeople:*:101:david,conan,tom,joan

For this reason, _gr_init looks to see if a group name or group-id number has already been seen. If it has, then the user names are simply concatenated onto the previous list of users. (There is actually a subtle problem with the code presented above. Suppose that the first time there were no names. This code adds the names with a leading comma. It also doesn’t check that there is a $4.)

Finally, _gr_init closes the pipeline to grcat, restores FS, RS, and $0, initializes _gr_count to zero (it is used later), and makes _gr_inited nonzero.

    function getgrnam(group)
    {
        _gr_init()
        if (group in _gr_byname)
            return _gr_byname[group]
        return ""
    }

The getgrnam function takes a group name as its argument, and if that group exists, it is returned. Otherwise, getgrnam returns the null string.

function getgrgid(gid)
{
    _gr_init()
    if (gid in _gr_bygid)
        return _gr_bygid[gid]
    return ""
}

The getgrgid function is similar, it takes a numeric group-id, and looks up the information associated with that group-id.

function getgruser(user)
{
    _gr_init()
    if (user in _gr_groupsbyuser)
        return _gr_groupsbyuser[user]
    return ""
}

The getgruser function does not have a C counterpart. It takes a user name, and returns the list of groups that have the user as a member.

function getgrent()
{
    _gr_init()
    if (++gr_count in _gr_bycount)
        return _gr_bycount[_gr_count]
    return ""
}

The getgrent function steps through the database one entry at a time. It uses _gr_count to track its position in the list.

function endgrent()
{
    _gr_count = 0
}

endgrent resets _gr_count to zero so that getgrent can start over again. As with the user database routines, each function calls _gr_init to ini-

tialize the arrays. Doing so only incurs the extra overhead of running grcat if these functions are used (as opposed to moving the body of _gr_init into a BEGIN rule).

Most of the work is in scanning the database and building the various associative arrays. The functions that the user calls are themselves very simple, relying on awk’s associative arrays to do work.

The id program in Section 16.1.3 [Printing Out User Information], page 213, uses these functions.

15.13 Naming Library Function Global Variables

Due to the way the awk language evolved, variables are either global (usable by the entire program), or local (usable just by a specific function). There is no intermediate state analogous to static variables in C.

Library functions often need to have global variables that they can use to preserve state information between calls to the function. For example, getopt’s variable _opti (see Section 15.10 [Processing Command Line Options], page 184), and the _tm_months array used by mktime (see Section 15.7 [Turning Dates Into Timestamps], page 175). Such variables are called private, since the only functions that need to use them are the ones in the library.

When writing a library function, you should try to choose names for your private variables so that they will not conflict with any variables used by either another library function or a user’s main program. For example, a name like ‘i’ or ‘j’ is not a good choice, since user programs often use variable names like these for their own purposes.

The example programs shown in this chapter all start the names of their private variables with an underscore (‘_’). Users generally don’t use leading underscores in their variable names, so this convention immediately decreases the chances that the variable name will be accidentally shared with the user’s program.

In addition, several of the library functions use a prefix that helps indicate what function or set of functions uses the variables. For example, _tm_months in mktime (see Section 15.7 [Turning Dates Into Timestamps], page 175), and _pw_byname in the user data base routines (see Section 15.11 [Reading the User Database], page 190). This convention is recommended, since it even further decreases the chance of inadvertent conflict among variable names. Note that this convention can be used equally well both for variable names and for private function names too.

While I could have re-written all the library routines to use this convention, I did not do so, in order to show how my own awk programming style has evolved, and to provide some basis for this discussion.

As a final note on variable naming, if a function makes global variables available for use by a main program, it is a good convention to start that variable’s name with a capital letter. For example, getopt’s Opterr and Optind variables (see Section 15.10 [Processing Command Line Options], page 184). The leading capital letter indicates that it is global, while the fact that the variable name is not all capital letters indicates that the variable is not one of awk’s built-in variables, like FS.

It is also important that all variables in library functions that do not need to save state are in fact declared local. If this is not done, the variable could accidentally be used in the user’s program, leading to bugs that are very difficult to track down.

    function lib_func(x, y,    l1, l2)
    {

…

        use variable some_var  # some_var could be local
        . . .  # but is not by oversight
    }

A different convention, common in the Tcl community, is to use a single associative array to hold the values needed by the library function(s), or “package.” This significantly decreases the number of actual global names in use. For example, the functions described in Section 15.11 [Reading the User Database], page 190, might have used PW_data["inited"],

PW_data["total"], PW_data["count"] and PW_data["awklib"], instead
of _pw_inited, _pw_awklib, _pw_total, and _pw_count.

The conventions presented in this section are exactly that, conventions. You are not required to write your programs this way, we merely recommend that you do so.

202 Effective AWK Programming

Chapter 16: Practical awk Programs 203

16 Practical awk Programs

This chapter presents a potpourri of awk programs for your reading enjoyment. There are two sections. The first presents awk versions of several common POSIX utilities. The second is a grab-bag of interesting programs.

Many of these programs use the library functions presented in Chapter 15 [A Library of awk Functions], page 167.

16.1 Re-inventing Wheels for Fun and Profit

This section presents a number of POSIX utilities that are implemented in awk. Re-inventing these programs in awk is often enjoyable, since the algorithms can be very clearly expressed, and usually the code is very concise and simple. This is true because awk does so much for you.

It should be noted that these programs are not necessarily intended to replace the installed versions on your system. Instead, their purpose is to illustrate awk language programming for “real world” tasks.

The programs are presented in alphabetical order.

16.1.1 Cutting Out Fields and Columns

The cut utility selects, or “cuts,” either characters or fields from its standard input and sends them to its standard output. cut can cut out either a list of characters, or a list of fields. By default, fields are separated by tabs, but you may supply a command line option to change the field delimiter, i.e. the field separator character. cut’s definition of fields is less general than awk’s.

A common use of cut might be to pull out just the login name of logged-on users from the output of who. For example, the following pipeline generates a sorted, unique list of the logged on users:

    who | cut -c1-8 | sort | uniq

The options for cut are:

-c list Use list as the list of characters to cut out. Items within the list may be separated by commas, and ranges of characters can be separated with dashes. The list ‘1-8,15,22-35’ specifies characters one through eight, 15, and 22 through 35.

-f list Use list as the list of fields to cut out.

-d delim Use delim as the field separator character instead of the tab character.

-s Suppress printing of lines that do not contain the field delimiter.

The awk implementation of cut uses the getopt library function (see Section 15.10 [Processing Command Line Options], page 184), and the join library function (see Section 15.6 [Merging an Array Into a String], page 174).

The program begins with a comment describing the options and a usage function which prints out a usage message and exits. usage is called if invalid arguments are supplied.

# cut.awk --- implement cut in awk
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
# May 1993
# Options:
#    -f list        Cut fields
#    -d c           Field delimiter character
#    -c list        Cut characters
#
#    -s        Suppress lines without the delimiter character
function usage(    e1, e2)
{
    e1 = "usage: cut [-f list] [-d c] [-s] [files...]"
    e2 = "usage: cut [-c list] [files...]"
    print e1 > "/dev/stderr"
    print e2 > "/dev/stderr"
    exit 1
}

The variables e1 and e2 are used so that the function fits nicely on the page.

Next comes a BEGIN rule that parses the command line options. It sets FS to a single tab character, since that is cut’s default field separator. The output field separator is also set to be the same as the input field separator. Then getopt is used to step through the command line options. One or the other of the variables by_fields or by_chars is set to true, to indicate that processing should be done by fields or by characters respectively. When cutting by characters, the output field separator is set to the null string.

BEGIN    \
{
    FS = "\t"    # default
    OFS = FS
    while ((c = getopt(ARGC, ARGV, "sf:c:d:")) != -1) {
        if (c == "f") {
            by_fields = 1
            fieldlist = Optarg

} else if (c == "c") {
by_chars = 1
fieldlist = Optarg
OFS = ""
} else if (c == "d") {
if (length(Optarg) > 1) {
    printf("Using first character of %s" \
    " for delimiter\n", Optarg) > "/dev/stderr"
    Optarg = substr(Optarg, 1, 1)
}
FS = Optarg
OFS = FS
if (FS == " ")    # defeat awk semantics
    FS = "[ ]"
} else if (c == "s")
suppress++
else
usage()
}
for (i = 1; i < Optind; i++)
ARGV[i] = ""

Special care is taken when the field delimiter is a space. Using " " (a single space) for the value of FS is incorrect—awk would separate fields with runs of spaces, tabs and/or newlines, and we want them to be separated with individual spaces. Also, note that after getopt is through, we have to clear out all the elements of ARGV from one to Optind, so that awk will not try to process the command line options as file names.

After dealing with the command line options, the program verifies that the options make sense. Only one or the other of ‘-c’ and ‘-f’ should be used, and both require a field list. Then either set_fieldlist or set_charlist is called to pull apart the list of fields or characters.

if (by_fields && by_chars)
usage()
if (by_fields == 0 && by_chars == 0)
by_fields = 1    # default
if (fieldlist == "") {
print "cut: needs list for -c or -f" > "/dev/stderr"
exit 1
}

        if (by_fields)
            set_fieldlist()
        else
            set_charlist()
    }

Here is set_fieldlist. It first splits the field list apart at the commas, into an array. Then, for each element of the array, it looks to see if it is actually a range, and if so splits it apart. The range is verified to make sure the first number is smaller than the second. Each number in the list is added to the flist array, which simply lists the fields that will be printed. Normal field splitting is used. The program lets awk handle the job of doing the field splitting.

    function set_fieldlist(        n, m, i, j, k, f, g)
    {
        n = split(fieldlist, f, ",")
        j = 1    # index in flist
        for (i = 1; i <= n; i++) {
            if (index(f[i], "-") != 0) { # a range
                m = split(f[i], g, "-")
                if (m != 2 || g[1] >= g[2]) {
                    printf("bad field list: %s\n",
                                      f[i]) > "/dev/stderr"
                    exit 1
                }
                for (k = g[1]; k <= g[2]; k++)
                    flist[j++] = k
            } else
                flist[j++] = f[i]
        }
        nfields = j - 1
    }

The set_charlist function is more complicated than set_fieldlist. The idea here is to use gawk’s FIELDWIDTHS variable (see Section 5.6 [Reading Fixed-width Data], page 49), which describes constant width input. When using a character list, that is exactly what we have.

Setting up FIELDWIDTHS is more complicated than simply listing the fields that need to be printed. We have to keep track of the fields to be printed, and also the intervening characters that have to be skipped. For example, suppose you wanted characters one through eight, 15, and 22 through 35. You would use ‘-c 1-8,15,22-35’. The necessary value for FIELDWIDTHS would be "8 6 1 6 14". This gives us five fields, and what should be printed are $1, $3, and $5. The intermediate fields are “filler,” stuff in between the desired data.

flist lists the fields to be printed, and t tracks the complete field list, including filler fields.

function set_charlist(    field, i, j, f, g, t,
              filler, last, len)
{
field = 1   # count total fields
n = split(fieldlist, f, ",")
j = 1       # index in flist
for (i = 1; i <= n; i++) {
if (index(f[i], "-") != 0) { # range
m = split(f[i], g, "-")
if (m != 2 || g[1] >= g[2]) {
    printf("bad character list: %s\n",
                   f[i]) > "/dev/stderr"
    exit 1
}
len = g[2] - g[1] + 1
if (g[1] > 1)  # compute length of filler
    filler = g[1] - last - 1
else
    filler = 0
if (filler)
    t[field++] = filler
t[field++] = len  # length of field
last = g[2]
flist[j++] = field - 1
} else {
if (f[i] > 1)
    filler = f[i] - last - 1
else
    filler = 0
if (filler)
    t[field++] = filler
t[field++] = 1
last = f[i]
flist[j++] = field - 1
}
}
FIELDWIDTHS = join(t, 1, field - 1)
nfields = j - 1
}

Here is the rule that actually processes the data. If the ‘-s’ option was given, then suppress will be true. The first if statement makes sure that the input record does have the field separator. If cut is processing fields, suppress is true, and the field separator character is not in the record, then the record is skipped.

If the record is valid, then at this point, gawk has split the data into fields, either using the character in FS or using fixed-length fields and FIELDWIDTHS. The loop goes through the list of fields that should be printed. If the corresponding field has data in it, it is printed. If the next field also has data, then the separator character is written out in between the fields.

    {
        if (by_fields && suppress && $0 !~ FS)
            next
        for (i = 1; i <= nfields; i++) {
            if ($flist[i] != "") {
                printf "%s", $flist[i]
                if (i < nfields && $flist[i+1] != "")
                    printf "%s", OFS
            }
        }
        print ""
    }

This version of cut relies on gawk’s FIELDWIDTHS variable to do the character-based cutting. While it would be possible in other awk implementations to use substr (see Section 12.3 [Built-in Functions for String Manipulation], page 135), it would also be extremely painful to do so. The FIELDWIDTHS variable supplies an elegant solution to the problem of picking the input line apart by characters.

16.1.2 Searching for Regular Expressions in Files

The egrep utility searches files for patterns. It uses regular expressions that are almost identical to those available in awk (see Section 7.1.2 [Regular Expression Constants], page 77). It is used this way:

egrep [ options ] 'pattern' files … The pattern is a regexp. In typical usage, the regexp is quoted to prevent

the shell from expanding any of the special characters as file name wildcards. Normally, egrep prints the lines that matched. If multiple file names are provided on the command line, each output line is preceded by the name of the file and a colon.

The options are:

-c Print out a count of the lines that matched the pattern, instead of the lines themselves.

-s Be silent. No output is produced, and the exit value indicates whether or not the pattern was matched.

-v Invert the sense of the test. egrep prints the lines that do not match the pattern, and exits successfully if the pattern was not matched.

-i Ignore case distinctions in both the pattern and the input data.

-l Only print the names of the files that matched, not the lines that matched.

-e pattern Use pattern as the regexp to match. The purpose of the ‘-e’ option is to allow patterns that start with a ‘-’.

This version uses the getopt library function (see Section 15.10 [Processing Command Line Options], page 184), and the file transition library program (see Section 15.9 [Noting Data File Boundaries], page 183).

The program begins with a descriptive comment, and then a BEGIN rule that processes the command line arguments with getopt. The ‘-i’ (ignore case) option is particularly easy with gawk; we just use the IGNORECASE built in variable (see Chapter 10 [Built-in Variables], page 115).

    # egrep.awk --- simulate egrep in awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # Options:
    #    -c    count of lines
    #    -s    silent - use exit value
    #    -v    invert test, success if no match
    #    -i    ignore case
    #    -l    print filenames only
    #    -e    argument is pattern
    BEGIN {
        while ((c = getopt(ARGC, ARGV, "ce:svil")) != -1) {
            if (c == "c")
                count_only++
            else if (c == "s")
                no_print++
            else if (c == "v")
                invert++

            else if (c == "i")
                IGNORECASE = 1
            else if (c == "l")
                filenames_only++
            else if (c == "e")
                pattern = Optarg
            else
                usage()
        }

Next comes the code that handles the egrep specific behavior. If no pattern was supplied with ‘-e’, the first non-option on the command line is used. The awk command line arguments up to ARGV[Optind] are cleared, so that awk won’t try to process them as files. If no files were specified, the standard input is used, and if multiple files were specified, we make sure to note this so that the file names can precede the matched lines in the output.

The last two lines are commented out, since they are not needed in gawk. They should be uncommented if you have to use another version of awk.

        if (pattern == "")
            pattern = ARGV[Optind++]
        for (i = 1; i < Optind; i++)
            ARGV[i] = ""
        if (Optind >= ARGC) {
            ARGV[1] = "-"
            ARGC = 2
        } else if (ARGC - Optind > 1)
            do_filenames++
    #    if (IGNORECASE)
    #        pattern = tolower(pattern)
    }

The next set of lines should be uncommented if you are not using gawk. This rule translates all the characters in the input line into lower-case if the ‘-i’ option was specified. The rule is commented out since it is not necessary with gawk.

    #{
    #    if (IGNORECASE)
    #        $0 = tolower($0)
    #}

The beginfile function is called by the rule in ‘ftrans.awk’ when each new file is processed. In this case, it is very simple; all it does is initialize a variable fcount to zero. fcount tracks how many lines in the current file matched the pattern.

    function beginfile(junk)
    {
        fcount = 0
    }

The endfile function is called after each file has been processed. It is used only when the user wants a count of the number of lines that matched. no_print will be true only if the exit status is desired. count_only will be true if line counts are desired. egrep will therefore only print line counts if printing and counting are enabled. The output format must be adjusted depending upon the number of files to be processed. Finally, fcount is added to total, so that we know how many lines altogether matched the pattern.

    function endfile(file)
    {
        if (! no_print && count_only)
            if (do_filenames)
                print file ":" fcount
            else
                print fcount
        total += fcount
    }

This rule does most of the work of matching lines. The variable matches will be true if the line matched the pattern. If the user wants lines that did not match, the sense of the matches is inverted using the ‘!’ operator. fcount is incremented with the value of matches, which will be either one or zero, depending upon a successful or unsuccessful match. If the line did not match, the next statement just moves on to the next record.

There are several optimizations for performance in the following few lines of code. If the user only wants exit status (no_print is true), and we don’t have to count lines, then it is enough to know that one line in this file matched, and we can skip on to the next file with nextfile. Along similar lines, if we are only printing file names, and we don’t need to count lines, we can print the file name, and then skip to the next file with nextfile.

Finally, each line is printed, with a leading filename and colon if necessary.

    {
        matches = ($0 ~ pattern)
        if (invert)
            matches = ! matches
        fcount += matches    # 1 or 0

if (! matches)
    next
if (no_print && ! count_only)
    nextfile
if (filenames_only && ! count_only) {
    print FILENAME
    nextfile
}
if (do_filenames && ! count_only)
    print FILENAME ":" $0
else if (! count_only)
    print
}

The END rule takes care of producing the correct exit status. If there were no matches, the exit status is one, otherwise it is zero.

END    \
{
if (total == 0)
    exit 1
exit 0
}

The usage function prints a usage message in case of invalid options and then exits.

function usage(    e)
{
e = "Usage: egrep [-csvil] [-e pat] [files ...]"
print e > "/dev/stderr"
exit 1
}

The variable e is used so that the function fits nicely on the printed page.

Just a note on programming style. You may have noticed that the END rule uses backslash continuation, with the open brace on a line by itself. This is so that it more closely resembles the way functions are written. Many of the examples in this chapter use this style. You can decide for yourself if you like writing your BEGIN and END rules this way, or not.

16.1.3 Printing Out User Information

The id utility lists a user’s real and effective user-id numbers, real and effective group-id numbers, and the user’s group set, if any. id will only print the effective user-id and group-id if they are different from the real ones. If possible, id will also supply the corresponding user and group names. The output might look like this:

    $ id
    ⊣ uid=2076(arnold) gid=10(staff) groups=10(staff),4(tty)

This information is exactly what is provided by gawk’s ‘/dev/user’ special file (see Section 6.7 [Special File Names in gawk], page 72). However, the id utility provides a more palatable output than just a string of numbers.

Here is a simple version of id written in awk. It uses the user database library functions (see Section 15.11 [Reading the User Database], page 190), and the group database library functions (see Section 15.12 [Reading the Group Database], page 195).

The program is fairly straightforward. All the work is done in the BEGIN rule. The user and group id numbers are obtained from ‘/dev/user’. If there is no support for ‘/dev/user’, the program gives up.

The code is repetitive. The entry in the user database for the real user-id number is split into parts at the ‘:’. The name is the first field. Similar code is used for the effective user-id number, and the group numbers.

    # id.awk --- implement id in awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # output is:
    # uid=12(foo) euid=34(bar) gid=3(baz) \
    #             egid=5(blat) groups=9(nine),2(two),1(one)
    BEGIN    \
    {
        if ((getline < "/dev/user") < 0) {
            err = "id: no /dev/user support - cannot run"
            print err > "/dev/stderr"
            exit 1
        }
        close("/dev/user")
        uid = $1
        euid = $2
        gid = $3
        egid = $4

printf("uid=%d", uid)
pw = getpwuid(uid)
if (pw != "") {
split(pw, a, ":")
printf("(%s)", a[1])
}
if (euid != uid) {
printf(" euid=%d", euid)
pw = getpwuid(euid)
if (pw != "") {
    split(pw, a, ":")
    printf("(%s)", a[1])
}
}
printf(" gid=%d", gid)
pw = getgrgid(gid)
if (pw != "") {
split(pw, a, ":")
printf("(%s)", a[1])
}
if (egid != gid) {
printf(" egid=%d", egid)
pw = getgrgid(egid)
if (pw != "") {
    split(pw, a, ":")
    printf("(%s)", a[1])
}
}
if (NF > 4) {
printf(" groups=");
for (i = 5; i <= NF; i++) {
    printf("%d", $i)
    pw = getgrgid($i)
    if (pw != "") {
        split(pw, a, ":")
        printf("(%s)", a[1])
    }

                if (i < NF) printf(",")
            }
        }
        print ""
    }

16.1.4 Splitting a Large File Into Pieces

The split program splits large text files into smaller pieces. By default, the output files are named ‘xaa’, ‘xab’, and so on. Each file has 1000 lines in it, with the likely exception of the last file. To change the number of lines in each file, you supply a number on the command line preceded with a minus, e.g., ‘-500’ for files with 500 lines in them instead of 1000. To change the name of the output files to something like ‘myfileaa’, ‘myfileab’, and so on, you supply an additional argument that specifies the filename.

Here is a version of split in awk. It uses the ord and chr functions presented in Section 15.5 [Translating Between Characters and Numbers], page 172.

The program first sets its defaults, and then tests to make sure there are not too many arguments. It then looks at each argument in turn. The first argument could be a minus followed by a number. If it is, this happens to look like a negative number, so it is made positive, and that is the count of lines. The data file name is skipped over, and the final argument is used as the prefix for the output file names.

    # split.awk --- do split in awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # usage: split [-num] [file] [outname]
    BEGIN {
        outfile = "x"    # default
        count = 1000
        if (ARGC > 4)
            usage()
        i = 1
        if (ARGV[i] ~ /^-[0-9]+$/) {
            count = -ARGV[i]
            ARGV[i] = ""
            i++
        }

# test argv in case reading from stdin instead of file
if (i in ARGV)
    i++    # skip data file name
if (i in ARGV) {
    outfile = ARGV[i]
    ARGV[i] = ""
}
s1 = s2 = "a"
out = (outfile s1 s2)
}

The next rule does most of the work. tcount (temporary count) tracks how many lines have been printed to the output file so far. If it is greater than count, it is time to close the current file and start a new one. s1 and s2 track the current suffixes for the file name. If they are both ‘z’, the file is just too big. Otherwise, s1 moves to the next letter in the alphabet and s2 starts over again at ‘a’.

{
if (++tcount > count) {
    close(out)
    if (s2 == "z") {
        if (s1 == "z") {
            printf("split: %s is too large to split\n", \
                   FILENAME) > "/dev/stderr"
            exit 1
        }
        s1 = chr(ord(s1) + 1)
        s2 = "a"
    } else
        s2 = chr(ord(s2) + 1)
    out = (outfile s1 s2)
    tcount = 1
}
print > out
}

The usage function simply prints an error message and exits.

function usage(   e)
{
e = "usage: split [-num] [file] [outname]"
print e > "/dev/stderr"
exit 1
}

The variable e is used so that the function fits nicely on the page.

This program is a bit sloppy; it relies on awk to close the last file for it automatically, instead of doing it in an END rule.

16.1.5 Duplicating Output Into Multiple Files

The tee program is known as a “pipe fitting.” tee copies its standard input to its standard output, and also duplicates it to the files named on the command line. Its usage is:

    tee [-a] file . . .

The ‘-a’ option tells tee to append to the named files, instead of truncating them and starting over.

The BEGIN rule first makes a copy of all the command line arguments, into an array named copy. ARGV[0] is not copied, since it is not needed. tee cannot use ARGV directly, since awk will attempt to process each file named in ARGV as input data.

If the first argument is ‘-a’, then the flag variable append is set to true, and both ARGV[1] and copy[1] are deleted. If ARGC is less than two, then no file names were supplied, and tee prints a usage message and exits. Finally, awk is forced to read the standard input by setting ARGV[1] to "-", and ARGC to two.

    # tee.awk --- tee in awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # Revised December 1995
    BEGIN    \
    {
        for (i = 1; i < ARGC; i++)
            copy[i] = ARGV[i]
        if (ARGV[1] == "-a") {
            append = 1
            delete ARGV[1]
            delete copy[1]
            ARGC--
        }
        if (ARGC < 2) {
            print "usage: tee [-a] file ..." > "/dev/stderr"
            exit 1
        }
        ARGV[1] = "-"
        ARGC = 2
    }

The single rule does all the work. Since there is no pattern, it is executed for each line of input. The body of the rule simply prints the line into each file on the command line, and then to the standard output.

    {
        # moving the if outside the loop makes it run faster
        if (append)
            for (i in copy)
                print >> copy[i]
        else
            for (i in copy)
                print > copy[i]
        print
    }

It would have been possible to code the loop this way:

    for (i in copy)
        if (append)
            print >> copy[i]
        else
            print > copy[i]

This is more concise, but it is also less efficient. The ‘if’ is tested for each record and for each output file. By duplicating the loop body, the ‘if’ is only tested once for each input record. If there are N input records and M input files, the first method only executes N ‘if’ statements, while the second would execute N*M ‘if’ statements.

Finally, the END rule cleans up, by closing all the output files.

    END    \
    {
        for (i in copy)
            close(copy[i])
    }

16.1.6 Printing Non-duplicated Lines of Text

The uniq utility reads sorted lines of data on its standard input, and (by default) removes duplicate lines. In other words, only unique lines are printed, hence the name. uniq has a number of options. The usage is:

uniq [-udc [-n]] [+n] [ input file [ output file ]] The option meanings are:

-d Only print repeated lines.

-u Only print non-repeated lines.

-c Count lines. This option overrides ‘-d’ and ‘-u’. Both repeated and non-repeated lines are counted.

-n Skip n fields before comparing lines. The definition of fields is similar to awk’s default: non-whitespace characters separated by runs of spaces and/or tabs.

+n Skip n characters before comparing lines. Any fields specified with ‘-n’ are skipped first.

input file Data is read from the input file named on the command line, instead of from the standard input.

output file The generated output is sent to the named output file, instead of to the standard output.

Normally uniq behaves as if both the ‘-d’ and ‘-u’ options had been provided.

Here is an awk implementation of uniq. It uses the getopt library function (see Section 15.10 [Processing Command Line Options], page 184), and the join library function (see Section 15.6 [Merging an Array Into a String], page 174).

The program begins with a usage function and then a brief outline of the options and their meanings in a comment.

The BEGIN rule deals with the command line arguments and options. It uses a trick to get getopt to handle options of the form ‘-25’, treating such an option as the option letter ‘2’ with an argument of ‘5’. If indeed two or more digits were supplied (Optarg looks like a number), Optarg is concatenated with the option digit, and then result is added to zero to make it into a number. If there is only one digit in the option, then Optarg is not needed, and Optind must be decremented so that getopt will process it next time. This code is admittedly a bit tricky.

If no options were supplied, then the default is taken, to print both repeated and non-repeated lines. The output file, if provided, is assigned to outputfile. Earlier, outputfile was initialized to the standard output,

‘/dev/stdout’.
    # uniq.awk --- do uniq in awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    function usage(    e)
    {
        e = "Usage: uniq [-udc [-n]] [+n] [ in [ out ]]"
        print e > "/dev/stderr"
        exit 1
    }

# -c    count lines. overrides -d and -u
# -d    only repeated lines
# -u    only non-repeated lines
# -n    skip n fields
# +n    skip n characters, skip fields first
BEGIN    \
{
count = 1
outputfile = "/dev/stdout"
opts = "udc0:1:2:3:4:5:6:7:8:9:"
while ((c = getopt(ARGC, ARGV, opts)) != -1) {
    if (c == "u")
        non_repeated_only++
    else if (c == "d")
        repeated_only++
    else if (c == "c")
        do_count++
    else if (index("0123456789", c) != 0) {
        # getopt requires args to options
        # this messes us up for things like -5
        if (Optarg ~ /^[0-9]+$/)
            fcount = (c Optarg) + 0
        else {
            fcount = c + 0
            Optind--
        }
    } else
        usage()
}
if (ARGV[Optind] ~ /^\+[0-9]+$/) {
    charcount = substr(ARGV[Optind], 2) + 0
    Optind++
}
for (i = 1; i < Optind; i++)
    ARGV[i] = ""
if (repeated_only == 0 && non_repeated_only == 0)
    repeated_only = non_repeated_only = 1

if (ARGC - Optind == 2) {
outputfile = ARGV[ARGC - 1]
ARGV[ARGC - 1] = ""
}
}

The following function, are_equal, compares the current line, $0, to the previous line, last. It handles skipping fields and characters.

If no field count and no character count were specified, are_equal simply returns one or zero depending upon the result of a simple string comparison of last and $0. Otherwise, things get more complicated.

If fields have to be skipped, each line is broken into an array using split (see Section 12.3 [Built-in Functions for String Manipulation], page 135), and then the desired fields are joined back into a line using join. The joined lines are stored in clast and cline. If no fields are skipped, clast and cline are set to last and $0 respectively.

Finally, if characters are skipped, substr is used to strip off the leading charcount characters in clast and cline. The two strings are then compared, and are_equal returns the result.

function are_equal(    n, m, clast, cline, alast, aline)
{
if (fcount == 0 && charcount == 0)
return (last == $0)
if (fcount > 0) {
n = split(last, alast)
m = split($0, aline)
clast = join(alast, fcount+1, n)
cline = join(aline, fcount+1, m)
} else {
clast = last
cline = $0
}
if (charcount) {
clast = substr(clast, charcount + 1)
cline = substr(cline, charcount + 1)
}
return (clast == cline)
}

The following two rules are the body of the program. The first one is executed only for the very first line of data. It sets last equal to $0, so that subsequent lines of text have something to be compared to.

The second rule does the work. The variable equal will be one or zero depending upon the results of are_equal’s comparison. If uniq is counting repeated lines, then the count variable is incremented if the lines are equal. Otherwise the line is printed and count is reset, since the two lines are not equal.

If uniq is not counting, count is incremented if the lines are equal. Otherwise, if uniq is counting repeated lines, and more than one line has been seen, or if uniq is counting non-repeated lines, and only one line has been seen, then the line is printed, and count is reset.

Finally, similar logic is used in the END rule to print the final line of input data.

    NR == 1 { last = $0 next }

    {
        equal = are_equal()
        if (do_count) {    # overrides -d and -u
            if (equal)
                count++
            else {
                printf("%4d %s\n", count, last) > outputfile
                last = $0
                count = 1    # reset
            }
            next
        }
        if (equal)
            count++
        else {
            if ((repeated_only && count > 1) ||
                (non_repeated_only && count == 1))
                    print last > outputfile
            last = $0
            count = 1
        }
    }

    END {
        if (do_count)
            printf("%4d %s\n", count, last) > outputfile
        else if ((repeated_only && count > 1) ||
                (non_repeated_only && count == 1))
            print last > outputfile
    }

16.1.7 Counting Things

The wc (word count) utility counts lines, words, and characters in one or more input files. Its usage is:

wc [-lwc] [ files …]

If no files are specified on the command line, wc reads its standard input. If there are multiple files, it will also print total counts for all the files. The options and their meanings are:

-l Only count lines.

-w Only count words. A “word” is a contiguous sequence of nonwhitespace characters, separated by spaces and/or tabs. Happily, this is the normal way awk separates fields in its input data.

-c Only count characters.

Implementing wc in awk is particularly elegant, since awk does a lot of the work for us; it splits lines into words (i.e. fields) and counts them, it counts lines (i.e. records) for us, and it can easily tell us how long a line is.

This version uses the getopt library function (see Section 15.10 [Processing Command Line Options], page 184), and the file transition functions (see Section 15.9 [Noting Data File Boundaries], page 183).

This version has one major difference from traditional versions of wc. Our version always prints the counts in the order lines, words, and characters. Traditional versions note the order of the ‘-l’, ‘-w’, and ‘-c’ options on the command line, and print the counts in that order.

The BEGIN rule does the argument processing. The variable print_total will be true if more than one file was named on the command line.

    # wc.awk --- count lines, words, characters
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # Options:
    #    -l    only count lines
    #    -w    only count words
    #    -c    only count characters

#
# Default is to count lines, words, characters
BEGIN {
# let getopt print a message about
# invalid options. we ignore them
while ((c = getopt(ARGC, ARGV, "lwc")) != -1) {
    if (c == "l")
        do_lines = 1
    else if (c == "w")
        do_words = 1
    else if (c == "c")
        do_chars = 1
}
for (i = 1; i < Optind; i++)
    ARGV[i] = ""
# if no options, do all
if (! do_lines && ! do_words && ! do_chars)
    do_lines = do_words = do_chars = 1
print_total = (ARGC - i > 2)
}

The beginfile function is simple; it just resets the counts of lines, words, and characters to zero, and saves the current file name in fname.

The endfile function adds the current file’s numbers to the running totals of lines, words, and characters. It then prints out those numbers for the file that was just read. It relies on beginfile to reset the numbers for the following data file.

function beginfile(file)
{
chars = lines = words = 0
fname = FILENAME
}
function endfile(file)
{
tchars += chars
tlines += lines
twords += words
if (do_lines)
    printf "\t%d", lines

if (do_words)
    printf "\t%d", words
if (do_chars)
    printf "\t%d", chars
printf "\t%s\n", fname
}

There is one rule that is executed for each line. It adds the length of the record to chars. It has to add one, since the newline character separating records (the value of RS) is not part of the record itself. lines is incremented for each line read, and words is incremented by the value of NF, the number of “words” on this line.^{1}

Finally, the END rule simply prints the totals for all the files.

# do per line
{
chars += length($0) + 1    # get newline
lines++
words += NF
}
END {
if (print_total) {
    if (do_lines)
        printf "\t%d", tlines
    if (do_words)
        printf "\t%d", twords
    if (do_chars)
        printf "\t%d", tchars
    print "\ttotal"
}
}

16.2 A Grab Bag of awk Programs

This section is a large “grab bag” of miscellaneous programs. We hope you find them both interesting and enjoyable.

1 Examine the code in Section 15.9 [Noting Data File Boundaries], page 183. Why must wc use a separate lines variable, instead of using the value of FNR in endfile?

16.2.1 Finding Duplicated Words in a Document

A common error when writing large amounts of prose is to accidentally duplicate words. Often you will see this in text as something like “the the program does the following . . ..” When the text is on-line, often the duplicated words occur at the end of one line and the beginning of another, making them very difficult to spot.

This program, ‘dupword.awk’, scans through a file one line at a time, and looks for adjacent occurrences of the same word. It also saves the last word on a line (in the variable prev) for comparison with the first word on the next line.

The first two statements make sure that the line is all lower-case, so that, for example, “The” and “the” compare equal to each other. The second statement removes all non-alphanumeric and non-whitespace characters from the line, so that punctuation does not affect the comparison either. This sometimes leads to reports of duplicated words that really are different, but this is unusual.

    # dupword --- find duplicate words in text
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # December 1991
    {
        $0 = tolower($0)
        gsub(/[^A-Za-z0-9 \t]/, "");
        if ($1 == prev)
            printf("%s:%d: duplicate %s\n",
                FILENAME, FNR, $1)
        for (i = 2; i <= NF; i++)
            if ($i == $(i-1))
                printf("%s:%d: duplicate %s\n",
                    FILENAME, FNR, $i)
        prev = $NF
    }

16.2.2 An Alarm Clock Program

The following program is a simple “alarm clock” program. You give it a time of day, and an optional message. At the given time, it prints the message on the standard output. In addition, you can give it the number of times to repeat the message, and also a delay between repetitions.

This program uses the gettimeofday function from Section 15.8 [Managing the Time of Day], page 181.

All the work is done in the BEGIN rule. The first part is argument checking and setting of defaults; the delay, the count, and the message to print. If the user supplied a message, but it does not contain the ASCII BEL character (known as the “alert” character, ‘\a’), then it is added to the message. (On many systems, printing the ASCII BEL generates some sort of audible alert. Thus, when the alarm goes off, the system calls attention to itself, in case the user is not looking at their computer or terminal.)

# alarm --- set an alarm
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
# May 1993
# usage: alarm time [ "message" [ count [ delay ] ] ]
BEGIN    \
{
# Initial argument sanity checking
usage1 = "usage: alarm time ['message' [count [delay]]]"
usage2 = sprintf("\t(%s) time ::= hh:mm", ARGV[1])
if (ARGC < 2) {
print usage > "/dev/stderr"
exit 1
} else if (ARGC == 5) {
delay = ARGV[4] + 0
count = ARGV[3] + 0
message = ARGV[2]
} else if (ARGC == 4) {
count = ARGV[3] + 0
message = ARGV[2]
} else if (ARGC == 3) {
message = ARGV[2]
} else if (ARGV[1] !~ /[0-9]?[0-9]:[0-9][0-9]/) {
print usage1 > "/dev/stderr"
print usage2 > "/dev/stderr"
exit 1
}
# set defaults for once we reach the desired time
if (delay == 0)
delay = 180    # 3 minutes
if (count == 0)
count = 5

if (message == "")
    message = sprintf("\aIt is now %s!\a", ARGV[1])
else if (index(message, "\a") == 0)
    message = "\a" message "\a"

The next section of code turns the alarm time into hours and minutes, and converts it if necessary to a 24-hour clock. Then it turns that time into a count of the seconds since midnight. Next it turns the current time into a count of seconds since midnight. The difference between the two is how long to wait before setting off the alarm.

# split up dest time
split(ARGV[1], atime, ":")
hour = atime[1] + 0    # force numeric
minute = atime[2] + 0  # force numeric
# get current broken down time
gettimeofday(now)
# if time given is 12-hour hours and it's after that
# hour, e.g., `alarm 5:30' at 9 a.m. means 5:30 p.m.,
# then add 12 to real hour
if (hour < 12 && now["hour"] > hour)
    hour += 12
# set target time in seconds since midnight
target = (hour * 60 * 60) + (minute * 60)
# get current time in seconds since midnight
current = (now["hour"] * 60 * 60) + \
           (now["minute"] * 60) + now["second"]
# how long to sleep for
naptime = target - current
if (naptime <= 0) {
    print "time is in the past!" > "/dev/stderr"
    exit 1
}

Finally, the program uses the system function (see Section 12.4 [Built-in Functions for Input/Output], page 144) to call the sleep utility. The sleep utility simply pauses for the given number of seconds. If the exit status is not zero, the program assumes that sleep was interrupted, and exits. If sleep exited with an OK status (zero), then the program prints the message in a loop, again using sleep to delay for however many seconds are necessary.

        # zzzzzz..... go away if interrupted
        if (system(sprintf("sleep %d", naptime)) != 0)
            exit 1
        # time to notify!
        command = sprintf("sleep %d", delay)
        for (i = 1; i <= count; i++) {
            print message
            # if sleep command interrupted, go away
            if (system(command) != 0)
                break
        }
        exit 0
    }

16.2.3 Transliterating Characters

The system tr utility transliterates characters. For example, it is often used to map upper-case letters into lower-case, for further processing.

generate data | tr '[A-Z]' '[a-z]' | process data … You give tr two lists of characters enclosed in square brackets. Usually,

the lists are quoted to keep the shell from attempting to do a filename expansion.^{2} When processing the input, the first character in the first list is replaced with the first character in the second list, the second character in the first list is replaced with the second character in the second list, and so on. If there are more characters in the “from” list than in the “to” list, the last character of the “to” list is used for the remaining characters in the “from” list.

Some time ago, a user proposed to us that we add a transliteration function to gawk. Being opposed to “creeping featurism,” I wrote the following program to prove that character transliteration could be done with a userlevel function. This program is not as complete as the system tr utility, but it will do most of the job.

The translate program demonstrates one of the few weaknesses of standard awk: dealing with individual characters is very painful, requiring repeated use of the substr, index, and gsub built-in functions (see Section 12.3 [Built-in Functions for String Manipulation], page 135).^{3}

2 On older, non-POSIX systems, tr often does not require that the lists be

enclosed in square brackets and quoted. This is a feature. 3 This program was written before gawk acquired the ability to split each

character in a string into separate array elements. How might this ability simplify the program?

There are two functions. The first, stranslate, takes three arguments.

from A list of characters to translate from.

to A list of characters to translate to.

target The string to do the translation on.

Associative arrays make the translation part fairly easy. t_ar holds the “to” characters, indexed by the “from” characters. Then a simple loop goes through from, one character at a time. For each character in from, if the character appears in target, gsub is used to change it to the corresponding to character.

The translate function simply calls stranslate using $0 as the target. The main program sets two global variables, FROM and TO, from the command line, and then changes ARGV so that awk will read from the standard input.

Finally, the processing rule simply calls translate for each record.

    # translate --- do tr like stuff
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # August 1989
    # bugs: does not handle things like: tr A-Z a-z, it has
    # to be spelled out. However, if `to' is shorter than `from',
    # the last character in `to' is used for the rest of `from'.
    function stranslate(from, to, target,     lf, lt, t_ar, i, c)
    {
        lf = length(from)
        lt = length(to)
        for (i = 1; i <= lt; i++)
            t_ar[substr(from, i, 1)] = substr(to, i, 1)
        if (lt < lf)
            for (; i <= lf; i++)
                t_ar[substr(from, i, 1)] = substr(to, lt, 1)
        for (i = 1; i <= lf; i++) {
            c = substr(from, i, 1)
            if (index(target, c) > 0)
                gsub(c, t_ar[c], target)
        }
        return target
    }

    function translate(from, to)
    {
        return $0 = stranslate(from, to, $0)
    }
    # main program
    BEGIN {
        if (ARGC < 3) {
            print "usage: translate from to" > "/dev/stderr"
            exit
        }
        FROM = ARGV[1]
        TO = ARGV[2]
        ARGC = 2
        ARGV[1] = "-"
    }
    {
        translate(FROM, TO)
        print
    }

While it is possible to do character transliteration in a user-level function, it is not necessarily efficient, and we started to consider adding a builtin function. However, shortly after writing this program, we learned that the System V Release 4 awk had added the toupper and tolower functions. These functions handle the vast majority of the cases where character transliteration is necessary, and so we chose to simply add those functions to gawk as well, and then leave well enough alone.

An obvious improvement to this program would be to set up the t_ar array only once, in a BEGIN rule. However, this assumes that the “from” and “to” lists will never change throughout the lifetime of the program.

16.2.4 Printing Mailing Labels

Here is a “real world”^{4} program. This script reads lists of names and addresses, and generates mailing labels. Each page of labels has 20 labels on it, two across and ten down. The addresses are guaranteed to be no more than five lines of data. Each address is separated from the next by a blank line.

4 “Real world” is defined as “a program actually used to get something done.”

The basic idea is to read 20 labels worth of data. Each line of each label is stored in the line array. The single rule takes care of filling the line array and printing the page when 20 labels have been read.

The BEGIN rule simply sets RS to the empty string, so that awk will split records at blank lines (see Section 5.1 [How Input is Split into Records], page 37). It sets MAXLINES to 100, since MAXLINE is the maximum number of lines on the page (20 * 5 = 100).

Most of the work is done in the printpage function. The label lines are stored sequentially in the line array. But they have to be printed horizontally; line[1] next to line[6], line[2] next to line[7], and so on. Two loops are used to accomplish this. The outer loop, controlled by i, steps through every 10 lines of data; this is each row of labels. The inner loop, controlled by j, goes through the lines within the row. As j goes from zero to four, ‘i+j’ is the j’th line in the row, and ‘i+j+5’ is the entry next to it. The output ends up looking something like this:

    line 1          line 6
    line 2          line 7
    line 3          line 8
    line 4          line 9
    line 5          line 10

As a final note, at lines 21 and 61, an extra blank line is printed, to keep the output lined up on the labels. This is dependent on the particular brand of labels in use when the program was written. You will also note that there are two blank lines at the top and two blank lines at the bottom.

The END rule arranges to flush the final page of labels; there may not have been an even multiple of 20 labels in the data.

    # labels.awk
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # June 1992
    # Program to print labels.  Each label is 5 lines of data
    # that may have blank lines.  The label sheets have 2
    # blank lines at the top and 2 at the bottom.
    BEGIN    { RS = "" ; MAXLINES = 100 }
    function printpage(    i, j)
    {
        if (Nlines <= 0)
            return
        printf "\n\n"        # header

for (i = 1; i <= Nlines; i += 10) {
    if (i == 21 || i == 61)
        print ""
    for (j = 0; j < 5; j++) {
        if (i + j > MAXLINES)
            break
        printf "   %-41s %s\n", line[i+j], line[i+j+5]
    }
    print ""
}
printf "\n\n"        # footer
for (i in line)
    line[i] = ""
}
# main rule
{
if (Count >= 20) {
    printpage()
    Count = 0
    Nlines = 0
}
n = split($0, a, "\n")
for (i = 1; i <= n; i++)
    line[++Nlines] = a[i]
for (; i <= 5; i++)
    line[++Nlines] = ""
Count++
}
END    \
{
printpage()
}

16.2.5 Generating Word Usage Counts

The following awk program prints the number of occurrences of each word in its input. It illustrates the associative nature of awk arrays by using strings as subscripts. It also demonstrates the ‘for x in array’ construction.

Finally, it shows how awk can be used in conjunction with other utility programs to do a useful task of some complexity with a minimum of effort. Some explanations follow the program listing.

    awk '
    # Print list of word frequencies
    {
        for (i = 1; i <= NF; i++)
            freq[$i]++
    }
    END {
        for (word in freq)
            printf "%s\t%d\n", word, freq[word]
    }'

The first thing to notice about this program is that it has two rules. The first rule, because it has an empty pattern, is executed on every line of the input. It uses awk’s field-accessing mechanism (see Section 5.2 [Examining Fields], page 40) to pick out the individual words from the line, and the built-in variable NF (see Chapter 10 [Built-in Variables], page 115) to know how many fields are available.

For each input word, an element of the array freq is incremented to reflect that the word has been seen an additional time.

The second rule, because it has the pattern END, is not executed until the input has been exhausted. It prints out the contents of the freq table that has been built up inside the first action.

This program has several problems that would prevent it from being useful by itself on real text files:

• Words are detected using the awk convention that fields are separated by whitespace and that other characters in the input (except newlines) don’t have any special meaning to awk. This means that punctuation characters count as part of words.

• The awk language considers upper- and lower-case characters to be distinct. Therefore, ‘bartender’ and ‘Bartender’ are not treated as the same word. This is undesirable since, in normal text, words are capitalized if they begin sentences, and a frequency analyzer should not be sensitive to capitalization.

• The output does not come out in any useful order. You’re more likely to be interested in which words occur most frequently, or having an alphabetized table of how frequently each word occurs.

The way to solve these problems is to use some of the more advanced features of the awk language. First, we use tolower to remove case distinctions. Next, we use gsub to remove punctuation characters. Finally, we use the system sort utility to process the output of the awk script. Here is the new version of the program:

    # Print list of word frequencies
    {
        $0 = tolower($0)    # remove case distinctions
        gsub(/[^a-z0-9_\t]/, "", $0)  # remove punctuation
        for (i = 1; i <= NF; i++)
            freq[$i]++
    }
    END {
        for (word in freq)
            printf "%s\t%d\n", word, freq[word]
    }

Assuming we have saved this program in a file named ‘wordfreq.awk’, and that the data is in ‘file1’, the following pipeline

    awk -f wordfreq.awk file1 | sort +1 -nr

produces a table of the words appearing in ‘file1’ in order of decreasing frequency.

The awk program suitably massages the data and produces a word frequency table, which is not ordered.

The awk script’s output is then sorted by the sort utility and printed on the terminal. The options given to sort in this example specify to sort using the second field of each input line (skipping one field), that the sort keys should be treated as numeric quantities (otherwise ‘15’ would come before ‘5’), and that the sorting should be done in descending (reverse) order.

We could have even done the sort from within the program, by changing the END action to:

    END {
        sort = "sort +1 -nr"
        for (word in freq)
            printf "%s\t%d\n", word, freq[word] | sort
        close(sort)
    }

You would have to use this way of sorting on systems that do not have true pipes.

See the general operating system documentation for more information on how to use the sort program.

16.2.6 Removing Duplicates from Unsorted Text

The uniq program (see Section 16.1.6 [Printing Non-duplicated Lines of Text], page 218), removes duplicate lines from sorted data.

Suppose, however, you need to remove duplicate lines from a data file, but that you wish to preserve the order the lines are in? A good example of this might be a shell history file. The history file keeps a copy of all the commands you have entered, and it is not unusual to repeat a command several times in a row. Occasionally you might wish to compact the history by removing duplicate entries. Yet it is desirable to maintain the order of the original commands.

This simple program does the job. It uses two arrays. The data array is indexed by the text of each line. For each line, data[$0] is incremented.

If a particular line has not been seen before, then data[$0] will be zero. In that case, the text of the line is stored in lines[count]. Each element of lines is a unique command, and the indices of lines indicate the order in which those lines were encountered. The END rule simply prints out the lines, in order.

    # histsort.awk --- compact a shell history file
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    # Thanks to Byron Rakitzis for the general idea
    {
        if (data[$0]++ == 0)
            lines[++count] = $0
    }
    END {
        for (i = 1; i <= count; i++)
            print lines[i]
    }

This program also provides a foundation for generating other useful information. For example, using the following print satement in the END rule would indicate how often a particular command was used.

    print data[lines[i]], lines[i]

This works because data[$0] was incremented each time a line was seen.

16.2.7 Extracting Programs from Texinfo Source Files

Both this chapter and the previous chapter (Chapter 15 [A Library of awk Functions], page 167), present a large number of awk programs. If you wish to experiment with these programs, it is tedious to have to type them in by hand. Here we present a program that can extract parts of a Texinfo input file into separate files.

This book is written in Texinfo, the GNU project’s document formatting language. A single Texinfo source file can be used to produce both printed and on-line documentation. Texinfo is fully documented in Texinfo—The GNU Documentation Format, available from the Free Software Foundation.

For our purposes, it is enough to know three things about Texinfo input files.

• The “at” symbol, ‘@’, is special in Texinfo, much like ‘\’ in C or awk. Literal ‘@’ symbols are represented in Texinfo source files as ‘@@’.

• Comments start with either ‘@c’ or ‘@comment’. The file extraction program will work by using special comments that start at the beginning of a line.

• Example text that should not be split across a page boundary is bracketed between lines containing ‘@group’ and ‘@end group’ commands.

The following program, ‘extract.awk’, reads through a Texinfo source file, and does two things, based on the special comments. Upon seeing ‘@c system …’, it runs a command, by extracting the command text from the control line and passing it on to the system function (see Section 12.4 [Built-in Functions for Input/Output], page 144). Upon seeing ‘@c file filename’, each subsequent line is sent to the file filename, until ‘@c endfile’ is encountered. The rules in ‘extract.awk’ will match either ‘@c’ or ‘@comment’ by letting the ‘omment’ part be optional. Lines containing ‘@group’ and ‘@end group’ are simply removed. ‘extract.awk’ uses the join library function (see Section 15.6 [Merging an Array Into a String], page 174).

The example programs in the on-line Texinfo source for Effective AWK Programming (‘gawk.texi’) have all been bracketed inside ‘file’, and ‘endfile’ lines. The gawk distribution uses a copy of ‘extract.awk’ to extract the sample programs and install many of them in a standard directory, where gawk can find them.

‘extract.awk’ begins by setting IGNORECASE to one, so that mixed uppercase and lower-case letters in the directives won’t matter.

The first rule handles calling system, checking that a command was given (NF is at least three), and also checking that the command exited with a zero exit status, signifying OK.

    # extract.awk --- extract files and run programs
    #                 from texinfo files
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # May 1993
    BEGIN    { IGNORECASE = 1 }

    /^@c(omment)?[ \t]+system/    \
    {
        if (NF < 3) {
            e = (FILENAME ":" FNR)
            e = (e  ": badly formed `system' line")
            print e > "/dev/stderr"
            next
        }
        $1 = ""
        $2 = ""
        stat = system($0)
        if (stat != 0) {
            e = (FILENAME ":" FNR)
            e = (e ": warning: system returned " stat)
            print e > "/dev/stderr"
        }
    }

The variable e is used so that the function fits nicely on the page.

The second rule handles moving data into files. It verifies that a file name was given in the directive. If the file named is not the current file, then the current file is closed. This means that an ‘@c endfile’ was not given for that file. (We should probably print a diagnostic in this case, although at the moment we do not.)

The ‘for’ loop does the work. It reads lines using getline (see Section 5.8 [Explicit Input with getline], page 53). For an unexpected end of file, it calls the unexpected_eof function. If the line is an “endfile” line, then it breaks out of the loop. If the line is an ‘@group’ or ‘@end group’ line, then it ignores it, and goes on to the next line.

Most of the work is in the following few lines. If the line has no ‘@’ symbols, it can be printed directly. Otherwise, each leading ‘@’ must be stripped off.

To remove the ‘@’ symbols, the line is split into separate elements of the array a, using the split function (see Section 12.3 [Built-in Functions for String Manipulation], page 135). Each element of a that is empty indicates two successive ‘@’ symbols in the original line. For each two empty elements (‘@@’ in the original file), we have to add back in a single ‘@’ symbol.

When the processing of the array is finished, join is called with the value of SUBSEP, to rejoin the pieces back into a single line. That line is then printed to the output file.

/^@c(omment)?[ \t]+file/    \
{
if (NF != 3) {
e = (FILENAME ":" FNR ": badly formed `file' line")
print e > "/dev/stderr"
next
}
if ($3 != curfile) {
if (curfile != "")
    close(curfile)
curfile = $3
}
for (;;) {
if ((getline line) <= 0)
    unexpected_eof()
if (line ~ /^@c(omment)?[ \t]+endfile/)
    break
else if (line ~ /^@(end[ \t]+)?group/)
    continue
if (index(line, "@") == 0) {
    print line > curfile
    continue
}
n = split(line, a, "@")
# if a[1] == "", means leading @,
# don't add one back in.
for (i = 2; i <= n; i++) {
    if (a[i] == "") { # was an @@
        a[i] = "@"
        if (a[i+1] == "")
            i++
    }
}
print join(a, 1, n, SUBSEP) > curfile
}
}

An important thing to note is the use of the ‘>’ redirection. Output done with ‘>’ only opens the file once; it stays open and subsequent output is appended to the file (see Section 6.6 [Redirecting Output of print and printf], page 70). This allows us to easily mix program text and explanatory prose for the same sample source file (as has been done here!) without any hassle. The file is only closed when a new data file name is encountered, or at the end of the input file.

Finally, the function unexpected_eof prints an appropriate error message and then exits.

The END rule handles the final cleanup, closing the open file.

    function unexpected_eof()
    {
        printf("%s:%d: unexpected EOF or error\n", \
            FILENAME, FNR) > "/dev/stderr"
        exit 1
    }
    END {
        if (curfile)
            close(curfile)
    }

16.2.8 A Simple Stream Editor

The sed utility is a “stream editor,” a program that reads a stream of data, makes changes to it, and passes the modified data on. It is often used to make global changes to a large file, or to a stream of data generated by a pipeline of commands.

While sed is a complicated program in its own right, its most common use is to perform global substitutions in the middle of a pipeline:

    command1 < orig.data | sed 's/old/new/g' | command2 > result

Here, the ‘s/old/new/g’ tells sed to look for the regexp ‘old’ on each input line, and replace it with the text ‘new’, globally (i.e. all the occurrences on a line). This is similar to awk’s gsub function (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

The following program, ‘awksed.awk’, accepts at least two command line arguments; the pattern to look for and the text to replace it with. Any additional arguments are treated as data file names to process. If none are provided, the standard input is used.

    # awksed.awk --- do s/foo/bar/g using just print
    #    Thanks to Michael Brennan for the idea
    # Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
    # August 1995

    function usage()
    {
        print "usage: awksed pat repl [files...]" > "/dev/stderr"
        exit 1
    }
    BEGIN {
        # validate arguments
        if (ARGC < 3)
            usage()
        RS = ARGV[1]
        ORS = ARGV[2]
        # don't use arguments as files
        ARGV[1] = ARGV[2] = ""
    }
    # look ma, no hands!
    {
        if (RT == "")
            printf "%s", $0
        else
            print
    }

The program relies on gawk’s ability to have RS be a regexp and on the setting of RT to the actual text that terminated the record (see Section 5.1 [How Input is Split into Records], page 37).

The idea is to have RS be the pattern to look for. gawk will automatically set $0 to the text between matches of the pattern. This is text that we wish to keep, unmodified. Then, by setting ORS to the replacement text, a simple print statement will output the text we wish to keep, followed by the replacement text.

There is one wrinkle to this scheme, which is what to do if the last record doesn’t end with text that matches RS? Using a print statement unconditionally prints the replacement text, which is not correct.

However, if the file did not end in text that matches RS, RT will be set to the null string. In this case, we can print $0 using printf (see Section 6.5 [Using printf Statements for Fancier Printing], page 64).

The BEGIN rule handles the setup, checking for the right number of arguments, and calling usage if there is a problem. Then it sets RS and ORS from the command line arguments, and sets ARGV[1] and ARGV[2] to the null string, so that they will not be treated as file names (see Section 10.3 [Using ARGC and ARGV], page 120).

The usage function prints an error message and exits.

Finally, the single rule handles the printing scheme outlined above, using print or printf as appropriate, depending upon the value of RT.

16.2.9 An Easy Way to Use Library Functions

Using library functions in awk can be very beneficial. It encourages code re-use and the writing of general functions. Programs are smaller, and therefore clearer. However, using library functions is only easy when writing awk programs; it is painful when running them, requiring multiple ‘-f’ options. If gawk is unavailable, then so too is the AWKPATH environment variable and the ability to put awk functions into a library directory (see Section 14.1 [Command Line Options], page 159).

It would be nice to be able to write programs like so:

    # library functions
    @include getopt.awk
    @include join.awk

…

    # main program
    BEGIN {
        while ((c = getopt(ARGC, ARGV, "a:b:cde")) != -1)

… …

    }

The following program, ‘igawk.sh’, provides this service. It simulates gawk’s searching of the AWKPATH variable, and also allows nested includes; i.e. a file that has been included with ‘@include’ can contain further ‘@include’ statements. igawk will make an effort to only include files once, so that nested includes don’t accidentally include a library function twice.

igawk should behave externally just like gawk. This means it should accept all of gawk’s command line arguments, including the ability to have multiple source files specified via ‘-f’, and the ability to mix command line and library source files.

The program is written using the POSIX Shell (sh) command language. The way the program works is as follows:

1. Loop through the arguments, saving anything that doesn’t represent awk source code for later, when the expanded program is run.

2. For any arguments that do represent awk text, put the arguments into a temporary file that will be expanded. There are two cases. a. Literal text, provided with ‘--source’ or ‘--source=’. This text is just echoed directly. The echo program will automatically supply a trailing newline.

b. File names provided with ‘-f’. We use a neat trick, and echo ‘@include filename’ into the temporary file. Since the file inclusion program will work the way gawk does, this will get the text of the file included into the program at the correct point.

3. Run an awk program (naturally) over the temporary file to expand ‘@include’ statements. The expanded program is placed in a second temporary file.

4. Run the expanded program with gawk and any other original command line arguments that the user supplied (such as the data file names). The initial part of the program turns on shell tracing if the first argument

was ‘debug’. Otherwise, a shell trap statement arranges to clean up any temporary files on program exit or upon an interrupt.

The next part loops through all the command line arguments. There are several cases of interest.

-- This ends the arguments to igawk. Anything else should be passed on to the user’s awk program without being evaluated.

-W This indicates that the next option is specific to gawk. To make argument processing easier, the ‘-W’ is appended to the front of the remaining arguments and the loop continues. (This is an sh programming trick. Don’t worry about it if you are not familiar with sh.)

-v

-F These are saved and passed on to gawk.

-f
--file
--file=

-Wfile= The file name is saved to the temporary file ‘/tmp/ig.s.$$’ with an ‘@include’ statement. The sed utility is used to remove the leading option part of the argument (e.g., ‘--file=’).

--source
--source=
-Wsource=

The source text is echoed into ‘/tmp/ig.s.$$’.

--version
--version
-Wversion

igawk prints its version number, and runs ‘gawk --version’ to get the gawk version information, and then exits.

If none of ‘-f’, ‘--file’, ‘-Wfile’, ‘--source’, or ‘-Wsource’, were supplied, then the first non-option argument should be the awk program. If there are no command line arguments left, igawk prints an error message and exits. Otherwise, the first argument is echoed into ‘/tmp/ig.s.$$’.

In any case, after the arguments have been processed, ‘/tmp/ig.s.$$’ contains the complete text of the original awk program.

The ‘$$’ in sh represents the current process ID number. It is often used in shell programs to generate unique temporary file names. This allows multiple users to run igawk without worrying that the temporary file names will clash.

Here’s the program:

#! /bin/sh
# igawk --- like gawk but do @include processing
# Arnold Robbins, arnold@gnu.ai.mit.edu, Public Domain
# July 1993
if [ "$1" = debug ]
then
    set -x
    shift
else
    # cleanup on exit, hangup, interrupt, quit, termination
    trap 'rm -f /tmp/ig.[se].$$' 0 1 2 3 15
fi
while [ $# -ne 0 ] # loop over arguments
do
    case $1 in
    --)     shift; break;;
    -W)     shift
            set -- -W"$@"
            continue;;
    -[vF])  opts="$opts $1 '$2'"
            shift;;
    -[vF]*) opts="$opts '$1'" ;;
    -f)     echo @include "$2" >> /tmp/ig.s.$$
            shift;;

-f*)    f=`echo "$1" | sed 's/-f//'`
        echo @include "$f" >> /tmp/ig.s.$$ ;;
-?file=*)    # -Wfile or --file
        f=`echo "$1" | sed 's/-.file=//'`
        echo @include "$f" >> /tmp/ig.s.$$ ;;
-?file)    # get arg, $2
        echo @include "$2" >> /tmp/ig.s.$$
        shift;;
-?source=*)    # -Wsource or --source
        t=`echo "$1" | sed 's/-.source=//'`
        echo "$t" >> /tmp/ig.s.$$ ;;
-?source)  # get arg, $2
        echo "$2" >> /tmp/ig.s.$$
        shift;;
-?version)
        echo igawk: version 1.0 1>&2
        gawk --version
        exit 0 ;;
-[W-]*)    opts="$opts '$1'" ;;
*)      break;;
esac
shift
done
if [ ! -s /tmp/ig.s.$$ ]
then
if [ -z "$1" ]
then
     echo igawk: no program! 1>&2
     exit 1
else
    echo "$1" > /tmp/ig.s.$$
    shift
fi
fi

    # at this point, /tmp/ig.s.$$ has the program

The awk program to process ‘@include’ directives reads through the program, one line at a time using getline (see Section 5.8 [Explicit Input with getline], page 53). The input file names and ‘@include’ statements are managed using a stack. As each ‘@include’ is encountered, the current file name is “pushed” onto the stack, and the file named in the ‘@include’ directive becomes the current file name. As each file is finished, the stack is “popped,” and the previous input file becomes the current input file again. The process is started by making the original file the first one on the stack.

The pathto function does the work of finding the full path to a file. It simulates gawk’s behavior when searching the AWKPATH environment variable (see Section 14.3 [The AWKPATH Environment Variable], page 164). If a file name has a ‘/’ in it, no path search is done. Otherwise, the file name is concatenated with the name of each directory in the path, and an attempt is made to open the generated file name. The only way in awk to test if a file can be read is to go ahead and try to read it with getline; that is what pathto does.^{5} If the file can be read, it is closed, and the file name is returned.

    gawk -- '
    # process @include directives
    function pathto(file,    i, t, junk)
    {
        if (index(file, "/") != 0)
            return file
        for (i = 1; i <= ndirs; i++) {
            t = (pathlist[i] "/" file)
            if ((getline junk < t) > 0) {
                # found it
                close(t)
                return t
            }
        }
        return ""
    }

The main program is contained inside one BEGIN rule. The first thing it does is set up the pathlist array that pathto uses. After splitting the path

5 On some very old versions of awk, the test ‘getline junk < t’ can loop forever if the file exists but is empty. Caveat Emptor. on ‘:’, null elements are replaced with ".", which represents the current directory.

    BEGIN {
        path = ENVIRON["AWKPATH"]
        ndirs = split(path, pathlist, ":")
        for (i = 1; i <= ndirs; i++) {
            if (pathlist[i] == "")
                pathlist[i] = "."
        }

The stack is initialized with ARGV[1], which will be ‘/tmp/ig.s.$$’. The main loop comes next. Input lines are read in succession. Lines that do not start with ‘@include’ are printed verbatim.

If the line does start with ‘@include’, the file name is in $2. pathto is called to generate the full path. If it could not, then we print an error message and continue.

The next thing to check is if the file has been included already. The processed array is indexed by the full file name of each included file, and it tracks this information for us. If the file has been seen, a warning message is printed. Otherwise, the new file name is pushed onto the stack and processing continues.

Finally, when getline encounters the end of the input file, the file is closed and the stack is popped. When stackptr is less than zero, the program is done.

        stackptr = 0
        input[stackptr] = ARGV[1] # ARGV[1] is first file
        for (; stackptr >= 0; stackptr--) {
            while ((getline < input[stackptr]) > 0) {
                if (tolower($1) != "@include") {
                    print
                    continue
                }
                fpath = pathto($2)
                if (fpath == "") {
                    printf("igawk:%s:%d: cannot find %s\n", \
                        input[stackptr], FNR, $2) > "/dev/stderr"
                    continue
                }

                if (! (fpath in processed)) {
                    processed[fpath] = input[stackptr]
                    input[++stackptr] = fpath
                } else
                    print $2, "included in", input[stackptr], \
                        "already included in", \
                        processed[fpath] > "/dev/stderr"
            }
            close(input[stackptr])
        }
    }' /tmp/ig.s.$$ > /tmp/ig.e.$$

The last step is to call gawk with the expanded program and the original options and command line arguments that the user supplied. gawk’s exit status is passed back on to igawk’s calling program.

    eval gawk -f /tmp/ig.e.$$ $opts -- "$@"
    exit $?

This version of igawk represents my third attempt at this program. There are three key simplifications that made the program work better.

1. Using ‘@include’ even for the files named with ‘-f’ makes building the initial collected awk program much simpler; all the ‘@include’ processing can be done once.

2. The pathto function doesn’t try to save the line read with getline when testing for the file’s accessibility. Trying to save this line for use with the main program complicates things considerably.

3. Using a getline loop in the BEGIN rule does it all in one place. It is not necessary to call out to a separate loop for processing nested ‘@include’ statements.

Also, this program illustrates that it is often worthwhile to combine sh and awk programming together. You can usually accomplish quite a lot, without having to resort to low-level programming in C or C++, and it is frequently easier to do certain kinds of string and argument manipulation using the shell than it is in awk.

Finally, igawk shows that it is not always necessary to add new features to a program; they can often be layered on top. With igawk, there is no real reason to build ‘@include’ processing into gawk itself.

As an additional example of this, consider the idea of having two files in a directory in the search path.

‘default.awk’

This file would contain a set of default library functions, such as

          getopt and assert.

‘site.awk’

This file would contain library functions that are specific to a site or installation, i.e. locally developed functions. Having a separate file allows ‘default.awk’ to change with new gawk releases, without requiring the system administrator to update it each time by adding the local functions.

One user suggested that gawk be modified to automatically read these files upon startup. Instead, it would be very simple to modify igawk to do this. Since igawk can process nested ‘@include’ directives, ‘default.awk’ could simply contain ‘@include’ statements for the desired library functions.

250 Effective AWK Programming

Chapter 17: The Evolution of the awk Language 251

17 The Evolution of the awk Language

This book describes the GNU implementation of awk, which follows the POSIX specification. Many awk users are only familiar with the original awk implementation in Version 7 Unix. (This implementation was the basis for awk in Berkeley Unix, through 4.3–Reno. The 4.4 release of Berkeley Unix uses gawk 2.15.2 for its version of awk.) This chapter briefly describes the evolution of the awk language, with cross references to other parts of the book where you can find more information.

17.1 Major Changes between V7 and SVR3.1

The awk language evolved considerably between the release of Version 7 Unix (1978) and the new version first made generally available in System V Release 3.1 (1987). This section summarizes the changes, with crossreferences to further details.

• The requirement for ‘;’ to separate rules on a line (see Section 2.6 [awk Statements Versus Lines], page 16).

• User-defined functions, and the return statement (see Chapter 13 [Userdefined Functions], page 151).

• The delete statement (see Section 11.6 [The delete Statement], page 128).

• The do-while statement (see Section 9.3 [The do-while Statement], page 106).

• The built-in functions atan2, cos, sin, rand and srand (see Section 12.2 [Numeric Built-in Functions], page 134).

• The built-in functions gsub, sub, and match (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

• The built-in functions close, and system (see Section 12.4 [Built-in Functions for Input/Output], page 144).

• The ARGC, ARGV, FNR, RLENGTH, RSTART, and SUBSEP built-in variables

(see Chapter 10 [Built-in Variables], page 115).

• The conditional expression using the ternary operator ‘?:’ (see Section 7.12 [Conditional Expressions], page 92).

• The exponentiation operator ‘^’ (see Section 7.5 [Arithmetic Operators], page 82) and its assignment operator form ‘^=’ (see Section 7.7 [Assignment Expressions], page 84).

• C-compatible operator precedence, which breaks some old awk programs (see Section 7.14 [Operator Precedence (How Operators Nest)], page 94).

• Regexps as the value of FS (see Section 5.5 [Specifying How Fields are Separated], page 44), and as the third argument to the split function (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

• Dynamic regexps as operands of the ‘~’ and ‘!~’ operators (see Section 4.1 [How to Use Regular Expressions], page 23).

• The escape sequences ‘\b’, ‘\f’, and ‘\r’ (see Section 4.2 [Escape Sequences], page 24). (Some vendors have updated their old versions of awk to recognize ‘\r’, ‘\b’, and ‘\f’, but this is not something you can rely on.)

• Redirection of input for the getline function (see Section 5.8 [Explicit Input with getline], page 53).

• Multiple BEGIN and END rules (see Section 8.1.5 [The BEGIN and END Special Patterns], page 100).

• Multi-dimensional arrays (see Section 11.9 [Multi-dimensional Arrays], page 130).

17.2 Changes between SVR3.1 and SVR4

The System V Release 4 version of Unix awk added these features (some of which originated in gawk):

• The ENVIRON variable (see Chapter 10 [Built-in Variables], page 115).

• Multiple ‘-f’ options on the command line (see Section 14.1 [Command Line Options], page 159).

• The ‘-v’ option for assigning variables before program execution begins (see Section 14.1 [Command Line Options], page 159).

• The ‘--’ option for terminating command line options.

• The ‘\a’, ‘\v’, and ‘\x’ escape sequences (see Section 4.2 [Escape Sequences], page 24).

• A defined return value for the srand built-in function (see Section 12.2 [Numeric Built-in Functions], page 134).

• The toupper and tolower built-in string functions for case translation (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

• A cleaner specification for the ‘%c’ format-control letter in the printf function (see Section 6.5.2 [Format-Control Letters], page 65).

• The ability to dynamically pass the field width and precision ("%*.*d") in the argument list of the printf function (see Section 6.5.2 [Format-Control Letters], page 65).

• The use of regexp constants such as /foo/ as expressions, where they are equivalent to using the matching operator, as in ‘$0 ~ /foo/’ (see Section 7.2 [Using Regular Expression Constants], page 78).

17.3 Changes between SVR4 and POSIX awk

The POSIX Command Language and Utilities standard for awk introduced the following changes into the language:

• The use of ‘-W’ for implementation-specific options.

• The use of CONVFMT for controlling the conversion of numbers to strings (see Section 7.4 [Conversion of Strings and Numbers], page 81).

• The concept of a numeric string, and tighter comparison rules to go with it (see Section 7.10 [Variable Typing and Comparison Expressions], page 88).

• More complete documentation of many of the previously undocumented features of the language.

The following common extensions are not permitted by the POSIX standard:

• \x escape sequences are not recognized (see Section 4.2 [Escape Sequences], page 24).

• Newlines do not act as whitespace to separate fields when FS is equal to a single space.

• The synonym func for the keyword function is not recognized (see Section 13.1 [Function Definition Syntax], page 151).

• The operators ‘**’ and ‘**=’ cannot be used in place of ‘^’ and ‘^=’ (see Section 7.5 [Arithmetic Operators], page 82, and also see Section 7.7 [Assignment Expressions], page 84).

• Specifying ‘-Ft’ on the command line does not set the value of FS to be a single tab character (see Section 5.5 [Specifying How Fields are Separated], page 44).

• The fflush built-in function is not supported (see Section 12.4 [Built-in Functions for Input/Output], page 144).

17.4 Extensions in the Bell Laboratories awk

Brian Kernighan, one of the original designers of Unix awk, has made his version available via anonymous ftp (see Section B.8 [Other Freely Available awk Implementations], page 291). This section describes extensions in his version of awk that are not in POSIX awk.

• The ‘-mf NNN’ and ‘-mr NNN’ command line options to set the maximum number of fields, and the maximum record size, respectively (see Section 14.1 [Command Line Options], page 159).

• The fflush built-in function for flushing buffered output (see Section 12.4 [Built-in Functions for Input/Output], page 144).

17.5 Extensions in gawk Not in POSIX awk

The GNU implementation, gawk, adds a number of features. This sections lists them in the order they were added to gawk. They can all be disabled with either the ‘--traditional’ or ‘--posix’ options (see Section 14.1 [Command Line Options], page 159).

Version 2.10 of gawk introduced these features:

• The AWKPATH environment variable for specifying a path search for the ‘-f’ command line option (see Section 14.1 [Command Line Options], page 159).

• The IGNORECASE variable and its effects (see Section 4.5 [Case-sensitivity in Matching], page 33).

• The ‘/dev/stdin’, ‘/dev/stdout’, ‘/dev/stderr’, and ‘/dev/fd/n’ file

name interpretation (see Section 6.7 [Special File Names in gawk], page 72). Version 2.13 of gawk introduced these features:

• The FIELDWIDTHS variable and its effects (see Section 5.6 [Reading Fixed-width Data], page 49).

• The systime and strftime built-in functions for obtaining and printing time stamps (see Section 12.5 [Functions for Dealing with Time Stamps], page 146).

• The ‘-W lint’ option to provide source code and run time error and portability checking (see Section 14.1 [Command Line Options], page 159).

• The ‘-W compat’ option to turn off these extensions (see Section 14.1 [Command Line Options], page 159).

• The ‘-W posix’ option for full POSIX compliance (see Section 14.1 [Command Line Options], page 159). Version 2.14 of gawk introduced these features:

• The next file statement for skipping to the next data file (see Section 9.8 [The nextfile Statement], page 112). Version 2.15 of gawk introduced these features:

• The ARGIND variable, that tracks the movement of FILENAME through ARGV (see Chapter 10 [Built-in Variables], page 115).

• The ERRNO variable, that contains the system error message when getline returns −1, or when close fails (see Chapter 10 [Built-in Variables], page 115).

• The ability to use GNU-style long named options that start with ‘--’ (see Section 14.1 [Command Line Options], page 159).

• The ‘--source’ option for mixing command line and library file source code (see Section 14.1 [Command Line Options], page 159).

• The ‘/dev/pid’, ‘/dev/ppid’, ‘/dev/pgrpid’, and ‘/dev/user’ file

name interpretation (see Section 6.7 [Special File Names in gawk], page 72). Version 3.0 of gawk introduced these features:

• The next file statement became nextfile (see Section 9.8 [The nextfile Statement], page 112).

• The ‘--lint-old’ option to warn about constructs that are not available in the original Version 7 Unix version of awk (see Section 17.1 [Major Changes between V7 and SVR3.1], page 251).

• The ‘--traditional’ option was added as a better name for ‘--compat’ (see Section 14.1 [Command Line Options], page 159).

• The ability for FS to be a null string, and for the third argument to split to be the null string (see Section 5.5.3 [Making Each Character a Separate Field], page 46).

• The ability for RS to be a regexp (see Section 5.1 [How Input is Split into Records], page 37).

• The RT variable (see Section 5.1 [How Input is Split into Records], page 37).

• The gensub function for more powerful text manipulation (see Section 12.3 [Built-in Functions for String Manipulation], page 135).

• The strftime function acquired a default time format, allowing it to be called with no arguments (see Section 12.5 [Functions for Dealing with Time Stamps], page 146).

• Full support for both POSIX and GNU regexps (see Chapter 4 [Regular Expressions], page 23).

• The ‘--re-interval’ option to provide interval expressions in regexps (see Section 4.3 [Regular Expression Operators], page 26).

• IGNORECASE changed, now applying to string comparison as well as regexp operations (see Section 4.5 [Case-sensitivity in Matching], page 33).

• The ‘-m’ option and the fflush function from the Bell Labs research version of awk (see Section 14.1 [Command Line Options], page 159; also see Section 12.4 [Built-in Functions for Input/Output], page 144).

• The use of GNU Autoconf to control the configuration process (see Section B.2.1 [Compiling gawk for Unix], page 282).

• Amiga support (see Section B.6 [Installing gawk on an Amiga], page 289).

256 Effective AWK Programming

Appendix A: gawk Summary 257

Appendix A gawk Summary

This appendix provides a brief summary of the gawk command line and the awk language. It is designed to serve as “quick reference.” It is therefore terse, but complete.

A.1 Command Line Options Summary

The command line consists of options to gawk itself, the awk program text (if not supplied via the ‘-f’ option), and values to be made available in the ARGC and ARGV predefined awk variables:

gawk [POSIX or GNU style options] -f source-file [--] file … gawk [POSIX or GNU style options] [--] 'program' file … The options that gawk accepts are:

-F fs
--field-separator fs

Use fs for the input field separator (the value of the FS predefined variable).

-f program-file --file program-file

Read the awk program source from the file program-file, instead of from the first command line argument.

-mf NNN

-mr NNN The ‘f’ flag sets the maximum number of fields, and the ‘r’ flag sets the maximum record size. These options are ignored by gawk, since gawk has no predefined limits; they are only for compatibility with the Bell Labs research version of Unix awk.

-v var=val

--assign var=val

Assign the variable var the value val before program execution begins.

-W traditional
-W compat
--traditional

--compat Use compatibility mode, in which gawk extensions are turned off.

-W copyleft
-W copyright
--copyleft
--copyright

Print the short version of the General Public License on the standard output, and exit. This option may disappear in a future version of gawk.

-W help
-W usage
--help

--usage Print a relatively short summary of the available options on the standard output, and exit.

-W lint

--lint Give warnings about dubious or non-portable awk constructs.

-W lint-old
--lint-old

Warn about constructs that are not available in the original Version 7 Unix version of awk.

-W posix

--posix Use POSIX compatibility mode, in which gawk extensions are turned off and additional restrictions apply.

-W re-interval
--re-interval

Allow interval expressions (see Section 4.3 [Regular Expression Operators], page 26), in regexps.

-W source=program-text --source program-text

Use program-text as awk program source code. This option allows mixing command line source code with source code from files, and is particularly useful for mixing command line programs with library functions.

-W version
--version

Print version information for this particular copy of gawk on the error output.

-- Signal the end of options. This is useful to allow further arguments to the awk program itself to start with a ‘-’. This is mainly for consistency with POSIX argument parsing conventions.

Any other options are flagged as invalid, but are otherwise ignored. See Section 14.1 [Command Line Options], page 159, for more details.

A.2 Language Summary

An awk program consists of a sequence of zero or more pattern-action statements and optional function definitions. One or the other of the pattern and action may be omitted.

pattern { action statements }

pattern { action statements }

function name(parameter list) { action statements } gawk first reads the program source from the program-file(s), if specified,

or from the first non-option argument on the command line. The ‘-f’ option may be used multiple times on the command line. gawk reads the program text from all the program-file files, effectively concatenating them in the order they are specified. This is useful for building libraries of awk functions, without having to include them in each new awk program that uses them. To use a library function in a file from a program typed in on the command line, specify ‘--source 'program'’, and type your program in between the single quotes. See Section 14.1 [Command Line Options], page 159.

The environment variable AWKPATH specifies a search path to use when finding source files named with the ‘-f’ option. The default path, which is ‘.:/usr/local/share/awk’^{1} is used if AWKPATH is not set. If a file name given to the ‘-f’ option contains a ‘/’ character, no path search is performed. See Section 14.3 [The AWKPATH Environment Variable], page 164.

gawk compiles the program into an internal form, and then proceeds to read each file named in the ARGV array. The initial values of ARGV come from the command line arguments. If there are no files named on the command line, gawk reads the standard input.

If a “file” named on the command line has the form ‘var=val’, it is treated as a variable assignment: the variable var is assigned the value val. If any of the files have a value that is the null string, that element in the list is skipped.

For each record in the input, gawk tests to see if it matches any pattern in the awk program. For each pattern that the record matches, the associated action is executed.

A.3 Variables and Fields

awk variables are not declared; they come into existence when they are first used. Their values are either floating-point numbers or strings. awk also

1 The path may use a directory other than ‘/usr/local/share/awk’, depending upon how gawk was built and installed. has one-dimensional arrays; multiple-dimensional arrays may be simulated. There are several predefined variables that awk sets as a program runs; these are summarized below.

A.3.1 Fields

As each input line is read, gawk splits the line into fields, using the value of the FS variable as the field separator. If FS is a single character, fields are separated by that character. Otherwise, FS is expected to be a full regular expression. In the special case that FS is a single space, fields are separated by runs of spaces, tabs and/or newlines.^{2} If FS is the null string (""), then each individual character in the record becomes a separate field. Note that the value of IGNORECASE (see Section 4.5 [Case-sensitivity in Matching], page 33) also affects how fields are split when FS is a regular expression.

Each field in the input line may be referenced by its position, $1, $2, and so on. $0 is the whole line. The value of a field may be assigned to as well. Field numbers need not be constants:

    n = 5
    print $n

prints the fifth field in the input line. The variable NF is set to the total number of fields in the input line.

References to non-existent fields (i.e. fields after $NF) return the null string. However, assigning to a non-existent field (e.g., $(NF+2) = 5) increases the value of NF, creates any intervening fields with the null string as their value, and causes the value of $0 to be recomputed, with the fields being separated by the value of OFS. Decrementing NF causes the values of fields past the new value to be lost, and the value of $0 to be recomputed, with the fields being separated by the value of OFS. See Chapter 5 [Reading Input Files], page 37.

A.3.2 Built-in Variables

gawk’s built-in variables are:

ARGC The number of elements in ARGV. See below for what is actually included in ARGV.

ARGIND The index in ARGV of the current file being processed. When gawk is processing the input data files, it is always true that

          ‘FILENAME == ARGV[ARGIND]’.

ARGV The array of command line arguments. The array is indexed from zero to ARGC − 1. Dynamically changing ARGC and the

2 In POSIX awk, newline does not separate fields. contents of ARGV can control the files used for data. A nullvalued element in ARGV is ignored. ARGV does not include the options to awk or the text of the awk program itself.

CONVFMT The conversion format to use when converting numbers to strings.

FIELDWIDTHS

A space separated list of numbers describing the fixed-width input data.

ENVIRON An array of environment variable values. The array is indexed by variable name, each element being the value of that variable. Thus, the environment variable HOME is ENVIRON["HOME"]. One possible value might be ‘/home/arnold’. Changing this array does not affect the environment seen by programs which gawk spawns via redirection or the system function. (This may change in a future version of gawk.) Some operating systems do not have environment variables. The ENVIRON array is empty when running on these systems.

ERRNO The system error message when an error occurs using getline

or close.

FILENAME The name of the current input file. If no files are specified on the command line, the value of FILENAME is the null string.

FNR The input record number in the current input file.

FS The input field separator, a space by default.

IGNORECASE

The case-sensitivity flag for string comparisons and regular expression operations. If IGNORECASE has a non-zero value, then pattern matching in rules, record separating with RS, field splitting with FS, regular expression matching with ‘~’ and ‘!~’, and

the gensub, gsub, index, match, split and sub built-in func-

tions all ignore case when doing regular expression operations, and all string comparisons are done ignoring case. The value of IGNORECASE does not affect array subscripting.

NF The number of fields in the current input record.

NR The total number of input records seen so far.

OFMT The output format for numbers for the print statement, "%.6g" by default.

OFS The output field separator, a space by default.

ORS The output record separator, by default a newline.

RS The input record separator, by default a newline. If RS is set to the null string, then records are separated by blank lines. When RS is set to the null string, then the newline character always acts as a field separator, in addition to whatever value FS may have. If RS is set to a multi-character string, it denotes a regexp; input text matching the regexp separates records.

RT The input text that matched the text denoted by RS, the record separator.

RSTART The index of the first character last matched by match; zero if no match.

RLENGTH The length of the string last matched by match; −1 if no match.

SUBSEP The string used to separate multiple subscripts in array elements, by default "\034".

See Chapter 10 [Built-in Variables], page 115, for more information.

A.3.3 Arrays

Arrays are subscripted with an expression between square brackets (‘[’ and ‘]’). Array subscripts are always strings; numbers are converted to strings as necessary, following the standard conversion rules (see Section 7.4 [Conversion of Strings and Numbers], page 81).

If you use multiple expressions separated by commas inside the square brackets, then the array subscript is a string consisting of the concatenation of the individual subscript values, converted to strings, separated by the subscript separator (the value of SUBSEP).

The special operator in may be used in a conditional context to see if an array has an index consisting of a particular value.

    if (val in array)
            print array[val]

If the array has multiple subscripts, use ‘(i, j, …) in array’ to test for existence of an element.

The in construct may also be used in a for loop to iterate over all the elements of an array. See Section 11.5 [Scanning All Elements of an Array], page 127.

You can remove an element from an array using the delete statement. You can clear an entire array using ‘delete array’. See Chapter 11 [Arrays in awk], page 123.

A.3.4 Data Types

The value of an awk expression is always either a number or a string.

Some contexts (such as arithmetic operators) require numeric values. They convert strings to numbers by interpreting the text of the string as a number. If the string does not look like a number, it converts to zero.

Other contexts (such as concatenation) require string values. They convert numbers to strings by effectively printing them with sprintf. See Section 7.4 [Conversion of Strings and Numbers], page 81, for the details.

To force conversion of a string value to a number, simply add zero to it. If the value you start with is already a number, this does not change it.

To force conversion of a numeric value to a string, concatenate it with the null string.

Comparisons are done numerically if both operands are numeric, or if one is numeric and the other is a numeric string. Otherwise one or both operands are converted to strings and a string comparison is performed. Fields, getline input, FILENAME, ARGV elements, ENVIRON elements and the elements of an array created by split are the only items that can be numeric strings. String constants, such as "3.1415927" are not numeric strings, they are string constants. The full rules for comparisons are described in Section 7.10 [Variable Typing and Comparison Expressions], page 88.

Uninitialized variables have the string value "" (the null, or empty, string). In contexts where a number is required, this is equivalent to zero.

See Section 7.3 [Variables], page 79, for more information on variable naming and initialization; see Section 7.4 [Conversion of Strings and Numbers], page 81, for more information on how variable values are interpreted.

A.4 Patterns

An awk program is mostly composed of rules, each consisting of a pattern followed by an action. The action is enclosed in ‘{’ and ‘}’. Either the pattern may be missing, or the action may be missing, but not both. If the pattern is missing, the action is executed for every input record. A missing action is equivalent to ‘{ print }’, which prints the entire line.

Comments begin with the ‘#’ character, and continue until the end of the line. Blank lines may be used to separate statements. Statements normally end with a newline; however, this is not the case for lines ending in a ‘,’, ‘{’, ‘?’, ‘:’, ‘&&’, or ‘||’. Lines ending in do or else also have their statements automatically continued on the following line. In other cases, a line can be continued by ending it with a ‘\’, in which case the newline is ignored.

Multiple statements may be put on one line by separating each one with a ‘;’. This applies to both the statements within the action part of a rule (the usual case), and to the rule statements.

See Section 2.2.5 [Comments in awk Programs], page 13, for information on awk’s commenting convention; see Section 2.6 [awk Statements Versus Lines], page 16, for a description of the line continuation mechanism in awk.

A.4.1 Pattern Summary

awk patterns may be one of the following:

/regular expression/ relational expression pattern && pattern pattern || pattern pattern ? pattern : pattern (pattern) ! pattern pattern1, pattern2

    BEGIN
    END

BEGIN and END are two special kinds of patterns that are not tested against the input. The action parts of all BEGIN rules are concatenated as if all the statements had been written in a single BEGIN rule. They are executed before any of the input is read. Similarly, all the END rules are concatenated, and executed when all the input is exhausted (or when an exit statement is executed). BEGIN and END patterns cannot be combined with other patterns in pattern expressions. BEGIN and END rules cannot have missing action parts.

For /regular-expression/ patterns, the associated statement is executed for each input record that matches the regular expression. Regular expressions are summarized below.

A relational expression may use any of the operators defined below in the section on actions. These generally test whether certain fields match certain regular expressions.

The ‘&&’, ‘||’, and ‘!’ operators are logical “and,” logical “or,” and logical “not,” respectively, as in C. They do short-circuit evaluation, also as in C, and are used for combining more primitive pattern expressions. As in most languages, parentheses may be used to change the order of evaluation.

The ‘?:’ operator is like the same operator in C. If the first pattern matches, then the second pattern is matched against the input record; otherwise, the third is matched. Only one of the second and third patterns is matched.

The ‘pattern1, pattern2’ form of a pattern is called a range pattern. It matches all input lines starting with a line that matches pattern1, and continuing until a line that matches pattern2, inclusive. A range pattern cannot be used as an operand of any of the pattern operators.

See Section 8.1 [Pattern Elements], page 97.

A.4.2 Regular Expressions

Regular expressions are based on POSIX EREs (extended regular expressions). The escape sequences allowed in string constants are also valid in regular expressions (see Section 4.2 [Escape Sequences], page 24). Regexps are composed of characters as follows:

c matches the character c (assuming c is none of the characters listed below).

\c matches the literal character c.

. matches any character, including newline. In strict POSIX mode, ‘.’ does not match the nul character, which is a character with all bits equal to zero.

^matches the beginning of a string.

$ matches the end of a string.

[abc…] matches any of the characters abc… (character list).

[[:class:]]

matches any character in the character class class. Allowable

          classes are alnum, alpha, blank, cntrl, digit, graph, lower,
          print, punct, space, upper, and xdigit.
[[.symbol.]]

matches the multi-character collating symbol symbol. gawk does not currently support collating symbols.

[[=classname=]] matches any of the equivalent characters in the current locale named by the equivalence class classname. gawk does not currently support equivalence classes.

[^abc…] matches any character except abc… (negated character list).

r1|r2 matches either r1 or r2 (alternation).

r1r2 matches r1, and then r2 (concatenation).

r+ matches one or more r’s.

r* matches zero or more r’s.

r? matches zero or one r’s.

(r) matches r (grouping).

r{n}
r{n,}

r{n,m} matches at least n, n to any number, or n to m occurrences of r (interval expressions).

\y matches the empty string at either the beginning or the end of a word.

\B matches the empty string within a word.

\< matches the empty string at the beginning of a word.

\> matches the empty string at the end of a word.

\w matches any word-constituent character (alphanumeric characters and the underscore).

\W matches any character that is not word-constituent.

\` matches the empty string at the beginning of a buffer (same as a string in gawk).

\' matches the empty string at the end of a buffer.

The various command line options control how gawk interprets characters in regexps.

No options In the default case, gawk provide all the facilities of POSIX regexps and the GNU regexp operators described above. However, interval expressions are not supported.

--posix Only POSIX regexps are supported, the GNU operators are not special (e.g., ‘\w’ matches a literal ‘w’). Interval expressions are allowed.

--traditional

Traditional Unix awk regexps are matched. The GNU operators are not special, interval expressions are not available, and neither are the POSIX character classes ([[:alnum:]] and so on). Characters described by octal and hexadecimal escape sequences are treated literally, even if they represent regexp metacharacters.

--re-interval

Allow interval expressions in regexps, even if ‘--traditional’ has been provided.

See Chapter 4 [Regular Expressions], page 23.

A.5 Actions

Action statements are enclosed in braces, ‘{’ and ‘}’. A missing action statement is equivalent to ‘{ print }’.

Action statements consist of the usual assignment, conditional, and looping statements found in most languages. The operators, control statements, and Input/Output statements available are similar to those in C.

Comments begin with the ‘#’ character, and continue until the end of the line. Blank lines may be used to separate statements. Statements normally end with a newline; however, this is not the case for lines ending in a ‘,’, ‘{’, ‘?’, ‘:’, ‘&&’, or ‘||’. Lines ending in do or else also have their statements automatically continued on the following line. In other cases, a line can be continued by ending it with a ‘\’, in which case the newline is ignored.

Multiple statements may be put on one line by separating each one with a ‘;’. This applies to both the statements within the action part of a rule (the usual case), and to the rule statements.

See Section 2.2.5 [Comments in awk Programs], page 13, for information on awk’s commenting convention; see Section 2.6 [awk Statements Versus Lines], page 16, for a description of the line continuation mechanism in awk.

A.5.1 Operators

The operators in awk, in order of decreasing precedence, are: (…) Grouping. $ Field reference. ++ -- Increment and decrement, both prefix and postfix. ^Exponentiation (‘**’ may also be used, and ‘**=’ for the assign-

ment operator, but they are not specified in the POSIX standard).

+ - ! Unary plus, unary minus, and logical negation.

* / % Multiplication, division, and modulus. + - Addition and subtraction. space String concatenation.

< <= > >= != ==

The usual relational operators.

~ !~ Regular expression match, negated match. in Array membership. && Logical “and”. || Logical “or”. ?: A conditional expression. This has the form ‘expr1 ? expr2 :

expr3’. If expr1 is true, the value of the expression is expr2; otherwise it is expr3. Only one of expr2 and expr3 is evaluated.

= += -= *= /= %= ^=

Assignment. Both absolute assignment (var=value) and operator assignment (the other forms) are supported.

See Chapter 7 [Expressions], page 77.

A.5.2 Control Statements

The control statements are as follows: if (condition) statement [ else statement ] while (condition) statement do statement while (condition) for (expr1; expr2; expr3) statement for (var in array) statement

break
continue

delete array[index]

delete array

exit [ expression ] { statements } See Chapter 9 [Control Statements in Actions], page 105.

A.5.3 I/O Statements

The Input/Output statements are as follows:

getline Set $0 from next input record; set NF, NR, FNR. See Section 5.8 [Explicit Input with getline], page 53.

getline <file

Set $0 from next record of file; set NF.

getline var

Set var from next input record; set NR, FNR.

getline var <file

Set var from next record of file.

command | getline

Run command, piping its output into getline; sets $0, NF, NR.

command | getline var

Run command, piping its output into getline; sets var.

next Stop processing the current input record. The next input record is read and processing starts over with the first pattern in the awk program. If the end of the input data is reached, the END rule(s), if any, are executed. See Section 9.7 [The next Statement], page 111.

nextfile Stop processing the current input file. The next input record read comes from the next input file. FILENAME is updated, FNR is set to one, ARGIND is incremented, and processing starts over with the first pattern in the awk program. If the end of the input data is reached, the END rule(s), if any, are executed. Earlier versions of gawk used ‘next file’; this usage is still supported, but is considered to be deprecated. See Section 9.8 [The nextfile Statement], page 112.

print Prints the current record. See Chapter 6 [Printing Output], page 61.

print expr-list Prints expressions.

print expr-list > file Prints expressions to file. If file does not exist, it is created. If it does exist, its contents are deleted the first time the print is executed.

print expr-list >> file Prints expressions to file. The previous contents of file are retained, and the output of print is appended to the file.

print expr-list | command Prints expressions, sending the output down a pipe to command. The pipeline to the command stays open until the close function is called.

printf fmt, expr-list Format and print.

printf fmt, expr-list > file Format and print to file. If file does not exist, it is created. If it does exist, its contents are deleted the first time the printf is executed.

printf fmt, expr-list >> file Format and print to file. The previous contents of file are retained, and the output of printf is appended to the file.

printf fmt, expr-list | command Format and print, sending the output down a pipe to command. The pipeline to the command stays open until the close function is called.

getline returns zero on end of file, and −1 on an error. In the event of an error, getline will set ERRNO to the value of a system-dependent string that describes the error.

A.5.4 printf Summary

Conversion specification have the form %[flag][width][.prec]format. Items in brackets are optional.

The awk printf statement and sprintf function accept the following conversion specification formats:

%c An ASCII character. If the argument used for ‘%c’ is numeric, it is treated as a character and printed. Otherwise, the argument is assumed to be a string, and the only first character of that string is printed.

%d

%i A decimal number (the integer part).

%e

%E A floating point number of the form ‘[-]d.dddddde[+-]dd’. The ‘%E’ format uses ‘E’ instead of ‘e’.

%f A floating point number of the form [-]ddd.dddddd.

%g

%G Use either the ‘%e’ or ‘%f’ formats, whichever produces a shorter string, with non-significant zeros suppressed. ‘%G’ will use ‘%E’ instead of ‘%e’.

%o An unsigned octal number (again, an integer).

%s A character string.

%x

%X An unsigned hexadecimal number (an integer). The ‘%X’ format uses ‘A’ through ‘F’ instead of ‘a’ through ‘f’ for decimal 10 through 15.

%% A single ‘%’ character; no argument is converted.

There are optional, additional parameters that may lie between the ‘%’ and the control letter:

- The expression should be left-justified within its field.

space For numeric conversions, prefix positive values with a space, and negative values with a minus sign.

+ The plus sign, used before the width modifier (see below), says to always supply a sign for numeric conversions, even if the data to be formatted is positive. The ‘+’ overrides the space modifier.

# Use an “alternate form” for certain control letters. For ‘o’, supply a leading zero. For ‘x’, and ‘X’, supply a leading ‘0x’ or ‘0X’ for a non-zero result. For ‘e’, ‘E’, and ‘f’, the result will always contain a decimal point. For ‘g’, and ‘G’, trailing zeros are not removed from the result.

0 A leading ‘0’ (zero) acts as a flag, that indicates output should be padded with zeros instead of spaces. This applies even to non-numeric output formats. This flag only has an effect when the field width is wider than the value to be printed.

width The field should be padded to this width. The field is normally padded with spaces. If the ‘0’ flag has been used, it is padded with zeros.

.prec A number that specifies the precision to use when printing. For the ‘e’, ‘E’, and ‘f’ formats, this specifies the number of digits you want printed to the right of the decimal point. For the ‘g’, and ‘G’ formats, it specifies the maximum number of significant digits. For the ‘d’, ‘o’, ‘i’, ‘u’, ‘x’, and ‘X’ formats, it specifies the minimum number of digits to print. For the ‘s’ format, it specifies the maximum number of characters from the string that should be printed.

Either or both of the width and prec values may be specified as ‘*’. In that case, the particular value is taken from the argument list.

See Section 6.5 [Using printf Statements for Fancier Printing], page 64.

A.5.5 Special File Names

When doing I/O redirection from either print or printf into a file, or via getline from a file, gawk recognizes certain special file names internally. These file names allow access to open file descriptors inherited from gawk’s parent process (usually the shell). The file names are:

‘/dev/stdin’

The standard input.

‘/dev/stdout’

The standard output.

‘/dev/stderr’

The standard error output.

‘/dev/fd/n’

The file denoted by the open file descriptor n. In addition, reading the following files provides process related informa-

tion about the running gawk program. All returned records are terminated with a newline.

‘/dev/pid’

Returns the process ID of the current process.

‘/dev/ppid’

Returns the parent process ID of the current process.

‘/dev/pgrpid’

Returns the process group ID of the current process.

‘/dev/user’

At least four space-separated fields, containing the return values

          of the getuid, geteuid, getgid, and getegid system calls. If

there are any additional fields, they are the group IDs returned by getgroups system call. (Multiple groups may not be supported on all systems.)

These file names may also be used on the command line to name data files. These file names are only recognized internally if you do not actually have files with these names on your system.

See Section 6.7 [Special File Names in gawk], page 72, for a longer description that provides the motivation for this feature.

A.5.6 Built-in Functions

awk provides a number of built-in functions for performing numeric operations, string related operations, and I/O related operations.

The built-in arithmetic functions are:

atan2(y, x)

the arctangent of y/x in radians.

cos(expr) the cosine of expr, which is in radians. exp(expr) the exponential function (e ^expr). int(expr) truncates to integer. log(expr) the natural logarithm of expr. rand() a random number between zero and one. sin(expr) the sine of expr, which is in radians.

sqrt(expr)

the square root function.

srand([expr])

use expr as a new seed for the random number generator. If no expr is provided, the time of day is used. The return value is the previous seed for the random number generator.

awk has the following built-in string functions: gensub(regex, subst, how [, target])

If how is a string beginning with ‘g’ or ‘G’, then replace each match of regex in target with subst. Otherwise, replace the how’th occurrence. If target is not supplied, use $0. The return value is the changed string; the original target is not modified. Within subst, ‘\n’, where n is a digit from one to nine, can be used to indicate the text that matched the n’th parenthesized subexpression. This function is gawk-specific. gsub(regex, subst [, target])

for each substring matching the regular expression regex in the string target, substitute the string subst, and return the number of substitutions. If target is not supplied, use $0.

index(str, search)

returns the index of the string search in the string str, or zero if search is not present.

length([str])

returns the length of the string str. The length of $0 is returned if no argument is supplied.

match(str, regex)

returns the position in str where the regular expression regex occurs, or zero if regex is not present, and sets the values of

RSTART and RLENGTH.

split(str, arr [, regex]) splits the string str into the array arr on the regular expression regex, and returns the number of elements. If regex is omitted, FS is used instead. regex can be the null string, causing each character to be placed into its own array element. The array arr is cleared first.

sprintf(fmt, expr-list)

prints expr-list according to fmt, and returns the resulting string.

sub(regex, subst [, target]) just like gsub, but only the first matching substring is replaced.

substr(str, index [, len]) returns the len-character substring of str starting at index. If len is omitted, the rest of str is used.

tolower(str)

returns a copy of the string str, with all the upper-case characters in str translated to their corresponding lower-case counterparts. Non-alphabetic characters are left unchanged.

toupper(str)

returns a copy of the string str, with all the lower-case characters in str translated to their corresponding upper-case counterparts. Non-alphabetic characters are left unchanged.

The I/O related functions are:

close(expr)

Close the open file or pipe denoted by expr.

fflush([expr])

Flush any buffered output for the output file or pipe denoted by expr. If expr is omitted, standard output is flushed. If expr is the null string (""), all output buffers are flushed.

system(cmd-line)

Execute the command cmd-line, and return the exit status. If your operating system does not support system, calling it will generate a fatal error. ‘system("")’ can be used to force awk to flush any pending output. This is more portable, but less obvious, than calling

fflush.

A.5.7 Time Functions

The following two functions are available for getting the current time of day, and for formatting time stamps. They are specific to gawk.

systime()

returns the current time of day as the number of seconds since a particular epoch (Midnight, January 1, 1970 UTC, on POSIX systems).

strftime([format[, timestamp]])

formats timestamp according to the specification in format. The current time of day is used if no timestamp is supplied. A default format equivalent to the output of the date utility is used if no format is supplied. See Section 12.5 [Functions for Dealing with Time Stamps], page 146, for the details on the conversion specifiers that strftime accepts.

See Chapter 12 [Built-in Functions], page 133, for a description of all of awk’s built-in functions.

A.5.8 String Constants

String constants in awk are sequences of characters enclosed in double quotes ("). Within strings, certain escape sequences are recognized, as in C. These are:

\\ A literal backslash.

\a The “alert” character; usually the ASCII BEL character.

\b Backspace.

\f Formfeed.

\n Newline.

\r Carriage return.

\t Horizontal tab.

\v Vertical tab.

\xhex digits The character represented by the string of hexadecimal digits following the ‘\x’. As in ANSI C, all following hexadecimal digits are considered part of the escape sequence. E.g., "\x1B" is a string containing the ASCII ESC (escape) character. (The ‘\x’ escape sequence is not in POSIX awk.)

\ddd The character represented by the one, two, or three digit sequence of octal digits. Thus, "\033" is also a string containing the ASCII ESC (escape) character.

\c The literal character c, if c is not one of the above.

The escape sequences may also be used inside constant regular expressions (e.g., the regexp /[ \t\f\n\r\v]/ matches whitespace characters).

See Section 4.2 [Escape Sequences], page 24.

A.6 User-defined Functions

Functions in awk are defined as follows:

function name(parameter list) { statements } Actual parameters supplied in the function call are used to instantiate the

formal parameters declared in the function. Arrays are passed by reference, other variables are passed by value.

If there are fewer arguments passed than there are names in parameterlist, the extra names are given the null string as their value. Extra names have the effect of local variables.

The open-parenthesis in a function call of a user-defined function must immediately follow the function name, without any intervening white space. This is to avoid a syntactic ambiguity with the concatenation operator.

The word func may be used in place of function (but not in POSIX

awk).

Use the return statement to return a value from a function.

See Chapter 13 [User-defined Functions], page 151.

A.7 Historical Features

There are two features of historical awk implementations that gawk supports.

First, it is possible to call the length built-in function not only with no arguments, but even without parentheses!

    a = length

is the same as either of

    a = length()
    a = length($0)

For example:

    $ echo abcdef | awk '{ print length }'
    ⊣ 6

This feature is marked as “deprecated” in the POSIX standard, and gawk will issue a warning about its use if ‘--lint’ is specified on the command line. (The ability to use length this way was actually an accident of the original Unix awk implementation. If any built-in function used $0 as its default argument, it was possible to call that function without the parentheses. In particular, it was common practice to use the length function in this fashion, and this usage was documented in the awk manual page.)

The other historical feature is the use of either the break statement, or the continue statement outside the body of a while, for, or do loop. Traditional awk implementations have treated such usage as equivalent to the next statement. More recent versions of Unix awk do not allow it. gawk supports this usage if ‘--traditional’ has been specified.

See Section 14.1 [Command Line Options], page 159, for more information about the ‘--posix’ and ‘--lint’ options.

Appendix B: Installing gawk 277

Appendix B Installing gawk

This appendix provides instructions for installing gawk on the various platforms that are supported by the developers. The primary developers support Unix (and one day, GNU), while the other ports were contributed. The file ‘ACKNOWLEDGMENT’ in the gawk distribution lists the electronic mail addresses of the people who did the respective ports, and they are also provided in Section B.7 [Reporting Problems and Bugs], page 290.

B.1 The gawk Distribution

This section first describes how to get the gawk distribution, how to extract it, and then what is in the various files and subdirectories.

B.1.1 Getting the gawk Distribution

There are three ways you can get GNU software.

1. You can copy it from someone else who already has it.

2. You can order gawk directly from the Free Software Foundation. Software distributions are available for Unix, MS-DOS, and VMS, on tape and CD-ROM. The address is:

Free Software Foundation 59 Temple Place—Suite 330 Boston, MA 02111-1307 USA Phone: +1-617-542-5942 Fax (including Japan): +1-617-542-2652

    E-mail: gnu@prep.ai.mit.edu

Ordering from the FSF directly contributes to the support of the foundation and to the production of more free software.

3. You can get gawk by using anonymous ftp to the Internet host

ftp.gnu.ai.mit.edu, in the directory ‘/pub/gnu’.

Here is a list of alternate ftp sites from which you can obtain GNU software. When a site is listed as “site:directory” the directory indicates the directory where GNU software is kept. You should use a site that is geographically close to you.

Asia:

          cair-archive.kaist.ac.kr:/pub/gnu
          ftp.cs.titech.ac.jp
          ftp.nectec.or.th:/pub/mirrors/gnu
          utsun.s.u-tokyo.ac.jp:/ftpsync/prep

Australia:

archie.au:/gnu
          (archie.oz or archie.oz.au for ACSnet)

Africa:

ftp.sun.ac.za:/pub/gnu

Middle East:

ftp.technion.ac.il:/pub/unsupported/gnu

Europe:

archive.eu.net
ftp.denet.dk
ftp.eunet.ch
ftp.funet.fi:/pub/gnu
ftp.ieunet.ie:pub/gnu
ftp.informatik.rwth-aachen.de:/pub/gnu
ftp.informatik.tu-muenchen.de
ftp.luth.se:/pub/unix/gnu
ftp.mcc.ac.uk
ftp.stacken.kth.se
ftp.sunet.se:/pub/gnu
ftp.univ-lyon1.fr:pub/gnu
ftp.win.tue.nl:/pub/gnu
irisa.irisa.fr:/pub/gnu
isy.liu.se
nic.switch.ch:/mirror/gnu
src.doc.ic.ac.uk:/gnu
unix.hensa.ac.uk:/pub/uunet/systems/gnu

South America:

ftp.inf.utfsm.cl:/pub/gnu
ftp.unicamp.br:/pub/gnu

Western Canada:

ftp.cs.ubc.ca:/mirror2/gnu

USA:

col.hp.com:/mirrors/gnu
f.ms.uky.edu:/pub3/gnu
ftp.cc.gatech.edu:/pub/gnu
ftp.cs.columbia.edu:/archives/gnu/prep
ftp.digex.net:/pub/gnu
ftp.hawaii.edu:/mirrors/gnu
ftp.kpc.com:/pub/mirror/gnu

USA (continued):

              ftp.uu.net:/systems/gnu
              gatekeeper.dec.com:/pub/GNU
              jaguar.utah.edu:/gnustuff
              labrea.stanford.edu
              mrcnext.cso.uiuc.edu:/pub/gnu
              vixen.cso.uiuc.edu:/gnu
              wuarchive.wustl.edu:/systems/gnu

B.1.2 Extracting the Distribution

gawk is distributed as a tar file compressed with the GNU Zip program,

gzip.

Once you have the distribution (for example, ‘gawk-3.0.3.tar.gz’), first use gzip to expand the file, and then use tar to extract it. You can use the following pipeline to produce the gawk distribution:

    # Under System V, add 'o' to the tar flags
    gzip -d -c gawk-3.0.3.tar.gz | tar -xvpf -

This will create a directory named ‘gawk-3.0.3’ in the current directory.

The distribution file name is of the form ‘gawk-V.R.n.tar.gz’. The V represents the major version of gawk, the R represents the current release of version V, and the n represents a patch level, meaning that minor bugs have been fixed in the release. The current patch level is 3, but when retrieving distributions, you should get the version with the highest version, release, and patch level. (Note that release levels greater than or equal to 90 denote “beta,” or non-production software; you may not wish to retrieve such a version unless you don’t mind experimenting.)

If you are not on a Unix system, you will need to make other arrangements for getting and extracting the gawk distribution. You should consult a local expert.

B.1.3 Contents of the gawk Distribution

The gawk distribution has a number of C source files, documentation files, subdirectories and files related to the configuration process (see Section B.2 [Compiling and Installing gawk on Unix], page 282), and several subdirectories related to different, non-Unix, operating systems.

various ‘.c’, ‘.y’, and ‘.h’ files These files are the actual gawk source code.

‘README’
‘README_d/README.*’

Descriptive files: ‘README’ for gawk under Unix, and the rest for the various hardware and software combinations.

‘INSTALL’ A file providing an overview of the configuration and installation process.

‘PORTS’ A list of systems to which gawk has been ported, and which have successfully run the test suite.

‘ACKNOWLEDGMENT’

A list of the people who contributed major parts of the code or documentation.

‘ChangeLog’

A detailed list of source code changes as bugs are fixed or improvements made.

‘NEWS’ A list of changes to gawk since the last release or patch.

‘COPYING’ The GNU General Public License.

‘FUTURES’ A brief list of features and/or changes being contemplated for future releases, with some indication of the time frame for the feature, based on its difficulty.

‘LIMITATIONS’

A list of those factors that limit gawk’s performance. Most of these depend on the hardware or operating system software, and are not limits in gawk itself.

‘POSIX.STD’

A description of one area where the POSIX standard for awk is incorrect, and how gawk handles the problem.

‘PROBLEMS’

A file describing known problems with the current release.

‘doc/awkforai.txt’

A short article describing why gawk is a good language for AI (Artificial Intelligence) programming.

‘doc/README.card’
‘doc/ad.block’
‘doc/awkcard.in’
‘doc/cardfonts’
‘doc/colors’
‘doc/macros’
‘doc/no.colors’
‘doc/setter.outline’

The troff source for a five-color awk reference card. A modern version of troff, such as GNU Troff (groff) is needed to produce the color version. See the file ‘README.card’ for instructions if you have an older troff.

‘doc/gawk.1’

The troff source for a manual page describing gawk. This is distributed for the convenience of Unix users.

‘doc/gawk.texi’

The Texinfo source file for this book. It should be processed with TEX to produce a printed document, and with makeinfo to produce an Info file.

‘doc/gawk.info’

The generated Info file for this book.

‘doc/igawk.1’

The troff source for a manual page describing the igawk program presented in Section 16.2.9 [An Easy Way to Use Library Functions], page 242.

‘doc/Makefile.in’

The input file used during the configuration process to generate the actual ‘Makefile’ for creating the documentation.

‘Makefile.in’
‘acconfig.h’
‘aclocal.m4’
‘configh.in’
‘configure.in’
‘configure’
‘custom.h’
‘missing/*’

These files and subdirectory are used when configuring gawk for various Unix systems. They are explained in detail in Section B.2 [Compiling and Installing gawk on Unix], page 282.

‘awklib/extract.awk’
‘awklib/Makefile.in’

The ‘awklib’ directory contains a copy of ‘extract.awk’ (see Section 16.2.7 [Extracting Programs from Texinfo Source Files], page 236), which can be used to extract the sample programs from the Texinfo source file for this book, and a ‘Makefile.in’ file, which configure uses to generate a ‘Makefile’. As part of the process of building gawk, the library functions from Chapter 15 [A Library of awk Functions], page 167, and the igawk program from Section 16.2.9 [An Easy Way to Use Library Functions], page 242, are extracted into ready to use files. They are installed as part of the installation process.

‘atari/*’ Files needed for building gawk on an Atari ST. See Section B.5 [Installing gawk on the Atari ST], page 287, for details.

‘pc/*’ Files needed for building gawk under MS-DOS and OS/2. See Section B.4 [MS-DOS and OS/2 Installation and Compilation], page 286, for details.

‘vms/*’ Files needed for building gawk under VMS. See Section B.3 [How to Compile and Install gawk on VMS], page 283, for details.

‘test/*’ A test suite for gawk. You can use ‘make check’ from the top level gawk directory to run your version of gawk against the test suite. If gawk successfully passes ‘make check’ then you can be confident of a successful port.

B.2 Compiling and Installing gawk on Unix

Usually, you can compile and install gawk by typing only two commands. However, if you do use an unusual system, you may need to configure gawk for your system yourself.

B.2.1 Compiling gawk for Unix

After you have extracted the gawk distribution, cd to ‘gawk-3.0.3’. Like most GNU software, gawk is configured automatically for your Unix system by running the configure program. This program is a Bourne shell script that was generated automatically using GNU autoconf. (The autoconf software is described fully in Autoconf—Generating Automatic Configuration Scripts, which is available from the Free Software Foundation.)

To configure gawk, simply run configure:

    sh ./configure

This produces a ‘Makefile’ and ‘config.h’ tailored to your system. The ‘config.h’ file describes various facts about your system. You may wish to edit the ‘Makefile’ to change the CFLAGS variable, which controls the command line options that are passed to the C compiler (such as optimization levels, or compiling for debugging).

Alternatively, you can add your own values for most make variables, such as CC and CFLAGS, on the command line when running configure:

    CC=cc CFLAGS=-g sh ./configure

See the file ‘INSTALL’ in the gawk distribution for all the details. After you have run configure, and possibly edited the ‘Makefile’, type:

    make

and shortly thereafter, you should have an executable version of gawk. That’s all there is to it! (If these steps do not work, please send in a bug report; see Section B.7 [Reporting Problems and Bugs], page 290.)

B.2.2 The Configuration Process

(This section is of interest only if you know something about using the C language and the Unix operating system.)

The source code for gawk generally attempts to adhere to formal standards wherever possible. This means that gawk uses library routines that are specified by the ANSI C standard and by the POSIX operating system interface standard. When using an ANSI C compiler, function prototypes are used to help improve the compile-time checking.

Many Unix systems do not support all of either the ANSI or the POSIX standards. The ‘missing’ subdirectory in the gawk distribution contains replacement versions of those subroutines that are most likely to be missing.

The ‘config.h’ file that is created by the configure program contains definitions that describe features of the particular operating system where you are attempting to compile gawk. The three things described by this file are what header files are available, so that they can be correctly included, what (supposedly) standard functions are actually available in your C libraries, and other miscellaneous facts about your variant of Unix. For example, there may not be an st_blksize element in the stat structure. In this case ‘HAVE_ST_BLKSIZE’ would be undefined.

It is possible for your C compiler to lie to configure. It may do so by not exiting with an error when a library function is not available. To get around this, you can edit the file ‘custom.h’. Use an ‘#ifdef’ that is appropriate for your system, and either #define any constants that configure should have defined but didn’t, or #undef any constants that configure defined and should not have. ‘custom.h’ is automatically included by ‘config.h’.

It is also possible that the configure program generated by autoconf will not work on your system in some other fashion. If you do have a problem, the file ‘configure.in’ is the input for autoconf. You may be able to change this file, and generate a new version of configure that will work on your system. See Section B.7 [Reporting Problems and Bugs], page 290, for information on how to report problems in configuring gawk. The same mechanism may be used to send in updates to ‘configure.in’ and/or ‘custom.h’.

B.3 How to Compile and Install gawk on VMS

This section describes how to compile and install gawk under VMS.

B.3.1 Compiling gawk on VMS

To compile gawk under VMS, there is a DCL command procedure that will issue all the necessary CC and LINK commands, and there is also a ‘Makefile’ for use with the MMS utility. From the source directory, use either

    $ @[.VMS]VMSBUILD.COM

or

    $ MMS/DESCRIPTION=[.VMS]DESCRIP.MMS GAWK

Depending upon which C compiler you are using, follow one of the sets of instructions in this table:

VAX C V3.x Use either ‘vmsbuild.com’ or ‘descrip.mms’ as is. These use CC/OPTIMIZE=NOLINE, which is essential for Version 3.0.

VAX C V2.x You must have Version 2.3 or 2.4; older ones won’t work. Edit either ‘vmsbuild.com’ or ‘descrip.mms’ according to the comments in them. For ‘vmsbuild.com’, this just entails removing two ‘!’ delimiters. Also edit ‘config.h’ (which is a copy of file ‘[.config]vms-conf.h’) and comment out or delete the two

          lines ‘#define __STDC__0’ and ‘#define VAXC_BUILTINS’ near

the end.

GNU C Edit ‘vmsbuild.com’ or ‘descrip.mms’; the changes are different from those for VAX C V2.x, but equally straightforward. No changes to ‘config.h’ should be needed.

DEC C Edit ‘vmsbuild.com’ or ‘descrip.mms’ according to their comments. No changes to ‘config.h’ should be needed.

gawk has been tested under VAX/VMS 5.5-1 using VAX C V3.2, GNU C 1.40 and 2.3. It should work without modifications for VMS V4.6 and up.

B.3.2 Installing gawk on VMS

To install gawk, all you need is a “foreign” command, which is a DCL symbol whose value begins with a dollar sign. For example:

    $ GAWK :== $disk1:[gnubin]GAWK

(Substitute the actual location of gawk.exe for ‘$disk1:[gnubin]’.) The symbol should be placed in the ‘login.com’ of any user who wishes to run gawk, so that it will be defined every time the user logs on. Alternatively, the symbol may be placed in the system-wide ‘sylogin.com’ procedure, which will allow all users to run gawk.

Optionally, the help entry can be loaded into a VMS help library:

    $ LIBRARY/HELP SYS$HELP:HELPLIB [.VMS]GAWK.HLP

(You may want to substitute a site-specific help library rather than the standard VMS library ‘HELPLIB’.) After loading the help text,

    $ HELP GAWK

will provide information about both the gawk implementation and the awk programming language.

The logical name ‘AWK_LIBRARY’ can designate a default location for awk program files. For the ‘-f’ option, if the specified filename has no device or directory path information in it, gawk will look in the current directory first, then in the directory specified by the translation of ‘AWK_LIBRARY’ if the file was not found. If after searching in both directories, the file still is not found, then gawk appends the suffix ‘.awk’ to the filename and the file search will be re-tried. If ‘AWK_LIBRARY’ is not defined, that portion of the file search will fail benignly.

B.3.3 Running gawk on VMS

Command line parsing and quoting conventions are significantly different on VMS, so examples in this book or from other sources often need minor changes. They are minor though, and all awk programs should run correctly.

Here are a couple of trivial tests:

    $ gawk -- "BEGIN {print ""Hello, World!""}"
    $ gawk -"W" version
    ! could also be -"W version" or "-W version"

Note that upper-case and mixed-case text must be quoted.

The VMS port of gawk includes a DCL-style interface in addition to the original shell-style interface (see the help entry for details). One side-effect of dual command line parsing is that if there is only a single parameter (as in the quoted string program above), the command becomes ambiguous. To work around this, the normally optional ‘--’ flag is required to force Unix style rather than DCL parsing. If any other dash-type options (or multiple parameters such as data files to be processed) are present, there is no ambiguity and ‘--’ can be omitted.

The default search path when looking for awk program files specified by the ‘-f’ option is "SYS$DISK:[],AWK_LIBRARY:". The logical name ‘AWKPATH’ can be used to override this default. The format of ‘AWKPATH’ is a comma-separated list of directory specifications. When defining it, the value should be quoted so that it retains a single translation, and not a multi-translation RMS searchlist.

B.3.4 Building and Using gawk on VMS POSIX

Ignore the instructions above, although ‘vms/gawk.hlp’ should still be made available in a help library. The source tree should be unpacked into a container file subsystem rather than into the ordinary VMS file system. Make sure that the two scripts, ‘configure’ and ‘vms/posix-cc.sh’, are executable; use ‘chmod +x’ on them if necessary. Then execute the following two commands:

    psx> CC=vms/posix-cc.sh configure
    psx> make CC=c89 gawk

The first command will construct files ‘config.h’ and ‘Makefile’ out of templates, using a script to make the C compiler fit configure’s expectations. The second command will compile and link gawk using the C compiler directly; ignore any warnings from make about being unable to redefine CC. configure will take a very long time to execute, but at least it provides incremental feedback as it runs.

This has been tested with VAX/VMS V6.2, VMS POSIX V2.0, and DEC C V5.2.

Once built, gawk will work like any other shell utility. Unlike the normal VMS port of gawk, no special command line manipulation is needed in the VMS POSIX environment.

B.4 MS-DOS and OS/2 Installation and Compilation

If you have received a binary distribution prepared by the DOS maintainers, then gawk and the necessary support files will appear under the ‘gnu’ directory, with executables in ‘gnu/bin’, libraries in ‘gnu/lib/awk’, and manual pages under ‘gnu/man’. This is designed for easy installation to a ‘/gnu’ directory on your drive, but the files can be installed anywhere provided AWKPATH is set properly. Regardless of the installation directory, the first line of ‘igawk.cmd’ and ‘igawk.bat’ (in ‘gnu/bin’) may need to be edited.

The binary distribution will contain a separate file describing the contents. In particular, it may include more than one version of the gawk executable. OS/2 binary distributions may have a different arrangement, but installation is similar.

The OS/2 and MS-DOS versions of gawk search for program files as described in Section 14.3 [The AWKPATH Environment Variable], page 164. However, semicolons (rather than colons) separate elements in the AWKPATH variable. If AWKPATH is not set or is empty, then the default search path is

".;c:/lib/awk;c:/gnu/lib/awk".

An sh-like shell (as opposed to command.com under MS-DOS or cmd.exe under OS/2) may be useful for awk programming. Ian Stewartson has written an excellent shell for MS-DOS and OS/2, and a ksh clone and GNU Bash are available for OS/2. The file ‘README_d/README.pc’ in the gawk distribution contains information on these shells. Users of Stewartson’s shell on DOS should examine its documentation on handling of command-lines. In particular, the setting for gawk in the shell configuration may need to be changed, and the ignoretype option may also be of interest.

gawk can be compiled for MS-DOS and OS/2 using the GNU development tools from DJ Delorie (DJGPP, MS-DOS-only) or Eberhard Mattes (EMX, MS-DOS and OS/2). Microsoft C can be used to build 16-bit versions for MS-DOS and OS/2. The file ‘README_d/README.pc’ in the gawk distribution contains additional notes, and ‘pc/Makefile’ contains important notes on compilation options.

To build gawk, copy the files in the ‘pc’ directory (except for ‘ChangeLog’) to the directory with the rest of the gawk sources. The ‘Makefile’ contains a configuration section with comments, and may need to be edited in order to work with your make utility.

The ‘Makefile’ contains a number of targets for building various MS-DOS and OS/2 versions. A list of targets will be printed if the make command is given without a target. As an example, to build gawk using the DJGPP tools, enter ‘make djgpp’.

Using make to run the standard tests and to install gawk requires additional Unix-like tools, including sh, sed, and cp. In order to run the tests, the ‘test/*.ok’ files may need to be converted so that they have the usual DOS-style end-of-line markers. Most of the tests will work properly with Stewartson’s shell along with the companion utilities or appropriate GNU utilities. However, some editing of ‘test/Makefile’ is required. It is recommended that the file ‘pc/Makefile.tst’ be copied to ‘test/Makefile’ as a replacement. Details can be found in ‘README_d/README.pc’.

B.5 Installing gawk on the Atari ST

There are no substantial differences when installing gawk on various Atari models. Compiled gawk executables do not require a large amount of memory with most awk programs and should run on all Motorola processor based models (called further ST, even if that is not exactly right).

In order to use gawk, you need to have a shell, either text or graphics, that does not map all the characters of a command line to upper-case. Maintaining case distinction in option flags is very important (see Section 14.1 [Command Line Options], page 159). These days this is the default, and it may only be a problem for some very old machines. If your system does not preserve the case of option flags, you will need to upgrade your tools. Support for I/O redirection is necessary to make it easy to import awk programs from other environments. Pipes are nice to have, but not vital.

B.5.1 Compiling gawk on the Atari ST

A proper compilation of gawk sources when sizeof(int) differs from sizeof(void *) requires an ANSI C compiler. An initial port was done with gcc. You may actually prefer executables where ints are four bytes wide, but the other variant works as well.

You may need quite a bit of memory when trying to recompile the gawk sources, as some source files (‘regex.c’ in particular) are quite big. If you run out of memory compiling such a file, try reducing the optimization level for this particular file; this may help.

With a reasonable shell (Bash will do), and in particular if you run Linux, MiNT or a similar operating system, you have a pretty good chance that the configure utility will succeed. Otherwise sample versions of ‘config.h’ and ‘Makefile.st’ are given in the ‘atari’ subdirectory and can be edited and copied to the corresponding files in the main source directory. Even if configure produced something, it might be advisable to compare its results with the sample versions and possibly make adjustments.

Some gawk source code fragments depend on a preprocessor define ‘atarist’. This basically assumes the TOS environment with gcc. Modify these sections as appropriate if they are not right for your environment. Also see the remarks about AWKPATH and envsep in Section B.5.2 [Running gawk on the Atari ST], page 288.

As shipped, the sample ‘config.h’ claims that the system function is missing from the libraries, which is not true, and an alternative implementation of this function is provided in ‘atari/system.c’. Depending upon your particular combination of shell and operating system, you may wish to change the file to indicate that system is available.

B.5.2 Running gawk on the Atari ST

An executable version of gawk should be placed, as usual, anywhere in your PATH where your shell can find it.

While executing, gawk creates a number of temporary files. When using gcc libraries for TOS, gawk looks for either of the environment variables TEMP or TMPDIR, in that order. If either one is found, its value is assumed to be a directory for temporary files. This directory must exist, and if you can spare the memory, it is a good idea to put it on a RAM drive. If neither TEMP nor TMPDIR are found, then gawk uses the current directory for its temporary files.

The ST version of gawk searches for its program files as described in Section 14.3 [The AWKPATH Environment Variable], page 164. The default value for the AWKPATH variable is taken from DEFPATH defined in ‘Makefile’. The sample gcc/TOS ‘Makefile’ for the ST in the distribution sets DEFPATH to ".,c:\lib\awk,c:\gnu\lib\awk". The search path can be modified by explicitly setting AWKPATH to whatever you wish. Note that colons cannot be used on the ST to separate elements in the AWKPATH variable, since they have another, reserved, meaning. Instead, you must use a comma to separate elements in the path. When recompiling, the separating character can be modified by initializing the envsep variable in ‘atari/gawkmisc.atr’ to another value.

Although awk allows great flexibility in doing I/O redirections from within a program, this facility should be used with care on the ST running under TOS. In some circumstances the OS routines for file handle pool processing lose track of certain events, causing the computer to crash, and requiring a reboot. Often a warm reboot is sufficient. Fortunately, this happens infrequently, and in rather esoteric situations. In particular, avoid having one part of an awk program using print statements explicitly redirected to "/dev/stdout", while other print statements use the default standard output, and a calling shell has redirected standard output to a file.

When gawk is compiled with the ST version of gcc and its usual libraries, it will accept both ‘/’ and ‘\’ as path separators. While this is convenient, it should be remembered that this removes one, technically valid, character (‘/’) from your file names, and that it may create problems for external programs, called via the system function, which may not support this convention. Whenever it is possible that a file created by gawk will be used by some other program, use only backslashes. Also remember that in awk, backslashes in strings have to be doubled in order to get literal backslashes (see Section 4.2 [Escape Sequences], page 24).

B.6 Installing gawk on an Amiga

You can install gawk on an Amiga system using a Unix emulation environment available via anonymous ftp from ftp.ninemoons.com in the directory ‘pub/ade/current’. This includes a shell based on pdksh. The primary component of this environment is a Unix emulation library, ‘ixemul.lib’.

A more complete distribution for the Amiga is available on the Geek Gadgets CD-ROM from:

CRONUS

1840 E. Warner Road #105-265

Tempe, AZ 85284 USA

US Toll Free: (800) 804-0833 Phone: +1-602-491-0442 FAX: +1-602-491-0048

    Email: info@ninemoons.com
    WWW: http://www.ninemoons.com
    Anonymous ftp site: ftp.ninemoons.com

Once you have the distribution, you can configure gawk simply by running

configure:
    configure -v m68k-amigaos

Then run make, and you should be all set! (If these steps do not work, please send in a bug report; see Section B.7 [Reporting Problems and Bugs], page 290.)

B.7 Reporting Problems and Bugs

There is nothing more dangerous than a bored archeologist.

The Hitchhiker’s Guide to the Galaxy

If you have problems with gawk or think that you have found a bug, please report it to the developers; we cannot promise to do anything but we might well want to fix it.

Before reporting a bug, make sure you have actually found a real bug. Carefully reread the documentation and see if it really says you can do what you’re trying to do. If it’s not clear whether you should be able to do something or not, report that too; it’s a bug in the documentation!

Before reporting a bug or trying to fix it yourself, try to isolate it to the smallest possible awk program and input data file that reproduces the problem. Then send us the program and data file, some idea of what kind of Unix system you’re using, and the exact results gawk gave you. Also say what you expected to occur; this will help us decide whether the problem was really in the documentation.

Once you have a precise problem, there are two e-mail addresses you can send mail to.

Internet: ‘bug-gnu-utils@prep.ai.mit.edu’
UUCP:  ‘uunet!prep.ai.mit.edu!bug-gnu-utils’

Please include the version number of gawk you are using. You can get this information with the command ‘gawk --version’. You should send a carbon copy of your mail to Arnold Robbins, who can be reached at

‘arnold@gnu.ai.mit.edu’.

Important! Do not try to report bugs in gawk by posting to the Usenet/Internet newsgroup comp.lang.awk. While the gawk developers do occasionally read this newsgroup, there is no guarantee that we will see your posting. The steps described above are the official, recognized ways for reporting bugs.

Non-bug suggestions are always welcome as well. If you have questions about things that are unclear in the documentation or are just obscure features, ask Arnold Robbins; he will try to help you out, although he may not have the time to fix the problem. You can send him electronic mail at the Internet address above.

If you find bugs in one of the non-Unix ports of gawk, please send an electronic mail message to the person who maintains that port. They are listed below, and also in the ‘README’ file in the gawk distribution. Information in the ‘README’ file should be considered authoritative if it conflicts with this book.

The people maintaining the non-Unix ports of gawk are:

MS-DOS Scott Deifik, ‘scottd@amgen.com’, and Darrel Hankerson,

‘hankedr@mail.auburn.edu’.

OS/2 Kai Uwe Rommel, ‘rommel@ars.de’.

VMS  Pat Rankin, ‘rankin@eql.caltech.edu’.
Atari ST Michal Jaegermann, ‘michal@gortel.phys.ualberta.ca’.
Amiga  Fred Fish, ‘fnf@ninemoons.com’.

If your bug is also reproducible under Unix, please send copies of your report to the general GNU bug list, as well as to Arnold Robbins, at the addresses listed above.

B.8 Other Freely Available awk Implementations

It’s kind of fun to put comments like this in your awk code.

// Do C++ comments work? answer: yes! of course

Michael Brennan

There are two other freely available awk implementations. This section briefly describes where to get them.

Unix awk Brian Kernighan has been able to make his implementation of awk freely available. You can get it via anonymous ftp to the host netlib.att.com. Change directory to ‘/netlib/research’. Use “binary” or “image” mode, and re-

trieve ‘awk.bundle.Z’.

This is a shell archive that has been compressed with the compress utility. It can be uncompressed with either

uncompress or the GNU gunzip utility.

This version requires an ANSI C compiler; GCC (the GNU C compiler) works quite nicely.

mawk Michael Brennan has written an independent implementation of awk, called mawk. It is available under the GPL (see [GNU GENERAL PUBLIC LICENSE], page 307), just as gawk is. You can get it via anonymous ftp to the host ftp.whidbey.net. Change directory to ‘/pub/brennan’. Use “binary” or “image” mode, and retrieve ‘mawk1.3.3.tar.gz’ (or the latest version that is there). gunzip may be used to decompress this file. Installation is similar to gawk’s (see Section B.2 [Compiling and Installing gawk on Unix], page 282).

292 Effective AWK Programming

Appendix C: Implementation Notes 293

Appendix C Implementation Notes

This appendix contains information mainly of interest to implementors and maintainers of gawk. Everything in it applies specifically to gawk, and not to other implementations.

C.1 Downward Compatibility and Debugging

See Section 17.5 [Extensions in gawk Not in POSIX awk], page 254, for a summary of the GNU extensions to the awk language and program. All of these features can be turned off by invoking gawk with the ‘--traditional’ option, or with the ‘--posix’ option.

If gawk is compiled for debugging with ‘-DDEBUG’, then there is one more option available on the command line:

-W parsedebug
--parsedebug

Print out the parse stack information as the program is being parsed.

This option is intended only for serious gawk developers, and not for the casual user. It probably has not even been compiled into your version of gawk, since it slows down execution.

C.2 Making Additions to gawk

If you should find that you wish to enhance gawk in a significant fashion, you are perfectly free to do so. That is the point of having free software; the source code is available, and you are free to change it as you wish (see [GNU GENERAL PUBLIC LICENSE], page 307).

This section discusses the ways you might wish to change gawk, and any considerations you should bear in mind.

C.2.1 Adding New Features

You are free to add any new features you like to gawk. However, if you want your changes to be incorporated into the gawk distribution, there are several steps that you need to take in order to make it possible for me to include to your changes.

1. Get the latest version. It is much easier for me to integrate changes if they are relative to the most recent distributed version of gawk. If your version of gawk is very old, I may not be able to integrate them at all. See Section B.1.1 [Getting the gawk Distribution], page 277, for information on getting the latest version of gawk.

2. Follow the GNU Coding Standards. This document describes how GNU software should be written. If you haven’t read it, please do so, preferably before starting to modify gawk. (The GNU Coding Standards are available as part of the Autoconf distribution, from the FSF.)

3. Use the gawk coding style. The C code for gawk follows the instructions in the GNU Coding Standards, with minor exceptions. The code is formatted using the traditional “K&R” style, particularly as regards the placement of braces and the use of tabs. In brief, the coding rules for gawk are:

• Use old style (non-prototype) function headers when defining functions.

• Put the name of the function at the beginning of its own line.

• Put the return type of the function, even if it is int, on the line above the line with the name and arguments of the function.

• The declarations for the function arguments should not be indented.

• Put spaces around parentheses used in control structures (if,

  while, for, do, switch and return).

• Do not put spaces in front of parentheses used in function calls.

• Put spaces around all C operators, and after commas in function calls.

• Do not use the comma operator to produce multiple side-effects, except in for loop initialization and increment parts, and in macro bodies.

• Use real tabs for indenting, not spaces.

• Use the “K&R” brace layout style.

• Use comparisons against NULL and '\0' in the conditions of if, while and for statements, and in the cases of switch statements, instead of just the plain pointer or character value.

• Use the TRUE, FALSE, and NULL symbolic constants, and the character constant '\0' where appropriate, instead of 1 and 0.

• Provide one-line descriptive comments for each function.

• Do not use ‘#elif’. Many older Unix C compilers cannot handle it.

• Do not use the alloca function for allocating memory off the stack. Its use causes more portability trouble than the minor benefit of not having to free the storage. Instead, use malloc and free.

If I have to reformat your code to follow the coding style used in gawk, I may not bother.

4. Be prepared to sign the appropriate paperwork. In order for the FSF to distribute your changes, you must either place those changes in the public domain, and submit a signed statement to that effect, or assign the copyright in your changes to the FSF. Both of these actions are easy to do, and many people have done so already. If you have questions, please contact me (see Section B.7 [Reporting Problems and Bugs], page 290),

or gnu@prep.ai.mit.edu.

5. Update the documentation. Along with your new code, please supply new sections and or chapters for this book. If at all possible, please use real Texinfo, instead of just supplying unformatted ASCII text (although even that is better than no documentation at all). Conventions to be followed in Effective AWK Programming are provided after the ‘@bye’ at the end of the Texinfo source file. If possible, please update the man page as well. You will also have to sign paperwork for your documentation changes.

6. Submit changes as context diffs or unified diffs. Use ‘diff -c -r -N’ or ‘diff -u -r -N’ to compare the original gawk source tree with your version. (I find context diffs to be more readable, but unified diffs are more compact.) I recommend using the GNU version of diff. Send the output produced by either run of diff to me when you submit your changes. See Section B.7 [Reporting Problems and Bugs], page 290, for the electronic mail information. Using this format makes it easy for me to apply your changes to the master version of the gawk source code (using patch). If I have to apply the changes manually, using a text editor, I may not do so, particularly if there are lots of changes.

Although this sounds like a lot of work, please remember that while you may write the new code, I have to maintain it and support it, and if it isn’t possible for me to do that with a minimum of extra work, then I probably will not.

C.2.2 Porting gawk to a New Operating System

If you wish to port gawk to a new operating system, there are several steps to follow.

1. Follow the guidelines in Section C.2.1 [Adding New Features], page 293, concerning coding style, submission of diffs, and so on.

2. When doing a port, bear in mind that your code must co-exist peacefully with the rest of gawk, and the other ports. Avoid gratuitous changes to the system-independent parts of the code. If at all possible, avoid sprinkling ‘#ifdef’s just for your port throughout the code. If the changes needed for a particular system affect too much of the code, I probably will not accept them. In such a case, you will, of course, be able to distribute your changes on your own, as long as you comply with the GPL (see [GNU GENERAL PUBLIC LICENSE], page 307).

3. A number of the files that come with gawk are maintained by other people at the Free Software Foundation. Thus, you should not change them unless it is for a very good reason. I.e. changes are not out of the question, but changes to these files will be scrutinized extra carefully. The

files are ‘alloca.c’, ‘getopt.h’, ‘getopt.c’, ‘getopt1.c’, ‘regex.h’,
‘regex.c’, ‘dfa.h’, ‘dfa.c’, ‘install-sh’, and ‘mkinstalldirs’.

4. Be willing to continue to maintain the port. Non-Unix operating systems are supported by volunteers who maintain the code needed to compile and run gawk on their systems. If no-one volunteers to maintain a port, that port becomes unsupported, and it may be necessary to remove it from the distribution.

5. Supply an appropriate ‘gawkmisc.???’ file. Each port has its own ‘gawkmisc.???’ that implements certain operating system specific functions. This is cleaner than a plethora of ‘#ifdef’s scattered throughout the code. The ‘gawkmisc.c’ in the main source directory includes the appropriate ‘gawkmisc.???’ file from each subdirectory. Be sure to update it as well. Each port’s ‘gawkmisc.???’ file has a suffix reminiscent of the machine or operating system for the port. For example, ‘pc/gawkmisc.pc’ and ‘vms/gawkmisc.vms’. The use of separate suffixes, instead of plain ‘gawkmisc.c’, makes it possible to move files from a port’s subdirectory into the main subdirectory, without accidentally destroying the real ‘gawkmisc.c’ file. (Currently, this is only an issue for the MS-DOS and OS/2 ports.)

6. Supply a ‘Makefile’ and any other C source and header files that are necessary for your operating system. All your code should be in a separate subdirectory, with a name that is the same as, or reminiscent of, either your operating system or the computer system. If possible, try to structure things so that it is not necessary to move files out of the subdirectory into the main source directory. If that is not possible, then be sure to avoid using names for your files that duplicate the names of files in the main source directory.

7. Update the documentation. Please write a section (or sections) for this book describing the installation and compilation steps needed to install and/or compile gawk for your system.

8. Be prepared to sign the appropriate paperwork. In order for the FSF to distribute your code, you must either place your code in the public domain, and submit a signed statement to that effect, or assign the copyright in your code to the FSF.

Following these steps will make it much easier to integrate your changes into gawk, and have them co-exist happily with the code for other operating systems that is already there.

In the code that you supply, and that you maintain, feel free to use a coding style and brace layout that suits your taste.

C.3 Probable Future Extensions

AWK is a language similar to PERL, only considerably more elegant. Arnold Robbins

Hey! Larry Wall

This section briefly lists extensions and possible improvements that indicate the directions we are currently considering for gawk. The file ‘FUTURES’ in the gawk distributions lists these extensions as well.

This is a list of probable future changes that will be usable by the awk language programmer.

Localization The GNU project is starting to support multiple languages. It will at least be possible to make gawk print its warnings and error messages in languages other than English. It may be possible for awk programs to also use the multiple language facilities, separate from gawk itself.

Databases It may be possible to map a GDBM/NDBM/SDBM file into an awk array.

A PROCINFO Array

The special files that provide process-related information (see Section 6.7 [Special File Names in gawk], page 72) may be superseded by a PROCINFO array that would provide the same information, in an easier to access fashion.

More lint warnings There are more things that could be checked for portability.

Control of subprocess environment Changes made in gawk to the array ENVIRON may be propagated to subprocesses run by gawk.

This is a list of probable improvements that will make gawk perform better.

An Improved Version of dfa The dfa pattern matcher from GNU grep has some problems. Either a new version or a fixed one will deal with some important regexp matching issues.

Use of GNU malloc The GNU version of malloc could potentially speed up gawk, since it relies heavily on the use of dynamic memory allocation.

Use of the rx regexp library The rx regular expression library could potentially speed up all regexp operations that require knowing the exact location of matches. This includes record termination, field and array splitting, and the sub, gsub, gensub and match functions.

C.4 Suggestions for Improvements

Here are some projects that would-be gawk hackers might like to take on. They vary in size from a few days to a few weeks of programming, depending on which one you choose and how fast a programmer you are. Please send any improvements you write to the maintainers at the GNU project. See Section C.2.1 [Adding New Features], page 293, for guidelines to follow when adding new features to gawk. See Section B.7 [Reporting Problems and Bugs], page 290, for information on contacting the maintainers.

1. Compilation of awk programs: gawk uses a Bison (YACC-like) parser to convert the script given it into a syntax tree; the syntax tree is then executed by a simple recursive evaluator. This method incurs a lot of overhead, since the recursive evaluator performs many procedure calls to do even the simplest things. It should be possible for gawk to convert the script’s parse tree into a C program which the user would then compile, using the normal C compiler and a special gawk library to provide all the needed functions (regexps, fields, associative arrays, type coercion, and so on). An easier possibility might be for an intermediate phase of awk to convert the parse tree into a linear byte code form like the one used in GNU Emacs Lisp. The recursive evaluator would then be replaced by a straight line byte code interpreter that would be intermediate in speed between running a compiled program and doing what gawk does now.

2. The programs in the test suite could use documenting in this book.

3. See the ‘FUTURES’ file for more ideas. Contact us if you would seriously like to tackle any of the items listed there.

Appendix D: Glossary 299

Appendix D Glossary

Action A series of awk statements attached to a rule. If the rule’s pattern matches an input record, awk executes the rule’s action. Actions are always enclosed in curly braces. See Section 8.2 [Overview of Actions], page 102.

Amazing awk Assembler Henry Spencer at the University of Toronto wrote a retargetable assembler completely as awk scripts. It is thousands of lines long, including machine descriptions for several eight-bit microcomputers. It is a good example of a program that would have been better written in another language.

Amazingly Workable Formatter (awf)

Henry Spencer at the University of Toronto wrote a formatter that accepts a large subset of the ‘nroff -ms’ and ‘nroff -man’ formatting commands, using awk and sh.

ANSI The American National Standards Institute. This organization produces many standards, among them the standards for the C and C++ programming languages.

Assignment An awk expression that changes the value of some awk variable or data object. An object that you can assign to is called an lvalue. The assigned values are called rvalues. See Section 7.7 [Assignment Expressions], page 84.

awk Language The language in which awk programs are written.

awk Program An awk program consists of a series of patterns and actions, collectively known as rules. For each input record given to the program, the program’s rules are all processed in turn. awk programs may also contain function definitions.

awk Script Another name for an awk program.

Bash The GNU version of the standard shell (the Bourne-Again shell). See “Bourne Shell.”

BBS See “Bulletin Board System.”

Boolean Expression Named after the English mathematician Boole. See “Logical Expression.”

Bourne Shell The standard shell (‘/bin/sh’) on Unix and Unix-like systems, originally written by Steven R. Bourne. Many shells (Bash, ksh, pdksh, zsh) are generally upwardly compatible with the Bourne shell.

Built-in Function The awk language provides built-in functions that perform various numerical, time stamp related, and string computations. Examples are sqrt (for the square root of a number) and substr (for a substring of a string). See Chapter 12 [Built-in Functions], page 133.

Built-in Variable

ARGC, ARGIND, ARGV, CONVFMT, ENVIRON, ERRNO, FIELDWIDTHS,
FILENAME, FNR, FS, IGNORECASE, NF, NR, OFMT, OFS, ORS,

RLENGTH, RSTART, RS, RT, and SUBSEP, are the variables that have special meaning to awk. Changing some of them affects awk’s running environment. Several of these variables are specific to gawk. See Chapter 10 [Built-in Variables], page 115.

Braces See “Curly Braces.” Bulletin Board System

A computer system allowing users to log in and read and/or leave messages for other users of the system, much like leaving paper notes on a bulletin board.

C The system programming language that most GNU software is written in. The awk programming language has C-like syntax, and this book points out similarities between awk and C when appropriate.

Character Set The set of numeric codes used by a computer system to represent the characters (letters, numbers, punctuation, etc.) of a particular country or place. The most common character set in use today is ASCII (American Standard Code for Information Interchange). Many European countries use an extension of ASCII known as ISO-8859-1 (ISO Latin-1).

CHEM A preprocessor for pic that reads descriptions of molecules and produces pic input for drawing them. It was written in awk by Brian Kernighan and Jon Bentley, and is available from

netlib@research.att.com.

Compound Statement A series of awk statements, enclosed in curly braces. Compound statements may be nested. See Chapter 9 [Control Statements in Actions], page 105.

Concatenation Concatenating two strings means sticking them together, one after another, giving a new string. For example, the string ‘foo’ concatenated with the string ‘bar’ gives the string ‘foobar’. See Section 7.6 [String Concatenation], page 83.

Conditional Expression An expression using the ‘?:’ ternary operator, such as ‘expr1 ? expr2 : expr3’. The expression expr1 is evaluated; if the result is true, the value of the whole expression is the value of expr2, otherwise the value is expr3. In either case, only one of expr2 and expr3 is evaluated. See Section 7.12 [Conditional Expressions], page 92.

Comparison Expression A relation that is either true or false, such as ‘(a < b)’. Comparison expressions are used in if, while, do, and for statements, and in patterns to select which input records to process. See Section 7.10 [Variable Typing and Comparison Expressions], page 88.

Curly Braces The characters ‘{’ and ‘}’. Curly braces are used in awk for delimiting actions, compound statements, and function bodies.

Dark Corner An area in the language where specifications often were (or still are) not clear, leading to unexpected or undesirable behavior. Such areas are marked in this book with “(d.c.)” in the text, and are indexed under the heading “dark corner.”

Data Objects These are numbers and strings of characters. Numbers are converted into strings and vice versa, as needed. See Section 7.4 [Conversion of Strings and Numbers], page 81.

Double Precision An internal representation of numbers that can have fractional parts. Double precision numbers keep track of more digits than do single precision numbers, but operations on them are more expensive. This is the way awk stores numeric values. It is the

C type double.

Dynamic Regular Expression A dynamic regular expression is a regular expression written as an ordinary expression. It could be a string constant, such as "foo", but it may also be an expression whose value can vary. See Section 4.7 [Using Dynamic Regexps], page 35.

Environment A collection of strings, of the form name=val, that each program has available to it. Users generally place values into the environment in order to provide information to various programs. Typical examples are the environment variables HOME and PATH.

Empty String See “Null String.”

Escape Sequences A special sequence of characters used for describing non-printing characters, such as ‘\n’ for newline, or ‘\033’ for the ASCII ESC (escape) character. See Section 4.2 [Escape Sequences], page 24.

Field When awk reads an input record, it splits the record into pieces separated by whitespace (or by a separator regexp which you can change by setting the built-in variable FS). Such pieces are called fields. If the pieces are of fixed length, you can use the built-in variable FIELDWIDTHS to describe their lengths. See Section 5.5 [Specifying How Fields are Separated], page 44, and also see See Section 5.6 [Reading Fixed-width Data], page 49.

Floating Point Number Often referred to in mathematical terms as a “rational” number, this is just a number that can have a fractional part. See “Double Precision” and “Single Precision.”

Format Format strings are used to control the appearance of output in the printf statement. Also, data conversions from numbers to strings are controlled by the format string contained in the builtin variable CONVFMT. See Section 6.5.2 [Format-Control Letters], page 65.

Function A specialized group of statements used to encapsulate general or program-specific tasks. awk has a number of built-in functions, and also allows you to define your own. See Chapter 12 [Built-in Functions], page 133, and Chapter 13 [User-defined Functions], page 151.

FSF See “Free Software Foundation.”

Free Software Foundation A non-profit organization dedicated to the production and distribution of freely distributable software. It was founded by Richard M. Stallman, the author of the original Emacs editor. GNU Emacs is the most widely used version of Emacs today.

gawk The GNU implementation of awk.

General Public License This document describes the terms under which gawk and its source code may be distributed. (see [GNU GENERAL PUB-LIC LICENSE], page 307)

GNU “GNU’s not Unix”. An on-going project of the Free Software Foundation to create a complete, freely distributable, POSIXcompliant computing environment.

GPL See “General Public License.” Hexadecimal

Base 16 notation, where the digits are 0-9 and A-F, with ‘A’ representing 10, ‘B’ representing 11, and so on up to ‘F’ for 15. Hexadecimal numbers are written in C using a leading ‘0x’, to indicate their base. Thus, 0x12 is 18 (one times 16 plus 2).

I/O Abbreviation for “Input/Output,” the act of moving data into and/or out of a running program.

Input Record A single chunk of data read in by awk. Usually, an awk input record consists of one line of text. See Section 5.1 [How Input is Split into Records], page 37.

Integer A whole number, i.e. a number that does not have a fractional part.

Keyword In the awk language, a keyword is a word that has special meaning. Keywords are reserved and may not be used as variable names.

gawk’s keywords are: BEGIN, END, if, else, while, do. . .while,
for, for. . .in, break, continue, delete, next, nextfile,
function, func, and exit.

Logical Expression An expression using the operators for logic, AND, OR, and NOT, written ‘&&’, ‘||’, and ‘!’ in awk. Often called Boolean expressions, after the mathematician who pioneered this kind of mathematical logic.

Lvalue An expression that can appear on the left side of an assignment operator. In most languages, lvalues can be variables or array elements. In awk, a field designator can also be used as an lvalue.

Null String A string with no characters in it. It is represented explicitly in awk programs by placing two double-quote characters next to each other (""). It can appear in input data by having two successive occurrences of the field separator appear next to each other.

Number A numeric valued data object. The gawk implementation uses double precision floating point to represent numbers. Very old awk implementations use single precision floating point.

Octal Base-eight notation, where the digits are 0-7. Octal numbers are written in C using a leading ‘0’, to indicate their base. Thus, 013 is 11 (one times 8 plus 3).

Pattern Patterns tell awk which input records are interesting to which rules. A pattern is an arbitrary conditional expression against which input is tested. If the condition is satisfied, the pattern is said to match the input record. A typical pattern might compare the input record against a regular expression. See Section 8.1 [Pattern Elements], page 97.

POSIX The name for a series of standards being developed by the IEEE that specify a Portable Operating System interface. The “IX” denotes the Unix heritage of these standards. The main standard of interest for awk users is IEEE Standard for Information Technology, Standard 1003.2-1992, Portable Operating System Interface (POSIX) Part 2: Shell and Utilities. Informally, this standard is often referred to as simply “P1003.2.”

Private Variables and/or functions that are meant for use exclusively by library functions, and not for the main awk program. Special care must be taken when naming such variables and functions. See Section 15.13 [Naming Library Function Global Variables], page 200.

Range (of input lines)

A sequence of consecutive lines from the input file. A pattern can specify ranges of input lines for awk to process, or it can specify single lines. See Section 8.1 [Pattern Elements], page 97.

Recursion When a function calls itself, either directly or indirectly. If this isn’t clear, refer to the entry for “recursion.”

Redirection Redirection means performing input from other than the standard input stream, or output to other than the standard output stream. You can redirect the output of the print and printf statements to a file or a system command, using the ‘>’, ‘>>’, and ‘|’ operators. You can redirect input to the getline statement using the ‘<’ and ‘|’ operators. See Section 6.6 [Redirecting Output of print and printf], page 70, and Section 5.8 [Explicit Input with getline], page 53.

Regexp Short for regular expression. A regexp is a pattern that denotes a set of strings, possibly an infinite set. For example, the regexp ‘R.*xp’ matches any string starting with the letter ‘R’ and ending with the letters ‘xp’. In awk, regexps are used in patterns and in conditional expressions. Regexps may contain escape sequences. See Chapter 4 [Regular Expressions], page 23.

Regular Expression See “regexp.”

Regular Expression Constant A regular expression constant is a regular expression written within slashes, such as /foo/. This regular expression is chosen when you write the awk program, and cannot be changed doing its execution. See Section 4.1 [How to Use Regular Expressions], page 23.

Rule A segment of an awk program that specifies how to process single input records. A rule consists of a pattern and an action. awk reads an input record; then, for each rule, if the input record satisfies the rule’s pattern, awk executes the rule’s action. Otherwise, the rule does nothing for that input record.

Rvalue A value that can appear on the right side of an assignment operator. In awk, essentially every expression has a value. These values are rvalues.

sed See “Stream Editor.”

Short-Circuit The nature of the awk logical operators ‘&&’ and ‘||’. If the value of the entire expression can be deduced from evaluating just the left-hand side of these operators, the right-hand side will not be evaluated (see Section 7.11 [Boolean Expressions], page 91).

Side Effect A side effect occurs when an expression has an effect aside from merely producing a value. Assignment expressions, increment and decrement expressions and function calls have side effects. See Section 7.7 [Assignment Expressions], page 84.

Single Precision An internal representation of numbers that can have fractional parts. Single precision numbers keep track of fewer digits than do double precision numbers, but operations on them are less expensive in terms of CPU time. This is the type used by some very old versions of awk to store numeric values. It is the C type

float.

Space The character generated by hitting the space bar on the keyboard.

Special File A file name interpreted internally by gawk, instead of being handed directly to the underlying operating system. For example, ‘/dev/stderr’. See Section 6.7 [Special File Names in gawk], page 72.

Stream Editor A program that reads records from an input stream and processes them one or more at a time. This is in contrast with batch programs, which may expect to read their input files in entirety before starting to do anything, and with interactive programs, which require input from the user.

String A datum consisting of a sequence of characters, such as ‘I am a string’. Constant strings are written with double-quotes in the awk language, and may contain escape sequences. See Section 4.2 [Escape Sequences], page 24.

Tab The character generated by hitting the TAB key on the keyboard. It usually expands to up to eight spaces upon output.

Unix A computer operating system originally developed in the early 1970’s at AT&T Bell Laboratories. It initially became popular in universities around the world, and later moved into commercial evnironments as a software development system and network server system. There are many commercial versions of Unix, as well as several work-alike systems whose source code is freely available (such as Linux, NetBSD, and FreeBSD).

Whitespace A sequence of space, tab, or newline characters occurring inside an input record or a string.

GNU GENERAL PUBLIC LICENSE 307

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 59 Temple Place — Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating

that you changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in

whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License. c. If the modified program normally reads commands interactively

when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable

source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for non-commercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EX-TENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUAL-ITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR COR-RECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDIS-TRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUD-ING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPER-ATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

     Copyright (C) 19yy name of author
     This program is free software; you can redistribute it and/or
     modify it under the terms of the GNU General Public License
     as published by the Free Software Foundation; either version 2
     of the License, or (at your option) any later version.
     This program is distributed in the hope that it will be useful,
     but WITHOUT ANY WARRANTY; without even the implied warranty of
     MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
     GNU General Public License for more details.
     You should have received a copy of the GNU General Public License
     along with this program; if not, write to the Free Software
     Foundation, Inc., 59 Temple Place --- Suite 330, Boston, MA 02111-1307, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

     Gnomovision version 69, Copyright (C) 19yy name of author
     Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
     type `show w'.  This is free software, and you are welcome
     to redistribute it under certain conditions; type `show c'
     for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

     Yoyodyne, Inc., hereby disclaims all copyright
     interest in the program `Gnomovision'
     (which makes passes at compilers) written
     by James Hacker.

signature of Ty Coon, 1 April 1989

     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Index 315

Index

!

! operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 != operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 !~ operator . . . . . . . . . . . . . 23, 33, 35, 77, 89

#

# (comment) . . . . . . . . . . . . . . . . . . . . . . . . . . 13 #! (executable scripts) . . . . . . . . . . . . . . . . 12

$

$ (field operator) . . . . . . . . . . . . . . . . . . . . . . 40

&

&& operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

-

--assign option . . . . . . . . . . . . . . . . . . . . . 159 --compat option . . . . . . . . . . . . . . . . . . . . . 160 --copyleft option . . . . . . . . . . . . . . . . . . . 160 --copyright option . . . . . . . . . . . . . . . . . . 160 --field-separator option . . . . . . . . . . . 159 --file option. . . . . . . . . . . . . . . . . . . . . . . . 159 --help option. . . . . . . . . . . . . . . . . . . . . . . . 161 --lint option. . . . . . . . . . . . . . . . . . . . . . . . 161 --lint-old option . . . . . . . . . . . . . . . . . . . 161 --posix option . . . . . . . . . . . . . . . . . . . . . . 161 --source option . . . . . . . . . . . . . . . . . . . . . 162 --traditional option. . . . . . . . . . . . . . . . 160 --usage option . . . . . . . . . . . . . . . . . . . . . . 161 --version option . . . . . . . . . . . . . . . . . . . . 162 -f option. . . . . . . . . . . . . . . . . . . . . . . . . 11, 159 -F option. . . . . . . . . . . . . . . . . . . . . . . . . 47, 159 -v option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 -W option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160

/

‘/dev/fd’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ‘/dev/pgrpid’ . . . . . . . . . . . . . . . . . . . . . . . . . 73 ‘/dev/pid’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 ‘/dev/ppid’ . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 ‘/dev/stderr’ . . . . . . . . . . . . . . . . . . . . . . . . . 72 ‘/dev/stdin’ . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ‘/dev/stdout’ . . . . . . . . . . . . . . . . . . . . . . . . . 72 ‘/dev/user’ . . . . . . . . . . . . . . . . . . . . . . 73, 190

=

== operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

_gr_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 _pw_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 _tm_addup . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 _tm_isleap . . . . . . . . . . . . . . . . . . . . . . . . . . 176

                         |

|| operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

                         ~

~ operator . . . . . . . . . . . . . . 23, 33, 35, 77, 89

                         >

> operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 >= operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

                         \

\' regexp operator . . . . . . . . . . . . . . . . . . . . 32 \` regexp operator . . . . . . . . . . . . . . . . . . . . 32 \> regexp operator . . . . . . . . . . . . . . . . . . . . 32 \< regexp operator . . . . . . . . . . . . . . . . . . . . 32 \B regexp operator . . . . . . . . . . . . . . . . . . . . 32 \w regexp operator . . . . . . . . . . . . . . . . . . . . 32 \W regexp operator . . . . . . . . . . . . . . . . . . . . 32 \y regexp operator . . . . . . . . . . . . . . . . . . . . 32

                         <

< operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 <= operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

A

accessing fields . . . . . . . . . . . . . . . . . . . . . . . . 40 account information . . . . . . . . . . . . . 190, 195 acronym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 action, curly braces . . . . . . . . . . . . . . . . . . 102 action, default . . . . . . . . . . . . . . . . . . . . . . . . 14 action, definition of . . . . . . . . . . . . . . . . . . 102 action, empty . . . . . . . . . . . . . . . . . . . . . . . . . 14 action, separating statements. . . . . . . . . 102 adding new features . . . . . . . . . . . . . . . . . . 293 addition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Aho, Alfred . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 AI programming, using gawk . . . . . . . . . 280 alarm.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 amiga. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 anchors in regexps . . . . . . . . . . . . . . . . . . . . 27 and operator . . . . . . . . . . . . . . . . . . . . . . . . . . 91 anonymous ftp. . . . . . . . . . . . . . . . . . 277, 291 applications of awk . . . . . . . . . . . . . . . . . . . . 18 ARGC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 ARGIND . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 163 argument processing . . . . . . . . . . . . . . . . . 184 arguments in function call . . . . . . . . . . . . . 93 arguments, command line . . . . . . . . . . . . 159 ARGV . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 163 arithmetic operators. . . . . . . . . . . . . . . . . . . 82 array assignment . . . . . . . . . . . . . . . . . . . . . 125 array reference . . . . . . . . . . . . . . . . . . . . . . . 124 Array subscripts and IGNORECASE. . . . . 124 array subscripts, uninitialized variables

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 arrays, associative. . . . . . . . . . . . . . . . . . . . 123 arrays, definition of . . . . . . . . . . . . . . . . . . 123 arrays, deleting an element . . . . . . . . . . . 128 arrays, deleting entire contents . . . . . . . 128 arrays, multi-dimensional subscripts . . 130 arrays, presence of elements . . . . . . . . . . 125 arrays, sparse . . . . . . . . . . . . . . . . . . . . . . . . 124 arrays, special for statement . . . . . . . . . 127 arrays, the in operator . . . . . . . . . . . . . . . 125 artificial intelligence, using gawk . . . . . . 280 ASCII . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 assert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 assert, C version . . . . . . . . . . . . . . . . . . . . 170 assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 assignment operators . . . . . . . . . . . . . . . . . . 84 assignment to fields . . . . . . . . . . . . . . . . . . . 42 associative arrays . . . . . . . . . . . . . . . . . . . . 123 atan2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 atari . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 automatic initialization. . . . . . . . . . . . . . . . 16 awk language, POSIX version . . 25, 27, 28, 31, 48, 64, 68, 82, 86, 95, 96, 109, 110, 111, 115, 136, 142, 152

awk language, V.4 version . . . . . 24, 25, 252 AWKPATH environment variable . . . . . . . . 164 awksed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

B

backslash continuation. . . . . . . . . . . . 16, 212 backslash continuation and comments . . 17 backslash continuation in csh . . . . . . 15, 17 basic function of awk . . . . . . . . . . . . . . . . . . . 9 ‘BBS-list’ file. . . . . . . . . . . . . . . . . . . . . . . . . . 7 BEGIN special pattern . . . . . . . . . . . . . . . . 100 beginfile . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 body of a loop . . . . . . . . . . . . . . . . . . . . . . . 106 book, using this . . . . . . . . . . . . . . . . . . . . . . . . 5 boolean expressions . . . . . . . . . . . . . . . . . . . 91 boolean operators . . . . . . . . . . . . . . . . . . . . . 91 break statement . . . . . . . . . . . . . . . . . . . . . 109 break, outside of loops . . . . . . . . . . . . . . . 109 Brennan, Michael. . . . . . . . . . . 128, 240, 291 buffer matching operators . . . . . . . . . . . . . 32 buffering output. . . . . . . . . . . . . . . . . 144, 146 buffering, interactive vs. non-interactive

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 buffering, non-interactive vs. interactive

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 buffers, flushing . . . . . . . . . . . . . . . . . 144, 146 bugs, known in gawk . . . . . . . . . . . . . . . . . 165 built-in functions. . . . . . . . . . . . . . . . . . . . . 133 built-in variables . . . . . . . . . . . . . . . . . . . . . 115 built-in variables, convey information

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 built-in variables, user modifiable. . . . . 115

C

call by reference . . . . . . . . . . . . . . . . . 154, 155 call by value . . . . . . . . . . . . . . . . . . . . . . . . . 154 calling a function . . . . . . . . . . . . . . . . . 93, 154 case conversion. . . . . . . . . . . . . . . . . . . . . . . 141 case sensitivity . . . . . . . . . . . . . . . . . . . . . . . . 33 changing contents of a field . . . . . . . . . . . . 42 changing the record separator . . . . . . . . . 37 character classes . . . . . . . . . . . . . . . . . . . . . . 28 character encodings . . . . . . . . . . . . . . . . . . 173 character list . . . . . . . . . . . . . . . . . . . . . . . . . . 27 character list, complemented . . . . . . . . . . 29 character sets . . . . . . . . . . . . . . . . . . . . . . . . 173 chr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 144 closing input files and pipes . . . . . . . . . . . 74 closing output files and pipes . . . . . . . . . . 74 coding style used in gawk. . . . . . . . . . . . . 294 collating elements . . . . . . . . . . . . . . . . . . . . . 29 collating symbols. . . . . . . . . . . . . . . . . . . . . . 29 command line . . . . . . . . . . . . . . . . . . . . . . . . 159 command line formats. . . . . . . . . . . . . . . . . 10 command line, setting FS on. . . . . . . . . . . 47 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 comments and backslash continuation . . 17 common mistakes . . . . . . . . . . 35, 44, 62, 90 comp.lang.awk. . . . . . . . . . . . . . . . . . . . . . . 290 comparison expressions . . . . . . . . . . . . . . . . 88 comparisons, string vs. regexp. . . . . . . . . 90 compatibility mode. . . . . . . . . . . . . . 160, 254 complemented character list . . . . . . . . . . . 29 compound statement . . . . . . . . . . . . . . . . . 105 computed regular expressions. . . . . . . . . . 35 concatenation . . . . . . . . . . . . . . . . . . . . . . . . . 83 conditional expression . . . . . . . . . . . . . . . . . 92 configuring gawk . . . . . . . . . . . . . . . . . . . . . 283 constants, types of . . . . . . . . . . . . . . . . . . . . 77 continuation of lines. . . . . . . . . . . . . . . . . . . 16 continue statement . . . . . . . . . . . . . . . . . . 110 continue, outside of loops. . . . . . . . . . . . 110 control statement . . . . . . . . . . . . . . . . . . . . 105 conversion of case . . . . . . . . . . . . . . . . . . . . 141 conversion of strings and numbers . . . . . 81 conversions, during subscripting . . . . . . 129 converting dates to timestamps. . . . . . . 175 CONVFMT . . . . . . . . . . . . . . . . . . . . . 81, 115, 129 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 csh, backslash continuation . . . . . . . . 15, 17 curly braces. . . . . . . . . . . . . . . . . . . . . . . . . . 102 custom.h configuration file . . . . . . . . . . . 283 cut utility . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 cut.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

D

d.c., see “dark corner”. . . . . . . . . . . . . . . . . . 6 dark corner . . 6, 26, 39, 47, 48, 51, 55, 59, 64, 66, 67, 78, 80, 81, 88, 101, 109, 110, 113, 118, 119, 130, 137, 159, 163 data-driven languages . . . . . . . . . . . . . . . . . . 9 dates, converting to timestamps . . . . . . 175 decrement operators. . . . . . . . . . . . . . . . . . . 87 default action . . . . . . . . . . . . . . . . . . . . . . . . . 14 default pattern . . . . . . . . . . . . . . . . . . . . . . . . 14 defining functions . . . . . . . . . . . . . . . . . . . . 151 Deifik, Scott . . . . . . . . . . . . . . . . . . . . . . . 4, 291 delete statement . . . . . . . . . . . . . . . . . . . . 128 deleting elements of arrays . . . . . . . . . . . 128 deleting entire arrays. . . . . . . . . . . . . . . . . 128 deprecated features . . . . . . . . . . . . . . . . . . 165 deprecated options . . . . . . . . . . . . . . . . . . . 165 differences between gawk and awk . . 33, 39, 46, 54, 59, 71, 72, 75, 77, 78, 83, 93, 102, 112, 121, 128, 133, 137, 140, 164 directory search . . . . . . . . . . . . . . . . . . . . . . 164 division . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 documenting awk programs . . . . . . . 13, 200 dupword.awk . . . . . . . . . . . . . . . . . . . . . . . . . 226 dynamic regular expressions . . . . . . . . . . . 35

E

EBCDIC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 egrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 28 egrep utility . . . . . . . . . . . . . . . . . . . . . . . . . 208 egrep.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 element assignment . . . . . . . . . . . . . . . . . . 125 element of array . . . . . . . . . . . . . . . . . . . . . 124 empty action . . . . . . . . . . . . . . . . . . . . . . . . . . 14 empty pattern . . . . . . . . . . . . . . . . . . . . . . . 102 empty program . . . . . . . . . . . . . . . . . . . . . . 159 empty string . . . . . . . . . . . . . . . 39, 46, 81, 88 END special pattern . . . . . . . . . . . . . . . . . . . 100 endfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 endgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 endpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 ENVIRON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 environment variable, AWKPATH . . . . . . . 164 environment variable, POSIXLY_CORRECT

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 equivalence classes . . . . . . . . . . . . . . . . . . . . 29 ERRNO. . . . . . . . . . . . . . . . . . . . . . . . . 54, 75, 118 errors, common . . . . . . . . . . . . 35, 44, 62, 90 escape processing, sub et. al. . . . . . . . . . 141 escape sequence notation . . . . . . . . . . . . . . 24 evaluation, order of . . . . . . . . . . . . . . . . . . 133 examining fields . . . . . . . . . . . . . . . . . . . . . . . 40 executable scripts . . . . . . . . . . . . . . . . . . . . . 12 exit statement . . . . . . . . . . . . . . . . . . . . . . 112 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 explicit input . . . . . . . . . . . . . . . . . . . . . . . . . 53 exponentiation . . . . . . . . . . . . . . . . . . . . . . . . 82 expression. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 expression, assignment . . . . . . . . . . . . . . . . 84 expression, boolean. . . . . . . . . . . . . . . . . . . . 91 expression, comparison . . . . . . . . . . . . . . . . 88 expression, conditional . . . . . . . . . . . . . . . . 92 expression, matching . . . . . . . . . . . . . . . . . . 88 extract.awk . . . . . . . . . . . . . . . . . . . . . . . . . 237

F

features, adding . . . . . . . . . . . . . . . . . . . . . . 293 fflush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 field operator $. . . . . . . . . . . . . . . . . . . . . . . . 40 field separator, choice of . . . . . . . . . . . . . . . 45 field separator, FS . . . . . . . . . . . . . . . . . . . . . 44 field separator, on command line . . . . . . 47 field, changing contents of . . . . . . . . . . . . . 42 fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 fields, separating . . . . . . . . . . . . . . . . . . . . . . 44 FIELDWIDTHS . . . . . . . . . . . . . . . . . . . . . . . . . 115 file descriptors . . . . . . . . . . . . . . . . . . . . . . . . 72 file, awk program . . . . . . . . . . . . . . . . . . . . . . 11 FILENAME . . . . . . . . . . . . . . . . . . . . . 37, 59, 118 FILENAME, being set by getline . . . . . . . 59 Fish, Fred . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 flushing buffers . . . . . . . . . . . . . . . . . . 144, 146 FNR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 118 for (x in …) . . . . . . . . . . . . . . . . . . . . . . . 127 for statement. . . . . . . . . . . . . . . . . . . . . . . . 107 format specifier . . . . . . . . . . . . . . . . . . . . . . . 65 format string. . . . . . . . . . . . . . . . . . . . . . . . . . 65 format, numeric output . . . . . . . . . . . . . . . 64 formatted output. . . . . . . . . . . . . . . . . . . . . . 64 formatted timestamps . . . . . . . . . . . . . . . . 181 Free Software Foundation . . . . . . . . . . 1, 277 FreeBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Friedl, Jeffrey . . . . . . . . . . . . . . . . . . . . . . . . . . 4 FS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 115 ftp, anonymous . . . . . . . . . . . . . . . . . 277, 291 function call. . . . . . . . . . . . . . . . . . . . . . 93, 154 function definition . . . . . . . . . . . . . . . . . . . 151 function, recursive . . . . . . . . . . . . . . . . . . . 152 functions, undefined. . . . . . . . . . . . . . . . . . 155 functions, user-defined . . . . . . . . . . . . . . . 151

G

gawk coding style . . . . . . . . . . . . . . . . . . . . 294 gensub. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 getgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 getgrent, C version. . . . . . . . . . . . . . . . . . 195 getgrgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 getgrnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 getgruser . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 getline. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 getline, return values . . . . . . . . . . . . . . . . 54 getline, setting FILENAME. . . . . . . . . . . . . 59 getopt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 getopt, C version . . . . . . . . . . . . . . . . . . . . 184 getpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 getpwent, C version. . . . . . . . . . . . . . . . . . 190 getpwnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 getpwuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 gettimeofday . . . . . . . . . . . . . . . . . . . . . . . . 181 getting gawk . . . . . . . . . . . . . . . . . . . . . . . . . 277 GNU Project. . . . . . . . . . . . . . . . . . . . . . . . . . . 2 grcat program . . . . . . . . . . . . . . . . . . . . . . . 195 grcat.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 group file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 group information. . . . . . . . . . . . . . . . . . . . 195 gsub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 gsub, third argument of . . . . . . . . . . . . . . 139

H

Hankerson, Darrel . . . . . . . . . . . . . . . . . 4, 291 historical features . . 47, 109, 110, 136, 275 history of awk . . . . . . . . . . . . . . . . . . . . . . . . . . 1 histsort.awk . . . . . . . . . . . . . . . . . . . . . . . . 236 how awk works . . . . . . . . . . . . . . . . . . . . . . . . 14 Hughes, Phil . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

I

I/O from BEGIN and END . . . . . . . . . . . . . 102 id utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 id.awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 if-else statement . . . . . . . . . . . . . . . . . . . 105 igawk.sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 IGNORECASE. . . . . . . . . . . . . . . . . . 33, 116, 124 IGNORECASE and array subscripts . . . . . 124 ignoring case . . . . . . . . . . . . . . . . . . . . . . . . . . 33 implementation limits . . . . . . . . . . . . . 59, 71 in operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 increment operators . . . . . . . . . . . . . . . . . . . 87 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 initialization, automatic . . . . . . . . . . . . . . . 16 input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 input file, sample . . . . . . . . . . . . . . . . . . . . . . . 7 input files, skipping . . . . . . . . . . . . . . . . . . 168 input pipeline . . . . . . . . . . . . . . . . . . . . . . . . . 57 input redirection . . . . . . . . . . . . . . . . . . . . . . 56 input, explicit . . . . . . . . . . . . . . . . . . . . . . . . . 53 input, getline command . . . . . . . . . . . . . . 53 input, multiple line records . . . . . . . . . . . . 51 input, standard . . . . . . . . . . . . . . . . . . . . . . . 11 installation, amiga . . . . . . . . . . . . . . . . . . . 289 installation, atari . . . . . . . . . . . . . . . . . . . . 287 installation, MS-DOS and OS/2 . . . . . . 286 installation, unix . . . . . . . . . . . . . . . . . . . . . 282 installation, vms . . . . . . . . . . . . . . . . . . . . . 283 int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 interaction, awk and other programs . . 144 interactive buffering vs. non-interactive

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 interval expressions. . . . . . . . . . . . . . . . . . . . 31 ‘inventory-shipped’ file . . . . . . . . . . . . . . . 7 invocation of gawk . . . . . . . . . . . . . . . . . . . 159 ISO 8601 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 ISO 8859-1 . . . . . . . . . . . . . . . . . . . . . . . 34, 300 ISO Latin-1 . . . . . . . . . . . . . . . . . . . . . . 34, 300

J

Jaegermann, Michal . . . . . . . . . . . . . . . 4, 291 join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

K

Kernighan, Brian . . . . . . . 1, 4, 83, 253, 291 known bugs . . . . . . . . . . . . . . . . . . . . . . . . . . 165

L

labels.awk . . . . . . . . . . . . . . . . . . . . . . . . . . 232 language, awk . . . . . . . . . . . . . . . . . . . . . . . . . . 5 language, data-driven . . . . . . . . . . . . . . . . . . 9 language, procedural . . . . . . . . . . . . . . . . . . . 9 leftmost longest match. . . . . . . . . . . . . 34, 51 length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 limitations . . . . . . . . . . . . . . . . . . . . . . . . 59, 71 line break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 line continuation . . . . . . . . . . . 16, 63, 92, 93 Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 288 locale, definition of . . . . . . . . . . . . . . . . . . . 148 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 logical false . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 logical operations . . . . . . . . . . . . . . . . . . . . . 91 logical true. . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 login information. . . . . . . . . . . . . . . . . . . . . 190 long options . . . . . . . . . . . . . . . . . . . . . . . . . 159 loop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 loops, exiting . . . . . . . . . . . . . . . . . . . . . . . . 109 lvalue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

M

mark parity . . . . . . . . . . . . . . . . . . . . . . . . . . 173 match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 matching ranges of lines . . . . . . . . . . . . . . . 99 matching, leftmost longest . . . . . . . . . 34, 51 mawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 merging strings . . . . . . . . . . . . . . . . . . . . . . 174 metacharacters . . . . . . . . . . . . . . . . . . . . . . . . 26 mistakes, common . . . . . . . . . . 35, 44, 62, 90 mktime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 modifiers (in format specifiers) . . . . . . . . 66 multi-dimensional subscripts . . . . . . . . . 130 multiple line records. . . . . . . . . . . . . . . . . . . 51 multiple passes over data . . . . . . . . . . . . . 164 multiple statements on one line . . . . . . . . 18 multiplication . . . . . . . . . . . . . . . . . . . . . . . . . 82

N

names, use of . . . . . . . . . . . . . . . . . . . . . . . . 151 namespace issues in awk . . . . . . . . . . . . . . 200 namespaces . . . . . . . . . . . . . . . . . . . . . . . . . . 151 NetBSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 new awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 new awk vs. old awk . . . . . . . . . . . . . . . . . . . . 9 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 next file statement . . . . . . . . . . . . . . . . . 112 next statement . . . . . . . . . . . . . . . . . . . . . . 111 next, inside a user-defined function… 111 nextfile function . . . . . . . . . . . . . . . . . . . 168 nextfile statement . . . . . . . . . . . . . . . . . . 112 NF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 118 non-interactive buffering vs. interactive

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 not operator . . . . . . . . . . . . . . . . . . . . . . . . . . 91 NR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 119 null string . . . . . . . . . . . . . . . . . . . . . 46, 81, 88 null string, as array subscript . . . . . . . . 130 number of fields, NF . . . . . . . . . . . . . . . . . . . 40 number of records, NR, FNR . . . . . . . . . . . . 40 numbers, used as subscripts . . . . . . . . . . 129 numeric character values . . . . . . . . . . . . . 172 numeric constant . . . . . . . . . . . . . . . . . . . . . . 77 numeric output format . . . . . . . . . . . . . . . . 64 numeric string . . . . . . . . . . . . . . . . . . . . . . . . 88 numeric value . . . . . . . . . . . . . . . . . . . . . . . . . 77

O

obsolete features . . . . . . . . . . . . . . . . . . . . . 165 obsolete options . . . . . . . . . . . . . . . . . . . . . . 165 OFMT. . . . . . . . . . . . . . . . . . . . . . . . . . 64, 82, 116 OFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 116 old awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 old awk vs. new awk . . . . . . . . . . . . . . . . . . . . 9 one-liners . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 operations, logical . . . . . . . . . . . . . . . . . . . . . 91 operator precedence . . . . . . . . . . . . . . . . . . . 94 operators, arithmetic . . . . . . . . . . . . . . . . . . 82 operators, assignment . . . . . . . . . . . . . . . . . 84 operators, boolean . . . . . . . . . . . . . . . . . . . . 91 operators, decrement . . . . . . . . . . . . . . . . . . 87 operators, increment . . . . . . . . . . . . . . . . . . 87 operators, regexp matching . . . . . . . . . . . . 23 operators, relational . . . . . . . . . . . . . . . 88, 89 operators, short-circuit . . . . . . . . . . . . . . . . 91 operators, string . . . . . . . . . . . . . . . . . . . . . . 83 operators, string-matching. . . . . . . . . . . . . 23 options, command line . . . . . . . . . . . . . . . 159 options, long . . . . . . . . . . . . . . . . . . . . . . . . . 159 or operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 ord . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 order of evaluation . . . . . . . . . . . . . . . . . . . 133 ORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 116 output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 output field separator, OFS . . . . . . . . . . . . 63 output format specifier, OFMT . . . . . . . . . . 64 output record separator, ORS . . . . . . . . . . 63 output redirection . . . . . . . . . . . . . . . . . . . . . 70 output, buffering . . . . . . . . . . . . . . . . 144, 146 output, formatted . . . . . . . . . . . . . . . . . . . . . 64 output, piping . . . . . . . . . . . . . . . . . . . . . . . . 71

P

passes, multiple . . . . . . . . . . . . . . . . . . . . . . 164 password file . . . . . . . . . . . . . . . . . . . . . . . . . 190 path, search. . . . . . . . . . . . . . . . . . . . . . . . . . 164 pattern, BEGIN . . . . . . . . . . . . . . . . . . . . . . . 100 pattern, default . . . . . . . . . . . . . . . . . . . . . . . 14 pattern, definition of . . . . . . . . . . . . . . . . . . 97 pattern, empty . . . . . . . . . . . . . . . . . . . . . . . 102 pattern, END . . . . . . . . . . . . . . . . . . . . . . . . . 100 pattern, range. . . . . . . . . . . . . . . . . . . . . . . . . 99 pattern, regular expressions . . . . . . . . . . . 23 patterns, types of . . . . . . . . . . . . . . . . . . . . . 97 per file initialization and clean-up …. 183 PERL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 pipeline, input . . . . . . . . . . . . . . . . . . . . . . . . 57 pipes for output . . . . . . . . . . . . . . . . . . . . . . . 71 portability issues. . . . . 17, 26, 75, 128, 136,

144, 152, 167 porting gawk . . . . . . . . . . . . . . . . . . . . . . . . . 295 POSIX awk . . . . . 25, 27, 28, 31, 48, 64, 68,

82, 86, 95, 96, 109, 110, 111, 115, 136, 142, 152

POSIX mode . . . . . . . . . . . . . . . . . . . . . . . . 161 POSIXLY_CORRECT environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 precedence . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 precedence, regexp operators . . . . . . . . . . 31 print statement . . . . . . . . . . . . . . . . . . . . . . 61 printf statement, syntax of . . . . . . . . . . . 65 printf, format-control characters. . . . . . 65 printf, modifiers . . . . . . . . . . . . . . . . . . . . . 66 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 procedural languages . . . . . . . . . . . . . . . . . . . 9 process information . . . . . . . . . . . . . . . . . . . 73 processing arguments . . . . . . . . . . . . . . . . 184 program file . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 program, awk. . . . . . . . . . . . . . . . . . . . . . . . . . . 5 program, definition of . . . . . . . . . . . . . . . . . . 9 program, self contained. . . . . . . . . . . . . . . . 12 programs, documenting . . . . . . . . . . . 13, 200 pwcat program . . . . . . . . . . . . . . . . . . . . . . . 191 pwcat.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

Q

quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 quoting, shell . . . . . . . . . . . . . . . . . . . . . . 11, 12

R

Rakitzis, Byron . . . . . . . . . . . . . . . . . . . . . . 236 rand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 random numbers, seed of . . . . . . . . . . . . . 135 range pattern . . . . . . . . . . . . . . . . . . . . . . . . . 99 Rankin, Pat . . . . . . . . . . . . . . . . . . . 4, 85, 291 reading files . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 reading files, getline command . . . . . . . 53 reading files, multiple line records . . . . . 51 record separator, RS . . . . . . . . . . . . . . . . . . . 37 record terminator, RT. . . . . . . . . . . . . . . . . . 39 record, definition of . . . . . . . . . . . . . . . . . . . 37 records, multiple line . . . . . . . . . . . . . . . . . . 51 recursive function . . . . . . . . . . . . . . . . . . . . 152 redirection of input. . . . . . . . . . . . . . . . . . . . 56 redirection of output . . . . . . . . . . . . . . . . . . 70 reference to array . . . . . . . . . . . . . . . . . . . . 124 regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 regexp as expression. . . . . . . . . . . . . . . . . . . 91 regexp comparison vs. string comparison

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 regexp constant . . . . . . . . . . . . . . . . . . . . . . . 24 regexp constants, difference between

slashes and quotes . . . . . . . . . . . . . . . . 35 regexp match/non-match operators … 23,

88 regexp matching operators . . . . . . . . . . . . 23 regexp operators . . . . . . . . . . . . . . . . . . . . . . 26 regexp operators, GNU specific . . . . . . . . 31 regexp operators, precedence of. . . . . . . . 31 regexp, anchors . . . . . . . . . . . . . . . . . . . . . . . 27 regexp, dynamic . . . . . . . . . . . . . . . . . . . . . . 35 regexp, effect of command line options

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 regular expression . . . . . . . . . . . . . . . . . . . . . 23 regular expression metacharacters . . . . . 26 regular expressions as field separators . . 45 regular expressions as patterns . . . . . . . . 23 regular expressions as record separators

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 regular expressions, computed . . . . . . . . . 35 relational operators . . . . . . . . . . . . . . . . 88, 89 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 removing elements of arrays . . . . . . . . . . 128 return statement . . . . . . . . . . . . . . . . . . . . 156 RFC-1036. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 RFC-822. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 RLENGTH . . . . . . . . . . . . . . . . . . . . . . . . . 119, 136 Robbins, Miriam . . . . . . . . . . . . . . . . . . . . . . . 4 Rommel, Kai Uwe . . . . . . . . . . . . . . . . . 4, 291 round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 rounding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 RS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37, 116 RSTART . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 136 RT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39, 53, 119 rule, definition of . . . . . . . . . . . . . . . . . . . . . . . 9 running awk programs . . . . . . . . . . . . . . . . . 10 running long programs . . . . . . . . . . . . . . . . 11 rvalue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

S

sample input file . . . . . . . . . . . . . . . . . . . . . . . 7 scanning an array . . . . . . . . . . . . . . . . . . . . 127 script, definition of . . . . . . . . . . . . . . . . . . . . . 9 scripts, executable. . . . . . . . . . . . . . . . . . . . . 12 scripts, shell . . . . . . . . . . . . . . . . . . . . . . . . . . 13 search path . . . . . . . . . . . . . . . . . . . . . . . . . . 164 search path, for source files. . . . . . . . . . . 164 sed utility . . . . . . . . . . . . . . . . . . . 48, 240, 244 seed for random numbers. . . . . . . . . . . . . 135 self contained programs . . . . . . . . . . . . . . . 12 shell quoting . . . . . . . . . . . . . . . . . . . . . . 11, 12 shell scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 short-circuit operators. . . . . . . . . . . . . . . . . 91 side effect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 simple stream editor . . . . . . . . . . . . . . . . . 240 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 single character fields. . . . . . . . . . . . . . . . . . 46 single quotes, why needed . . . . . . . . . . . . . 10 skipping input files . . . . . . . . . . . . . . . . . . . 168 skipping lines between markers . . . . . . . 100 sparse arrays . . . . . . . . . . . . . . . . . . . . . . . . . 124 split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 split utility . . . . . . . . . . . . . . . . . . . . . . . . . 215 split.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 sprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 srand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Stallman, Richard . . . . . . . . . . . . . . . . . . . 1, 3 standard error output . . . . . . . . . . . . . . . . . 72 standard input . . . . . . . . . . . . . . . . . 11, 37, 72 standard output. . . . . . . . . . . . . . . . . . . . . . . 72 statement, compound . . . . . . . . . . . . . . . . 105 stream editor . . . . . . . . . . . . . . . . . . . . . . . . . 48 stream editor, simple . . . . . . . . . . . . . . . . . 240 strftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 string comparison vs. regexp comparison

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 string constants . . . . . . . . . . . . . . . . . . . . . . . 77 string operators . . . . . . . . . . . . . . . . . . . . . . . 83 string-matching operators . . . . . . . . . . . . . 23 sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 sub, third argument of . . . . . . . . . . . . . . . 139 subscripts in arrays . . . . . . . . . . . . . . . . . . 130 SUBSEP . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 130 substr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 subtraction . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 systime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

T

Tcl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 tee utility . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 tee.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 terminator, record. . . . . . . . . . . . . . . . . . . . . 39 time of day . . . . . . . . . . . . . . . . . . . . . . . . . . 146 timestamps . . . . . . . . . . . . . . . . . . . . . . . . . . 146 timestamps, converting from dates … 175 timestamps, formatted . . . . . . . . . . . . . . . 181 tolower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 toupper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 translate.awk. . . . . . . . . . . . . . . . . . . . . . . 230 Trueman, David . . . . . . . . . . . . . . . . . . . . . . . . 4 truth values . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 type conversion . . . . . . . . . . . . . . . . . . . . . . . 81 types of variables . . . . . . . . . . . . . . . . . . 84, 88

U

undefined functions . . . . . . . . . . . . . . . . . . 155 undocumented features. . . . . . . . . . . . . . . 165 uninitialized variables, as array subscripts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 uniq utility . . . . . . . . . . . . . . . . . . . . . . . . . . 218 uniq.awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 use of comments . . . . . . . . . . . . . . . . . . . . . . 13 user information . . . . . . . . . . . . . . . . . . . . . 190 user-defined functions . . . . . . . . . . . . . . . . 151 user-defined variables. . . . . . . . . . . . . . . . . . 79 uses of awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 using this book. . . . . . . . . . . . . . . . . . . . . . . . . 5

V

values of characters as numbers . . . . . . 172 variable shadowing . . . . . . . . . . . . . . . . . . . 152 variable typing . . . . . . . . . . . . . . . . . . . . . . . . 88 variables, user-defined . . . . . . . . . . . . . . . . . 79

W

Wall, Larry . . . . . . . . . . . . . . . . . . . . . . . . . . 297 wc utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 wc.awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 Weinberger, Peter . . . . . . . . . . . . . . . . . . . . . . 1 when to use awk. . . . . . . . . . . . . . . . . . . . . . . 18 while statement . . . . . . . . . . . . . . . . . . . . . 106 word boundaries, matching . . . . . . . . . . . . 32 word, regexp definition of . . . . . . . . . . . . . 31 wordfreq.sh . . . . . . . . . . . . . . . . . . . . . . . . . 235

324 Effective AWK Programming

Short Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 Getting Started with awk . . . . . . . . . . . . . . . . . . . . . . . . . 9 3 Useful One Line Programs . . . . . . . . . . . . . . . . . . . . . . . . 21 4 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5 Reading Input Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6 Printing Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 8 Patterns and Actions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 9 Control Statements in Actions. . . . . . . . . . . . . . . . . . . . . 105 10 Built-in Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 11 Arrays in awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12 Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 13 User-defined Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 151 14 Running awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 15 A Library of awk Functions . . . . . . . . . . . . . . . . . . . . . . 167 16 Practical awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . 203 17 The Evolution of the awk Language. . . . . . . . . . . . . . . . . 251 Appendix A gawk Summary . . . . . . . . . . . . . . . . . . . . . . . . 257 Appendix B Installing gawk . . . . . . . . . . . . . . . . . . . . . . . . 277 Appendix C Implementation Notes. . . . . . . . . . . . . . . . . . . . 293 Appendix D Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 GNU GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . . 307 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

ii Effective AWK Programming

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

History of awk and gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 The GNU Project and This Book . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.1 Using This Book . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Dark Corners . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.2 Typographical Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3 Data Files for the Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2 Getting Started with awk . . . . . . . . . . . . . . . . . . . 9

2.1 A Rose By Any Other Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2 How to Run awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.2.1 One-shot Throw-away awk Programs . . . . . . . . . . . . 10 2.2.2 Running awk without Input Files . . . . . . . . . . . . . . . 11 2.2.3 Running Long Programs . . . . . . . . . . . . . . . . . . . . . . . 11 2.2.4 Executable awk Programs . . . . . . . . . . . . . . . . . . . . . . 12 2.2.5 Comments in awk Programs . . . . . . . . . . . . . . . . . . . . 13

2.3 A Very Simple Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.4 An Example with Two Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5 A More Complex Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6 awk Statements Versus Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.7 Other Features of awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.8 When to Use awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3 Useful One Line Programs . . . . . . . . . . . . . . . . 21

4 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . 23

4.1 How to Use Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . 23 4.2 Escape Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.3 Regular Expression Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4 Additional Regexp Operators Only in gawk. . . . . . . . . . . . . . . 31 4.5 Case-sensitivity in Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.6 How Much Text Matches? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.7 Using Dynamic Regexps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 iv Effective AWK Programming

5 Reading Input Files . . . . . . . . . . . . . . . . . . . . . . . 37

5.1 How Input is Split into Records . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.2 Examining Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.3 Non-constant Field Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.4 Changing the Contents of a Field . . . . . . . . . . . . . . . . . . . . . . . . 42 5.5 Specifying How Fields are Separated . . . . . . . . . . . . . . . . . . . . . 44

5.5.1 The Basics of Field Separating. . . . . . . . . . . . . . . . . . 44 5.5.2 Using Regular Expressions to Separate Fields …. 45 5.5.3 Making Each Character a Separate Field . . . . . . . . 46 5.5.4 Setting FS from the Command Line . . . . . . . . . . . . . 47 5.5.5 Field Splitting Summary . . . . . . . . . . . . . . . . . . . . . . . 48

5.6 Reading Fixed-width Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.7 Multiple-Line Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.8 Explicit Input with getline . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

5.8.1 Introduction to getline . . . . . . . . . . . . . . . . . . . . . . . 54 5.8.2 Using getline with No Arguments . . . . . . . . . . . . . 54 5.8.3 Using getline Into a Variable . . . . . . . . . . . . . . . . . . 55 5.8.4 Using getline from a File . . . . . . . . . . . . . . . . . . . . . 56 5.8.5 Using getline Into a Variable from a File . . . . . . . 56 5.8.6 Using getline from a Pipe. . . . . . . . . . . . . . . . . . . . . 57 5.8.7 Using getline Into a Variable from a Pipe . . . . . . 58 5.8.8 Summary of getline Variants . . . . . . . . . . . . . . . . . . 58

6 Printing Output . . . . . . . . . . . . . . . . . . . . . . . . . . 61

6.1 The print Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 6.2 Examples of print Statements . . . . . . . . . . . . . . . . . . . . . . . . . . 62 6.3 Output Separators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.4 Controlling Numeric Output with print . . . . . . . . . . . . . . . . . 64 6.5 Using printf Statements for Fancier Printing . . . . . . . . . . . . 64

6.5.1 Introduction to the printf Statement . . . . . . . . . . . 65 6.5.2 Format-Control Letters . . . . . . . . . . . . . . . . . . . . . . . . 65 6.5.3 Modifiers for printf Formats. . . . . . . . . . . . . . . . . . . 66 6.5.4 Examples Using printf . . . . . . . . . . . . . . . . . . . . . . . . 68

6.6 Redirecting Output of print and printf . . . . . . . . . . . . . . . . 70 6.7 Special File Names in gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 6.8 Closing Input and Output Files and Pipes. . . . . . . . . . . . . . . . 74

7 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

7.1 Constant Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.1.1 Numeric and String Constants . . . . . . . . . . . . . . . . . . 77 7.1.2 Regular Expression Constants . . . . . . . . . . . . . . . . . . 77

7.2 Using Regular Expression Constants . . . . . . . . . . . . . . . . . . . . . 78 7.3 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

7.3.1 Using Variables in a Program. . . . . . . . . . . . . . . . . . . 79 7.3.2 Assigning Variables on the Command Line . . . . . . 80

7.4 Conversion of Strings and Numbers . . . . . . . . . . . . . . . . . . . . . . 81 7.5 Arithmetic Operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 7.6 String Concatenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 7.7 Assignment Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7.8 Increment and Decrement Operators . . . . . . . . . . . . . . . . . . . . . 87 7.9 True and False in awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 7.10 Variable Typing and Comparison Expressions . . . . . . . . . . . 88 7.11 Boolean Expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 7.12 Conditional Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 7.13 Function Calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.14 Operator Precedence (How Operators Nest) . . . . . . . . . . . . . 94

8 Patterns and Actions . . . . . . . . . . . . . . . . . . . . . 97

8.1 Pattern Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 8.1.1 Kinds of Patterns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 8.1.2 Regular Expressions as Patterns . . . . . . . . . . . . . . . . 97 8.1.3 Expressions as Patterns . . . . . . . . . . . . . . . . . . . . . . . . 98 8.1.4 Specifying Record Ranges with Patterns . . . . . . . . . 99 8.1.5 The BEGIN and END Special Patterns . . . . . . . . . . . 100

8.1.5.1 Startup and Cleanup Actions . . . . . . . . . 100 8.1.5.2 Input/Output from BEGIN and END Rules

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

8.1.6 The Empty Pattern. . . . . . . . . . . . . . . . . . . . . . . . . . . 102 8.2 Overview of Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

9 Control Statements in Actions. . . . . . . . . . . . 105

9.1 The if-else Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 9.2 The while Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 9.3 The do-while Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 9.4 The for Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 9.5 The break Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 9.6 The continue Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 9.7 The next Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 9.8 The nextfile Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 9.9 The exit Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 vi Effective AWK Programming

10 Built-in Variables. . . . . . . . . . . . . . . . . . . . . . . 115

10.1 Built-in Variables that Control awk . . . . . . . . . . . . . . . . . . . . 115 10.2 Built-in Variables that Convey Information. . . . . . . . . . . . . 117 10.3 Using ARGC and ARGV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

11 Arrays in awk . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

11.1 Introduction to Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 11.2 Referring to an Array Element . . . . . . . . . . . . . . . . . . . . . . . . 124 11.3 Assigning Array Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 11.4 Basic Array Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 11.5 Scanning All Elements of an Array . . . . . . . . . . . . . . . . . . . . 127 11.6 The delete Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 11.7 Using Numbers to Subscript Arrays . . . . . . . . . . . . . . . . . . . 129 11.8 Using Uninitialized Variables as Subscripts . . . . . . . . . . . . . 129 11.9 Multi-dimensional Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 11.10 Scanning Multi-dimensional Arrays . . . . . . . . . . . . . . . . . . . 132

12 Built-in Functions . . . . . . . . . . . . . . . . . . . . . . 133

12.1 Calling Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 12.2 Numeric Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 12.3 Built-in Functions for String Manipulation . . . . . . . . . . . . . 135 12.4 Built-in Functions for Input/Output . . . . . . . . . . . . . . . . . . . 144 12.5 Functions for Dealing with Time Stamps . . . . . . . . . . . . . . . 146

13 User-defined Functions. . . . . . . . . . . . . . . . . . 151

13.1 Function Definition Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 13.2 Function Definition Examples . . . . . . . . . . . . . . . . . . . . . . . . . 152 13.3 Calling User-defined Functions . . . . . . . . . . . . . . . . . . . . . . . . 154 13.4 The return Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

14 Running awk. . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

14.1 Command Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 14.2 Other Command Line Arguments . . . . . . . . . . . . . . . . . . . . . 163 14.3 The AWKPATH Environment Variable . . . . . . . . . . . . . . . . . . . 164 14.4 Obsolete Options and/or Features . . . . . . . . . . . . . . . . . . . . . 165 14.5 Undocumented Options and Features . . . . . . . . . . . . . . . . . . 165 14.6 Known Bugs in gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

15 A Library of awk Functions . . . . . . . . . . . . . . 167

15.1 Simulating gawk-specific Features. . . . . . . . . . . . . . . . . . . . . . 167 15.2 Implementing nextfile as a Function . . . . . . . . . . . . . . . . . 168 15.3 Assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 15.4 Rounding Numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 15.5 Translating Between Characters and Numbers . . . . . . . . . . 172 15.6 Merging an Array Into a String . . . . . . . . . . . . . . . . . . . . . . . 174 15.7 Turning Dates Into Timestamps . . . . . . . . . . . . . . . . . . . . . . . 175 15.8 Managing the Time of Day . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 15.9 Noting Data File Boundaries . . . . . . . . . . . . . . . . . . . . . . . . . . 183 15.10 Processing Command Line Options . . . . . . . . . . . . . . . . . . . 184 15.11 Reading the User Database . . . . . . . . . . . . . . . . . . . . . . . . . . 190 15.12 Reading the Group Database . . . . . . . . . . . . . . . . . . . . . . . . 195 15.13 Naming Library Function Global Variables . . . . . . . . . . . . 200

16 Practical awk Programs . . . . . . . . . . . . . . . . . 203

16.1 Re-inventing Wheels for Fun and Profit . . . . . . . . . . . . . . . . 203 16.1.1 Cutting Out Fields and Columns . . . . . . . . . . . . . 203 16.1.2 Searching for Regular Expressions in Files . . . . . 208 16.1.3 Printing Out User Information . . . . . . . . . . . . . . . 213 16.1.4 Splitting a Large File Into Pieces . . . . . . . . . . . . . 215 16.1.5 Duplicating Output Into Multiple Files. . . . . . . . 217 16.1.6 Printing Non-duplicated Lines of Text. . . . . . . . . 218 16.1.7 Counting Things . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

16.2 A Grab Bag of awk Programs . . . . . . . . . . . . . . . . . . . . . . . . . 225 16.2.1 Finding Duplicated Words in a Document . . . . . 226 16.2.2 An Alarm Clock Program . . . . . . . . . . . . . . . . . . . . 226 16.2.3 Transliterating Characters. . . . . . . . . . . . . . . . . . . . 229 16.2.4 Printing Mailing Labels . . . . . . . . . . . . . . . . . . . . . . 231 16.2.5 Generating Word Usage Counts. . . . . . . . . . . . . . . 233 16.2.6 Removing Duplicates from Unsorted Text . . . . . 236 16.2.7 Extracting Programs from Texinfo Source Files

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 16.2.8 A Simple Stream Editor. . . . . . . . . . . . . . . . . . . . . . 240 16.2.9 An Easy Way to Use Library Functions . . . . . . . 242

17 The Evolution of the awk Language . . . . . . 251

17.1 Major Changes between V7 and SVR3.1 . . . . . . . . . . . . . . . 251 17.2 Changes between SVR3.1 and SVR4 . . . . . . . . . . . . . . . . . . . 252 17.3 Changes between SVR4 and POSIX awk . . . . . . . . . . . . . . . 253 17.4 Extensions in the Bell Laboratories awk . . . . . . . . . . . . . . . . 253 17.5 Extensions in gawk Not in POSIX awk . . . . . . . . . . . . . . . . . 254 viii Effective AWK Programming

Appendix A gawk Summary . . . . . . . . . . . . . . . . 257

A.1 Command Line Options Summary . . . . . . . . . . . . . . . . . . . . . 257 A.2 Language Summary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 A.3 Variables and Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259

A.3.1 Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 A.3.2 Built-in Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 A.3.3 Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 A.3.4 Data Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262

A.4 Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 A.4.1 Pattern Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 A.4.2 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . 265

A.5 Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 A.5.1 Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 A.5.2 Control Statements . . . . . . . . . . . . . . . . . . . . . . . . . . 268 A.5.3 I/O Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 A.5.4 printf Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 A.5.5 Special File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 A.5.6 Built-in Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 A.5.7 Time Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 A.5.8 String Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274

A.6 User-defined Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 A.7 Historical Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275

Appendix B Installing gawk. . . . . . . . . . . . . . . . . 277

B.1 The gawk Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 B.1.1 Getting the gawk Distribution . . . . . . . . . . . . . . . . . 277 B.1.2 Extracting the Distribution . . . . . . . . . . . . . . . . . . . 279 B.1.3 Contents of the gawk Distribution . . . . . . . . . . . . . 279

B.2 Compiling and Installing gawk on Unix . . . . . . . . . . . . . . . . . 282 B.2.1 Compiling gawk for Unix. . . . . . . . . . . . . . . . . . . . . . 282 B.2.2 The Configuration Process . . . . . . . . . . . . . . . . . . . . 283

B.3 How to Compile and Install gawk on VMS . . . . . . . . . . . . . . 283 B.3.1 Compiling gawk on VMS. . . . . . . . . . . . . . . . . . . . . . 283 B.3.2 Installing gawk on VMS . . . . . . . . . . . . . . . . . . . . . . 284 B.3.3 Running gawk on VMS . . . . . . . . . . . . . . . . . . . . . . . 285 B.3.4 Building and Using gawk on VMS POSIX . . . . . . 285

B.4 MS-DOS and OS/2 Installation and Compilation . . . . . . . . 286 B.5 Installing gawk on the Atari ST . . . . . . . . . . . . . . . . . . . . . . . . 287

B.5.1 Compiling gawk on the Atari ST . . . . . . . . . . . . . . 287 B.5.2 Running gawk on the Atari ST . . . . . . . . . . . . . . . . 288

B.6 Installing gawk on an Amiga . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 B.7 Reporting Problems and Bugs . . . . . . . . . . . . . . . . . . . . . . . . . 290 B.8 Other Freely Available awk Implementations . . . . . . . . . . . . 291

Appendix C Implementation Notes . . . . . . . . . 293

C.1 Downward Compatibility and Debugging . . . . . . . . . . . . . . . 293 C.2 Making Additions to gawk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

C.2.1 Adding New Features. . . . . . . . . . . . . . . . . . . . . . . . . 293 C.2.2 Porting gawk to a New Operating System . . . . . . 295

C.3 Probable Future Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 C.4 Suggestions for Improvements . . . . . . . . . . . . . . . . . . . . . . . . . 298

Appendix D Glossary . . . . . . . . . . . . . . . . . . . . . . 299

GNU GENERAL PUBLIC LICENSE . . . . . . . . 307

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION

AND MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 How to Apply These Terms to Your New Programs . . . . . . . . . . . 313

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

x Effective AWK Programming